### v1

# Impact of the carbon price on credit portfolio’s loss with stochastic collateral

Abstract
In this work, we propose models to quantify the distortion a credit portfolio (expected and unexpected) losses, when the obligor companies as well as their guarantees belong to an economy subject to the climate transition. The economy is driven by its productivity which is a multidimensional Ornstein-Uhlenbeck process while the climate transition is declined thanks to the carbon price, a continuous deterministic process. We define each loan’s loss at default as the difference between Exposure at Default (EAD) and the liquidated collateral, which will help us to define the Loss Given Default (LGD) – the expected percentage of exposure that is lost if a debtor defaults. We consider two types of collateral: financial asset such as invoices, cash, or investments or physical asset such as real estate, business equipment, or inventory. First, if it is a financial asset, we model the later by the continuous time version of the discounted cash flows methodology, where the cash flows SDE is driven by the instantaneous output growth, the instantaneous growth of a carbon price function, and an arithmetic Brownian motion. Secondly, for physical asset, we focus on the example of a property in housing market. Therefore, we define, as Sopgoui, (2024), its value as the difference between the price of an equivalent efficient building following an exponential Ornstein-Uhlenbeck as well as the actualized renovation costs and the actualized sum of the future additional energy costs due to the inefficiency of the building, before an optimal renovation date which depends on the carbon price process. Finally, we obtain how the loss’ risk measures of a credit portfolio are skewed in the context of climate transition through carbon price and/or energy performance of buildings when both the obligors and their guarantees are affected. This work provides a methodology to calculate the (statistics of the) loss of a portfolio of secured loans, starting from a given climate transition scenario described by a carbon price.
keywords:
Credit risk , Climate risk , Collateral , Stochastic modelling , Transition risk , Carbon price , Firm valuation , Loss Given Default
journal: arXiv
This research is part of the PhD thesis in Mathematical Finance of Lionel Sopgoui whose works are funded by a CIFRE grant from BPCE S.A. The opinions expressed in this research are those of the authors and are not meant to represent the opinions or official positions of BPCE S.A.
We would like to thank Jean-François Chassagneux, Antoine Jacquier, Smail Ibbou, and Géraldine Bouveret for helpful comments on an earlier version of this work.
Notations
  • 1.
    $\mathbb{N}$ is the set of non-negative integers, :=$\mathbb{N}^{*}:=\mathbb{N}\setminus\{0\}$, and $\mathbb{Z}$ is the set of integers.
  • 2.
    $\mathbb{R}^{d}$ denotes the $d$-dimensional Euclidean space, $\mathbb{R}_{+}$ is the set of non-negative real numbers, +:=+$\mathbb{R}_{+}^{*}:=\mathbb{R}_{+}\setminus\{0\}$.
  • 3.
    $\mathbf{1}:=(1,\ldots,1)\in\mathbb{R}^{I}$.
  • 4.
    $\mathbb{R}^{n\times d}$ is the set of real-valued $n\times d$ matrices ($\mathbb{R}^{n\times 1}=\mathbb{R}^{n}$), $\mathbf{I}_{n}$ is the identity $n\times n$ matrix.
  • 5.
    $x^{i}$ denotes the $i$-th component of the vector $x\in\mathbb{R}^{d}$. For all $A:=(A^{ij})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$, we denote by $A^{\top}:=(A^{ji})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$ the transpose matrix, and $\lambda(A)$ denotes the spectrum of $A$.
  • 6.
    $\bigotimes$ is the Kronecker product.
  • 7.
    For a given finite set $S$, we define as the cardinal of $S$, #$\#S$.
  • 8.
    For all $x,y\in\mathbb{R}^{d}$, we denote the scalar product $x^{\top}y$, the Euclidean norm $|x|:=\sqrt{x^{\top}x}$ and for a matrix $M\in\mathbb{R}^{d\times d}$, we denote
    $|M|:=\sup_{a\in\mathbb{R}^{d},|a|\leq 1}|Ma|.$
  • 9.
    $(\Omega,\mathcal{H},\mathbb{P})$ is a complete probability space.
  • 10.
    For $p\in[1,\infty]$, ${E}$ is a finite dimensional Euclidian vector space and for a $\sigma$-field $\mathcal{H}$, $\mathcal{L}^{p}(\mathcal{H},{E})$, denotes the set of $\mathcal{H}$-meassurable random variable $X$ with values in ${E}$ such that $\|X\|_{p}:=\left(\mathbb{E}\left[|X|^{p}\right]\right)^{\frac{1}{p}}<\infty$ for $p<\infty$ and for $p=\infty$, $\|X\|_{\infty}:=\mathrm{esssup}|X(\omega)|<\infty$.
  • 11.
    For a filtration $\mathbb{G}$, $p\in[1,+\infty]$ and $I\in\mathbb{N}^{*}$, +$\mathscr{L}^{p}_{+}(\mathbb{G},(0,\infty)^{I})$ is the set of continuous-time processes that are $\mathbb{G}$-adapted valued in $(0,\infty)^{I}$ and which satisfy
    $\lVert X_{t}\rVert_{p}<\infty\text{ for all }t\in\mathbb{R}_{+}.$
  • 12.
    If $X$ and $Y$ are two random variables $\mathbb{R}^{d}$-valued, for $x\in\mathbb{R}^{d}$, we note $Y|X=x$ the conditional distribution of $Y$ given $X=x$, and $Y|\mathcal{F}$ the conditional distribution of $Y$ given the filtration $\mathcal{F}$.
  • 13.
    If $f:\mathbb{R}\to\mathbb{R},t\mapsto f(t)$ is a differentiable function, we note $\dot{f}$ its first derivative.
Introduction
When an obligor (firm, government, or individual) defaults, the creditor (bank) stands to lose its money. One way to ensure the stability of the banks’ business and more generally the soundness of the whole financial system, is ideally, to anticipate when the default will happen and how much a bank could lose. In order to achieve that, the Basel Committee on Banking Supervision, (2017) introduces four parameters: the probability of default (PD) which measures the default risk associated with each borrower, the exposure at default (EAD) which quantifies the outstanding debt at the time of default, the loss given default (LGD) which captures the expected percentage of EAD that is lost if the debtor defaults, and the effective maturity T which represents the duration of the credit. By using, among others, these parameters, banks can compute various risk measures (such as Expected, Unexpected and Stressed Losses) which help later on to determine provisions, as well as economic and regulatory capital. An essential part of a bank risk division is to estimate how these risk measures change with various factors such as time and economic conditions.
Let us focus for example on LGD. When a debtor defaults, banks can lose part or whole of its exposure. The fraction of the loss relative to EAD is LGD while the recovery rate is the fraction of EAD recovered so that $LGD=1-Recovery$. So modelling LGD or modelling recovery are equivalent. According to Chalupka and Kopecsni, (2008), there are three ways to handle LGD: ”Market LGD is observed from market prices of defaulted bonds or marketable loans soon after the actual default event. Workout LGD is derived from a set of estimated cash flows resulting from a workout and collection process, properly discounted to a date of default. Thirdly, implied market LGD is derived from risky but not defaulted bond prices using a theoretical asset pricing model”. In the IRB approach, LGD refers to Workout LGD and there are several techniques to model it. In economic modeling, as detailed by Bastos, (2010); Roncalli, (2020), LGD is a (linear or non linear) function of many factors which can be factors external to the issuer, specific to the issuer or specific to the debt issuance. That function can be obtained/calibrated through logistic regression, regression trees, or neural networks. In stochastic modeling, it is assumed that LGD follows a given distribution (parametric or non-parametric). In this case, LGD is commonly modeled by a Beta distribution as Roncalli, (2020)[Page 193] and Chalupka and Kopecsni, (2008). Its parameter are then estimated on historical data. Fermanian, (2020), for his part, proposes a joint modelling of PD and LGD by writing the potential loss at default as the difference between the debt amount (EAD) and the assets at the default date.
There are secured and unsecured loans. In these approaches, even for secured loans, there is one parameter essential that is sometimes overlooked: the collateral. However, not all borrowers put up collateral when taking out loans. It is even worse, there is even some evidence that loans with collateral attached may be riskier for lenders (see Berger and Udell, (1990)). If the loan is secured, when the counterpart defaults, the bank liquidates the collateral (guarantee), and if the EAD is not reached, it can recover the remainder by liquidating other assets (called residual recovery). These guarantees can be tangible assets (buildings, business equipments, inventories, etc.) or intangible assets (cash deposits, public bonds, securities, etc.) as noted by Berger and Udell, (1990), Blazy and Weill, (2013). In the presence of collateral, the recovery (i.e. $1-\mathrm{LGD}$) is therefore made up of the value of the collateral at the date of default and the value of the residual recovery Frontczak and Rostek, (2015), and Pelizza and Schenk-Hoppé, (2020). We will model here two examples of guarantees: either a security or a (commercial or residential) building, which both, will be affected by the climate transition at its liquidation.
A security can represent ownership in a corporation in the form of stock, a creditor relationship with a governmental body or a corporation represented by owning that entity’s bond; or rights to ownership as represented by an option. A security generate a stream of cash flows. The proxy of the security value is the infinite sum of the present value of the future cash flows. As we already know from the literature and from what we propose in Bouveret et al., (2023), the security (notably if it is a firm) value will be affected by transition risk. We will therefore revisit the results of Bouveret et al., (2023) where carbon price dynamics affect the firm value and credit risk measures such as probability of default, expected and unexpected losses. In particular, we redesign the multisectoral model with carbon price, the firm valuation model, and the credit risk model proposed in continuous time.
In the same way, a commercial or residential building price will be affected by transition for example through the Energy Performance (or Energy Efficiency) as mentioned in Aydin et al., (2020); Franke and Nadler, (2019). Ter Steege and Vogel, (2021) quantify the depreciation by writing the price difference per square meter between two properties with different energy efficiency as the sum of the discounted value of (expected) energy cost differences. Sopgoui, (2024) enhance this work by assuming that initially, the building’s owner incurs additional energy costs (writing as a function of the carbon price) due to the inefficiency of their property. Then, he may decide to spend money on renovations to make their building energy efficient. After the renovation, they no longer incur additional energy costs. Therefore, he obtains the value of a building as the difference between the price of an equivalent efficient building following an exponential Ornstein-Uhlenbeck as well as the actualized renovation costs and the actualized sum of the future additional energy costs.
The rest of the present work is organized as follows. We revisit in Section 2 the results of Bouveret et al., (2023) in a continuous time setting, namely a multisectoral economic model with carbon price, a firm valuation model, and a credit risk model. In section 3, we define the loss at default as the difference between EAD and the liquidated collateral, which will help us to define LGD. If the collateral is a financial asset, we model it in Subsection 3.2 by the continuous time version of the discounted cash flows, where the cash flows SDE is driven by the instantaneous consumption growth, the instantaneous growth of a carbon price’s function and a Brownian motion. If the collateral is a building, we will use the housing valuation under climate transition proposed in Sopgoui, (2024) to compute the portfolio’s loss. The last section is dedicated to estimations, simulations, and discussion. Our simulations show that (1) expected and unexpected losses increase when the price of carbon rises, (2) the presence of collateral significantly reduces expected and unexpected losses, (3) the positive effect of collateral on losses is reduced if the collateral is energy inefficient (for a building) or depends on a polluting sector (financial asset).
1 The problem
We consider a bank credit portfolio composed of $N\in\mathbb{N}^{*}$ firms in a closed economy (in other words no import and no export). In credit risk assessment, one of the first steps is to create homogeneous sub-portfolios of firms. As we are dealing here with climate transition risk, we would like to classify firms by carbon intensity: firms with similar carbon intensities belong to a same homogeneous sub-portfolio. It should be noted that in the absence of a climate transition, firms are traditionally clustered in terms of industry, geography, size, and credit rating, for example.
We thus assume $I\in\mathbb{N}^{*}$ ($I\leq N$) homogeneous carbon emission sectors in the economy. Nevertheless, as we rarely have the firm individual carbon emissions/intensities, we assume that each company has the carbon intensity of its industry sector. This amounts to grouping ”industry sectors” into $I$ ”carbon emission sectors”. From now on, sectors are to be interpreted as carbon emission sectors.
Definition 1.1.
We divide our portfolio into $I$ disjunct sub-portfolios $g_{1},\ldots,g_{I}$ so that each sub-portfolio represents a single risk class and the firms in each sub-portfolio belong to a single carbon emission sectors. From now on, we denote $\mathcal{I}$ the set of sectors with cardinal $I\in\mathbb{N}^{*}$. We also fix $n_{i}:=\min{\{n\in\{1,\ldots,N\}\text{ such that }n\in g_{i}\}}$ for each $i\in\mathcal{I}$. Therefore, firm $n_{i}$ is a representative of the group $i$.
We would like to know how the whole portfolio loss and sub-portfolios losses would be affected should the regulator introduce a carbon price in the economy, in order to mitigate the effects of climate change. This precisely amounts to quantifying the distortion over time of credit risk measures created by the introduction of a carbon price. For example, if the government decides to charge firms and households GHG emissions between 2025 and 2035, a bank would like to estimate today how the probability of a company to default in 2030 is impacted.
The bank’s potential loss caused by a firm depends essentially on the default date and on the liquidation of the guarantees if they exist. The firm as well as the guarantee belong to the same economy subject to the climate transition. Thus, we build in the first stage a dynamic, stochastic, and multisectoral economic model in which direct and indirect GHG emissions from companies as well as direct GHG emissions from households are charged. We choose a representative firm in each sector and a representative household for the whole economy. By observing that each firm belongs to a sector and its cash flows are a proportion of its sales. The latter are themselves a proportion of the sectoral output, we obtain the cash flows dynamics that we use to model the value of firms in an environment where GHG emissions are charged. Then, starting from a default model in which a company defaults if its value falls below its debt, we calculate the probability of default of each firm. Finally, we compute the distortion of the (associated statistics of) loss – defined as the difference between the exposure and the liquidated collateral if exists – by the introduction of a carbon price.
2 Main assumptions and results of Bouveret et al., (2023) in continuous time
In this section, we revisit the framework developed in Bouveret et al., (2023) in continuous time. Precisely, we decline, in continuous time, the two standing assumptions as well as the three main results respectively on the dynamic stochastic multisectoral model with carbon emissions costs, on the firm valuation model, and on the structural credit risk model. Most of the proofs can be derived from the discrete time so that we will either skip them or detail them in appendix.
2.1 A Multisectoral Model with Carbon price
Each sector $i\in\mathcal{I}$ has a representative firm which produces a single good, so that we can associate sector, firm and good. We introduce the following standing assumption which describes the productivity, which is considered to have stationary Ornstein-Uhlenbeck dynamics.
Standing Assumption 2.1.
We define the $\mathbb{R}^{I}$-valued process $\mathcal{A}$ which evolves according to
$\left\{\begin{array}[]{rl}\mathrm{d}\mathcal{Z}_{t}&=-\Gamma\mathcal{Z}_{t}\mathrm{d}t+\Sigma\mathrm{d}B_{t}^{\mathcal{Z}}\\ \mathrm{d}\mathcal{A}_{t}&=\left(\mu+\varsigma\mathcal{Z}_{t}\right)\mathrm{d}t\end{array}\quad\textrm{for all }t\in\mathbb{R}_{+},\right.$ (2.1)
where $(B_{t}^{\mathcal{Z}})_{t\in\mathbb{R}^{*}}$ is a $I$-dimensional Brownian Motion, and where the constants $\mu,\mathcal{A}_{0}\in\mathbb{R}^{I}$, the matrices $\Gamma,\Sigma\in\mathbb{R}^{I\times I}$, $\mathcal{Z}_{0}\sim\mathcal{N}\left(0,\Sigma\Sigma^{\top}\right)$, and $0<\varsigma\leq 1$ is an intensity of noise parameter that is fixed: it will be used later to obtain a tractable proxy of the firm value. Moreover, $\Sigma$ is a positive definite matrix and $-\Gamma$ is a Hurwitz matrix i.e. its eigenvalues have strictly negative real parts.
The processes $\mathcal{Z}^{i}$ and $\mathcal{A}^{i}$ play a major role in our factor productivity model since, for any $i\in\mathcal{I}$, the total factor productivity of sector $i$ is defined as
$A^{i}:=\exp{(\mathcal{A}^{i})},$ (2.2)
so that $\mathcal{Z}^{i}$ is the log-productivity growth and $\mathcal{A}^{i}$ is the cumulative log-productivity growth. In the rest of the paper, the terminology ”productivity” will be used within a context that will allow the reader to understand if the term refers to $\mathcal{Z}^{i}$, $\mathcal{A}^{i}$, or $A^{i}$.
We also introduce the following filtration $\mathbb{G}:=(\mathcal{G}_{t})_{t\in\mathbb{R}^{*}}$ with $\mathcal{G}_{0}:=\sigma(\mathcal{Z}_{0})$ and for $t>0$, $\mathcal{G}_{t}:=\sigma\left(\left\{\mathcal{Z}_{0},B_{s}^{\mathcal{Z}}:s\leq t\right\}\right)$.
Remark 2.2 (O.U. process).
We have the following results on O.U. that we will use later:
  1. 1.
    According to Gobet and She, (2016)[Proposition 1], if one assumes that $\mathcal{Z}_{0}$ and $B^{\mathcal{Z}}$ are independent and $\mathcal{Z}_{0}$ is square integrable, then, there exists a unique square integrable solution to the $I$-dimentional Ornstein-Uhlenbeck process $\mathcal{Z}$ satisfying $\mathrm{d}\mathcal{Z}_{t}=-\Gamma\mathcal{Z}_{t}\mathrm{d}t+\Sigma\mathrm{d}B_{t}^{\mathcal{Z}}$, represented as
    $\mathcal{Z}_{t}=e^{-\Gamma t}\left(\mathcal{Z}_{0}+\int_{0}^{t}e^{\Gamma u}\Sigma\mathrm{d}B_{u}^{\mathcal{Z}}\right),\quad\textrm{for all }t\in\mathbb{R}_{+}.$
    Additionally, for any $t,h\geq 0$, the distribution of $\mathcal{Z}_{t+h}$ conditional on $\mathcal{G}_{t}$ is Gaussian $\mathcal{N}\left(M^{\mathcal{Z},h}_{t},\Sigma^{\mathcal{Z},h}_{t}\right)$, with the mean vector
    $M^{\mathcal{Z},h}_{t}:=\mathbb{E}[\mathcal{Z}_{t+h}|\mathcal{G}_{t}]=e^{-\Gamma h}\mathcal{Z}_{t},$ (2.3)
    and the covariance matrix
    $\Sigma^{\mathcal{Z},h}_{t}:=\mathbb{V}[\mathcal{Z}_{t+h}|\mathcal{G}_{t}]=\int_{0}^{h}e^{-\Gamma u}\Sigma\Sigma^{\top}e^{-\Gamma^{\top}u}\mathrm{d}u.$ (2.4)
  2. 2.
    Since $-\Gamma$ is a Hurwitz matrix, then if we note $\lambda_{\Gamma}:=\max_{\lambda\in\lambda(\Gamma)}Re(\lambda)$, there exists $c_{\Gamma}>0$ so that $\lVert e^{-\Gamma t}\rVert<c_{\Gamma}e^{-\lambda_{\Gamma}t}$ for all $t\geq 0$. Therefore, according to Gobet and She, (2016)[Proposition 2], $\mathcal{Z}$ has a unique stationary distribution which is Gaussian with mean 00 and covariance $\int_{0}^{+\infty}e^{-\Gamma u}\Sigma\Sigma^{\top}e^{-\Gamma^{\top}u}\mathrm{d}u$.
  3. 3.
    We show in Sopgoui, (2024)[Appendix A] that for any $t,h\geq 0$, we have
    $\mathcal{A}_{t+h}=\mathcal{A}_{t}+\int_{t}^{t+h}(\mu+\varsigma\mathcal{Z}_{s})\mathrm{d}s=\mu h+\varsigma\int_{t}^{t+h}\mathcal{Z}_{s}\mathrm{d}s,$
    and conditionally on $\mathcal{G}_{t}$, $\mathcal{A}_{t+h}$ has an $I$-dimensional normal distribution with the mean vector
    $M^{\mathcal{A},h}_{t}:=\mu h+\varsigma\Upsilon_{h}\mathcal{Z}_{t}+\mathcal{A}_{t},$ (2.5)
    with
    $\Upsilon_{h}:=\int_{0}^{h}e^{-\Gamma s}\mathrm{d}s=\Gamma^{-1}(\mathbf{I}_{I}-e^{-\Gamma h}),$ (2.6)
    and the covariance matrix
    $\Sigma^{\mathcal{A},h}_{t}:=\varsigma^{2}\Gamma^{-1}\left(\int_{0}^{h}\left(e^{-\Gamma u}-\mathbf{I}_{I}\right)\Sigma\Sigma^{\top}\left(e^{-\Gamma u}-\mathbf{I}_{I}\right)\mathrm{d}u\right)(\Gamma^{-1})^{\top}=\varsigma^{2}\int_{0}^{h}\Upsilon_{u}\Sigma\Sigma^{\top}\Upsilon_{u}^{\top}\mathrm{d}u.$ (2.7)
  4. 4.
    For later use, we define
    $\mathcal{A}^{\circ}_{t}:=\mathcal{A}_{t}-\mathcal{A}_{0},$ (2.8)
    and observe that $(\mathcal{A}^{\circ}_{t},\mathcal{Z}_{t})_{t\geq 0}$ is a Markov process.
Firms emit GHG when they consume intermediary input from other sectors and when they produce output. Likewise, households emit GHG when they consume. All these emissions are charged through a carbon price dynamics. For the whole economy, we introduce a deterministic and exogenous carbon price in euro/dollar per ton. It allows us to model the impact of the transition pathways on the whole economy. We will note $\delta$ the complete carbon price process. We shall then assume the following setting.
Standing Assumption 2.3.
We introduce the carbon price and the carbon intensities (the quantity of GHG in tons emits for each unit of production/consumption) processes:
  1. 1.
    Let $0\leq t_{\circ}<t_{\star}$ be given. The sequence $\delta$ satisfies
    • (a)
      for $t\in[0;t_{\circ}]$, $\delta_{t}=\delta_{0}\in(\mathbb{R}_{+})^{I}$, namely the carbon price is constant;
    • (b)
      for $t\in(t_{\circ},t_{\star})$, $\delta_{t}\in(\mathbb{R}_{+})^{I}$, the carbon price may evolve;
    • (c)
      for $t\geq t_{\star}$, $\delta_{t}=\delta_{t_{\star}}\in(\mathbb{R}_{+})^{I}$, namely the carbon price is constant.
    We assume moreover that $t\mapsto\delta_{t}$ is $\mathcal{C}^{1}(\mathbb{R}_{+},\mathbb{R}_{+})$.
  2. 2.
    We also introduce carbon intensities as the sequences $\tau$, $\zeta$, and $\kappa$ being respectively $\mathbb{R}_{+}^{I}$, $\mathbb{R}_{+}^{I\times I}$, and $\mathbb{R}_{+}^{I}$-processes, and representing respectively carbon intensities on firm’s output, on firm’s intermediary consumption, and on household’s consumption, and satisfying for all $t\in\mathbb{R}_{+}$ and
    $\mathfrak{y}\in\{\tau^{1},\ldots,\tau^{I},\zeta^{11},\zeta^{12},\ldots,\zeta^{II-1},\zeta^{II},\kappa^{1},\ldots,\kappa^{I}\}$,
    $\mathfrak{y}_{t}=\left\{\begin{array}[]{ll}&\mathfrak{y}_{0}\exp{\left(g_{\mathfrak{y},0}\frac{1-\exp{(-\theta_{\mathfrak{y}}t)}}{\theta_{\mathfrak{y}}}\right)}\qquad\text{if }0\leq t\leq t_{\star}\\ &\mathfrak{y}_{0}\exp{\left(g_{\mathfrak{y},0}\frac{1-\exp{(-\theta_{\mathfrak{y}}t_{\star})}}{\theta_{\mathfrak{y}}}\right)}\qquad\text{else},\end{array}\right.$ (2.9)
    with $\mathfrak{y}_{0},g_{\mathfrak{y},0},\theta_{\mathfrak{y}}>0$. For each $t\geq 0$, we call $\mathfrak{y}_{t}\delta_{t}$ the emissions cost rate at time $t$.
  3. 3.
    For each $i\in\mathcal{I}$ and for each $t\in\mathbb{R}_{+}$,
    $\delta_{t}\max_{i\in\mathcal{I}}{\tau^{i}_{0}}<1.$ (2.10)
In the following, we will note for all $t\geq 0$,
$\mathfrak{d}_{t}:=(\tau_{t}\delta_{t},\zeta_{t}\delta_{t},\kappa_{t}\delta_{t}).$ (2.11)
An example of carbon price process
We assume the regulator fixes $t_{\circ}\geq 0$ when the transition starts and the transition horizon time $t_{\star}>t_{\circ}$, the carbon price at the beginning of the transition $P_{carbon}>0$, at the end of the transition $\delta_{t_{\star}}>P_{carbon}$, and the annual growth rate $\eta_{\delta}>0$. Then, for all $t\geq 0$,
$\delta_{t}=\left\{\begin{array}[]{ll}\displaystyle P_{carbon},&\mbox{if }t\leq t_{\circ},\\ \displaystyle P_{carbon}e^{\eta_{\delta}(t-t_{\circ})},&\mbox{if }t\in(t_{\circ},t_{\star}],\\ \displaystyle\delta_{t_{\star}}=P_{carbon}e^{\eta_{\delta}(t_{\star}-t_{\circ})},&\mbox{otherwise}.\end{array}\right.$ (2.12)
In the example above that will be used in the rest of this work, we assume that the carbon price increases. However, there are several scenarios that could be considered, including a carbon price that would increase until a certain year before leveling off or even decreasing. We also assume an unique carbon price for the entire economy whereas we could proceed differently. For example, the carbon price could increase for production when stabilize or disappear on households in order to avoid social movements and so on. The framework can be adapted to various sectors as well as scenarios.
In our framework, a representative firm in each sector which maximizes its profits by choosing, at each time and for a given productivity, the quantities of labor and intermediary inputs, while, a representative household solves a dynamic optimization problem to decide how to allocate its consumption expenditures among the different goods and hours worked and among the different sectors. We assume that the utility function $U(x,y):=\log{x}-\frac{y^{1+\varphi}}{1+\varphi}$ with $\varphi\geq 0$. Moreover, $\boldsymbol{\lambda}$ (respectively $\psi$) are matrix in $(\mathbb{R}_{+}^{*})^{I\times I}$ (respectively vector in $(\mathbb{R}_{+}^{*})^{I}$) of the elasticities of intermediary inputs (respectively labor). We also assume a constant return to scale, namely
$\psi^{i}+\sum_{j\in\mathcal{I}}\boldsymbol{\lambda}^{ji}=1,\qquad\text{for each }i\in\mathcal{I}.$ (2.13)
Since the productivity and the carbon price processes are continuous, the firms and households problems are well posed and their solutions exist. More details are given in B. In the following proposition, we give the expression of the output.
Proposition 2.4.
For $(\overline{\tau},\overline{\zeta},\overline{\kappa},\overline{\delta})\in\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}^{I\times I}\times\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}$, let us note
$\Psi(\overline{\mathfrak{d}}):=\left(\psi^{i}\frac{1-\overline{\tau}^{i}\overline{\delta}}{1+\overline{\kappa}^{i}\overline{\delta}}\right)_{i\in\mathcal{I}}\ \quad\text{and}\quad\Lambda(\overline{\mathfrak{d}}):=\left(\boldsymbol{\lambda}^{ji}\frac{1-\overline{\tau}^{i}\overline{\delta}}{1+\overline{\zeta}_{t}^{ji}\overline{\delta}}\frac{1+\overline{\kappa}^{j}\overline{\delta}}{1+\overline{\kappa}^{i}\overline{\delta}}\right)_{j,i\in\mathcal{I}},$ (2.14)
with $\overline{\mathfrak{d}}:=(\overline{\tau}\overline{\delta},\overline{\zeta}\overline{\delta},\overline{\kappa}\overline{\delta})$. Assume that
  1. 1.
    $\mathbf{I}_{I}-\boldsymbol{\lambda}$ is not singular,
  2. 2.
    $\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{t})^{\top}$ is not singular for all $t\in\mathbb{R}_{+}$.
Then, for all $t\in\mathbb{R}_{+}$, there exists an unique couple of consumption and output $({C}_{t},{Y}_{t})$ solving the (dynamic stochastic) multisectoral model. Moreover, for all $t\in\mathbb{R}_{+}$.
  1. 1.
    if $\mathfrak{e}_{t}^{i}:=\frac{Y^{i}_{t}}{C^{i}_{t}}$ for $i\in\mathcal{I}$, we have
    $\mathfrak{e}_{t}=\mathfrak{e}(\mathfrak{d}_{t}):=(\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{t})^{\top})^{-1}\mathbf{1},$ (2.15)
  2. 2.
    Using $\mathcal{B}_{t}=(\mathcal{B}^{i}_{t})_{i\in\mathcal{I}}:=\left[\mathcal{A}_{t}^{i}+v^{i}(\mathfrak{d}_{t})\right]_{i\in\mathcal{I}}$ with for $(\overline{\tau},\overline{\zeta},\overline{\kappa},\overline{\delta})\in\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}^{I\times I}\times\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}$,
    $v^{i}(\overline{\mathfrak{d}}):=\log\left((\mathfrak{e}(\overline{\mathfrak{d}})^{i})^{-\frac{\varphi\psi^{i}}{1+\varphi}}\left(\Psi^{i}(\overline{\mathfrak{d}})\right)^{\frac{\psi^{i}}{1+\varphi}}\prod_{j\in\mathcal{I}}\left(\Lambda^{ji}(\overline{\mathfrak{d}})\right)^{\boldsymbol{\lambda}^{ji}}\right)_{i\in\mathcal{I}}+((\mathbf{I}_{I}-\boldsymbol{\lambda})\log{(\mathfrak{e}(\overline{\mathfrak{d}}))}),$ (2.16)
    We obtain
    ${Y}_{t}=\exp{\left((\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}{\mathcal{B}_{t}}\right)}.$ (2.17)
  3. 3.
    Furthermore, since $\mathfrak{d}\in\mathcal{C}^{1}(\mathbb{R}_{+},[0,1)^{I}\times(\mathbb{R}_{+})^{I\times I}\times(\mathbb{R}_{+})^{I})$, we directly have $\Psi(\overline{\mathfrak{d}}_{\cdot}),\Lambda(\overline{\mathfrak{d}}_{\cdot})\in\mathcal{C}^{1}(\mathbb{R}_{+},\mathbb{R})$. Moreover, $\bar{\mathfrak{d}}\mapsto(\mathbf{I}_{I}-{\Lambda}(\bar{\mathfrak{d}})^{\top})^{-1}$ on $\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}^{I\times I}\times\mathbb{R}_{+}^{I}$) is differentiable, then $(\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{\cdot})^{\top})^{-1})\in\mathcal{C}^{1}(\mathbb{R}_{+},\mathbb{R})$.
The output ${Y}$ is also positive, we can then introduce, from the third item, for all $t\geq 0$, $\mathrm{d}\log{({Y}_{t})}$ representing the instantaneous consumption growth. This proposition is an equivalent of Bouveret et al., (2023)[Theorem 1.10.] in continuous time.
2.2 A Firm Valuation Model
Consider a fix $n\in\{1,\ldots,N\}$. For any time $t\in\mathbb{R}_{+}$ and firm $n$, we note $F^{n}_{t}$ the free cash flows of $n$ at $t$, and $r>0$ the discount rate, we introduce the following assumption:
Assumption 2.5.
The $\mathbb{R}$-valued process on the instantaneous growth of the cash flows of firm $n$ denoted by $\mathrm{d}\log{F^{n}_{t,\mathfrak{d}}}$ is linear in the economic factors (output growth by sector), specifically we set for all $t\in\mathbb{R}_{+}$,
$\mathrm{d}\log{F^{n}_{t,\mathfrak{d}}}=\tilde{\mathfrak{a}}^{n\cdot}\mathrm{d}\log{{Y}_{t}}+\sigma_{n}\mathrm{d}\mathcal{W}^{n}_{t}=\mathfrak{a}^{n\cdot}(\mathrm{d}\mathcal{A}_{t}+\mathrm{d}v(\mathfrak{d}_{t}))+\sigma_{n}\mathrm{d}\mathcal{W}^{n}_{t},$ (2.18)
for $\tilde{\mathfrak{a}}^{n\cdot}\in\mathbb{R}^{I}$ and $\mathfrak{a}^{n\cdot}=\tilde{\mathfrak{a}}^{n\cdot}(\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}$, where $(\mathcal{W}^{n}_{t})_{t\in\mathbb{R}_{+}}$ is a $\mathbb{R}^{N}$-Brownian motion with $\sigma_{n}>0$. Moreover, $B^{\mathcal{Z}}$ and $\mathcal{W}^{n}$ are independent.
We define the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ by $\mathcal{F}_{t}=\sigma\left(\mathcal{G}_{t}\cup\sigma\left\{\mathfrak{b}_{s}:s\in[0,t]\right\}\right)$ for $t\geq 0$, and we denote $\mathbb{E}_{t}[\cdot]:=\mathbb{E}[\cdot|\mathcal{F}_{t}]$.
Recall that the economic motivation behind (2.18) comes from the fact that if firm $n$ belongs to sector $i$, then its production is proportional to the sectoral output and its cash flows are proportional to its production (as in the Dechow-Kothari-Watts model in Barth et al., (2001)). Thus, we obtain a relation between the cash flows of firm $n$ and the total output of sector $i$. The assumption $\tilde{\mathfrak{a}}^{n\cdot}\in\mathbb{R}^{I}$ stems from the fact that a company is not restricted to one sector only in general. However, since we are considering the emission sector here, we expect that each firm $n$ only belongs to one sector ($i$ for example). Therefore $\mathfrak{a}^{nj}=0$ for all $i\neq j$ and hence $|\mathfrak{a}^{ni}|=\max_{j\in\mathcal{I}}|\mathfrak{a}^{nj}|$.
Let $r\geq 0$ representing the interest rate, by the continuous form of the discounted cash flows valuation, the value $V^{n}_{t,\mathfrak{d}}$ of the firm $n$, at time $t$, is
$V^{n}_{t,\mathfrak{d}}:=\mathbb{E}_{t}\left[\int_{t}^{+\infty}e^{-r(s-t)}F^{n}_{s,\mathfrak{d}}\mathrm{d}s\right].$ (2.19)
For $n\in\{1,\ldots,N\}$ and $t\geq 0$, describing $V^{n}_{t,\mathfrak{d}}$ as a function of the underlying processes driving the economy does not lead to an easily tractable formula. To facilitate the forthcoming credit risk analysis, when $\varsigma$ (introduced in Standing Assumption 2.1) is closed to 00, we approach $V^{n}_{t,\mathfrak{d}}$ by the quantity
$\mathcal{V}_{t,\mathfrak{d}}^{n}:=F^{n}_{t,\mathfrak{d}}\int_{t}^{+\infty}e^{-r(s-t)}\mathbb{E}_{t}\left[\exp\left((s-t)\mathfrak{a}^{n\cdot}\mu+\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)+\sigma_{n}(\mathcal{W}^{n}_{s}-\mathcal{W}^{n}_{t})\right)\mathrm{d}s\right],$ (2.20)
that we describe as a proxy the firm $n$ value at time $t$. We will work directly with $\mathcal{V}^{n}_{t,\mathfrak{d}}$ instead of $V^{n}_{t,\mathfrak{d}}$. We have the following proposition, whose proof is given in B.4.
Proposition 2.6.
For any $n\in\{1,\ldots,N\}$ and for all $t\in\mathbb{R}_{+}$.
  1. 1.
    Assume that $\varrho_{n}:=\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu-r<0$, then $\mathcal{V}_{t,\mathfrak{d}}^{n}$ is well defined and
    $\mathcal{V}^{n}_{t,\mathfrak{d}}=F^{n}_{0}\mathfrak{R}^{n}_{t}(\mathfrak{d})\exp{\left(\mathfrak{a}^{n\cdot}(\mathcal{A}_{t}-v(\mathfrak{d}_{0}))\right)}\exp\left(\sigma_{n}\mathcal{W}^{n}_{t}\right),$ (2.21)
    where
    $\mathfrak{R}^{n}_{t}(\mathfrak{d}):=\int_{0}^{\infty}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s.$ (2.22)
  2. 2.
    Moreover, with $t_{\circ}$ and $t_{\star}$ defined in Standing Assumption 2.3, we obtain the following explicit form,
    $\mathfrak{R}^{n}_{t}(\mathfrak{d})=\left\{\begin{array}[]{ll}\displaystyle-\frac{e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})}}{\varrho_{n}},&\mbox{if }t\geq t_{\star},\\ \displaystyle\int_{0}^{t_{\star}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s-\frac{e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})+\varrho_{n}(t_{\star}-t)}}{\varrho_{n}},&\mbox{if }0\leq t<t_{\star},.\end{array}\right.$ (2.23)
  3. 3.
    Assume that
    $\rho_{n}:=\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu+\frac{1}{2}\varsigma^{2}\frac{c_{\Gamma}^{2}}{\lambda_{\Gamma}^{2}}\lVert\mathfrak{a}^{n\cdot}\rVert^{2}\lVert\Sigma\rVert^{2}<r.$ (2.24)
    therefore $V_{t,\mathfrak{d}}^{n}$ is well defined and there exists a constant $C$ such that $\mathbb{E}\left[\left|\frac{V^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}-\frac{\mathcal{V}^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}\right|\right]\leq C\varsigma$, for all $\varsigma>0$.
The following corollary gives (conditional) laws of the (proxy) of the firm value $\mathcal{V}_{\cdot,\mathfrak{d}}^{n}$.
Corollary 2.7.
For all $t,T\geq 0$.
  1. 1.
    We note $\mathfrak{m}^{n}(\mathfrak{d},t,\mathcal{A}_{t}^{\circ}):=\log{(F^{n}_{0}\mathfrak{R}^{n}_{t}(\mathfrak{d}))}+\left(\mathfrak{a}^{n\cdot}(\mathcal{A}_{t}^{\circ}-v(\mathfrak{d}_{0}))\right)$ and we have
    $\log{\mathcal{V}_{t,\mathfrak{d}}^{n}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\mathfrak{m}^{n}(\mathfrak{d},t,\mathcal{A}_{t}^{\circ}),t\sigma_{n}^{2}\right).$ (2.25)
  2. 2.
    We note $\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t}):=\log{(F^{n}_{0}\mathfrak{R}^{n}_{t+T}(\mathfrak{d}))}+\mathfrak{a}^{n\cdot}(\mu T+\varsigma\Upsilon_{T}\mathcal{Z}_{t}+\mathcal{A}_{t}^{\circ}-v(\mathfrak{d}_{0}))$ and $\mathcal{L}^{n}(t,T):=\mathfrak{a}^{n\cdot}\Sigma^{\mathcal{A},T}_{t}\mathfrak{a}^{n\cdot}+(t+T)\sigma_{n}^{2}$, we have
    $\log{\mathcal{V}_{t+T,\mathfrak{d}}^{n}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t}),\mathcal{L}^{n}(t,T)\right).$ (2.26)
2.3 A Credit Risk Model without collateral
To conclude this section, we present the probability of default. As Basel Committee on Banking Supervision, (2017), we introduce four credit risk parameters: the probability of default (PD) measures the default risk associated with each borrower, the exposure at default (EAD) measures the outstanding debt at the time of default, the loss given default (LGD) denotes the expected percentage of EAD that is lost if the debtor defaults, and the effective maturity $T$ represents the duration of the credit. With these four parameters, we can compute the portfolio loss $L$, with some assumptions:
Assumption 2.8.
Consider a portfolio of $N\in\mathbb{N}^{*}$ credits. For $1\leq n\leq N$,
  1. (1)
    Firm $n$ has issued two classes of securities: equity and debt.
  2. (2)
    $(\mathrm{EAD}_{t}^{n})_{t\geq 0}$ is a $\mathbb{R}_{+}^{*}$-valued continuous and deterministic process, and for all $t\geq 0$, the family $(\mathrm{EAD}_{t}^{n})_{n=1,\ldots,N}$ is a sequence of positive constants such that
    1. (a)
      $\sum_{n\geq 1}\mathrm{EAD}_{t}^{n}=+\infty$;
    2. (b)
      there exists $\upsilon>0$ such that $\frac{\mathrm{EAD}_{t}^{n}}{\sum_{n=1}^{N}\mathrm{EAD}_{t}^{n}}=\mathcal{O}(N^{-(\frac{1}{2}+\upsilon)})$, as $N$ tends to infinity.
  3. (3)
    $(\mathrm{LGD}_{t}^{n})_{t\geq 0}$ is a $(0,1]$-valued continuous and deterministic process;
  4. (4)
    $(\mathcal{D}^{n}_{t})_{t\geq 0}$ is a $\mathbb{R}_{+}$-valued continuous and deterministic process, representing the debt of firm $n$ at time $t$. We will also denote $D^{n}_{t}:=\displaystyle\frac{\mathcal{D}_{t}^{n}}{\mathbb{E}[F_{t,0}^{n}]}$ representing the debt to cash flows ratio.
  5. (5)
    The value of the firm $n$ at time $t$ is assumed to be a tradable asset given by $V^{n}_{t,\mathfrak{d}}$ defined in (2.21).
According to Kruschwitz and Löffler, (2020), there are two ways to handle the default of a company: for a given financing policy, a levered firm is
  • 1.
    in danger of illiquidity if the cash flows do not suffice to fulfill the creditors’ payment claims (interest and net redemption) as contracted,
  • 2.
    over-indebted if the market value of debt exceeds the firm’s market value.
We follow in the present work the second definition of default proposed: a firm default when it is over-indebted, that is in fact the same approach used in the structural credit risk models. We retain this definition in this continuous setting in the same way as Bourgey et al., (2022). Actually, the term default may even be considered an abuse of language. So, to avoid any confusion in the following, we will use the terms default and over-indebted without distinction. Therefore, the over-indebtedness of entity $n$ occurs at time $t$ when the firm value $\mathcal{V}^{n}_{t,\mathfrak{d}}$ falls below a given barrier $\mathcal{D}_{t}^{n}$, related to the net debt, namely on the event $\left\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\right\}$.
However, it should be noted that in a continuous time setting, it can be interesting to work in the Black and Cox, (1976) model. Here, the default event depends on the trajectory of the firm value process $\mathcal{V}$. Therefore, at a given time $t$, the firm defaults if it has been over-indebted at least one time during the period $[0,t]$, that is $\left\{\exists s\in[0,t]\text{ such that }\mathcal{V}^{n}_{s,\mathfrak{d}}<\mathcal{D}^{n}_{s}\right\}$. Thus, the default time is given by
$\tau^{n}:=\inf\left\{t\geq 0,\quad\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\right\}.$ (2.27)
Then, if we are interested in the probability of the firm $n$ defaulting before $t$ conditionally to $\mathcal{G}_{t}$ that is noted $PD^{n}_{t,\mathfrak{d}}$, we have
$PD^{n}_{t,T,\mathfrak{d}}=\mathbb{P}\left(\tau^{n}\leq t|\mathcal{G}_{t}\right)=\mathbb{P}\left(\inf_{0\leq s\leq t}\mathcal{V}^{n}_{s,\mathfrak{d}}<\mathcal{D}^{n}_{s}\middle|\mathcal{G}_{t}\right)=\mathbb{P}\left(\inf_{0\leq s\leq t}\log{\mathcal{V}^{n}_{s,\mathfrak{d}}}<\log{\mathcal{D}^{n}_{s}}\middle|\mathcal{G}_{t}\right).$
But for $0\leq s\leq t$ and from (2.21),
$\log{\mathcal{V}_{s,\mathfrak{d}}^{n}}=\log{\left(F_{0,\mathfrak{d}}^{n}\mathfrak{R}^{n}_{s}(\mathfrak{d})\right)}+\mathfrak{a}^{n\cdot}(\mathcal{A}_{s}^{\circ}-v(\mathfrak{d}_{0}))+\sigma_{n}\mathcal{W}^{n}_{s},$
therefore $\log{\mathcal{V}^{n}_{\cdot,\mathfrak{d}}}$ is a Gaussian process. However, as Azaıs and Wschebor, (2000) summarizes, the computation of the distribution function of the random variable $\inf_{0\leq s\leq t}\log{\mathcal{V}^{n}_{s,\mathfrak{d}}}$ is by means of a closed formula is known only for a very restricted number of stochastic processes as the Brownian Motion, the Brownian Bridge, the Brownian Motion with a linear drift, and the stationary Gaussian processes with relatively simple with covariance. This is not the case here.
At each time $t\geq 0$, we are interested in the probability that firm $n$ is over-indebted at a certain date $t+T$, we note $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ and we have
$\mathrm{PD}^{n}_{t,T,\mathfrak{d}}:=\mathbb{P}\left(\mathcal{V}^{n}_{t+T,\mathfrak{d}}\leq\mathcal{D}^{n}_{t+T}\middle|\mathcal{G}_{t}\right).$ (2.28)
We have the following proposition whose proof is a direct application of Corollary 2.7.
Proposition 2.9 (Probability of default).
For $t\geq 0$, $T\geq 0$, and $n\in\{1,\ldots,N\}$, the (conditional) probability of default of the entity $n$ at time $t$ over the horizon $T$ is
$\mathrm{PD}^{n}_{t,T,\mathfrak{d}}=\Phi\left(\frac{\log(\mathcal{D}_{t+T}^{n})-\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})}{\sqrt{\mathcal{L}^{n}(t,T)}}\right),$ (2.29)
where $\mathcal{K}^{n}(\mathfrak{d},t,T,a,\theta)$ and $\mathcal{L}^{n}(t,T)$ are defined in Corollary 2.7.
The previous results tell us that the probability of the over-indebtedness depends on:
  1. 1.
    parameters specific to the climate transition
    • (a)
      the carbon price $\delta$,
    • (b)
      the carbon intensities $\tau,\zeta,\kappa$,
  2. 2.
    parameters specific to the company (the contract),
    • (a)
      the factors loading $\mathfrak{a}^{n\cdot}$ and the standard deviation of the cash flows $\sigma_{n}$,
    • (b)
      the time $t$ when it is computed,
    • (c)
      the potential date of the over-indebtedness $t+T$,
    • (d)
      the over-indebtedness’s barrier $\mathcal{D}$,
  3. 3.
    parameters specific to the economy to which the company belongs to:
    • (a)
      the productivity $\mathcal{Z}$ and $\mathcal{A}$ (and their parameters) of the economy,
    • (b)
      the interest rate $r$.
Therefore, by assuming that EAD and LGD are deterministic and independent of the carbon price, we could obtain the expressions of the EL and UL. In the next section, we will express LGD as a function of some guarantees which are affected by the climate transition.
3 LGD with stochastic collaterals in continuous time
We are in the same framework as in the previous section, but Assumption 2.8(2) is not satisfied anymore, therefore the bank could require from each counterpart $1\leq n\leq N$ a (single) collateral $C^{n}$ to secure its debt. Collateral can take the form of a physical asset such as real estate, business equipment, or inventory, or it can be a financial asset such as invoices, cash, or investments. If a firm is over-indebted at time $t$, we assume that the liquidation ends at $t+a$ with $a\in\mathbb{R}_{+}$ where $a$ is the liquidation delay. Moreover, $k\in[0,1)$ represents the fraction of the collateral used to cover liquidations auctions, as well as other legal and administrative procedures.
A firm is over-indebted at time $t\geq 0$ if the market value of its debt $\mathcal{D}^{n}_{t}$ exceed its market value $\mathcal{V}^{n}_{t,\mathfrak{d}}$, namely $\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\}$. At time $t$, if the company $n$ in the portfolio defaults i.e. $\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}~{}$, the bank recovers $(1-k)e^{-ra}C^{n}_{t+a}$ after the collateral liquidation. In general, the liquidations do not cover all the debt, i.e. $\mathrm{EAD}_{t}^{n}\geq(1-k)e^{-ra}C^{n}_{t+a}$, the bank deploys further actions to recover an additional fraction. We note that fraction $\gamma\in[0,1)$. Therefore, the bank recovers $\gamma(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})_{+}$ by other tools.
The potential loss that would be recorded due to the firm default event is the difference between the debt amount $\mathrm{EAD}_{t}^{n}$ and the amount gets after the recovery processes at the time horizon. Consequently, if there is not default ($\mathcal{V}^{n}_{t,\mathfrak{d}}\geq\mathcal{D}^{n}_{t}$) or there is default and the collateral liquidated exceed the exposure ($(1-k)e^{-ra}C^{n}_{t+a}\geq\mathrm{EAD}_{t}^{n}$), the loss noted $L_{n,t}$ is zero, and if there is default and if the exposure exceed the collateral liquidated, the loss is
$L_{n,t}=\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a}-\gamma(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})=(1-\gamma)\left(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a}\right),$ (3.1)
where the constant $r$ is the discount rate. The loss of the portfolio at time $t$, is in fact, defined as
$L^{N}_{t}:=\sum_{n=1}^{N}L_{n,t}=\sum_{n=1}^{N}(1-\gamma)(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\}}.$ (3.2)
The following result is similar with the one introduced in Bouveret et al., (2023)[Theorem 3.5]. It gives a proxy of the loss of the portfolio.
Theorem 3.1 (Definition of PD and LGD).
For all $t\in\mathbb{R}_{+}$, define
$\mathrm{L}^{\mathbb{G},N}_{t}:=\sum_{n=1}^{N}L_{n,t}^{\mathbb{G}}\quad\text{with}\quad L_{n,t}^{\mathbb{G}}=\mathbb{E}\left[L_{n,t}\middle|\mathcal{G}_{t}\right]=\mathrm{EAD}_{t}^{n}\cdot\mathrm{LGD}_{t,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t,\mathfrak{d}}^{n},$ (3.3)
where
$\displaystyle\mathrm{PD}_{t,\mathfrak{d}}^{n}$ $:=\mathbb{P}\left(\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}|\mathcal{G}_{t}\right)=\Phi\left(\frac{\log(\mathcal{D}_{t}^{n})-\mathfrak{m}^{n}(\mathfrak{d},t,\mathcal{A}_{t})}{\sigma^{n}\sqrt{t}}\right),$ (3.4a)
$\displaystyle\mathrm{LGD}_{t,\mathfrak{d}}^{n}$ $:=(1-\gamma)\mathbb{E}\left[\left(1-(1-k)e^{-ra}\frac{C^{n}_{t+a}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right].$ (3.4b)
Under Assumptions 2.8, we have $L^{N}_{t}-\mathrm{L}^{\mathbb{G},N}_{t}$ converges to zero almost surely as $N$ tends to infinity, for all $t\in\mathbb{R}_{+}$.
Proof.
Let $t\in\mathbb{R}_{+}$,
$\begin{split}\mathrm{L}^{\mathbb{G},N}_{t}&=\mathbb{E}\left[L^{N}_{t}\middle|\mathcal{G}_{t}\right]=\mathbb{E}\left[\sum_{n=1}^{N}(1-\gamma)(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\}}\middle|\mathcal{G}_{t}\right].\end{split}$
$(\mathrm{EAD}_{t}^{n})_{n\in\{1,\ldots,N\}}$ is deterministic, we have:
$\begin{split}\mathrm{L}^{\mathbb{G},N}_{t}&=\sum_{n=1}^{N}\mathrm{EAD}_{t}^{n}\cdot\mathbb{E}\left[(1-\gamma)\left(1-(1-k)e^{-ra}\frac{C^{n}_{t+a}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\}}\middle|\mathcal{G}_{t}\right]\\ &=\sum_{n=1}^{N}\mathrm{EAD}_{t}^{n}\cdot\mathbb{E}\left[(1-\gamma)\left(1-(1-k)e^{-ra}\frac{C^{n}_{t+a}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right]\cdot\mathbb{P}\left[\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\middle|\mathcal{G}_{t}\right].\\ \end{split}$
The rest of the proof requires a version of the strong law of large numbers (Appendix of (Gordy, , 2003, Propositions 1, 2)), where the systematic risk factor is $\mathcal{G}_{t}$. ∎
Explicitly, in the above theorem, we assume that our portfolio is perfectly fine grained, so that we can approximate $L_{t}^{N}$ – the portfolio loss – by $L_{t}^{\mathbb{G},N}$ – the conditional expectation of loss given the systemic factor–. By construction, the loss given default noted $\mathrm{LGD}$ is the percentage of the total exposure that is lost when a over-indebtedness occurs. The literature on LGD modeling is fairly extensive. We can distinguish namely, economic modeling Bastos, (2010); Roncalli, (2020) and stochastic modeling Roncalli, (2020)[Page 193], Chalupka and Kopecsni, (2008). As the definition of $\mathrm{PD}$ does not change compare to what we did in Section 2 and as $\mathrm{EAD}$ is given, we will focus on $\mathrm{LGD}$ modeling. We can first remark that $0\leq\mathrm{LGD}_{t,\mathfrak{d}}^{n}\leq 1-\gamma$, then the presence of a collateral necessarily reduces $\mathrm{LGD}$.
Key quantities for the bank to understand the (dynamics of the) risk in the portfolio are the (expected and unexpected) losses and probability of default conditionally to the (information generated by the) risk factors. Precisely, for a date $t$ and a horizon $T$, a bank would like to know some risk measures at $t$ of its portfolio maturing at horizon $T$.
Definition 3.2 (Projected losses).
Let $t\geq 0$ be the time when the risk measure is computed for a period $T\geq 0$. As classically done, the potential loss is separated into three components:
  • 1.
    The (conditional) Expected Loss (EL) reads
    $\mathrm{EL}^{N,T}_{t}:=\mathbb{E}\left[\mathrm{L}^{\mathbb{G},N}_{t+T}\middle|\mathcal{G}_{t}\right].$ (3.5)
  • 2.
    The Unexpected Loss (UL) reads for $\alpha\in(0,1)$,
    $\mathrm{UL}^{N,T}_{t}\!(\alpha):=\mathrm{VaR}^{\alpha,N,T}_{t}-\mathrm{EL}^{N,T}_{t},\quad\text{where}\quad 1-\alpha=\mathbb{P}\left[L_{t+T}^{\mathbb{G},N}\leq\mathrm{VaR}^{\alpha,N,T}_{t}\middle|\mathcal{G}_{t}\right].$ (3.6)
  • 3.
    The Stressed Loss (or Expected Shortfall or ES) reads $\mathrm{VaR}^{\alpha}_{t}(L^{N}_{s})$:
    $\mathrm{ES}^{N,T}_{t}(\alpha):=\mathbb{E}\left[L^{N}_{t+T}\middle|L^{N}_{t+T}\geq\mathrm{VaR}^{\alpha,N,T}_{t},\mathcal{G}_{t}\right],\qquad\text{for }\alpha\in(0,1).$ (3.7)
From now, if the collateral exists, we focus on two types: a financial asset and a property in housing market. We introduce the two following lemmas which will help later on to set up explicit formulas for LGD. Let $n\in\{1,\ldots,N\}$.
Lemma 3.3.
Assume that a stochastic process $K^{n}$ satisfies for all $t\in\mathbb{R}_{+}$,
  1. 1.
    $\log{K^{n}}_{t}|\mathcal{G}_{t}\sim\mathcal{N}(m_{t}^{n},(\sigma^{n}_{t})^{2})$ with $m^{n}_{t}\in\mathbb{R}$ and $\sigma^{n}_{t}>0$,
  2. 2.
    and $K^{n}_{t}|\mathcal{G}_{t}$ and $\mathcal{V}^{n}_{t,\mathfrak{d}}|\mathcal{G}_{t}$ are independent.
Therefore, for $t,u\in\mathbb{R}_{+}$,
$\mathbb{E}\left[\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right]=u\Phi\left(\frac{w_{t}^{n}}{\sigma^{n}_{t}}\right)-\exp{\left(-w_{t}^{n}+\frac{1}{2}(\sigma^{n}_{t})^{2}\right)}\Phi\left(\frac{w_{t}^{n}}{\sigma^{n}_{t}}-\sigma^{n}_{t}\right),$ (3.8)
where
$w_{t}^{n}:=\log{\left(u\frac{\mathrm{EAD}_{t}^{n}}{1-k}\right)}-m_{t}^{n}.$ (3.9)
Proof.
Let $t,u\in\mathbb{R}_{+}$, we have
$\mathbb{E}\left[\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right]$
$\qquad=\mathbb{E}\left[\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{G}_{t}\right]\quad\text{because}\quad K^{n}_{t}|\mathcal{G}_{t}\text{ and }\mathcal{V}_{t,\mathfrak{d}}^{n}|\mathcal{G}_{t}\text{ are independent}$
$\qquad=\mathbb{E}\left[\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\right)\mathbf{1}_{\{u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\geq 0\}}\middle|\mathcal{G}_{t}\right]$
$\qquad=u\mathbb{P}\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\geq 0\middle|\mathcal{G}_{t}\right)-\frac{(1-k)}{\mathrm{EAD}_{t}^{n}}\mathbb{E}\left[K^{n}_{t}\mathbf{1}_{\{u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\geq 0\}}\middle|\mathcal{G}_{t}\right].$
However, $\log{K^{n}_{t}}|\mathcal{G}_{t}\sim\mathcal{N}(m_{t}^{n},(\sigma^{n}_{t})^{2})$. We also consider $w_{t}^{n}$ defined in (3.9), therefore
$\mathbb{P}\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\geq 0\middle|\mathcal{G}_{t}\right)=\Phi\left(\frac{w_{t}^{n}}{\sigma^{n}_{t}}\right).$
We also have
$\mathbb{E}\left[K^{n}_{t}\mathbf{1}_{\{u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\geq 0\}}\middle|\mathcal{G}_{t}\right]=\exp{\left(-w_{t}^{n}+\frac{1}{2}(\sigma^{n}_{t})^{2}\right)}\Phi\left(\frac{w_{t}^{n}}{\sigma^{n}_{t}}-\sigma^{n}_{t}\right).$
The conclusion follows. ∎
Lemma 3.4.
Assume that a stochastic process $K^{n}$ satisfies, for each $t,T\in\mathbb{R}_{+}$,
  1. 1.
    $\log{K^{n}}_{t+T}|\mathcal{G}_{t}\sim\mathcal{N}(m_{t,T}^{n},(\sigma^{n}_{t,T})^{2})$ with $m^{n}_{t,T}\in\mathbb{R}$ and $\sigma^{n}_{t,T}>0$,
  2. 2.
    and $\begin{bmatrix}\log{\mathcal{V}_{t+T}^{n}}\\ \log{K_{t+T+a}^{n}}\end{bmatrix}|\mathcal{G}_{t}\sim\mathcal{N}\left(\begin{bmatrix}\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})\\ \overline{\mathcal{K}}^{n}_{t,T+a}\end{bmatrix},\begin{bmatrix}\mathcal{L}^{n}(t,T)&cv_{t,T,a}^{n}\\ cv_{t,T,a}^{n}&\overline{\mathcal{L}}_{t,T+a}^{n}\end{bmatrix}\right)$, where ¯$\overline{\mathcal{L}}_{t,T+a}^{n}>0$, $cv_{t,T,a}^{n},\overline{\mathcal{K}}^{n}_{t,T+a}\in\mathbb{R}$, and $\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})$ as well as $\mathcal{L}^{n}(t,T)$ are defined in Corollary 2.7.
Therefore, for $t,T,u\in\mathbb{R}_{+}$, we have
$\begin{split}&\mathbb{E}\left[\left(1-(1-k)e^{-ra}\frac{\mathcal{C}^{n}_{t+T+a,\mathfrak{d}}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right]\\ &\qquad=u\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}});\rho_{t,T,a}^{n}\right)-\exp{\left(\frac{1}{2}\overline{\mathcal{L}}_{t,T+a}^{n}-\sqrt{\overline{\mathcal{L}}_{t,T+a}^{n}}\overline{\omega}_{t,T,a}^{n}\right)}\times\\ &\qquad\qquad\qquad\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n}-\sqrt{\overline{\mathcal{L}}_{t,T+a}^{n}},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}})-\rho_{t,T,a}^{n}\sqrt{\overline{\mathcal{L}}_{t,T+a}^{n}};\rho_{t,T,a}^{n}\right),\end{split}$ (3.10)
where $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ is defined in (2.29) and where
$\rho_{t,T,a}^{n}:=\frac{cv_{t,T,a}^{n}}{\sqrt{\mathcal{L}^{n}(t,T)\overline{\mathcal{L}}^{n}_{t,T+a}}},$
and
$\overline{\omega}_{t,T,a}^{n}:=\frac{\log{\left(u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\right)}-\overline{\mathcal{K}}^{n}_{\mathfrak{d},t,T+a}}{\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}}}.$
Proof.
Let $t,T,u\in\mathbb{R}_{+}$, we have
$\begin{split}&\mathbb{E}\left[\left(u-(1-k)e^{-ra}\frac{K_{t+T+a}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right]\\ &\qquad=\mathbb{E}\left[\left(u-(1-k)e^{-ra}\frac{K_{t+T+a}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)\cdot\mathbf{1}_{u-(1-k)e^{-ra}\frac{K_{t+T+a}^{n}}{\mathrm{EAD}_{t+T}^{n}}\geq 0}\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right]\\ &\qquad=\mathbb{E}\left[\left(u-(1-k)e^{-ra}\frac{K_{t+T+a}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},\quad K_{t+T+a}^{n}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}}\middle|\mathcal{G}_{t}\right]\\ &\qquad=u\mathbb{P}\left[\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},K_{t+T+a}^{n}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}\middle|\mathcal{G}_{t}\right]\\ &\qquad\qquad-\frac{(1-k)e^{-ra}}{\mathrm{EAD}_{t+T}^{n}}\mathbb{E}\left[K_{t+T+a}^{n}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},K_{t+T+a}^{n}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}}\middle|\mathcal{G}_{t}\right].\end{split}$
However $\begin{bmatrix}\log{\mathcal{V}_{t+T}^{n}}\\ \log{K_{t+T+a}^{n}}\end{bmatrix}|\mathcal{G}_{t}\sim\mathcal{N}\left(\begin{bmatrix}\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})\\ \overline{\mathcal{K}}^{n}_{t,T+a}\end{bmatrix},\begin{bmatrix}\mathcal{L}^{n}(t,T)&cv_{t,T,a}^{n}\\ cv_{t,T,a}^{n}&\overline{\mathcal{L}}_{t,T+a}^{n}\end{bmatrix}\right)$, therefore we have
$\mathbb{P}\left[\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},K^{n}_{t+T+a}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}\middle|\mathcal{G}_{t}\right]$
$\qquad=\Phi_{2}\left(\frac{\log{u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}}-\overline{\mathcal{K}}^{n}_{t,T+a}}{\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}}},\frac{\log{\mathcal{D}^{n}_{t+T}}-\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})}{\sqrt{\mathcal{L}^{n}(t,T)}};\frac{cv_{t,T,a}}{\sqrt{\mathcal{L}^{n}(t,T)\overline{\mathcal{L}}^{n}_{t,T+a}}}\right),$
and
$\begin{split}&\frac{(1-k)e^{-ra}}{\mathrm{EAD}_{t+T}^{n}}\mathbb{E}\left[K^{n}_{t+T+a}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},K^{n}_{t+T+a}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}}\middle|\mathcal{G}_{t}\right]\\ &\qquad=\mathbb{E}\left[e^{\log{K^{n}_{t+T+a}}}\cdot\mathbf{1}_{\left\{\frac{\log{K^{n}_{t+T+a}}-\overline{\mathcal{K}}^{n}_{t,T+a}}{\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}}}\leq\overline{\omega}_{t,T,a}^{n},\frac{\log{\mathcal{V}^{n}_{t+T,\mathfrak{d}}}-\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})}{\sqrt{\mathcal{L}^{n}(t,T)}}<\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}})\right\}}\middle|\mathcal{G}_{t}\right]\end{split}$
However, according to (A.3), $\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]=e^{\frac{1}{2}\sigma^{2}}\Phi_{2}\left(x-\sigma,y-\rho\sigma;\rho\right)$, therefore,
$\frac{(1-k)e^{-ra}}{\mathrm{EAD}_{t+T}^{n}}\mathbb{E}\left[e^{\log{K^{n}_{t+T+a}}}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},K^{n}_{t+T+a}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}}\middle|\mathcal{G}_{t}\right]$
$\qquad=\exp{\left(\frac{1}{2}\overline{\mathcal{L}}^{n}_{t,T+a}-\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}}\overline{\omega}_{t,T,a}^{n}\right)}\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n}-\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}})-\frac{cv_{t,T,a}}{\sqrt{\mathcal{L}^{n}(t,T)}};\frac{cv_{t,T,a}}{\sqrt{\mathcal{L}^{n}(t,T)\overline{\mathcal{L}}^{n}_{t,T+a}}}\right);$
Moreover, from Proposition 2.9,
$\mathbb{P}\left[\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\middle|\mathcal{G}_{t}\right]=\Phi\left(\frac{\log(\mathcal{D}_{t+T}^{n})-\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})}{\sqrt{\mathcal{L}^{n}(t,T)}}\right).$
This concludes the proof. ∎
3.1 When there is not collateral
When firm $n$ does not have a collateral, therefore $C^{n}=0$ and from (3.4b), LGD is
$\mathrm{LGD}_{t,\mathfrak{d}}^{n}=1-\gamma.$ (3.11)
3.2 When collateral is a financial asset
Here we assume that the collateral of the firm $n$ is an investment in a financial asset. Precisely, we assume that that investment is a proportion $\alpha^{n}\in(0,1]$ of a given firm located in the economy described in Section 2. Consequently, it is subjected to the same constraints in terms of productivity and of carbon transition scenarios as firm $n$. As any investment, it should generate a stream of cash flows so that at each time, we can compute its value by using the discounted cash flows model introduced in (2.19).
Let note the collateral cash flows $(\overline{F}_{t}^{n})_{t\in\mathbb{R}_{+}}$, its dynamics is similar to the firm cash flows introduced in Assumption 2.5. We have for all $t\in\mathbb{R}_{+}$,
$\mathrm{d}\overline{F}^{n}_{t,\mathfrak{d}}=\mathfrak{\overline{a}}^{n\cdot}((\mu+\varsigma\mathcal{Z}_{t})\mathrm{d}t+\mathrm{d}v(\mathfrak{d}_{t}))+\overline{\sigma}_{n}\mathrm{d}\overline{\mathcal{W}}^{n}_{t},$ (3.12)
where $\mathfrak{\overline{a}}^{n\cdot}\in\mathbb{R}^{I}$ and where $(\overline{\mathcal{W}}_{t})_{t\in\mathbb{R}_{+}}$ is a $\mathbb{R}^{N}$-Brownian motion with $\overline{\sigma}_{n}>0$. Moreover, $B^{\mathcal{Z}}$ (noise of productivity), $\overline{\mathcal{W}}^{n}$ (noise of collateral), and $(\mathcal{W}^{n})_{n\in\{1,\ldots,N\}}$ (noise of debtors) are independent. We also note $\overline{\tilde{\mathfrak{a}}}^{n\cdot}=\overline{\mathfrak{a}}^{n\cdot}(\mathbf{I}_{I}-\boldsymbol{\lambda})$.
Remark 3.5.
We have assumed that $\overline{\mathcal{W}}^{n}$ and $\mathcal{W}^{n}$ are not correlated, but this is not always the case. For example, if the depreciation of the firm value heading to its over-indebtedness implies the depreciation of the collateral value, then we should have a positive correlation.
Inspired by (2.19), the collateral value at time $t$ is
$C^{n}_{t,\mathfrak{d}}:=\alpha^{n}\mathbb{E}_{t}\left[\int_{s=t}^{+\infty}e^{-rs}\overline{F}^{n}_{s,\mathfrak{d}}\mathrm{d}s\right],$
and by (2.20), the approached collateral value as
$\mathcal{C}_{t,\mathfrak{d}}^{n}:=\alpha^{n}\overline{F}^{n}_{t,\mathfrak{d}}\int_{t}^{+\infty}e^{-r(s-t)}\mathbb{E}_{t}\left[\exp\left((s-t)\overline{\mathfrak{a}}^{n\cdot}\mu+\overline{\mathfrak{a}}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)+\sigma_{n}(\overline{\mathcal{W}}^{n}_{s}-\overline{\mathcal{W}}^{n}_{t})\right)\mathrm{d}s\right].$ (3.13)
Therefore, the following proposition and its proof are inspired by Lemma 2.6 and corollary 2.7, gives a proxy of the collateral value.
Proposition 3.6.
For any $n\in\{1,\ldots,N\}$ and for all $t\in\mathbb{R}_{+}$
  1. 1.
    Assume that $\overline{\varrho}_{n}:=\frac{1}{2}\overline{\sigma}_{\mathfrak{b}_{n}}^{2}+\mathfrak{\overline{a}}^{n\cdot}\mu-r<0$. Given the carbon emissions costs sequence $\mathfrak{d}$, the proxy of collateral value defined in (3.13), is well defined and
[…]
    where
[…]
  2. 2.
[…]
    $\log{\mathcal{C}^{n}_{t,\mathfrak{d}}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\mathfrak{\overline{m}}^{n}(\mathfrak{d},t,\mathcal{A}_{t}^{\circ}),t\overline{\sigma}_{n}^{2}\right),$ (3.16)
    and we note $\overline{\mathcal{K}}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t}):=\log{(\alpha^{n}\overline{F}^{n}_{0}\overline{\mathfrak{R}}^{n}_{t+T}(\mathfrak{d}))}+\overline{\mathfrak{a}}^{n\cdot}(\mu T+\varsigma\Upsilon_{T}\mathcal{Z}_{t}+\mathcal{A}_{t}^{\circ}-v(\mathfrak{d}_{0}))$ and ¯$\overline{\mathcal{L}}^{n}(t,T):=\overline{\mathfrak{a}}^{n\cdot}\Sigma^{\mathcal{A}}_{t,t+T}\overline{\mathfrak{a}}^{n\cdot}+(t+T)\overline{\sigma}_{n}^{2}$, and we have
    $\log{\mathcal{C}_{t+T}^{n}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\overline{\mathcal{K}}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t}),\overline{\mathcal{L}}^{n}(t,T)\right).$ (3.17)
  3. 3.
    Assume that
    $\overline{\rho}_{n}:=\frac{1}{2}\overline{\sigma}_{n}^{2}+\overline{\mathfrak{a}}^{n\cdot}\mu+\frac{1}{2}\varsigma^{2}\frac{c_{\Gamma}^{2}}{\lambda_{\Gamma}^{2}}\lVert\overline{\mathfrak{a}}^{n\cdot}\rVert^{2}\lVert\Sigma\rVert^{2}<r,$ (3.18)
    therefore $C_{t,\mathfrak{d}}^{n}$ is well defined and there exists a constant $\overline{C}$ such that $\mathbb{E}\left[\left|\frac{C^{n}_{t,\mathfrak{d}}}{\overline{F}^{n}_{t,\mathfrak{d}}}-\frac{\mathcal{C}^{n}_{t,\mathfrak{d}}}{\overline{F}^{n}_{t,\mathfrak{d}}}\right|\right]\leq\overline{C}\varsigma$, for all $\varsigma>0$.
Proof.
Let $n\in\{1,\ldots,N\}$ and $t\in\mathbb{R}_{+}$, (3.14) directly comes from Proposition 2.6. The proofs of the three points are equivalent to B.4. Let us develop the conditional laws. From (3.14), we have
$\log{\mathcal{C}^{n}_{t}}=\log{\alpha^{n}\overline{F}^{n}_{0}\overline{\mathfrak{R}}^{n}_{t}(\mathfrak{d})}+\overline{\mathfrak{a}}^{n\cdot}(\mathcal{A}_{t}^{\circ}-v(\mathfrak{d}_{0}))+\overline{\sigma}_{n}\overline{\mathcal{W}}^{n}_{t}.$
Because $\overline{\mathcal{W}}^{n}$ is a Brownian motion, $\overline{\mathcal{W}}^{n}_{t}\sim\mathcal{N}\left(0,t\right)$ and, $\overline{\mathcal{W}}^{n}$ and $B^{\mathcal{Z}}$ are independent, we obtain $\log{\mathcal{C}_{t}^{n}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\overline{\mathfrak{m}}^{n}(\mathfrak{d},t,\mathcal{A}_{t}^{\circ}),t\overline{\sigma}_{n}^{2}\right)$. Let also $T\in\mathbb{R}_{+}$, we have
$\begin{split}\log{\mathcal{C}^{n}_{t+T}}&=\log{\overline{F}^{n}_{0}\overline{\mathfrak{R}}^{n}_{t+T}(\mathfrak{d})}+\overline{\mathfrak{a}}^{n\cdot}(\mathcal{A}_{t+T}^{\circ}-v(\mathfrak{d}_{0}))+\overline{\mathcal{W}}^{n}_{t+T}.\end{split}$
From Remark 2.2, $\mathcal{A}_{t+T}|\mathcal{G}_{t}\sim\mathcal{N}\left(M^{\mathcal{A},T}_{t},\Sigma^{\mathcal{A},T}_{t}\right)$ and because $\overline{\mathcal{W}}^{n}$ is a Brownian motion, $\overline{\mathcal{W}}^{n}_{t+T}\sim\mathcal{N}\left(0,t+T\right)$. Moreover, $\overline{\mathcal{W}}^{n}$ and $B^{\mathcal{Z}}$ are independent. We have
$\log{\mathcal{C}^{n}_{t+T}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\log{\alpha^{n}\overline{F}^{n}_{0}\overline{\mathfrak{R}}^{n}_{t+T}(\mathfrak{d})}+\overline{\mathfrak{a}}^{n\cdot}(M^{\mathcal{A},T}_{t}-v(\mathfrak{d}_{0})),\overline{\mathfrak{a}}^{n\cdot}\Sigma^{\mathcal{A},T}_{t}\overline{\mathfrak{a}}^{n\cdot}+(t+T)\overline{\sigma}_{n}^{2}\right).$
The conclusion follows. ∎
From the (proxy of the) collateral value $\mathcal{C}$, we can then derive a precised expression of $\mathrm{LGD}$ based on Theorem 3.1. We have:
Theorem 3.7.
When $a=0$ (no liquidation delay), the Loss Given Default of the obligor $n$ over-indebted at time $t\in\mathbb{R}_{+}$, conditional on $\mathcal{G}_{t}$ is
$\mathrm{LGD}_{t,\mathfrak{d}}^{n}=(1-\gamma)\left[\Phi\left(\frac{w_{t}^{n}}{\overline{\sigma}_{\mathfrak{b}_{n}}\sqrt{t}}\right)-\exp{\left(-w_{t}^{n}+\frac{1}{2}t\overline{\sigma}_{n}^{2}\right)}\Phi\left(\frac{w_{t}^{n}}{\overline{\sigma}_{n}\sqrt{t}}-\overline{\sigma}_{n}\sqrt{t}\right)\right],$ (3.19)
where
$w_{t}^{n}:=\log{\left(\frac{\mathrm{EAD}_{t}^{n}}{1-k}\right)}-\mathfrak{\overline{m}}^{n}(\mathfrak{d},t,\mathcal{A}_{t}).$ (3.20)
Proof.
Let $t\in\mathbb{R}_{+}$. By remarking that
  1. 1.
    $\log{\mathcal{C}^{n}_{t,\mathfrak{d}}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\mathfrak{\overline{m}}^{n}(\mathfrak{d},t,\mathcal{A}_{t}),t\overline{\sigma}_{n}^{2}\right)$,
  2. 2.
    and $\mathcal{C}_{t,\mathfrak{d}}^{n}|\mathcal{G}_{t}$ and $\mathcal{V}_{t,\mathfrak{d}}^{n}|\mathcal{G}_{t}$ are independent,
we can simply apply Lemma 3.3 with $u=1$. ∎
We can also remark that the situation where there is no collateral corresponds to $\overline{F}^{n}_{0}=0$. We then have
$\begin{split}\overline{F}^{n}_{0}\to 0&\implies\log{(\overline{F}^{n}_{0})}\to-\infty\implies\mathfrak{\overline{m}}^{n}(\mathfrak{d},t,\mathcal{A}_{t}^{\circ})\to-\infty\implies w_{t}^{n}\to+\infty\implies\mathrm{LGD}_{t,\mathfrak{d}}^{n}\to 1-\gamma.\end{split}$
For each $t,T\in\mathbb{R}_{+}$, we introduce now the (conditional) $\mathrm{LGD}$ of the entity $n$ at time $t$ on the horizon $T$, namely
$\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}:=(1-\gamma)\mathbb{E}\left[\left(1-(1-k)e^{-ra}\frac{\mathcal{C}^{n}_{t+T+a,\mathfrak{d}}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},\mathcal{G}_{t}\right].$
It is precisely about calculating at date $t$ the proportion of the exposure that the bank would lose if the counterpart $n$ is over-indebted at date $t+T$.
For $x,y\in\mathbb{R}$, we note $\Phi_{2}(x,y;\rho)$ is the cumulative distribution function of the bi-variate Gaussian vector $(X,Y)$ with correlation $\rho$ on the space $[-\infty,x]\times[-\infty,y]$.
Proposition 3.8 (Projected PD and LGD).
For each $t,T\in\mathbb{R}_{+}$, the (conditional) $\mathrm{LGD}$ of the entity $n$ at time $t$ on the horizon $T$, reads
$\begin{split}\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}&=\frac{1-\gamma}{\mathrm{PD}^{n}_{t,T,\mathfrak{d}}}\left[\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}});\rho_{t,T,a}^{n}\right)-\exp{\left(\frac{1}{2}\overline{\mathcal{L}}^{n}(t,T+a)-\sqrt{\overline{\mathcal{L}}^{n}(t,T+a)}\overline{\omega}_{t,T,a}^{n}\right)}\times\right.\\ &\qquad\qquad\qquad\left.\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n}-\sqrt{\overline{\mathcal{L}}^{n}(t,T+a)},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}})-\rho_{t,T,a}^{n}\sqrt{\overline{\mathcal{L}}^{n}(t,T+a)};\rho_{t,T,a}^{n}\right)\right],\end{split}$ (3.21)
where
$\rho_{t,T,a}^{n}:=\frac{\varsigma^{2}\mathfrak{a}^{n\cdot}\Gamma^{-1}\left(\int_{0}^{T}\left(e^{-\Gamma u}-\mathbf{I}_{I}\right)\Sigma\Sigma^{\top}\left(e^{-\Gamma(u+a)}-\mathbf{I}_{I}\right)\mathrm{d}u\right)(\overline{\mathfrak{a}}^{n\cdot}\Gamma^{-1})^{\top}}{\sqrt{\mathcal{L}^{n}(t,T)\overline{\mathcal{L}}^{n}(t,T+a)}},$
and
$\overline{\omega}_{t,T,a}^{n}:=\frac{\log{\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}}-\overline{\mathcal{K}}^{n}(\mathfrak{d},t,T+a,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})}{\sqrt{\overline{\mathcal{L}}^{n}(t,T+a)}},$
and where $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ defined in Proposition 2.9.
Proof.
Let $t,T\in\mathbb{R}_{+}$, from (3.3) and (3.5),
$\mathrm{EL}^{N,T}_{t}:=\mathbb{E}\left[\mathrm{L}^{\mathbb{G},N}_{t+T}\middle|\mathcal{G}_{t}\right]=\mathbb{E}\left[\sum_{n=1}^{N}L_{n,t+T}^{\mathbb{G}}\middle|\mathcal{G}_{t}\right]=\sum_{n=1}^{N}\mathbb{E}\left[L_{n,t+T}^{\mathbb{G}}\middle|\mathcal{G}_{t}\right].$
But for $1\leq n\leq N$, we have
$\begin{split}\mathbb{E}\left[L_{n,t+T}^{\mathbb{G}}\middle|\mathcal{G}_{t}\right]&=\mathbb{E}\left[\mathrm{EAD}_{t+T}^{n}\mathrm{LGD}^{n}_{t+T,\mathfrak{d}}\mathrm{PD}^{n}_{t+T,\mathfrak{d}}\middle|\mathcal{G}_{t}\right]\\ &=\mathrm{EAD}_{t+T}^{n}\mathbb{E}\left[\mathrm{LGD}^{n}_{t+T,\mathfrak{d}}\mathrm{PD}_{t+T,\mathfrak{d}}^{n}\middle|\mathcal{G}_{t}\right]\quad\text{as }\mathrm{EAD}_{t+T}^{n}\text{ is deterministic}\\ &=\mathrm{EAD}_{t+T}^{n}\mathbb{E}\left[\mathbb{E}\left[(1-\gamma)\left(1-(1-k)e^{-ra}\frac{\mathcal{C}^{n}_{t+T+a,\mathfrak{d}}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t+T}\right]\middle|\mathcal{G}_{t}\right]\\ &=(1-\gamma)\mathrm{EAD}_{t+T}^{n}\mathbb{E}\left[\left(1-(1-k)e^{-ra}\frac{\mathcal{C}^{n}_{t+T+a,\mathfrak{d}}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right].\end{split}$
[…]
$cov(\log{F_{t+T}^{n}},\log{\mathcal{C}_{t+T+a}^{n}}|\mathcal{G}_{t})=\mathbb{E}[\log{F_{t+T}^{n}}\log{\mathcal{C}_{t+T+a}^{n}}|\mathcal{G}_{t}]-\mathbb{E}[\log{F_{t+T}^{n}}|\mathcal{G}_{t}]\mathbb{E}[\log{\mathcal{C}_{t+T+a}^{n}}|\mathcal{G}_{t}].$
However,
$\mathbb{E}[\log{F_{t+T}^{n}}\log{\mathcal{C}_{t+T+a}^{n}}|\mathcal{G}_{t}]$
$\qquad=\mathbb{E}\left[\left(\log{\left(F^{n}_{0}\mathfrak{R}^{n}_{t+T}(\mathfrak{d})e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{0})}\right)}+\mathfrak{a}^{n\cdot}\mathcal{A}_{t+T}^{\circ}+\sigma_{n}\mathcal{W}^{n}_{t+T}\right)\right.$
[…]
$\qquad\qquad\left.+\log{\left(F^{n}_{0}\mathfrak{R}^{n}_{t+T}(\mathfrak{d})e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{0})}\right)}(\overline{\mathfrak{a}}^{n\cdot}\mathcal{A}_{t+T+a}^{\circ}+\overline{\sigma}_{n}\overline{\mathcal{W}}_{t+T+a}^{n})+\overline{\sigma}_{n}\overline{\mathcal{W}}_{t+T+a}^{n}\sigma_{n}\mathcal{W}^{n}_{t+T}\right.$
[…]
$\qquad\qquad+\log{\left(F^{n}_{0}\mathfrak{R}^{n}_{t+T}(\mathfrak{d})e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{0})}\right)}\overline{\mathfrak{a}}^{n\cdot}\mathbb{E}[\mathcal{A}_{t+T+a}^{\circ}|\mathcal{G}_{t}]$
[…]
By also developing $\mathbb{E}[\log{F_{t+T}^{n}}|\mathcal{G}_{t}]\mathbb{E}[\log{\mathcal{C}_{t+T}^{n}}|\mathcal{G}_{t}]$, we obtain
$\begin{split}&cov(\log{F_{t+T}^{n}},\log{\mathcal{C}_{t+T}^{n}}|\mathcal{G}_{t})=cov(\mathfrak{a}^{n\cdot}\mathcal{A}_{t+T}^{\circ},\overline{\mathfrak{a}}^{n\cdot}\mathcal{A}_{t+T}^{\circ}|\mathcal{G}_{t})\\ &\qquad=\varsigma^{2}\mathfrak{a}^{n\cdot}\Gamma^{-1}\left(\int_{0}^{T}\left(e^{-\Gamma u}-\mathbf{I}_{I}\right)\Sigma\Sigma^{\top}\left(e^{-\Gamma(u+a)}-\mathbf{I}_{I}\right)\mathrm{d}u\right)(\overline{\mathfrak{a}}^{n\cdot}\Gamma^{-1})^{\top}:=cv_{t,T,a}.\end{split}$
We obtain
$\begin{bmatrix}\log{\mathcal{V}_{t+T}^{n}}\\ \log{\mathcal{C}_{t+T}^{n}}\end{bmatrix}|\mathcal{G}_{t}\sim\mathcal{N}\left(\begin{bmatrix}\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})\\ \overline{\mathcal{K}}^{n}(\mathfrak{d},t,T+a,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})\end{bmatrix},\begin{bmatrix}\mathcal{L}^{n}(t,T)&cv_{t,T,a}\\ cv_{t,T,a}&\overline{\mathcal{L}}^{n}(t,T+a)\end{bmatrix}\right).$ (3.22)
Then, we use the Lemma 3.4 with $u=1$ to conclude the proof. ∎
We remark that the carbon price introduced in our economy affect both PD through the obligor cash flows and LGD through the collateral cash flows. See more remarks in Section 3.5.
3.3 When collateral is commercial or residential property
In this section, we assume that loans are backed by either residential or commercial building. The problem here is then to model the real estate market in the presence of the climate transition risk. The latter is represented by energy efficiency as well as the carbon price. We would like to compute the value of a dwelling at time time $t$. We use exactly as in Sopgoui, (2024) the actualized sum of the cash flows before the renovation date (taking into account the additional energy costs due to inefficiency of the building), at the renovation date, and after the renovation date (when the building becomes efficient). Moreover, the agent chooses rationally the date of renovation which maximizes the value of his property. Therefore, according to Sopgoui, (2024)[Theorem 2.4], we have the following proposition.
Theorem 3.9.
Assume that the following conditions are satisfied:
  1. 1.
    the carbon price function $\delta:t\mapsto\delta_{t}$ is non decreasing on $\mathbb{R}_{+}$ and deterministic;
  2. 2.
    the energy price $\mathfrak{f}(\cdot,\mathfrak{p})$ is non decreasing on $\mathbb{R}_{+}$ for all $\mathfrak{p}$.
Then, the market value of the building serving as the collateral to firm $n$ at $t\geq 0$, given the carbon price sequence $\delta$, is given by
$\mathcal{C}^{n}_{t,\delta}=C^{n}_{t}-R_{n}X_{t,\delta}^{n},$ (3.23)
where
$X_{t,\delta}^{n}:=\mathfrak{c}(\alpha^{n},\alpha^{\star})e^{-\bar{r}(\mathfrak{t}_{n}-t)}+(\alpha^{n}-\alpha^{\star})\int_{t}^{\mathfrak{t}_{n}}\mathfrak{f}(\delta_{u},\mathfrak{p})e^{-\bar{r}(u-t)}\mathrm{d}u,$ (3.24)
and where the optimal date of renovations $\mathfrak{t}_{n}\in[t,+\infty]$ is given by
$\mathfrak{t}_{n}=$ $t$$ if $\mathfrak{f}(\delta_{\theta},\mathfrak{p})-\bar{r}\frac{\mathfrak{c}(\alpha^{n},\alpha^{\star})}{\alpha^{n}-\alpha^{\star}}>0$ for all $\theta\in[t,\infty)$ (3.25)
$\mathfrak{t}_{n}=$ +$\infty$+ ∞ if $\mathfrak{f}(\delta_{\theta},\mathfrak{p})-\bar{r}\frac{\mathfrak{c}(\alpha^{n},\alpha^{\star})}{\alpha^{n}-\alpha^{\star}}<0$ for all $\theta\in[t,\infty)$ (3.26)
$\mathfrak{t}_{n}=$ $\theta^{\star}$ the unique solution of $\mathfrak{f}(\delta_{\theta},\mathfrak{p})=\bar{r}\frac{\mathfrak{c}(\alpha^{n},\alpha^{\star})}{\alpha^{n}-\alpha^{\star}}$ on $\theta\in[t,\infty)$. (3.27)
Moreover,
$C^{n}_{t}$ $:=R_{n}C^{n}_{0}e^{K_{t}},$ (3.28)
where
$\displaystyle\mathrm{d}K_{t}$ $=\left(\dot{\chi}_{t}+\nu(\chi_{t}-K_{t})\right)\mathrm{d}t+\overline{\sigma}\mathrm{d}\overline{B}_{t},$ (3.29a)
$\mathrm{d}\overline{B}_{t}$ $=\rho^{\top}\mathrm{d}B^{\mathcal{Z}}_{t}+\sqrt{1-\lVert\rho\rVert^{2}}\mathrm{d}\overline{\mathcal{W}}_{t},$ (3.29b)
with $(\overline{\mathcal{W}}_{t})_{t\in\mathbb{R}_{+}}$ is a standard Brownian motion independent to $B^{\mathcal{Z}}$ introduced in Standing Assumption 2.1 and driving the productivity of the economy. Moreover, $C^{n}_{0}$, $r$, $R_{n},\overline{\sigma}>0$, $\rho\in\mathbb{R}_{+}^{I}$, and $\chi\in\mathcal{C}^{1}(\mathbb{R}_{+},\mathbb{R}_{+})$. We introduce the following filtration $\mathbb{U}:=(\mathcal{U}_{t})_{t\in\mathbb{R}^{*}}$ with for $t\geq 0$, $\mathcal{U}_{t}:=\sigma\left(\left\{\overline{W}_{s},B^{\mathcal{Z}}_{s}:s\leq t\right\}\right)$.
The following corollary gives the conditional distribution of the collateral. Its proof is straightforward and is detailed in Sopgoui, (2024).
Corollary 3.10.
For $0\leq t\leq T$, the law of $C^{n}_{t+T}=R_{n}C_{0}^{n}\exp{(K_{t+T})}$ conditional on $\mathcal{G}_{t}$ is log-Normal $\mathcal{LN}(m_{t,T}^{n},v_{t,T}^{n})$ with
$m_{t,T}^{n}:=\log{(R_{n}C_{0}^{n})}+\chi_{t+T}-\left(\chi_{0}-K_{0}\right)e^{-\nu(t+T)}+\overline{\sigma}\rho^{\top}\int_{0}^{t}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s},$ (3.30)
and
$v_{t,T}^{n}:=\frac{(\overline{\sigma}\lVert\rho\rVert)^{2})}{2\nu}\left(1-e^{-2\nu T}\right)+\frac{(\overline{\sigma})^{2}(1-\lVert\rho\rVert^{2})}{2\nu}\left(1-e^{-2\nu(t+T)}\right).$ (3.31)
An example of the energy price function
We can assume that the price of each type of energy $\mathfrak{p}$ is a linear function of the carbon price, therefore
$\mathfrak{f}:(\delta_{t},\mathfrak{p})\mapsto\mathfrak{f}^{\mathfrak{p}}_{1}\delta_{t}+\mathfrak{f}^{\mathfrak{p}}_{0}\qquad t\geq 0,$ (3.32)
with $\mathfrak{f}^{\mathfrak{p}}_{1},\mathfrak{f}^{\mathfrak{p}}_{0}>0$ and $\delta$ is the carbon price defined in the Standing Assumption 2.3 or an example given in (2.12).
An example of the renovation costs function
We can consider that the costs of renovation of a dwelling $\mathfrak{c}$, to move its energy efficiency from $x$ to $y$, is
$\mathfrak{c}:(x,y)\mapsto c_{0}|x-y|^{1+c_{1}},$ (3.33)
with $c_{0}>0$ and $c_{1}\geq-1$. This choice of $\mathfrak{c}$ allows us to model that when a building has a bad energy efficiency, its renovation is costly.
An example of the optimal renovation time
With the example of the carbon price in (2.12), the example of the energy price in (3.32), and the example of the renovation costs in (3.33), the optimal renovation time, solution of (3.27) is given by
$\mathfrak{t}_{n}=t_{\circ}+\frac{1}{\eta_{\delta}}\log{\left(\frac{c_{0}r|\alpha^{n}-\alpha^{\star}|^{c_{1}}-\mathfrak{f}^{\mathfrak{p}}_{0}}{\mathfrak{f}^{\mathfrak{p}}_{1}P_{carbon}}\right)}.$ (3.34)
We can clearly remark that the optimal renovation date depends on the climate transition policy ($P_{carbon}$ and $\eta_{\delta}$), on the energy prices ($\mathfrak{f}^{\mathfrak{p}}_{0}$ and $\mathfrak{f}^{\mathfrak{p}}_{1}$), on the renovation costs ($c_{0}$ and $c_{1}$), and on the energy efficiencies ($\alpha^{n}$ and $\alpha^{\star}$).
By using the housing price under the climate transition as given in Proposition 3.9, we can then derive a precised expression of $\mathrm{LGD}$ when the collateral exists and is a building. We have:
Theorem 3.11.
Let $1\leq n\leq N$. When $a=0$ (no liquidation delay), the Loss Given Default of the obligor $n$ is over-indebted at time $t\in\mathbb{R}_{+}$, conditional on $\mathcal{G}_{t}$, is
$\mathrm{LGD}_{t,\delta}^{n}=(1-\gamma)\left[\left(1+(1-k)\frac{R_{n}X_{t,\delta}^{n}}{\mathrm{EAD}_{t}^{n}}\right)\Phi\left(\frac{w_{t}^{n}}{\sqrt{v_{t,0}^{n}}}\right)-\exp{\left(-w_{t}^{n}+\frac{1}{2}v_{t,t}^{n}\right)}\Phi\left(\frac{w_{t}^{n}}{\sqrt{v_{t,0}^{n}}}-\sqrt{v_{t,0}^{n}}\right)\right],$ (3.35)
where
$w_{t}^{n}:=\log{\left(\frac{\mathrm{EAD}_{t}^{n}}{(1-k)R_{n}}+X_{t,\delta}^{n}\right)}-m_{t,0}^{n},$ (3.36)
and with $m_{t,0}^{n}$ and $v_{t,t}^{n}$ defined in Corollary 3.10, and $X_{t,\delta}^{n}$ defined in (3.24).
We can also verify that when there is not collateral corresponding to $C^{n}_{0}=0$. We then have
$\begin{split}C^{n}_{0}\to 0&\implies\log{(R_{n}C^{n}_{0})}\to-\infty\implies m_{t,0}^{n}\to-\infty\implies w_{t}^{n}\to+\infty\implies\mathrm{LGD}_{t,\delta}^{n}\to 1-\gamma.\end{split}$
It is even worse when the costs associated with the transition explode, LGD also explodes as
$X_{t,\delta}^{n}\to+\infty\implies w_{t}^{n}\to+\infty\implies\mathrm{LGD}_{t,\delta}^{n}\to+\infty.$
Proof.
Let $t\geq 0$ and $1\leq n\leq N$. By remarking that
  1. 1.
    $\log{C^{n}_{t,\mathfrak{d}}}|\mathcal{G}_{t}\sim\mathcal{N}\left(m_{t,0}^{n},v_{t,0}^{n}\right)$, and
  2. 2.
    $C_{t,\mathfrak{d}}^{n}|\mathcal{G}_{t}$ and $\mathcal{V}_{t,\mathfrak{d}}^{n}|\mathcal{G}_{t}$ are independent,
  3. 3.
    and from (3.4b) when $a=0$, we have
    $\mathrm{LGD}_{t,\delta}^{n}$ $=(1-\gamma)\mathbb{E}\left[\left(1-(1-k)\frac{\mathcal{C}^{n}_{t,\delta}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right]$
    $=(1-\gamma)\mathbb{E}\left[\left(1+(1-k)\frac{R_{n}X^{n}_{t,\delta}}{\mathrm{EAD}_{t}^{n}}-(1-k)\frac{C^{n}_{t,\delta}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right],$
we can simply apply Lemma 3.3 with $u=1+(1-k)\frac{R_{n}X^{n}_{t,\delta}}{\mathrm{EAD}_{t}^{n}}$. ∎
Once again, we want to compute the (conditional) Loss Given Default of the entity $n$ at time $t$ on the horizon $T$. We can formalize that in the following proposition:
Proposition 3.12 (Projected LGD).
For each $t,T\geq 0$ and $1\leq n\leq N$, the (conditional) Loss Given Default of the entity $n$ at time $t$ on the horizon $T$, reads
$\begin{split}\mathrm{LGD}^{n}_{t,T,\delta}&=\frac{1-\gamma}{\mathrm{PD}^{n}_{t,T,\mathfrak{d}}}\left[\left(1+(1-k)e^{-ra}\frac{R_{n}X_{t+T+a,\delta}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}});\rho_{t,T,a}^{n}\right)\right.\\ &\qquad\qquad\quad\left.-\exp{\left(\frac{1}{2}v_{t,T+a}^{n}-\sqrt{v_{t,T+a}^{n}}\overline{\omega}_{t,T,a}^{n}\right)}\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n}-\sqrt{v_{t,T+a}^{n}},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}})-\rho_{t,T,a}^{n}\sqrt{v_{t,T+a}^{n}};\rho_{t,T,a}^{n}\right)\right],\end{split}$ (3.37)
where
$\rho_{t,T,a}^{n}:=\overline{\sigma}\varsigma\frac{\mathfrak{a}^{n\cdot}\Gamma^{-1}\left(\int_{0}^{T}e^{-\nu(u+a)}\left(\mathbf{I}_{I}-e^{-\Gamma u}\right)\Sigma\mathrm{d}u\right)\rho}{\sqrt{\mathcal{L}^{n}(t,T)\times v_{t,T+a}^{n}}},$
$\overline{w}_{t,T,a}^{n}:=\frac{\log{\left(\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)R_{n}e^{-ra}}+X_{t+T+a,\delta}^{n}\right)}-m_{t,T+a}^{n}}{\sqrt{v_{t,T+a}^{n}}},$
and with $m_{t,t+T+a}^{n}$ and $v_{t,t+T+a}^{n}$ defined in Corollary 3.10 and $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ defined in Proposition 2.9.
Proof.
Let $t,T\geq 0$ and $1\leq n\leq N$. From the beginning of the proof of Proposition 3.8, we have
$\mathbb{E}\left[L_{n,t+T}^{\mathbb{G}}\middle|\mathcal{G}_{t}\right]$ $=(1-\gamma)\mathrm{EAD}_{t+T}^{n}\mathbb{E}\left[\left(1-(1-k)e^{-ra}\frac{\mathcal{C}^{n}_{t+T+a,\mathfrak{d}}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right],$
then
$\mathbb{E}\left[L_{n,t+T}^{\mathbb{G}}\middle|\mathcal{G}_{t}\right]$ $=(1-\gamma)\mathrm{EAD}_{t+T}^{n}\mathbb{E}\left[\left(1+(1-k)e^{-ra}\frac{R_{n}X_{t+T+a,\delta}^{n}}{\mathrm{EAD}_{t+T}^{n}}-(1-k)e^{-ra}\frac{\mathcal{C}_{t+T+a,\delta}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right].$
Remark that
$\begin{split}\log{C_{0}^{n}e^{K_{t+T+a}}}&=\log{(C_{0}^{n})}+\chi_{t+T+a}-(\chi_{0}-K_{0})e^{-\nu(t+T+a)}\\ &\qquad+\overline{\sigma}\rho^{\top}\int_{0}^{t+T+a}e^{-\nu(t+T+a-s)}\mathrm{d}B_{s}^{\mathcal{Z}}+\overline{\sigma}\sqrt{1-\lVert\rho\rVert^{2}}\int_{0}^{t+T+a}e^{-\nu(t+T+a-s)}\mathrm{d}\overline{\mathcal{W}}_{s},\end{split}$
and recall that $\log{(\mathcal{V}^{n}_{t+T,\mathfrak{d}})}=\log{(F^{n}_{0}\mathfrak{R}^{n}_{t+T}(\mathfrak{d}))}+\mathfrak{a}^{n\cdot}(\mathcal{A}_{t+T}^{\circ}-v(\mathfrak{d}_{0}))+\sigma_{n}\mathcal{W}^{n}_{t+T}$. Therefore,
$\begin{split}cov\left(\log{C_{0}^{n}e^{K_{t+T+a}}},\log{F^{n}_{t+T}}\middle|\mathcal{G}_{t}\right)&=cov\left(\mathfrak{a}^{n\cdot}\mathcal{A}_{t+T}^{\circ},\overline{\sigma}\rho^{\top}\int_{0}^{t+T+a}e^{-\nu(t+T+a-s)}\mathrm{d}B_{s}^{\mathcal{Z}}\middle|\mathcal{G}_{t}\right)\\ &=\overline{\sigma}\mathfrak{a}^{n\cdot}cov\left(\mathcal{A}_{t+T}^{\circ},\int_{0}^{t+T+a}e^{-\nu(t+T+a-s)}\mathrm{d}B_{s}^{\mathcal{Z}}\middle|\mathcal{G}_{t}\right)\rho\\ &=\overline{\sigma}\varsigma\mathfrak{a}^{n\cdot}\Gamma^{-1}\left(\int_{0}^{T}e^{-\nu(u+a)}\left(\mathbf{I}_{I}-e^{-\Gamma u}\right)\Sigma\mathrm{d}u\right)\rho:=cv_{t,T,a}.\end{split}$
Consequently, we can write
$\begin{bmatrix}\log{F_{t+T}^{n}}\\ \log{C_{0}^{n}e^{K_{t+T+a}}}\end{bmatrix}|\mathcal{G}_{t}\sim\mathcal{N}\left(\begin{bmatrix}\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})\\ m_{t,T+a}^{n}\end{bmatrix},\begin{bmatrix}\mathcal{L}^{n}(t,T)&cv_{t,T,a}\\ cv_{t,T,a}&v_{t,T+a}^{n}\end{bmatrix}\right).$ (3.38)
Then, we use the Lemma 3.4 with $u=1+(1-k)e^{-ra}\frac{R_{n}X_{t+T+a,\delta}^{n}}{\mathrm{EAD}_{t+T}^{n}}$ to conclude the proof. ∎
We can remark that $\mathrm{LGD}^{n}_{t,\delta}$ as well as $\mathrm{LGD}^{n}_{t,T,\delta}$ are also functions of the optimal renovation time $\mathfrak{t}_{n}$. Furthermore, if both the financial asset and the housing price are affected by the climate transition through their dependence on the carbon price sequence $\delta$, the financial asset depends also on the carbon price intensities (of firms production/consumption and of households consumption) $(\tau,\zeta,\kappa)$ which are not specific to a given company but to the economy as a whole. The housing price is clearly affected by specific climate factors, namely the energy efficiency $\alpha^{n}$ and the renovation date $\mathfrak{t}_{n}$.
3.4 Expected and Unexpected losses
Let us recall that we have a portfolio with N loans. We assume that loans from 1 to $N_{1}$ are unsecured, loans from $N_{1}+1$ to $N_{2}$ are secured by a financial asset as collateral, and loans from $N_{2}+1$ to $N$ are secured by a commercial or residential property as collateral.
We write the expression of the portfolio EL and UL as functions of the parameters and of the processes introduced above, and introduce the entity’s probability of default.
We can therefore give expressions of EL and UL. Let $t,T\geq 0$, the (conditional) Expected Loss of the portfolio at time $t$ on the horizon $T$ defined in (3.5), reads
$\begin{split}\mathrm{EL}^{N,T}_{t}&=\mathbb{E}\left[\mathrm{L}^{\mathbb{G},N}_{t+T}\middle|\mathcal{G}_{t}\right]=\sum_{n=1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}\\ &=\sum_{n=1}^{N_{1}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}+\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}\\ &\qquad\qquad+\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\delta}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}.\end{split}$ (3.39)
We can then compute each term conditionally to $\mathcal{G}_{t}$.
  1. 1.
    Given that $\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}=1-\gamma$ for $1\leq n\leq N_{1}$, to compute $\sum_{n=1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$, all you have to do is calculate $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$.
  2. 2.
    Given that $N_{1}+1\leq n\leq N_{2}$, the collaterals are financial assets, therefore, to compute $\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$, we compute first $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$. Then we compute $\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}$ directly through (3.21).
  3. 3.
    Given that $N_{1}+1\leq n\leq N_{2}$, the collaterals are properties, therefore, to compute $\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\delta}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$, we compute first $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$. Then we compute $\mathrm{LGD}^{n}_{t,T,\delta}$ directly through (3.37).
For $\alpha\in(0,1)$, the (conditional) Unexpected Loss of the portfolio at time $t$ on the horizon $T$, cannot be obtained in closed-form as EL. Precisely, there is not a closed-form expression neither of $\mathrm{UL}^{\alpha,N,T}_{t}$ nor of $\mathrm{VaR}^{\alpha,N,T}_{t}$. But we can describe how to compute $\mathrm{VaR}^{\alpha,N,T}_{t}$ given that $\mathbb{P}\left[L_{t+T}^{\mathbb{G},N}\leq\mathrm{VaR}^{\alpha,N,T}_{t}\middle|\mathcal{G}_{t}\right]$ as introduced in (3.6). First, let us note that from Theorem 3.1, we have
$\begin{split}L_{t+T}^{\mathbb{G},N}&=\sum_{n=1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}\\ &=\sum_{n=1}^{N_{1}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}+\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}\\ &\qquad\qquad+\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\delta}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}.\end{split}$
We can then describe each term’s law conditionally to $\mathcal{G}_{t}$.
  1. 1.
    $\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}=1-\gamma$ and from (2.28), we have $\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ which depends on $\mathcal{A}_{t+T}$. Then to simulate law of $\sum_{n=1}^{N_{1}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ conditional on $\mathcal{G}_{t}$, just simulate $\mathcal{A}_{t+T}|\mathcal{G}_{t}$.
  2. 2.
    From (3.19), we have $\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}$ which depends on $\mathcal{A}_{t+T}$ through $w_{t+T}^{n}$ defined in (3.20). We said in the previous item that $\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ depends on $\mathcal{A}_{t+T}$. Therefore, to simulate law of $\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ conditional on $\mathcal{G}_{t}$, just simulate $\mathcal{A}_{t+T}|\mathcal{G}_{t}$.
  3. 3.
    From (3.35), we have $\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}$ which depends on $\int_{0}^{t+T}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s}$ through $w_{t+T}^{n}$ defined in (3.36). We said in the previous item that $\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ depends on $\mathcal{A}_{t+T}$. Therefore, to simulate law of $\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\delta}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ conditional on $\mathcal{G}_{t}$, just simulate $\mathcal{A}_{t+T}|\mathcal{G}_{t}$ and $\int_{0}^{t+T}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s}|\mathcal{G}_{t}$ (which are in fact the same because both $\mathcal{A}_{t+T}$ and $\int_{0}^{t+T}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s}$ are $\mathcal{G}_{t+T}$-measurable).
3.5 Remarks on the determinants of LGD
The results  (3.19) and (3.37) tell us that, in the case the collateral is an investment, Loss Given Default depends on:
  1. 1.
    the carbon price $\delta$ for both (3.19) and (3.37),
  2. 2.
    parameters specific to the company (the contract),
    • (a)
      the time $t$ when it is computed,
    • (b)
      the date of default $t+T$,
    • (c)
      the Exposure at Default $EAD$,
  3. 3.
    parameters specific to the collateral,
    • (a)
      its liquidation time $t+T+a$,
    • (b)
      the liquidation costs $k$,
    • (c)
      the correlation of its cash flows with the environment $\overline{\mathfrak{a}}$,
    • (d)
      the standard deviation of its cash flows $\overline{\sigma}_{\mathfrak{b}}$,
    • (e)
      the fraction of recovery from other means $\gamma$,
  4. 4.
    the nature of the collateral:
    • (a)
      if it is a financial asset, then parameters related to the carbon intensities $\tau,\zeta,\kappa$,
    • (b)
      if it is a building, then parameters related to the energy efficiency $\alpha$, type of energy $\mathfrak{p}$, and renovation costs $\mathfrak{c}$,
  5. 5.
    parameters specific to the economy to which the colateral belongs to:
    • (a)
      the (cumulative) productivity $\mathcal{A}$ (and its parameters) of the economy,
    • (b)
      the interest rates $r$ and $\bar{r}$.
Some of these typical risk drivers are reported by Chalupka and Kopecsni, (2008).
We could also look at the sensitivities of the LGD to each of these variables and parameters. However, the expressions of LGD we obtained are not very tractable so that it would be difficult to get detailed expressions of theses sensitivities. If necessary, they can be calculated using numerical methods.
4 Numerical experiments, estimation and calibration
In this section, we describe how the parameters of multisectoral model, of the firm valuation model, and of the credit risk model are estimated given the historical macroeconomic variables (consumption, labour, output, GHG emissions, housing prices, etc.) as well as the historical credit portfolio data (firms rated and defaulted, collateral, etc.) In a second step, we give the expression of the risk measures (PD, LGD, EL, and UL) introduced in the previous sections, that we compute using Monte Carlo simulations.
4.1 Calibration and estimation
We will calibrate the model parameters on a set of data ranging from time $\mathfrak{t}_{0}$ to $\mathfrak{t}_{1}$. In practice, $\mathfrak{t}_{0}=1978$ and $\mathfrak{t}_{1}=t_{\circ}=2021$. From now on, we will discretize the observation interval into $M\in\mathbb{N}^{*}$ steps $t_{m}=\mathfrak{t}_{0}+\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}m$ for $0\leq m\leq M$. We note $\mathcal{T}^{M}:=\{t_{0},t_{1},\ldots,t_{M}\}$. We will not be interested in convergence results here.
4.1.1 Estimation of carbon intensities
For each sector $i\in\mathcal{I}$ and for $0\leq m\leq M$, we observe the output $Y_{t_{m}}^{i}$, the labor $N_{t_{m}}^{i}$, the intermediary input $(Z_{t_{m}}^{ji})_{j\in\mathcal{I}}$, and the consumption $C_{t_{m}}^{i}$ (recall that the transition starts at year $t_{\circ}$). For the sake of clarity, we will omit the dependence of each estimated parameter on $M$.
To calibrate each carbon intensity $\mathfrak{y}\in\{\tau^{1},\ldots,\tau^{I},\zeta^{11},\zeta^{12},\ldots,\zeta^{II-1},\zeta^{II},\kappa^{1},\ldots,\kappa^{I}\}$, we follow exactly the same process already presented in Bouveret et al., (2023). The main difference is that after calibration, we can compute $\mathfrak{y}$ for each $t\in\mathbb{R}_{+}$. Afterwards, if we consider the example of the carbon price introduces in (2.12), we can compute the emissions cost rate $\widehat{\mathfrak{d}}_{t}$.
4.1.2 Estimation of economic parameters
As in Galí, (2015), we assume a unitary Frisch elasticity of labor supply so $\varphi=1$ and the utility of consumption is logarithmic so $\sigma=1$, while we calibrate $({\boldsymbol{\lambda}}^{ij})_{i,j\in\mathcal{I}}$ and $({\chi}^{i})_{i\in\mathcal{I}}$in the same way as in Bouveret et al., (2023). We can then compute the functions $\chi$ and $\Lambda$ defined in Proposition 2.4, followed by the function $\widehat{v}^{i}$ as defined in (2.16). We can also compute the output growth $\left(\Delta^{{Y}}_{t_{m}}=(\log({Y}_{t_{m}}^{i})-\log({Y}_{t_{m-1}}^{i}))_{j\in\mathcal{I}}\right)_{1\leq m\leq M}$ directly from data.
Without carbon tax in any sector, it follows from (2.17) in Corollary 2.4 that, for each $1\leq m\leq M$, the computed consumption growth $\Delta^{{Y}}_{t_{m}}$ is equal to $\Delta^{{Y}}_{t_{m}}=\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}(\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}})^{-1}\widehat{\Theta}_{t_{m}}$ when $\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}}$ is not singular; hence $\widehat{\Theta}_{t_{m}}=\frac{M}{\mathfrak{t}_{1}-\mathfrak{t}_{0}}(\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}})\Delta^{{Y}}_{t_{m}}$. We can then compute the estimations $\widehat{\mu}$, $\widehat{\Gamma}$, $\widehat{\Sigma}$ and $\widehat{\varsigma}$, parameters $\mu$, $\Gamma$, $\Sigma$, and $\varsigma$ (all defined in Standing Assumption 2.1), as detailed in Sopgoui, (2024)[Section 3.1.1.].
4.1.3 Estimation of firm and of the credit model parameters
Recall that we have a portfolio with $N\in\mathbb{N}^{*}$ firms (or credit) at time $t_{\circ}$. For each firm $n\in\{1,\ldots,N\}$, we have its historical cash flows $(F_{t_{m}}^{n})_{1\leq m\leq M}$, hence its log-cash flow growths. For any $t\in\mathcal{T}^{M}$ and $1\leq i\leq I$, we denote by $r_{t}^{i}$ (resp. $d_{t}^{i}$) the number of firms in $g_{i}$ rated at the beginning of the year $t$ (resp. defaulted during the year $t$). In particular, $r_{\mathfrak{t}_{0}}=\#g_{i}$. Within each group $g_{i}$, all the firms behave in the same way as there is only one risk class. Since each sub-portfolio constitutes a single risk class, we have for each $n\in g_{m}$, $\mathfrak{a}^{n}=\mathfrak{a}^{n_{i}}$, $\sigma_{\mathfrak{b}^{n}}=\sigma_{\mathfrak{b}^{n_{i}}}$, and $B^{n}=B^{n_{i}}$. We then proceed as follows:
  1. 1.
    Knowing the output growth $\left(\Delta^{{Y}}_{t}\right)_{t\in\mathcal{T}^{M}}$, we calibrate the factor loading $\mathfrak{a}_{n_{i}}$ and the standard deviation $\sigma_{n_{i}}$, according to Assumption 2.5, appealing to the regression
    $\sum_{n\in g_{i}}\log{F_{t_{m}}^{n}}-\log{F_{t_{m-1}}^{n}}=(\#g_{i})\mathfrak{a}^{n_{i}}\Delta^{{Y}}_{t_{k}}+\sqrt{\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}\#g_{i}}\sigma_{\mathfrak{b}^{n_{i}}}\mathfrak{u}_{t_{m}}\quad\text{where}\quad\mathfrak{u}_{t_{m}}\sim\mathcal{N}(0,1),\quad\forall\quad 1\leq m\leq M.$ (4.1)
  2. 2.
    We then estimate the barrier $B^{n_{i}}$ by MLE as detailed in (Gordy and Heitfield, , 2002, Section 3):
    we compute
    $\widehat{B}^{n_{i}}:=\operatorname*{arg\,max}_{B^{n_{i}}\in\mathbb{R}^{+}}\mathcal{L}(B^{n_{i}}),$
    where $\mathcal{L}(B^{n_{i}})$ is the log-likelihood function defined by
[…]
    and where
    $\mathbb{P}[D^{n_{i}}=d_{t_{m}}^{i}|(\mathcal{A}_{t_{m}}^{\circ},\mathcal{Z}_{t_{m}})]={\binom{r_{t_{m}}^{i}}{d_{t_{m}}^{i}}}(\mathrm{PD}^{n_{i}}_{t_{m},1,0})^{d_{t_{m}}^{i}}\Big{(}1-\mathrm{PD}^{n_{i}}_{t_{m},1,0}\Big{)}^{r_{t_{m}}^{i}-d_{t_{m}}^{i}},$
    with $D^{n_{i}}$ the Binomial random variable standing for the conditional number of defaults, and $\mathrm{PD}^{n_{i}}_{t_{m},1,0}$ in Proposition 2.29, depending on $\sigma_{\mathfrak{b}^{n_{i}}}=\widehat{\sigma}_{\mathfrak{b}^{n_{i}}}$, $\mathfrak{a}^{n_{i}}=\widehat{\mathfrak{a}}^{n_{i}}$, for $1\leq m\leq M$, $\delta_{t_{m}}=0$ and on $B^{n_{i}}$.
4.1.4 Calibration of collateral
Recall that we have a portfolio with $N\in\mathbb{N}^{*}$ firms (or credit) at time $t_{\circ}$. For each firm $n\in\{1,\ldots,N\}$, if the collateral is
A financial asset
We have its historical cash flows $(\overline{F}_{t_{m}}^{n})_{0\leq m\leq M}$, hence its log-cash flow growths. Recall that, even if two firms belong to the same sub-portfolio, there is no reason that their collaterals behave in the same way. We also know the output growth $\left(\Delta^{{Y}}_{t_{m}}\right)_{1\leq m\leq M}$. We then have,
  1. 1.
    the proportion $\alpha^{n}$ of the investment representing the collateral is known.
  2. 2.
    we calibrate the factor loading $\widehat{\overline{\mathfrak{a}}}_{n}$ and the standard deviation $\widehat{\overline{\sigma}}_{n}$, according to (3.12), appealing to the regression
    $\log{\overline{F}_{t_{m}}^{n}}-\log{\overline{F}_{t_{m-1}}^{n}}={\overline{\mathfrak{a}}}^{n}\Delta^{{Y}}_{t_{m}}+\sqrt{\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}}\overline{\sigma}_{n}\mathfrak{u}_{t_{m}}\quad\text{where}\quad\mathfrak{u}_{t_{m}}\sim\mathcal{N}(0,1),\quad\text{for all}\quad 1\leq m\leq M.$ (4.2)
A commercial or residential property
We assume that in the past, carbon price did not have impact on the dwelling price so that for all $t\in\mathcal{T}^{M}$, $X_{t,\delta}^{n}$ defined in (3.24) is zero. Moreover, $C^{n}_{0}$ and $R_{n}$ defined in (3.28), the value of the collateral at 00 and the surface, are known. All that remains is to calibrate the parameters of the process $K$ defined in (3.29a) and (3.29b). Let us consider a real estate index $(REI_{t_{m}})_{0\leq m\leq M}$. We assume that the long-term average of the real estate index $\chi$, introduced in (3.29a) is linear and for $t\in\mathbb{R}_{+}$, $\chi_{t}=\varrho t+\vartheta$. We then estimate $\varrho$, $\vartheta$, $\nu$, $\overline{\sigma}$, and $\rho$ as Sopgoui, (2024)[Section 3.1.2].
4.2 Simulations
In this section as well, the idea here is not to (re)demonstrate or improve convergence results.
4.2.1 Of the productivities $\mathcal{Z}$ and $\mathcal{A}$
Let $K~{}\in\mathbb{N}$, for $0\leq k\leq K$, we note $u_{k}=t_{\circ}+\frac{t_{\star}-t_{\circ}}{K}k$ for $0\leq k\leq K$. We would like to simulate $\mathcal{Z}_{u_{k}}$ and $\mathcal{A}_{u_{k}}$. For $\mathcal{Z}$, we adopt the Euler-Maruyana Maruyama, (1955); Kanagawa, (1988) scheme as in Sopgoui, (2024)[Section 3.2.1].
4.2.2 Of the probability of over-indebtedness PD and of LGD
For $n\in\{1,\ldots,N\}$ and $t_{\circ}\leq t\leq t_{\star}$, We would like to compute $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ as defined in (2.29) as well as $\mathrm{LGD}_{t,T,\mathfrak{d}}^{n}$ defined in (3.21) and $\mathrm{LGD}_{t,T,\delta}^{n}$ in (3.37). After simulating $\mathcal{Z}_{t}$ and $\mathcal{A}_{t}$ as described in 4.2.1, we get $\widehat{\mathcal{Z}}_{t}$ and $\widehat{\mathcal{A}}_{t}$. Then, for each $1\leq i\leq I$ and for each $n\in g_{i}$, we have
  1. 1.
    from (2.29), the estimated probability of default of firm $n$ is
    $\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}=\Phi\left(\frac{\log(\mathcal{D}_{t+T}^{n})-\widehat{\mathcal{K}}^{n}(\widehat{\mathfrak{d}},t,T,\widehat{\mathcal{A}}_{t},\widehat{\mathcal{Z}}_{t})}{\widehat{\mathcal{L}}^{n}(t,T)}\right),$ (4.3)
    with
    $\widehat{\mathcal{K}}^{n}(\widehat{\mathfrak{d}},t,T,\widehat{\mathcal{A}}_{t},\widehat{\mathcal{Z}}_{t})=\log{(F^{n}_{0}\widehat{\mathfrak{R}}^{n}_{t+T}(\widehat{\mathfrak{d}}))}+\widehat{\mathfrak{a}}^{n\cdot}(\widehat{\mu}T+\widehat{\varsigma}\Upsilon_{T}\widehat{\mathcal{Z}}_{t}+\widehat{\mathcal{A}}_{t}-\widehat{v}(\widehat{\mathfrak{d}}_{0})),$ (4.4)
    and
    ^$\widehat{\mathcal{L}}^{n}(t,T):=\widehat{\varsigma}^{2}\widehat{\mathfrak{a}}^{n\cdot}\widehat{\Gamma}^{-1}\left(\frac{T}{L}\sum_{l=0}^{L}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\widehat{\Sigma}\widehat{\Sigma}^{\top}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\right)(\widehat{\mathfrak{a}}^{n\cdot}\widehat{\Gamma}^{-1})^{\top}+(t+T)\widehat{\sigma}_{n}^{2},$ (4.5)
    where $F^{n}_{0}$ and $\mathcal{D}_{t+T}^{n}$ are know, $\widehat{\mathfrak{d}}$ defined in Section 4.1.1, ^$\widehat{\mathfrak{R}}^{n}_{t+T}(\widehat{\mathfrak{d}}))$ in (2.22) in Theorem 2.6, $\widehat{\Gamma},\widehat{\varsigma},\widehat{v}$ in Section 4.1.2, $\widehat{\mathfrak{a}}^{n\cdot},\widehat{\sigma}^{n}$ in Section 4.1.3, $\widehat{\Upsilon}_{T}:=\widehat{\Gamma}^{-1}(\mathbf{I}_{I}-e^{-\widehat{\Gamma}T})$ and with $u_{l}:=\frac{Tl}{L},l=0,\ldots,L$.
  2. 2.
    If the collateral of loan $n$ is a financial asset, from (3.21),
    $\begin{split}\widehat{\mathrm{LGD}}^{n}_{t,T,\mathfrak{d}}&=\frac{1-\gamma}{\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}}\left[\Phi_{2}\left(\widehat{\overline{\omega}}_{t,T,a}^{n},\Phi^{-1}\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}});\widehat{\rho}_{t,T,a}^{n}\right)-\exp{\left(\frac{1}{2}\widehat{\overline{\mathcal{L}}}^{n}(t,T+a)\right)}\times\right.\\ &\qquad\qquad\qquad\left.\Phi_{2}\left(\widehat{\overline{\omega}}_{t,T,a}^{n}-\sqrt{\widehat{\overline{\mathcal{L}}}^{n}(t,T+a)},\Phi^{-1}(\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}})-\widehat{\rho}_{t,T,a}^{n}\sqrt{\widehat{\overline{\mathcal{L}}}^{n}(t,T+a)};\widehat{\rho}_{t,T,a}^{n}\right)\right],\end{split}$ (4.6)
    where $\widehat{\overline{\mathcal{K}}}$ and $\widehat{\overline{\mathcal{L}}}$ are computed in the same way $\widehat{{\mathcal{K}}}$ and $\widehat{{\mathcal{L}}}$ were in (4.4) and (4.5). Moreover,
    $\widehat{\rho}_{t,T,a}^{n}:=\frac{\widehat{\varsigma}^{2}\widehat{\mathfrak{a}}^{n\cdot}\widehat{\Gamma}^{-1}\left(\frac{T}{L}\sum_{l=0}^{L}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\widehat{\Sigma}\widehat{\Sigma}^{\top}\left(e^{-\widehat{\Gamma}(u_{l}+a)}-\mathbf{I}_{I}\right)\right)(\widehat{\overline{\mathfrak{a}}}^{n\cdot}\widehat{\Gamma}^{-1})^{\top}}{\sqrt{\widehat{\mathcal{L}}^{n}(t,T)\widehat{\overline{\mathcal{L}}}^{n}(t,T+a)}},$
    and
    $\widehat{\overline{\omega}}_{t,T,a}^{n}:=\frac{\log{\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}}-\widehat{\overline{\mathcal{K}}}^{n}(\widehat{\mathfrak{d}},t,T+a,\widehat{\mathcal{A}}_{t}^{\circ},\widehat{\mathcal{Z}}_{t})}{\sqrt{\widehat{\overline{\mathcal{L}}}^{n}(t,T+a)}},$
[…]
  3. 3.
    If the collateral of loan $n$ is a commercial or residential property, we compute in order (3.30), (3.31), (3.36), and (3.37). Since $\chi_{t}=\varrho t+\vartheta$ and $C_{0}^{n}$ are known
    $\widehat{m}_{t,T+a}^{n}:=\log{(R_{n}C_{0}^{n})}+(\widehat{\varrho}t+\widehat{\vartheta})-(\widehat{\vartheta}-K_{0})e^{-\widehat{\nu}(t+T+a)}+\widehat{\overline{\sigma}}\widehat{\rho}^{\top}\sum_{k=0}^{L}e^{-\widehat{\nu}((t+T+a)-\frac{kt}{L})}\eta_{u_{\frac{kT}{L}}},$
    where $\eta_{\frac{kt}{L}}\sim\mathcal{N}\left(0,\frac{t}{L}\mathbf{I}_{I}\right),k=0,\ldots,L$ with $L\in\mathbb{N}^{*}$, and
    $\widehat{v}_{t,T+a}^{n}:=\frac{(\widehat{\overline{\sigma}}\widehat{\rho})^{2})}{\widehat{\nu}}\left(1-e^{-2\widehat{\nu}(T+a)}\right)+\frac{(\widehat{\overline{\sigma}})^{2}(1-(\widehat{\rho})^{2})}{\widehat{\nu}}\left(1-e^{-2\widehat{\nu}(t+T+a)}\right).$
    Therefore, we have
    $\begin{split}\widehat{\mathrm{LGD}}^{n}_{t,T,\delta}&=\frac{1-\gamma}{\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}}\left[\left(1+(1-k)e^{-ra}\frac{R_{n}\widehat{X}_{t+T+a,\delta}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)\Phi_{2}\left(\widehat{\overline{w}}_{t,T,a}^{n},\Phi^{-1}(\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}});\widehat{\rho}_{t,T,a}^{n}\right)\right.\\ &\qquad\qquad\quad\left.-\exp{\left(\frac{1}{2}\widehat{v}_{t,T+a}^{n}-\sqrt{\widehat{v}_{t,T+a}^{n}}\widehat{\overline{w}}_{t,T,a}^{n}\right)}\Phi_{2}\left(\widehat{\overline{w}}_{t,T,a}^{n}-\sqrt{\widehat{v}_{t,T+a}^{n}},\Phi^{-1}(\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}})-\widehat{\rho}_{t,T,a}^{n}\sqrt{\widehat{v}_{t,T+a}^{n}};\widehat{\rho}_{t,T,a}^{n}\right)\right],\end{split}$ (4.7)
    where
    $\widehat{\rho}_{t,T,a}^{n}:=\widehat{\overline{\sigma}}\widehat{\varsigma}\frac{\widehat{\mathfrak{a}}^{n\cdot}\widehat{\Gamma}^{-1}\left(\frac{T}{L}\sum_{l=0}^{L}e^{-\widehat{\nu}(u_{l}+a)}\left(\mathbf{I}_{I}-e^{-\widehat{\Gamma}u_{l}}\right)\widehat{\Sigma}\right)\widehat{\rho}}{\sqrt{\widehat{\mathcal{L}}^{n}(t,T)\times\widehat{v}_{t,t+T+a}^{n}}},$
    and
    $\widehat{\overline{w}}_{t,T,a}^{n}:=\log{\left(\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)R_{n}e^{-ra}}+\widehat{X}_{t+T+a,\delta}^{n}\right)}-\widehat{m}_{t,T+a}^{n},$
    and $\widehat{X}$ is obtained by considering that from (3.24),
    $\widehat{X}_{t,\delta}^{n}=\mathfrak{c}(\alpha^{n},\alpha^{\star})e^{-r(\mathfrak{t}_{n}-t)}+(\alpha^{n}-\alpha^{\star})\frac{(\mathfrak{t}_{n}-t)}{P}\sum_{p=1}^{P}\mathfrak{f}(\delta_{v_{p}},\mathfrak{p})e^{-r(v_{p}-t)},$
    and where $\gamma$, $k$, $r$, $R_{n}$, and $\mathrm{EAD}_{t}^{n}$ are known, $\mathfrak{t}_{n}$ given by (3.34), $u_{l}:=\frac{(t_{\star}-t)l}{L},l=0,\ldots,L$, and $v_{p}:=\frac{(\mathfrak{t}_{n}-t)p}{P},l=0,\ldots,P$.
4.2.3 Of the (un)expected losses EL and UL
For EL, the result is direct by using $\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}$ in (4.3), $\widehat{\mathrm{LGD}}^{n}_{t,T,\mathfrak{d}}$ in (4.6), and $\widehat{\mathrm{LGD}}^{n}_{t,T,\delta}$ in (4.7), we have from (3.39),
$\begin{split}\widehat{\mathrm{EL}}^{N,T}_{t}&:=\sum_{n=1}^{N_{1}}(1-\gamma)\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}+\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}^{n}_{t,T,\mathfrak{d}}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}\\ &\qquad\qquad+\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}^{n}_{t,T,\delta}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}.\end{split}$ (4.8)
For UL, we use
$\begin{split}\widehat{L}_{t+T}^{\mathbb{G},N}&=\sum_{n=1}^{N_{1}}(1-\gamma)\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{PD}}_{t+T,\mathfrak{d}}^{n}+\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}_{t+T,\mathfrak{d}}^{n}\cdot\widehat{\mathrm{PD}}_{t+T,\mathfrak{d}}^{n}\\ &\qquad\qquad+\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}_{t+T,\delta}^{n}\cdot\widehat{\mathrm{PD}}_{t+T,\mathfrak{d}}^{n},\end{split}$ (4.9)
by noting that $\mathrm{PD}_{t+T,\mathfrak{d}}^{n}=\mathrm{PD}^{n}_{t+T,0,\mathfrak{d}}$, $\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}=\mathrm{LGD}^{n}_{t+T,0,\mathfrak{d}}$, and $\mathrm{LGD}_{t+T,\delta}^{n}=\mathrm{LGD}^{n}_{t+T,0,\delta}$. Therefore, as $\widehat{L}_{t+T}^{\mathbb{G},N}$ depends on $(\widehat{\mathrm{PD}}_{t+T,\mathfrak{d}}^{n},\widehat{\mathrm{LGD}}_{t+T,\mathfrak{d}}^{n},\widehat{\mathrm{LGD}}_{t+T,\delta}^{n})$ which depends on $(\widehat{\mathcal{A}}_{t+T},\widehat{\mathcal{Z}}_{t+T})$. However, we want to compute $\mathrm{VaR}^{\alpha,N,T}_{t}$ so that $\mathbb{P}\left[L_{t+T}^{\mathbb{G},N}\leq\mathrm{VaR}^{\alpha,N,T}_{t}\middle|\mathcal{G}_{t}\right]$. Then, we simulate $D\in\mathbb{N}^{*}$ couples noted $(\widehat{\mathcal{A}}_{t+T|t}^{p},\widehat{\mathcal{Z}}_{t+T|t}^{p})_{1\leq p\leq D}$ so that $\widehat{\mathcal{Z}}_{t+T|t}^{p}=^{d}\mathcal{Z}_{t+T}|\mathcal{G}_{t}$ and $\widehat{\mathcal{A}}_{t+T|t}^{p}=^{d}\mathcal{A}_{t+T}|\mathcal{G}_{t}$. That is straightforward and
$\widehat{\mathcal{Z}}_{t+T|t}^{p}|\mathcal{G}_{t}\sim\mathcal{N}\left(e^{-\widehat{\Gamma}T}\widehat{\mathcal{Z}}_{t},\frac{T}{L}\sum_{l=0}^{L}e^{-\widehat{\Gamma}u_{l}}\widehat{\Sigma}\widehat{\Sigma}^{\top}e^{-\widehat{\Gamma}^{\top}u_{l}}\right),$
and
$\widehat{\mathcal{A}}_{t+T|t}^{p}|\mathcal{G}_{t}\sim\mathcal{N}\left(\widehat{\mu}T+\widehat{\varsigma}\widehat{\Upsilon}_{T}\widehat{\mathcal{Z}}_{t}+\widehat{\mathcal{A}}_{t},\widehat{\varsigma}^{2}\widehat{\Gamma}^{-1}\left[\frac{T}{L}\sum_{l=0}^{L}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\widehat{\Sigma}\widehat{\Sigma}^{\top}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\right](\widehat{\Gamma}^{-1})^{\top}\right),$
with $u_{l}:=\frac{Tl}{L},l=0,\ldots,L$. We also need to simulate $\mathfrak{h}_{t+T}|\mathcal{G}_{t}:=\int_{0}^{t+T}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s}|\mathcal{G}_{t}$ (which comes from $m_{t+T,0}^{n}$ in (3.30)). As $\mathfrak{h}_{t+T}|\mathcal{G}_{t}\sim\mathcal{N}\left(\int_{0}^{t}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s},\frac{1-e^{-2\nu T}}{2\nu}\mathbf{I}_{I}\right)$, we have
$\widehat{\mathfrak{h}}_{t+T}^{p}|\mathcal{G}_{t}\sim\mathcal{N}\left(\sum_{k=0}^{L}e^{-\widehat{\nu}((t+T)-\frac{kt}{L})}\eta_{u_{\frac{kT}{L}}},\frac{1-e^{-2\widehat{\nu}T}}{2\widehat{\nu}}\mathbf{I}_{I}\right),\quad\eta_{\frac{kt}{L}}\sim\mathcal{N}\left(0,\frac{t}{L}\mathbf{I}_{I}\right).$
Then, the unexpected loss is
$\widehat{\mathrm{UL}}^{N,T}_{t,\delta,\alpha}:=q_{\alpha,D}\left(\left\{(\widehat{L}_{t+T}^{\mathbb{G},N})^{1},(\widehat{L}_{t+T}^{\mathbb{G},N})^{2},\ldots,(\widehat{L}_{t+T}^{\mathbb{G},N})^{D}\right\}\right)-\widehat{\mathrm{EL}}^{N,T}_{t,\mathfrak{d}},$ (4.10)
where $(\widehat{L}_{t+T}^{\mathbb{G},N})^{p}$ is obtained by replacing $(\widehat{\mathcal{A}}_{t+T},\widehat{\mathcal{Z}}_{t+T})$ in (4.9) by $(\widehat{\mathcal{A}}_{t+T|t}^{p},\widehat{\mathcal{Z}}_{t+T|t}^{p})$, and where $q_{\alpha,M}(\{Y^{1},\ldots,Y^{D}\})$ denotes the empirical $\alpha$-quantile of the distribution of $Y$.
5 Discussion
In this section, we describe the data used to calibrate the different parameters, we perform some simulations, and we comment the results.
5.1 Data
As in Bouveret et al., (2023) and Sopgoui, (2024), we work on data related to the French economy.
  1. 1.
    Due to data availability (precisely, we do not find public monthly/quaterly data for the intermediary inputs), we consider an annual frequency.
  2. 2.
    Annual consumption, labor, output, and intermediary inputs come from INSEE [^1] The French National Institute of Statistics and Economic Studies from 1978 to 2021 (see INSEE, (2023) for details) and are expressed in billion euros, therefore $\mathfrak{t}_{0}=1978$, $\mathfrak{t}_{1}=2021$, and $M=44$.
  3. 3.
    For the climate transition, we consider a time horizon of ten years with $t_{\circ}=2021$ as starting point, a time step of one year and $t_{\star}=2030$ as ending point. In addition, we will be extending the curves to 2034 to see what happens after the transition, even though the results will be calculated and analyzed during the transition.
  4. 4.
    The 38 INSEE sectors are grouped into four categories: Very High Emitting, Very Low Emitting, Low Emitting, and High Emitting, based on their carbon intensities.
  5. 5.
    The carbon intensities are calibrated on the realized emissions from Eurostat, (2023) (expressed in tonnes of CO2-equivalent) between 2008 and 2021.
  6. 6.
    Metropolitan France housing price index comes from OECD data and are from 1980 to 2021 (see OECD Stat, (2024) for details) in Base 2015. We renormalize in Base 2021.
5.2 Definition of the climate transition
We consider four deterministic transition scenarios giving four deterministic carbon price trajectories. The scenarios used come from the NGFS simulations, whose descriptions are given by NGFS, (2022) as follows:
  • 1.
    Net Zero 2050 is an ambitious scenario that limits global warming to $1.5^{\circ}C$ through stringent climate policies and innovation, reaching net zero $\mathrm{CO}_{2}$ emissions around 2050. Some jurisdictions such as the US, EU and Japan reach net zero for all GHG by this point.
  • 2.
    Divergent Net Zero reaches net-zero by 2050 but with higher costs due to divergent policies introduced across sectors and a quicker phase out of fossil fuels.
  • 3.
    Nationally Determined Contributions (NDCs) includes all pledged policies even if not yet implemented.
  • 4.
    Current Policies assumes that only currently implemented policies are preserved, leading to high physical risks.
For each scenario, we compute the carbon price $P_{carbon,0}$ in $t_{0}$ and the evolution rate $\eta_{\delta}$ as defined in (2.9).
Current Policies NDCs Divergent Net Zero Net Zero 2050
$P_{carbon,0}$ (in euro/ton) 30.957 33.321 32.963 34.315
$\eta_{\delta}$ (in %) 1.693 7.994 12.893 17.935
Table 1: Carbon price parameters
We can then compute the carbon price, whose evolution is plotted in Figure 1(a), at each date using (2.12).
For the energy price, we consider electricity as the unique source of energy. Then, we assume a linear relation between the electricity and the carbon price inspired by Abrell et al., (2023), where a variation of the carbon price is linked withe the variation of the electricity by a the pass-through rate noted $k$. This means that $\mathfrak{f}^{elec}_{1}$ and $\mathfrak{f}^{elec}_{2}$ define in (3.32) are respectively $k$ and $P_{elec,0}-k\times P_{carbon,0}$. For France, we take the electricity price $P_{elec,0}=0.2161$ euro per Kilowatt-hour and $k=0.55$ (see Abrell et al., (2023)) ton per Kilowatt-hour. Its evolution is plotted in Figure 1(b).
(a) Carbon price
(b) Energy price
Figure 1: per scenario and per year
For the renovation costs to improve a building for the energy efficiency $\alpha$ to $\alpha^{\star}$ as defined in (3.33), we take $c_{0}=0.01$ euro per kilowatt-hour and per square meter (€/KWh.m2) and $c_{1}=0.1$.
5.3 Estimations
5.3.1 The carbon intensities
We use the realized GHG emissions as well as the macroeconomic variables and their frequency being the same as in Bouveret et al., (2023), we use the same estimations. But after that, we can compute the carbon intensities at each date in $\mathbb{R}_{+}$ using (2.9).
5.3.2 Economic and housing pricing index (HPI) parameters
We keep the values of $\phi$, $\sigma$, $(\chi^{i})_{i\in\mathcal{I}}$, and $(\boldsymbol{\lambda}^{ji})_{i,j\in\mathcal{I}}$ already estimated. For the productivity process, we switched from a vector autoregressive model to an Ornstein-Uhlenbeck. We therefore calibrate ${\mu}$, ${\varsigma}$, ${\Sigma}$, and ${\Gamma}$ as detailed in Section 4.1.2 and we obtain the same results as in Sopgoui, (2024)[Section 4.3.1.].
We write the housing price index $K$ in Base 2021 and we apply the logarithm function. This means that $K_{t_{0}}=0$. We can therefore calibrate $\varrho,\vartheta,\nu,\overline{\sigma}$, and $\rho$. The values are presented in Sopgoui, (2024)[Table 5].
5.4 Simulations and discussions
In the previous work in discrete time, we simulate for different climate transition scenario between $t_{\circ}=2021$ and $t_{\star}=2030$, the annual evolution of (1) the output growth per sector (2) the output share per sector in the total output, (3) the firms direct GHG emissions per sector, (4) a given firm value and distribution, (5) the probabilities of default of fictive sub-portfolio of 4 firms each and of the resulting portfolio, (6) the expected and the unexpected losses of the previous (sub-)portfolios when the LGD are constant and deterministic, (7) the sensitivities of the losses to the carbon price.
In the current simulations, since we are keeping the same data at the same frequency (annual), the main change is then to replace the VAR process by the O.-U. process. Therefore, the comments already made for (1) to (5) concerning the trends, the impact of the carbon price, the difference of scenarios, the relation between sectors, etc. do not change. We will focus here on the LGD and on the losses, with different type of collateral.
5.4.1 Impact of the carbon price on Loss Given Default
When there is no guarantee, we assume as in the previous work that $\mathrm{LGD}$ is equal to $45\%$ so that $\gamma=0.55$. To illustrate the case where there is guarantee, we consider, both if the collateral is a financial asset and a building, $\mathrm{EAD}$ starts at $200$ and growths annually as the economic total output growth in the Current Policies scenario (see Table 2 below).
Year 2021 2022 2023 2024 2025 2026 2027 2028 2029 2030 2031 2032 2033 2034
EAD 200. 202.8 206.6 209.9 213.1 216.2 219.5 222.6 226.2 229.6 233.1 236.3 239.9 243.9
Table 2: EAD per year
If the collateral is a financial asset
We consider 4 firms so that firm $1$, $2$, $3$, and $4$ respectively belong to the Very High Emitting, High Emitting, Low Emitting, and Very Low Emitting groups. Each firm is characterized by its cash flows $F_{t_{\circ}-1}$ at $t_{\circ}-1$, the standard deviation of its cash flows $\sigma_{\mathfrak{b}}$, and the contribution $\mathfrak{a}$ of sectoral output growth to its cash flows growth as detailed in table 3. The chosen interest rate $r=5\%$.
Firm $1$ $2$ $3$ $4$
$\sigma_{\mathfrak{b}^{n}}$ 0.05 0.05 0.05 0.05
$F_{0}^{n}$ 1.0 1.0 1.0 1.0
$\mathfrak{a}^{n}$(Very High) 1.0 0.0 0.0 0.0
$\mathfrak{a}^{n}$(High) 0.0 1.0 0.0 0.0
$\mathfrak{a}^{n}$(Low) 0.0 0.0 1.0 0.0
$\mathfrak{a}^{n}$(Very Low) 0.0 0.0 0.0 1.0
Table 3: Characteristics of the firms
We compute here for $M=500$ simulations of the productivity processes $(\mathcal{Z},\mathcal{A})$, the loss given default of 4 loans with the same exposure but with 4 different financial assets collateral described in Table 3.
Emissions level No collateral Firm 1 Firm 2 Firm 3 Firm 4
Current Policies 45. 32.934 31.960 34.561 29.281
NDCs 45. 33.177 32.184 34.609 29.357
Divergent Net Zero 45. 33.485 32.471 34.673 29.459
Net Zero 2050 45. 33.995 32.940 34.784 29.640
Table 4: Average annual LGD per scenario between 2021 and 2030 (in %)
Both in Table 4 and in Figure 2, we can first see that the presence of guarantees reduce $\mathrm{LGD}$. Without collateral, we assume 45%, and with collateral, for all scenarios and for different characteristics of firms, $\mathrm{LGD}$ is less than 45%.
Figure 2: LGD with a financial asset as collateral
However, the decreasing of LGD depends on the scenarios. When the scenario becomes tougher, the impact of the presence of the collateral on LGD is lessened. This is logical and due to the fact that the value of the liquidated asset loses value when the price of carbon rises. The decreasing of the LGD also depends on the distinctive characteristics of the guarantees. Precisely, each firm in Table 3, serving as collateral, belongs to a unique and distinct sector (through $\mathfrak{a}$), which go from the more to the less polluting. Therefore the more the collateral is in a polluting sector, the less it reduces $\mathrm{LGD}$.
If the collateral is a building
We consider 5 apartments of 25 square meters whose price of the square meter fixed to 4000 euros in $t_{\circ}=2021$ is the same for all, but whose the energy efficiency are different. Moreover, we assume that the optimal energy efficiency equals to $\alpha^{\star}=70$ kilowatt hour per square meter per year (see Total Energies, (2024)) is reached.
Building $1$ $2$ $3$ $4$ $5$
$C_{n}^{0}$ 4000 4000 4000 4000 4000
$\alpha^{n}$ 320. 253. 187. 120. 70.
$R_{n}$ 25.0 25.0 25.0 25.0 25.
Table 5: Characteristics of the building
We use the $M=500$ trajectories of the productivity processes $(\mathcal{Z},\mathcal{A})$ simulated above. We compute the loss given default of 4 loans with the same exposure but with the 4 buildings described in Table 5 as collateral.
Emissions level No collateral Building 1 Building 2 Building 3 Building 4
Current Policies 45. 36.020 36.152 35.928 36.095
NDCs 45. 38.383 37.752 36.922 36.499
Divergent Net Zero 45. 38.939 38.303 37.377 36.751
Net Zero 2050 45. 39.102 38.524 37.615 36.908
Table 6: Average annual LGD per scenario between 2021 and 2030 (in %)
All the comments made for a financial asset as collateral are valid here: the presence of a collateral reduces LGD, that increases when the climate transition scenario becomes tougher.
Figure 3: LGD with a building as collateral
There are two main differences. First, the more the building is energetically inefficient, the more LGD increases (it is the same above when the financial asset belongs to a very polluting sector). Secondly, LGD decreases when time increases. This is a consequence of the dynamics of the impact of the carbon price of the housing market (as described in Sopgoui, (2024)): as we approach the optimal renovation date, the prices of energy-inefficient buildings rise and converge progressively towards the prices of energy-efficient buildings (we can see on Sopgoui, (2024)[Figure 3]). LGD follows the same behaviour logically but with an inverse monotony.
5.4.2 Expected and unexpected loss
To this aim, to keep things simple, we will consider a credit portfolio of $N=12$ loans contracted by the firms described in Table 7 below.
Loans 1 2 3 4 5 6 7 8 9 10 11 12
$\mathrm{EAD}_{n}$ 200. 200. 200. 200. 200. 200. 200. 200. 200. 200. 200. 200.
$F_{0}^{n}$ 1. 1. 1. 1. 1. 1. 1. 1. 1. 1. 1. 1.
$B^{n}$ 3.76 3.98 3.75 4.41 3.76 3.98 3.75 4.41 3.76 3.98 3.75 4.41
$\sigma_{\mathfrak{b}^{n}}$ 0.05 0.05 0.05 0.05 0.05 0.05 0.05 0.05 0.05 0.05 0.05 0.05
$\mathfrak{a}^{n}$(Very High) 1. 0. 0. 0. 1. 0. 0. 0. 1. 0. 0. 0.
$\mathfrak{a}^{n}$(High) 0. 1. 0. 0. 0. 1. 0. 0. 0. 1. 0. 0.
$\mathfrak{a}^{n}$(Low) 0. 0. 1. 0. 0. 0. 1. 0. 0. 0. 0. 0.
$\mathfrak{a}^{n}$(Very Low) 0. 0. 0. 1. 0. 0. 0. 1. 0. 0. 0. 0.
Collateral type No No No No Fa Fa Fa Fa Ho Ho Ho Ho
$\overline{F}_{0}^{n}$ 1. 1. 1. 1.
$\overline{\sigma}_{\mathfrak{b}^{n}}$ 0.05 0.05 0.05 0.05
$\overline{\mathfrak{a}}^{n}$(Very High) 1. 0. 0. 0.
$\overline{\mathfrak{a}}^{n}$(High) 0. 1. 0. 0.
$\overline{\mathfrak{a}}^{n}$(Low) 0. 0. 1. 0.
$\overline{\mathfrak{a}}^{n}$(Very Low) 0. 0. 0. 1.
$C_{n}^{0}$ 4000. 4000. 4000. 4000.
$R_{n}$ 25. 25. 25. 25.
$\alpha^{n}$ 320. 253. 187. 120.
Table 7: Characteristics of the portfolio (No = no collateral, Fa = Financial asset collateral, Ho = housing collateral)
We can remark that, for each $k=0,\ldots,2$, firms $14k+$ 1, $24k+$ 2, $34k+$ 3, and $44k+$ 4 respectively belong to the Very High Emitting, High Emitting, Low Emitting, and Very Low Emitting groups. Moreover, we assume that
  • 1.
    the loans of the firms $1$, $2$, $3$, and $4$ are not collateralized;
  • 2.
    the loans of the firms $5$, $6$, $7$, and $8$ are collateralized by financial assets described in Table 3;
  • 3.
    the loans of the firms $9$, $10$, $11$, and $12$ are collateralized by a building described in Table 5.
We want to calculate the expected (respectively unexpected) loss noted $\mathrm{EL}$ (respectively $\mathrm{UL}$) for each loan $n=1,\ldots,12$, by using (4.8) (respectively (4.10)).
Emissions level 1 2 3 4 5 6 7 8 9 10 11 12
Current Policies 1.00 1.00 1.00 1.00 0.79 0.77 0.81 0.71 0.58 0.59 0.58 0.58
NDCs 1.28 1.17 1.04 1.02 1.01 0.90 0.85 0.72 0.91 0.79 0.66 0.62
Divergent Net Zero 1.74 1.41 1.10 1.05 1.38 1.10 0.90 0.75 1.29 0.99 0.72 0.65
Net Zero 2050 2.85 1.91 1.21 1.11 2.27 1.47 0.98 0.79 2.13 1.37 0.81 0.70
Table 8: Average annual EL per scenario between 2021 and 2030 (in %)
Table 8 (respectively Table 9) shows average annual EL (respectively UL) normalized to the EL without collateral observed in the scenario Current Policies. We can make two key observations that were to be expected from the PD and LGD calculations:
  1. 1.
    Whether collateral is involved or not, we can see that EL and UL increase as the transition hardens. This is to be expected, since PD and LGD behave in the same way.
  2. 2.
    When a loan is collateralized, it significantly reduces the bank’s expected and unexpected losses. And for collateralized loans, these losses increase if the collateral has a high carbon footprint: in particular, if the collateral is a financial asset whose value growth is driven by a polluting sector or if it is a building that is not energy efficient.
Emissions level 1 2 3 4 5 6 7 8 9 10 11 12
Current Policies 1.00 1.00 1.00 1.00 0.81 0.81 0.84 0.82 0.59 0.57 0.56 0.62
NDCs 1.18 1.01 1.00 1.00 0.95 0.82 0.84 0.82 0.85 0.67 0.62 0.64
Divergent Net Zero 1.42 1.02 1.00 1.00 1.16 0.83 0.84 0.82 1.05 0.71 0.64 0.65
Net Zero 2050 1.80 1.02 1.0 0.99 1.49 0.85 0.85 0.82 1.34 0.73 0.65 0.66
Table 9: Average annual UL per scenario between 2021 and 2030 (in %)
Conclusion
Following Bouveret et al., (2023), we developed here a framework to quantify the impacts of the carbon price on a credit portfolio (expected and unexpected) losses, when the obligor companies as well as their guarantees belong to an economy subject to the climate transition declined by carbon price. We start by describing a closed economy, driven by a productivity following a multidimensional Ornstein-Uhlenbeck and subject to a climate transition modeled through a dynamic and deterministic carbon price, by a dynamic stochastic multisectoral. Then, by using the discounted cash flow methodology with the cash flows, following a stochastic differential equation, depending on the productivity as well as the carbon price, we evaluate the obligor value that helps us later on to compute its probability of over-indebtedness. We then turn to the bank’s loss in the event of a borrower’s over-indebtedness and if its loan is collateralized. When that is the case, the potential loss of the bank is written as the difference between the debt amount (EAD) and the collateral liquidated. We finally distinguish two types of collateral: either a financial asset or a building, both belonging to the economy so affected by the productivity and the carbon price. This work opens the door to many extensions as a finer modeling of the real estate market, taking into account other types of guarantees, modeling the unsecured loans that we assumed constant, modelling the impact of the carbon price on the exposure.
Appendix A Proofs
A.1 Hurwitz matrix
Assume that $-\Gamma$ is a Hurwitz matrix, then
  1. 1.
    if we note $\lambda_{\Gamma}:=\max_{\lambda\in\lambda(\Gamma)}Re(\lambda)\geq 0$, there exists $c_{\Gamma}>0$ so that $\lVert e^{-\Gamma t}\rVert<c_{\Gamma}e^{-\lambda_{\Gamma}t}$ for all $t\geq 0$.
  2. 2.
    Moreover, for $t\geq 0$n $\Upsilon_{t}$ defined in (2.6) is such that
    $\lVert\Upsilon_{t}\rVert=\left\lVert\int_{0}^{t}e^{-\Gamma s}\mathrm{d}s\right\rVert\leq\int_{0}^{t}\left\lVert e^{-\Gamma s}\right\rVert\mathrm{d}s\leq c_{\Gamma}\int_{0}^{t}e^{-\lambda_{\Gamma}s}\mathrm{d}s\leq c_{\Gamma}\min{\left\{\frac{1}{\lambda_{\Gamma}},t\right\}}.$ (A.1)
A.2 Bivariate Gaussian
Assume that $X$ and $Y$ are two standard Gaussian with correlation coefficient $\rho$. We then have for $(x,y)\in\mathbb{R}^{2}$, the cdf,
$\Phi_{2}(x,y):=\mathbb{P}[X\leq x,Y\leq y]=\frac{1}{2\pi(1-\rho^{2})}\int_{-\infty}^{x}\int_{-\infty}^{y}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left(u^{2}+v^{2}-2\rho uv\right)\right)}\mathrm{d}u\mathrm{d}v.$ (A.2)
Let $\sigma>0$, we want to compute $\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]$. We have
$\begin{split}\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]&=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\int_{-\infty}^{x}\int_{-\infty}^{y}e^{\sigma u}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left(u^{2}+v^{2}-2\rho uv\right)\right)}\mathrm{d}u\mathrm{d}v\\ &=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\int_{-\infty}^{x}e^{\sigma u-\frac{1}{2(1-\rho^{2})}u^{2}}\int_{-\infty}^{y}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left(v^{2}-2\rho uv\right)\right)}\mathrm{d}v\quad\mathrm{d}u\\ &=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\int_{-\infty}^{x}e^{\sigma u-\frac{1}{2(1-\rho^{2})}u^{2}}\int_{-\infty}^{y}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left((v-\rho u)^{2}-\rho^{2}u^{2}\right)\right)}\mathrm{d}v\quad\mathrm{d}u\\ &=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\int_{-\infty}^{x}e^{\sigma u-\frac{1}{2}u^{2}}\int_{-\infty}^{y}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left((v-\rho u)^{2}\right)\right)}\mathrm{d}v\quad\mathrm{d}u\\ \end{split}$
But
$\int_{-\infty}^{y}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left((v-\rho u)^{2}\right)\right)}\mathrm{d}v=\sqrt{2\pi(1-\rho^{2})}\Phi\left(\frac{y-\rho u}{\sqrt{1-\rho^{2}}}\right),$
therefore,
$\begin{split}\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]&=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{x}e^{\sigma u-\frac{1}{2}u^{2}}\Phi\left(\frac{y-\rho u}{\sqrt{1-\rho^{2}}}\right)\mathrm{d}u\\ &=\frac{e^{\frac{1}{2}\sigma^{2}}}{\sqrt{2\pi}}\int_{-\infty}^{x}e^{-\frac{1}{2}(u-\sigma)^{2}}\Phi\left(\frac{y-\rho u}{\sqrt{1-\rho^{2}}}\right)\mathrm{d}u\\ &=e^{\frac{1}{2}\sigma^{2}}\int_{-\infty}^{x-\sigma}\phi(u)\Phi\left(\frac{y-\rho\sigma}{\sqrt{1-\rho^{2}}}+\frac{-\rho}{\sqrt{1-\rho^{2}}}u\right)\mathrm{d}u.\end{split}$
However,
$\int_{-\infty}^{c}\Phi(a+bx)\phi(x)\mathrm{d}x=\Phi_{2}\left(c,\frac{a}{\sqrt{1+b^{2}}};\frac{-b}{\sqrt{1+b^{2}}}\right),$
we can then conclude that
$\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]=e^{\frac{1}{2}\sigma^{2}}\Phi_{2}\left(x-\sigma,y-\rho\sigma;\rho\right).$ (A.3)
Appendix B The multisectoral model in continuous time
For all $i\in\mathcal{I}$, let us consider the following $\mathbb{G}$-measurable and positive processes: $Y^{i}$ the production of sector $i$, $N^{i}$ the labor demand in sector $i$, and for all $j\in\mathcal{I}$, $Z^{ji}$ the consumption by sector $i$ of intermediate inputs produced by sector $j$.
B.1 The firm’s point of view
Aiming to work with a simple model, we follow (Galí, , 2015, Chapter 2). It then appears that the firm’s problem corresponds to an optimization performed at each period, depending on the state of the world. This problem will depend, in particular, on the productivity and the price processes introduced above. Moreover, it will also depend on $P^{i}$ and $W^{i}$, two $\mathbb{G}$-adapted positive stochastic processes representing respectively the price of good $i$ and the wage paid in sector $i\in\mathcal{I}$. We start by considering the associated deterministic problem below, when time and randomness are fixed.
Solution for the deterministic problem
We denote $\overline{a}\in(0,+\infty)^{I}$ the level of technology in each sector, $\overline{p}\in(0,\infty)^{I}$ the price of the goods produced by each sector, $\overline{w}\in(0,\infty)^{I}$ the nominal wage in each sector, $\overline{\tau}\in[0,1)^{I}$ and $\overline{\zeta}\in[0,1)^{I\times I}$ the price on production and consumption of goods. For $i\in\mathcal{I}$, we consider a representative firm of sector $i$, with technology described by the production function
+×$\mathbb{R}_{+}\times\mathbb{R}_{+}^{I}\ni(n,z)\mapsto F^{i}_{\overline{a}}(n,z)=\overline{a}^{i}n^{\psi^{i}}\prod_{j\in\mathcal{I}}(z^{j})^{\boldsymbol{\lambda}^{ji}}\in\mathbb{R}_{+},$ (B.1)
where $n$ represents the number of hours of work in the sector, and $z^{j}$ the firm’s consumption of intermediary input produced by sector $j$. The coefficients $\psi\in(\mathbb{R}_{+}^{*})^{I}$ and $\boldsymbol{\lambda}\in(\mathbb{R}_{+}^{*})^{I\times I}$ are elasticities satisfying (2.13). The management of firm $i$ then solves the classical problem of profit maximization
$\widehat{\Pi}^{i}_{(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta},\overline{\delta})}:=\sup_{(n,z)\in\mathbb{R}_{+}\times\mathbb{R}_{+}^{I}}\Pi^{i}(n,z),$ (B.2)
where, omitting the dependency in $(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})$,
$\Pi^{i}(n,z):=F^{i}_{\overline{a}}(n,z)\overline{p}^{i}-\overline{\tau}^{i}F^{i}_{\overline{a}}(n,z)\overline{p}^{i}\overline{\delta}-\overline{w}^{i}n-\sum_{j\in\mathcal{I}}z^{j}\overline{p}^{j}+z^{j}\overline{\zeta}^{ji}\overline{p}^{j}\overline{\delta}.$ (B.3)
Note that $F^{i}_{\overline{a}}(n,z)(1-\overline{\tau}^{i})\overline{p}^{i}$ represents the firm’s revenues after carbon price, that $\overline{w}^{i}n$ stands for the firm’s total compensations, and that $\sum_{j\in\mathcal{I}}z^{j}(1+\overline{\zeta}^{ji})\overline{p}^{j}$ is the firm’s total intermediary inputs. Now, we would like to solve the optimization problem for the firms, namely determine the optimal demands $\mathfrak{n}$ and $\mathfrak{z}$ as functions of $(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})$. Because we will lift these optimal quantities in a dynamical stochastic setting, we impose that they are expressed as measurable functions. We thus introduce:
Definition B.1.
An admissible solution to problem (B.2) is a pair of measurable functions
$(\mathfrak{n},\mathfrak{z}):(0,+\infty)^{I}\times(0,+\infty)^{I}\times(0,+\infty)^{I}\times[0,1)^{I}\times[0,1)^{I\times I}\rightarrow[0,+\infty)^{I}\times[0,+\infty)^{I\times I},$
such that, for each sector $i$, denoting $\overline{n}:=\mathfrak{n}^{i}(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})$ and $\overline{z}:=\mathfrak{z}^{\cdot i}(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})$,
$F^{i}_{\overline{a}}(\overline{n},\overline{z})(1-\overline{\tau}^{i}\overline{\delta})\overline{p}^{i}-\overline{w}^{i}\overline{n}-\sum_{j\in\mathcal{I}}\overline{z}^{j}(1+\overline{\zeta}^{ji}\overline{\delta})\overline{p}^{j}=\widehat{\Pi}^{i}_{(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta},\overline{\delta})},$
and $F^{i}_{\overline{a}}(\overline{n},\overline{z})>0$ (non-zero production), according to (B.2).
Remark B.2.
The solution obviously depends also on the coefficients $\psi$ and $\boldsymbol{\lambda}$. But these are fixed once and we will not study the dependence of the solution with respect to them.
Proposition B.3.
There exists admissible solutions in the sense of Definition B.1. Any admissible solution is given by for all $i\in\mathcal{I}$, $\mathfrak{n}^{i}>0$ and for all $(i,j)\in\mathcal{I}^{2}$,
$\mathfrak{z}^{ji}=\frac{\boldsymbol{\lambda}^{ji}}{\psi^{i}}\frac{\overline{w}^{i}}{(1+\overline{\zeta}^{ji}\overline{\delta})\overline{p}^{j}}\mathfrak{n}^{i}>0.$ (B.4)
Moreover, it holds that $\widehat{\Pi}^{i}_{(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta},\overline{\delta})}=0$ (according to (B.2)) and
$\displaystyle\mathfrak{n}^{i}$ $=\psi^{i}F^{i}_{\overline{a}}(\mathfrak{n}^{i},\mathfrak{z}^{\cdot i})\frac{(1-\overline{\tau}^{i}\overline{\delta})\overline{p}^{i}}{\overline{w}^{i}},$ (B.5a)
$\displaystyle\mathfrak{z}^{ji}$ $=\boldsymbol{\lambda}^{ji}F^{i}_{\overline{a}}(\mathfrak{n}^{i},\mathfrak{z}^{\cdot i})\frac{(1-\overline{\tau}^{i}\overline{\delta})\overline{p}^{i}}{(1+\overline{\zeta}^{ji}\overline{\delta})\overline{p}^{j}}.$ (B.5b)
Proof.
We study the optimization problem for the representative firm $i\in\mathcal{I}$. Since $\psi^{i}>0$ and $\boldsymbol{\lambda}^{ji}>0$, for all $j\in\mathcal{I}$, as soon as $n=0$ or $z^{j}=0$, for some $j\in\mathcal{I}$, the production is equal to 00. From problem (B.2), we obtain that necessarily $n\neq 0$ and $z^{j}\neq 0$ for all $j$ in this case. So an admissible solution, which has non-zero production, has positive components.
Setting $\overline{n}=\mathfrak{n}^{i}(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})>0$ and $\overline{z}=\mathfrak{z}^{\cdot i}(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})>0$, the optimality of $(\overline{n},\overline{z})$ yields
$\partial_{n}\Pi^{i}(\overline{n},\overline{z})=0\text{ and for any }j\in\mathcal{I},\quad\partial_{z^{j}}\Pi^{i}(\overline{n},\overline{z})=0.$
We then compute
$\psi^{i}\frac{F^{i}_{\overline{a}}(\overline{n},\overline{z})}{\overline{n}}(1-\overline{\tau}^{i}\overline{\delta})\overline{p}^{i}-\overline{w}^{i}=0\text{ and for any }j\in\mathcal{I},\quad\boldsymbol{\lambda}^{ji}\frac{F^{i}_{\overline{a}}(\overline{n},\overline{z})}{\overline{z}^{j}}(1-\overline{\tau}^{i}\overline{\delta})\overline{p}^{i}-(1+\overline{\zeta}^{ji}\overline{\delta})\overline{p}^{j}=0,$
which leads to (B.4), (B.5a), and (B.5b). ∎
Dynamic setting
In B.3 below, we characterize the dynamics of the output and consumption processes using market equilibrium arguments. There, the optimal demand by the firm for intermediary inputs and labor is lifted to the stochastic setting where the admissible solutions then write as functions of the productivity, carbon price, price of goods/services; and wage processes, see Definition B.5. For all $i\in\mathcal{I}$, $Y^{i}$ representing the production of sector $i$, $N^{i}$ representing the labor demand in sector $i$, and for all $j\in\mathcal{I}$, $Z^{ji}$ representing the consumption by sector $i$ of intermediate inputs produced by sector $j$ are therefore positive and $\mathbb{G}$-adapted processes.
B.2 The household’s point of view
Let $(r_{t})_{t\geq 0}$ be the (exogenous) deterministic interest rate, valued in $\mathbb{R}_{+}$. At each time $t\geq 0$ and for each sector $i\in\mathcal{I}$, we denote
  • 1.
    $C_{t}^{i}$ the quantity consumed of the single good in the sector $i$, valued in $\mathbb{R}_{+}^{*}$;
  • 2.
    $H_{t}^{i}$ the number of hours of work in sector $i$, valued in $\mathbb{R}_{+}^{*}$.
We also introduce a time preference parameter $\beta\in[0,1)$ and a utility function $U:(0,\infty)^{2}\to\mathbb{R}$ given, for $\varphi\geq 0$, by $U(x,y):=\frac{x^{1-\sigma}}{1-\sigma}-\frac{y^{1+\varphi}}{1+\varphi}$ if $\sigma\in[0,1)\cup(1,+\infty)$ and by $U(x,y):=\log(x)-\frac{y^{1+\varphi}}{1+\varphi}$, if $\sigma=1$. We also suppose that
$\mathfrak{P}:=\sup_{t\geq 0,i\in\mathcal{I}}\mathbb{E}\left[\left(\frac{P_{t}^{i}}{W_{t}^{i}}\right)^{1+\varphi}\right]<+\infty.$ (B.6)
For any ${C},{H}\in\mathscr{L}^{1}_{+}(\mathbb{G},(0,\infty)^{I})$, we introduce the wealth process
$\mathrm{d}Q_{t}=r_{t}Q_{t}\mathrm{d}t+\sum_{i\in\mathcal{I}}W_{t}^{i}H_{t}^{i}-\sum_{i\in\mathcal{I}}P_{t}^{i}C_{t}^{i}-\sum_{i\in\mathcal{I}}\kappa_{t}^{i}P_{t}^{i}C_{t}^{i}\delta_{t},\qquad\text{for any }t\geq 0,$ (B.7)
with the convention $Q_{0}:=0$ and $r_{0}:=0$. Note that we do not indicate the dependence of $Q$ upon $C$ and $H$ to alleviate the notations.
For $t\geq 0$ and $i\in\mathcal{I}$, $P_{t}^{i}C_{t}^{i}$ represents the household’s consumption in the sector $i$ and $\kappa_{t}^{i}P_{t}^{i}C_{t}^{i}\delta_{t}$ is the cost paid by households due to their emissions when they consume goods $i$, so $\sum_{i\in\mathcal{I}}P_{t}^{i}C_{t}^{i}(1+\kappa_{t}^{i}\delta_{t})$ is the household’s total expenses. Moreover, $W_{t}^{i}H_{t}^{i}$ is the household’s labor income in the sector $i$, $(1+r_{t-1})Q_{t-1}$ the household’s capital income, and $(1+r_{t-1})Q_{t-1}+\sum_{i\in\mathcal{I}}W_{t}^{i}H_{t}^{i}$ the household’s total revenue.
We define $\mathscr{A}$ as the set of all couples $({C},{H})$ with ${C},{H}\in\mathscr{L}^{1}_{+}(\mathbb{G},(0,\infty)^{I})$ such that
$\left\{\begin{array}[]{ll}&\displaystyle\mathbb{E}\left[\sum_{i\in\mathcal{I}}\int_{t=0}^{\infty}\beta^{t}|U(C_{t}^{i},H_{t}^{i})|\mathrm{d}t\right]<\infty,\\ &\lim_{T\uparrow\infty}\mathbb{E}[Q_{T}|\mathcal{G}_{t}]\geq 0,\qquad\text{forall }t\geq 0.\end{array}\right.$
The representative household consumes the $I$ goods of the economy and provides labor to all the sectors. For any $({C},{H})\in\mathscr{A}$, let
$\mathcal{J}({C},{H}):=\sum_{i\in\mathcal{I}}\mathcal{J}_{i}({C}^{i},{H}^{i}),\qquad\textrm{with}\qquad\mathcal{J}_{i}({C}^{i},{H}^{i}):=\mathbb{E}\left[\int_{t=0}^{\infty}\beta^{t}U(C_{t}^{i},H_{t}^{i})\mathrm{d}t\right],\quad\textrm{for all }i\in\mathcal{I}.$
The representative household seeks to maximize its objective function by solving
$\max_{({C},{H})\in\mathscr{A}}\quad\mathcal{J}({C},{H}).$ (B.8)
We choose above a separable utility function as Miranda-Pinto and Young, (2019) does, meaning that the representative household optimizes its consumption and hours of work for each sector independently but under a global budget constraint. The following proposition provides an explicit solution to (B.8).
Proposition B.4.
Assume that (B.8) has a solution $(C,H)\in\mathscr{A}$. Then, for all $i,j\in\mathcal{I}$, the household’s optimality condition reads, for any $t\geq 0$,
$\displaystyle\frac{P_{t}^{i}}{W_{t}^{i}}$ $=\displaystyle\frac{1}{1+\kappa_{t}^{i}\delta_{t}}(H_{t}^{i})^{-\varphi}(C_{t}^{i})^{-\sigma},$ (B.9a)
$\displaystyle\frac{P_{t}^{i}}{P_{t}^{j}}$ $=\displaystyle\frac{1+\kappa_{t}^{j}\delta_{t}}{1+\kappa_{t}^{i}\delta_{t}}\left(\frac{C_{t}^{i}}{C_{t}^{j}}\right)^{-\sigma}.$ (B.9b)
Note that the discrete-time processes $C$ and $H$ cannot hit zero by definition of $\mathscr{A}$, so that the quantities above are well defined.
Proof.
Suppose that $\sigma\neq 1$. We first check that $\mathscr{A}$ is non empty. Assume that, for all $t\geq 0$ and $i\in\mathcal{I}$, $\tilde{C}_{t}^{i}=1$ and $\tilde{H}_{t}^{i}=\frac{P_{t}^{i}(1+\kappa_{t}^{i})}{W_{t}^{i}}$, then
$\mathbb{E}\left[\sum_{i\in\mathcal{I}}\int_{t=0}^{\infty}\beta^{t}|U(\tilde{C}_{t}^{i},\tilde{H}_{t}^{i})|\mathrm{d}t\right]$ $\leq\sum_{i\in\mathcal{I}}\int_{t=0}^{\infty}\beta^{t}\left(\frac{1}{1-\sigma}+\frac{1}{1+\varphi}\mathbb{E}\left[\left(\frac{P_{t}^{i}(1+\kappa_{t}^{i}\delta_{t})}{W_{t}^{i}}\right)^{1+\varphi}\right]\right)\mathrm{d}t.$
$\leq\sum_{i\in\mathcal{I}}\int_{t=0}^{\infty}\beta^{t}\left(\frac{1}{1-\sigma}+\frac{\mathfrak{P}(1+\kappa_{t}^{i}\delta_{t})^{1+\varphi}}{1+\varphi}\right)\mathrm{d}t<+\infty,$
using (B.6). We also observe that $Q$ built from $\tilde{H},\tilde{C}$ satisfies $Q_{t}=0$, for $t\geq 0$. Thus $(\tilde{H},\tilde{C})\in\mathscr{A}$.
Let now $(\widehat{{C}},\widehat{{H}})\in\mathscr{A}$ be such that $\mathcal{J}(\widehat{{C}},\widehat{{H}})=\max_{({C},{H})\in\mathscr{A}}\mathcal{J}({C},{H})$.
[…]
$\overline{{C}}:=\widehat{{C}}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}}\Delta^{(i,s)}\text{ and }\overline{{H}}:=\widehat{{H}}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}}\Delta^{(i,s)}\frac{{P}^{i}(1+\kappa^{i}\delta_{s})}{{W}^{i}}.$ (B.10)
We observe that for $(j,t)\neq(i,s)$, $\overline{{C}}^{j}_{t}=\widehat{{C}}^{j}_{t}$ and $\overline{H}^{j}_{t}=\widehat{H}^{j}_{t}$ and we compute
$\overline{{C}}^{i}_{s}\geq\widehat{{C}}^{i}_{s}-\theta^{(i,s)}\geq\frac{1}{2}\widehat{{C}}^{i}_{s}>0.$
Similarly, we obtain $\overline{H}^{i}_{s}>0$. We also observe that $\overline{C}\leq\frac{3}{2}\widehat{C}$ and $\overline{H}\leq\frac{3}{2}\widehat{H}$. Finally, we have that
$\sum_{j\in\mathcal{I}}W_{t}^{j}\overline{H}_{t}^{j}-\sum_{j\in\mathcal{I}}P_{t}^{j}(1+\kappa_{t}^{j}\delta_{t})\overline{C}_{t}^{j}=\sum_{j\in\mathcal{I}}W_{t}^{j}\widehat{H}_{t}^{j}-\sum_{j\in\mathcal{I}}P_{t}^{j}(1+\kappa_{t}^{j}\delta_{t})\widehat{C}_{t}^{j}.$
This allows us to conclude that $(\overline{{C}},\overline{{H}})\in\mathscr{A}$.
We have, by optimality of $(\widehat{{C}},\widehat{{H}})$,
$\mathcal{J}(\widehat{{C}},\widehat{{H}})-\mathcal{J}(\overline{{C}},\overline{{H}})=\sum_{j\in\mathcal{I}}\mathcal{J}_{j}(\widehat{{C}}^{j},\widehat{{H}}^{j})-\sum_{j\in\mathcal{I}}\mathcal{J}_{j}(\overline{{C}}^{j},\overline{{H}}^{j})\geq 0.$
However, for all $(t,j)\neq(s,i)$, $\overline{{C}}^{j}_{t}=\widehat{{C}}^{j}_{t}$ and $\overline{{H}}^{j}_{t}=\widehat{{H}}^{j}_{t}$, then
$\mathbb{E}\left[\beta^{s}U(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})\right]-\mathbb{E}\left[\beta^{s}U\left(\widehat{{C}}_{s}^{i}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}},\widehat{{H}}_{s}^{i}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}}\frac{P_{s}^{i}(1+\kappa_{s}^{i}\delta_{s})}{W_{s}^{i}}\right)\right]\geq 0,$
i.e.
$\frac{1}{h}\mathbb{E}\left[U(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})-U\left(\widehat{{C}}_{s}^{i}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}},\widehat{{H}}_{s}^{i}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}}\frac{P_{s}^{i}(1+\kappa_{s}^{i}\delta_{s})}{W_{s}^{i}}\right)\right]\geq 0.$
Letting $h$ tend to 00, we obtain
$\mathbb{E}\left[\eta\theta^{(i,s)}\mathbf{1}_{A^{s}}\frac{\partial U}{\partial x}(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})+\eta\theta^{(i,s)}\mathbf{1}_{A^{s}}\frac{P_{s}^{i}(1+\kappa_{s}^{i}\delta_{s})}{W_{s}^{i}}\frac{\partial U}{\partial y}(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})\right]\geq 0.$
Since the above holds for all $A^{s}\in\mathcal{G}_{s}$, $\eta=\pm 1$ and since $\theta^{(i,s)}>0$, then
$\frac{\partial U}{\partial x}(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})+\frac{P_{s}^{i}(1+\kappa_{s}^{i}\delta_{s})}{W_{s}^{i}}\frac{\partial U}{\partial y}(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})=0,$
leading to (B.9a).
For $j\in\mathcal{I}\setminus\{i\}$ and $\theta^{(i,j,s)}:=\frac{1}{2}\left(1\wedge\frac{{P}^{j}_{s}(1+\kappa_{s}^{j}\delta_{s})}{{P}^{i}_{s}(1+\kappa_{s}^{i}\delta_{s})}\right)(1\wedge\widehat{C}^{i}_{s}\wedge\widehat{C}^{j}_{s})>0$, setting now
$\overline{{C}}:=\widehat{{C}}+\eta h\mathbf{1}_{A^{s}}\theta^{(i,j,s)}\left(\Delta^{(i,s)}-\Delta^{(j,s)}\frac{{P}^{i}(1+\kappa^{i}\delta_{s})}{{P}^{j}(1+\kappa^{j}\delta_{s})}\right)\quad\text{and}\quad\overline{{H}}:=\widehat{{H}},$
and using similar arguments as above, we obtain (B.9b).
When $\sigma=1$, we carry out an analogous proof. ∎
B.3 Markets equilibrium
We now consider that firms and households interact on the labor and goods markets.
Definition B.5.
A market equilibrium is a $\mathbb{G}$-adapted positive random process $(\overline{W},\overline{P})$ such that
  1. 1.
    Condition (B.6) holds true for $(\overline{W},\overline{P})$.
  2. 2.
    The goods’ and labor’s market clearing conditions are met, namely, for each sector $i\in\mathcal{I}$, and for all $t\geq 0$,
    $Y_{t}^{i}=C_{t}^{i}+\sum_{j\in\mathcal{I}}Z_{t}^{ij}\qquad\text{and}\qquad H_{t}^{i}=N_{t}^{i},$ (B.11)
    where $N_{t}=\overline{n}(A_{t},\overline{W}_{t},\overline{P}_{t},\kappa_{t},\zeta_{t})$, $Z_{t}=\overline{z}(A_{t},\overline{W}_{t},\overline{P}_{t},\kappa_{t},\zeta_{t})$, $Y=F_{A}(N,Z)$ with $(\overline{n},\overline{z})$ an admissible solution (B.5a)-(B.5b) to (B.2), from Proposition B.3 while $C$ and $H$ satisfy (B.9a)-(B.9b) for $(\overline{W},\overline{P})$.
In the case of the existence of a market equilibrium, we can derive equations that must be satisfied by the output production process $Y$ and the consumption process $C$.
Proposition B.6.
Assume that there exists a market equilibrium as in Definition B.5. Then, for $t\geq 0$, $i\in\mathcal{I}$, it must hold that
$\left\{\begin{array}[]{rl}Y_{t}^{i}&=\displaystyle C_{t}^{i}+\sum_{j\in\mathcal{I}}\Lambda^{ij}(\mathfrak{d}_{t})\left(\frac{C_{t}^{j}}{C_{t}^{i}}\right)^{-\sigma}Y_{t}^{j},\\ Y_{t}^{i}&=\displaystyle A^{i}_{t}\left[\Psi^{i}(\mathfrak{d}_{t})(C_{t}^{i})^{-\sigma}Y_{t}^{i}\right]^{\frac{\psi^{i}}{1+\varphi}}\prod_{j\in\mathcal{I}}\left[\Lambda^{ji}(\mathfrak{d}_{t})\left(\frac{C_{t}^{i}}{C_{t}^{j}}\right)^{-\sigma}Y_{t}^{i}\right]^{\boldsymbol{\lambda}^{ji}},\end{array}\right.$ (B.12)
where $\Psi$ and $\Lambda$ are defined in (2.14), and $\mathfrak{d}_{t}$ is defined in (2.11).
Proof.
Let $i,j\in\mathcal{I}$ and $t\geq 0$. Combining Proposition B.3 and Proposition B.4, we obtain
$Z_{t}^{ji}=\boldsymbol{\lambda}^{ji}\frac{1-\tau_{t}^{i}\delta_{t}}{1+\zeta^{ji}_{t}\delta_{t}}\frac{1+\kappa_{t}^{j}\delta_{t}}{1+\kappa_{t}^{i}\delta_{t}}\left(\frac{C_{t}^{i}}{C_{t}^{j}}\right)^{-\sigma}Y_{t}^{i}.$ (B.13)
From Propositions B.3 and B.4 again, we also have
$N_{t}^{i}=\psi^{i}\frac{1-\tau_{t}^{i}\delta_{t}}{1+\kappa_{t}^{i}\delta_{t}}(H_{t}^{i})^{-\varphi}(C_{t}^{i})^{-\sigma}Y_{t}^{i}.$
The labor market clearing condition in Definition B.5 yields
$N_{t}^{i}=\left[\psi^{i}\frac{1-\tau_{t}^{i}\delta_{t}}{1+\kappa_{t}^{i}\delta_{t}}(C_{t}^{i})^{-\sigma}Y_{t}^{i}\right]^{\frac{1}{1+\varphi}}.$ (B.14)
Then, by inserting the expression of $N_{t}^{i}$ given in (B.14)and $Z_{t}^{ji}$ given in (B.13) into the production function $F$, we obtain the second equation in (B.12). The first equation in (B.12) is obtained by combining the market clearing condition with (B.13) (at index $(i,j)$ instead of $(j,i)$). ∎
B.4 Output and consumption dynamics and associated growth
For each time $t\geq 0$ and noise realization, the system (B.12) is nonlinear with $2I$ equations and $2I$ variables, and its well-posedness is hence relatively involved. Moreover, it is computationally heavy to solve this system for each price trajectory and productivity scenario. We thus consider a special value for the parameter $\sigma$ which allows to derive a unique solution in closed form. From now on, and following (Golosov et al., , 2014, page 63), we assume that $\sigma=1$, namely $U(x,y):=\log(x)-\frac{y^{1+\varphi}}{1+\varphi}$ on $(0,\infty)^{2}$.
Theorem B.7.
Assume that
  1. 1.
    $\sigma=1$,
  2. 2.
    $\mathbf{I}_{I}-\boldsymbol{\lambda}$ is not singular,
  3. 3.
    $\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{t})^{\top}$ is not singular for all $t\in\mathbb{R}_{+}$.
Then for all $t\geq 0$, there exists a unique $({C}_{t},{Y}_{t})$ satisfying (B.12). Moreover, with $\mathfrak{e}_{t}^{i}:=\frac{Y^{i}_{t}}{C^{i}_{t}}$ for $i\in\mathcal{I}$, we have
$\mathfrak{e}_{t}=\mathfrak{e}(\mathfrak{d}_{t}):=(\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{t})^{\top})^{-1}\mathbf{1},$ (B.15)
and using $\mathcal{B}_{t}=(\mathcal{B}^{i}_{t})_{i\in\mathcal{I}}:=\left[\mathcal{A}_{t}^{i}+\mathfrak{v}^{i}(\mathfrak{d}_{t})\right]_{i\in\mathcal{I}}$ with
$\mathfrak{v}^{i}(\mathfrak{d}_{t}):=\log\left((\mathfrak{e}_{t}^{i})^{-\frac{\varphi\psi^{i}}{1+\varphi}}\left(\Psi^{i}(\mathfrak{d}_{t})\right)^{\frac{\psi^{i}}{1+\varphi}}\prod_{j\in\mathcal{I}}\left(\Lambda^{ji}(\mathfrak{d}_{t})\right)^{\boldsymbol{\lambda}^{ji}}\right),$ (B.16)
we obtain
${C}_{t}=\exp{\left((\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}{\mathcal{B}_{t}}\right)}.$ (B.17)
Proof.
Let $t\geq 0$. When $\sigma=1$, the system (B.12) becomes for all $i\in\mathcal{I}$,
$\left\{\begin{array}[]{rl}Y_{t}^{i}&=\displaystyle C_{t}^{i}+\sum_{j\in\mathcal{I}}\Lambda^{ij}(\mathfrak{d}_{t})\left(\frac{C_{t}^{i}}{C_{t}^{j}}\right)Y_{t}^{j},\\ Y_{t}^{i}&=\displaystyle A_{t}^{i}\left[\Psi^{i}(\mathfrak{d}_{t})\mathfrak{e}^{i}_{t}\right]^{\frac{\psi^{i}}{1+\varphi}}\prod_{j\in\mathcal{I}}\left[\Lambda^{ji}(\mathfrak{d}_{t})C_{t}^{j}\mathfrak{e}_{t}^{i}\right]^{\boldsymbol{\lambda}^{ji}}.\end{array}\right.$ (B.20)
For any $i\in\mathcal{I}$, dividing the first equation in (B.20) by $C_{t}^{i}$, we get
$\mathfrak{e}^{i}_{t}=1+\sum_{j\in\mathcal{I}}\Lambda^{ij}(\mathfrak{d}_{t})\mathfrak{e}^{j}_{t},$
which corresponds to (B.15), thanks to (2.13). Using $\sum_{j\in\mathcal{I}}\boldsymbol{\lambda}^{ji}=1-\psi^{i}$ and $Y_{t}^{i}=\mathfrak{e}_{t}^{i}C_{t}^{i}$ in the second equation in (B.20), we compute
$C_{t}^{i}=A_{t}^{i}(\mathfrak{e}_{t}^{i})^{-\frac{\varphi\psi^{i}}{1+\varphi}}\left[\Psi^{i}(\mathfrak{d}_{t})\right]^{\frac{\psi^{i}}{1+\varphi}}\prod_{j\in\mathcal{I}}\left[\Lambda^{ji}(\mathfrak{d}_{t})\right]^{\boldsymbol{\lambda}^{ji}}\prod_{j\in\mathcal{I}}(C_{t}^{j})^{\boldsymbol{\lambda}^{ji}}.$
Applying log and writing in matrix form, we obtain $(\mathbf{I}_{I}-\boldsymbol{\lambda})\log({C}_{t})={\mathcal{B}_{t}}$, implying (B.17). ∎
Remark B.8.
The matrix $\boldsymbol{\lambda}$ is generally not diagonal, and therefore, from (B.17), the sectors (in output and in consumption) are linked to each other through their respective productivity process. Similarly, an introduction of price in one sector affects the other ones.
Remark B.9.
For any $t\geq 0$, $i\in\mathcal{I}$, we observe that
$\mathcal{B}^{i}_{t}=\mathcal{A}^{i}_{t}+v^{i}(\mathfrak{d}_{t}),$ (B.21)
where $v^{i}(\cdot)$ is defined using (B.16). Namely, $\mathcal{B}_{t}$ is the sum of the (random) productivity term and a term involving the price. The economy is therefore subject to fluctuations of two different natures: the first one comes from the productivity process while the second one comes from the price processes.
We now look at the dynamics of production and consumption growth.
Theorem B.10.
For any $t\geq 0$ and for $\varpi\in\{{Y},{C}\}$. With the same assumptions as in Theorem B.7,
$\mathrm{d}\log{\varpi}_{t}\sim\mathcal{N}\left({m}_{t}^{\varpi},\widehat{\Sigma}_{t}\right),\qquad\text{for }\varpi\in\{{Y},{C}\},$ (B.22)
with
$\widehat{\Sigma}_{t}$ $=\varsigma^{2}(\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}\overline{\Sigma}(\mathbf{I}_{I}-\boldsymbol{\lambda}^{\top})^{-1}(\mathrm{d}t)^{2},$ (B.23)
$m^{C}_{t}$ $=(I-\boldsymbol{\lambda})^{-1}\left[\mu\mathrm{d}t+\mathrm{d}\mathfrak{v}(\mathfrak{d}_{t})\right],$ (B.24)
$m^{Y}_{t}$ $=(I-\boldsymbol{\lambda})^{-1}\left[\mu\mathrm{d}t+\mathrm{d}\mathfrak{v}(\mathfrak{d}_{t})\right],$ (B.25)
and
$v(\mathfrak{d}_{t}):=\mathfrak{v}(\mathfrak{d}_{t})+(\mathbf{I}_{I}-\boldsymbol{\lambda})\log(\mathfrak{e}(\mathfrak{d}_{t})),$ (B.26)
where $\overline{\mu}$ and $\varsigma^{2}\overline{\Sigma}$ are the mean and the variance of the stationary process $\mathcal{Z}$ (Remark 2.2), $v$ is defined in (B.16) and $\mathfrak{e}$ in (B.15).
Proof.
Let $t\geq 0^{*}$, from (B.21), we have, for $i\in\mathcal{I}$,
$\mathrm{d}\mathcal{B}^{i}_{t}=(\mu^{i}+\varsigma\mathcal{Z}_{t}^{i})\mathrm{d}t+\mathrm{d}v^{i}(\mathfrak{d}_{t}).$
Combining the previous equality with (B.17), we get
$\mathrm{d}\log{{C}}_{t}=(\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}\left[(\mu+\varsigma\mathcal{Z}_{t})\mathrm{d}t+\mathrm{d}v(\mathfrak{d}_{t})\right].$ (B.27)
Applying Remark 2.2 leads to $\mathrm{d}\log{{C}}_{t}\sim\mathcal{N}\left({m}_{t}^{{C}},\widehat{\Sigma}_{t}\right)$. Using (B.15), we observe that, for $i\in\mathcal{I}$,
$(\mathrm{d}\log{{Y}_{t}})^{i}=(\mathrm{d}\log{{C}_{t}})^{i}+\mathrm{d}\log(\mathfrak{e}^{i}(\mathfrak{d}_{t})),$ (B.28)
which, using the previous characterization of the law of $\mathrm{d}\log{{C}_{t}}$, allows to conclude. ∎
From the previous result, we observe that output and consumption growth processes have a stationary variance but a time-dependent mean.
Proof.
of Proposition 2.6.
Let $t\geq 0$, $n\in\{1,\ldots,N\}$, and $T>t_{\star}$.
  1. 1.
    we also introduce,
    $\mathcal{V}^{n,K}_{t,\mathfrak{d}}:=F^{n}_{t,\mathfrak{d}}\int_{t}^{+\infty}e^{-r(s-t)}\mathbb{E}_{t}\left[\exp\left((s-t)\mathfrak{a}^{n\cdot}\mu+\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)+\sigma_{n}(\mathcal{W}^{n}_{s}-\mathcal{W}^{n}_{t})\right)\mathrm{d}s\right].$ (B.29)
    Similar computations as (in fact easier than) the ones performed in the proof of Proposition 2.5. in Bouveret et al., (2023) show that $\mathcal{V}^{n}_{t,\mathfrak{d}}=\lim_{K\to+\infty}\mathcal{V}^{n,K}_{t}$ is well defined in $\mathcal{L}^{q}(\mathcal{H},\mathbb{E})$ for any $q\geq 1$. Furthermore,
    $\mathcal{V}^{n,K}_{t,\mathfrak{d}}=F^{n}_{t,\mathfrak{d}}\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{t+s})-v(\mathfrak{d}_{t})\right)\right)}\mathrm{d}s=F^{n}_{t,\mathfrak{d}}e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t})}\int_{s=0}^{K}e^{\varrho_{n}s}\exp{(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s}))}\mathrm{d}s,$
    where $\varrho_{n}$ is defined in the lemma, and from Assumption 2.5 and Corollary 2.4,
    $F^{n}_{t}=F^{n}_{0}\exp\left(\int_{u=0}^{t}\mathfrak{a}^{n\cdot}(\Theta_{u}\mathrm{d}u+\mathrm{d}v(\mathfrak{d}_{u}))+\sigma_{n}\mathrm{d}\mathcal{W}^{n}_{t}\mathrm{d}u\right)=F^{n}_{0}e^{\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{t})-v(\mathfrak{d}_{0})\right)}\exp\left(\mathfrak{a}^{n\cdot}\mathcal{A}^{\circ}_{t}+\sigma_{n}\mathcal{W}^{n}_{t}\right).$
    We then have
    $F^{n}_{t,\mathfrak{d}}e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t})}\int_{s=0}^{K}e^{\varrho_{n}s}\exp{(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s}))}\mathrm{d}s=F^{n}_{0}e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{0})}\exp\left(\mathfrak{a}^{n\cdot}\mathcal{A}^{\circ}_{t}+\sigma_{n}\mathcal{W}^{n}_{t}\right)\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s.$
  2. 2.
    Moreover,
    • (a)
      If $t<t_{\circ}$, then
      $\begin{split}\mathfrak{R}^{n,K}_{t}(\mathfrak{d})&:=\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s\\ &=\int_{s=0}^{t_{\circ}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s+\int_{s=t_{\circ}-t}^{t_{\star}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s+\int_{s=t_{\star}-t}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s\\ &=e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\circ}})}\frac{1-e^{\varrho_{n}(t_{\circ}-t)}}{-\varrho_{n}}+\int_{s=t_{\circ}-t}^{t_{\star}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s+e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})+\varrho_{n}(t_{\star}-t)}\frac{1-e^{\varrho_{n}(K-t_{\star}+t)}}{-\varrho_{n}}.\end{split}$
    • (b)
      If $t_{\circ}\leq t<t_{\star}$, then
      $\begin{split}\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s&=\int_{s=0}^{t_{\star}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s+\int_{s=t_{\star}-t+1}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s\\ &=\int_{s=0}^{t_{\star}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s+e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})+\varrho_{n}(t_{\star}-t+1)}\frac{1-e^{\varrho_{n}(K-t_{\star}+t)}}{-\varrho_{n}}.\end{split}$
    • (c)
      If $t\geq t_{\star}$, then
      $\begin{split}\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s&=\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})\right)}\mathrm{d}s=e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})}\frac{1-e^{\varrho_{n}(K+1)}}{-\varrho_{n}}.\end{split}$
    Finally, $e^{\varrho_{n}(K+1)}$ and $e^{\varrho_{n}(K-t_{\star}+t)}$ converge to 00 for $\varrho_{n}<0$ as $K$ tends to infinity, and the result follows.
  3. 3.
    We denote
    $V^{n,T}_{t,\mathfrak{d}}:=\mathbb{E}_{t}\left[\int_{t}^{T}e^{-r(s-t)}F^{n}_{s,\mathfrak{d}}\mathrm{d}s\right].$
    As we have from (2.18), $F^{n}_{s,\mathfrak{d}}=F^{n}_{t,\mathfrak{d}}\exp\left(\mathfrak{a}^{n\cdot}(\mathcal{A}_{s}-\mathcal{A}_{t})+\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)+\sigma_{n}(\mathcal{W}^{n}_{s}-\mathcal{W}^{n}_{t})\right)$, and given that for all $h,t\geq 0$,
    $\mathcal{A}_{t+h}=\mathcal{A}_{t}+\mu h+\varsigma\Upsilon_{h}\mathcal{Z}_{t}-\varsigma\Gamma^{-1}\int_{t}^{t+h}\left(e^{-\Gamma(t+h-s)}-\mathbf{I}_{I}\right)\Sigma\mathrm{d}B_{s}^{\mathcal{Z}}.$
    We obtain
    $V^{n,T}_{t,\mathfrak{d}}$ $=\mathbb{E}_{t}\left[\int_{t}^{T}e^{-r(s-t)}F^{n}_{t}\exp\left(\mathfrak{a}^{n\cdot}(\mathcal{A}_{s}-\mathcal{A}_{t})+\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)+\sigma_{n}(\mathcal{W}^{n}_{s}-\mathcal{W}^{n}_{t})\right)\mathrm{d}s\right]$
    $=F^{n}_{t,\mathfrak{d}}\int_{t}^{T}e^{\left(\frac{1}{2}\sigma_{n}^{2}-r\right)(s-t)}\exp\left(\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)\right)\mathbb{E}_{t}\left[\exp\left(\mathfrak{a}^{n\cdot}(\mathcal{A}_{s}-\mathcal{A}_{t})\right)\right]\mathrm{d}s$
    $=F^{n}_{t,\mathfrak{d}}\int_{t}^{T}e^{\left(\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu-r\right)(s-t)}\exp\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)\exp\left(\varsigma\mathfrak{a}^{n\cdot}\Upsilon_{s-t}\mathcal{Z}_{t}+\frac{1}{2}\mathfrak{a}^{n\cdot}\Sigma^{\mathcal{A},h}_{t}(\mathfrak{a}^{n\cdot})^{\top}\right)\mathrm{d}s.$
    Then using Hölder’s inequality (with $1=\frac{1}{p}+\frac{1}{q}$), we have
    $\lVert V^{n,T}_{t,\mathfrak{d}}\rVert_{1}$ $\leq\lVert F^{n}_{t,\mathfrak{d}}\rVert_{q}\left\lVert\int_{t}^{T}e^{\left(\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu-r\right)(s-t)}\exp\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)\exp\left(\varsigma\mathfrak{a}^{n\cdot}\Upsilon_{s-t}\mathcal{Z}_{t}+\frac{1}{2}\mathfrak{a}^{n\cdot}\Sigma^{\mathcal{A},s-t}_{t}(\mathfrak{a}^{n\cdot})^{\top}\right)\mathrm{d}s\right\rVert_{p}$
    $\leq\lVert F^{n}_{t,\mathfrak{d}}\rVert_{q}\int_{t}^{T}e^{\left(\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu-r\right)(s-t)}\exp\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)\exp\left(\frac{1}{2}\mathfrak{a}^{n\cdot}\Sigma^{\mathcal{A},s-t}_{t}(\mathfrak{a}^{n\cdot})^{\top}\right)\left\lVert\exp\left(\varsigma\mathfrak{a}^{n\cdot}\Upsilon_{s-t}\mathcal{Z}_{t}\right)\right\rVert_{p}\mathrm{d}s.$
    Observe that under Assumption 2.3, there exists a constant $\mathfrak{C}_{\mathfrak{d}}>0$ such that
    $\sup_{n,s,t}\exp\left(\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)\right)\leq\mathfrak{C}_{\mathfrak{d}}\,.$
    Given that $\mathcal{Z}$ is stationary and $\Upsilon_{s-t}$ is bounded ((A.1)), there exists $\mathfrak{C}_{n,p}>0$ so that $\leq\mathfrak{C}_{n,p}$
    $\left\lVert\exp\left(\varsigma\mathfrak{a}^{n\cdot}\Upsilon_{s-t}\mathcal{Z}_{t}\right)\right\rVert_{p}=\mathbb{E}\left[\exp\left(\varsigma p\mathfrak{a}^{n\cdot}\Upsilon_{s-t}\mathcal{Z}_{t}\right)\right]^{\frac{1}{p}}\leq\mathfrak{C}_{n,p}.$
    Moreover,
    $\exp\left(\frac{1}{2}\mathfrak{a}^{n\cdot}\Sigma^{\mathcal{A},h}_{t}(\mathfrak{a}^{n\cdot})^{\top}\right)$ $=\exp\left(\frac{1}{2}\varsigma^{2}\int_{0}^{s-t}\mathfrak{a}^{n\cdot}\Upsilon_{u}\Sigma\Sigma^{\top}\Upsilon_{u}^{\top}(\mathfrak{a}^{n\cdot})^{\top}\mathrm{d}u\right)$
    $\leq\exp\left(\frac{1}{2}\varsigma^{2}\int_{0}^{s-t}\lVert\mathfrak{a}^{n\cdot}\rVert^{2}\lVert\Sigma\rVert^{2}\lVert\Upsilon_{u}\rVert^{2}\mathrm{d}u\right)$
    $\leq\exp\left(\frac{1}{2}\varsigma^{2}\frac{c_{\Gamma}^{2}}{\lambda_{\Gamma}^{2}}\lVert\mathfrak{a}^{n\cdot}\rVert^{2}\lVert\Sigma\rVert^{2}(s-t)\right).$
    Next, we can write
    $\lVert V^{n,T}_{t,\mathfrak{d}}\rVert_{1}$ $\leq\mathfrak{C}_{\mathfrak{d}}\mathfrak{C}_{n,p}\lVert F^{n}_{t,\mathfrak{d}}\rVert_{q}\int_{t}^{T}\exp{\left(\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu+\frac{1}{2}\varsigma^{2}\frac{c_{\Gamma}^{2}}{\lambda_{\Gamma}^{2}}\lVert\mathfrak{a}^{n\cdot}\rVert^{2}\lVert\Sigma\rVert^{2}-r\right)(s-t)}\mathrm{d}s,$
    and if (2.24) is satisfied and $T\to+\infty$, then $V^{n,K}_{t,\mathfrak{d}}$ converges to $V^{n}_{t,\mathfrak{d}}$. Finally, similar methods must be used to show $\mathbb{E}\left[\left|\frac{V^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}-\frac{\mathcal{V}^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}\right|\right]\leq C\varsigma$.