### v2

# Impact of the carbon price on credit portfolio’s loss with stochastic collateral

Abstract
The aim of this work is to propose an end-by-end modeling framework to evaluate the risk measures of a bank’s portfolio of collateralized loans when the bank, the borrowers, as well as their guarantees operate in an economy subject to the climate transition. The economy, organized in sectors, is driven by its productivity which is a multidimensional Ornstein-Uhlenbeck process while the climate transition is declined thanks to the carbon price and carbon intensities, continuous deterministic processes. We thus derive the dynamics of macroeconomic variables for each climate transition scenario. By considering that a firm defaults if it is over-indebted i.e. its market value – depreciated due to the carbon price – becomes less than the market value of its debt, we define each loan’s loss at default as the difference between Exposure at Default (EAD) and the liquidated collateral, which will help us to define the Loss Given Default (LGD) – the expected percentage of exposure that is lost if a debtor defaults. We consider two types of collateral: financial asset such as invoices, cash, or investments or physical asset such as real estate, business equipment, or inventory. First, if it is a financial asset, we model the later by the continuous time version of the discounted cash flows methodology, where the cash flows growth is driven by the instantaneous output growth, the instantaneous growth of a carbon price function, and an arithmetic Brownian motion. Secondly, for physical asset, we focus on the example of a property in housing market. Therefore, we define, as Sopgoui, (2024), its value as the difference between the price of an equivalent efficient building following an exponential Ornstein-Uhlenbeck (OU) as well as the actualized renovation costs and the actualized sum of the future additional energy costs due to the inefficiency of the building, before an optimal renovation date which depends on the carbon price process. Finally, we obtain expressions for risk measures of a portfolio of collateralized loans as a function of various parameters and variables, mainly those linked to the climate transition, such as the carbon price and the energy efficiency of buildings. These risk measures will be used by banks, depending on the climate transition scenarios, to define their operating expenses, the fees applied to clients, as well as their economic and regulatory capital.
keywords:
Credit risk , Climate risk , Collateral , Stochastic modelling , Transition risk , Carbon price , Firm valuation , Loss Given Default
journal: arXiv
This research is part of the PhD thesis in Mathematical Finance of Lionel Sopgoui whose works are funded by a CIFRE grant from BPCE S.A. The opinions expressed in this research are those of the authors and are not meant to represent the opinions or official positions of BPCE S.A.
We would like to thank Jean-François Chassagneux, Antoine Jacquier, Smail Ibbou, and Géraldine Bouveret for helpful comments on an earlier version of this work.
Notations
  • 1.
    $\mathbb{N}$ is the set of non-negative integers, :=$\mathbb{N}^{*}:=\mathbb{N}\setminus\{0\}$, and $\mathbb{Z}$ is the set of integers.
  • 2.
    $\mathbb{R}^{d}$ denotes the $d$-dimensional Euclidean space, $\mathbb{R}_{+}$ is the set of non-negative real numbers, +:=+$\mathbb{R}_{+}^{*}:=\mathbb{R}_{+}\setminus\{0\}$.
  • 3.
    For $n,d\in\mathbb{N}^{*}$, $\mathbb{R}^{n\times d}$ is the set of real-valued $n\times d$ matrices ($\mathbb{R}^{n\times 1}=\mathbb{R}^{n}$), $\mathbf{I}_{n}$ is the identity $n\times n$ matrix and $\mathbf{1}:=(1,\ldots,1)\in\mathbb{R}^{n}$.
  • 4.
    $x^{i}$ denotes the $i$-th component of the vector $x\in\mathbb{R}^{d}$. For $A:=(A^{ij})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$, we denote by $A^{\top}:=(A^{ji})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$ the transpose matrix, and $\lambda(A)$ denotes the spectrum of $A$.
  • 5.
    For a given finite set $S$, we define as the cardinal of $S$, #$\#S$.
  • 6.
    For all $x,y\in\mathbb{R}^{d}$, we denote the scalar product $x^{\top}y$, the Euclidean norm $|x|:=\sqrt{x^{\top}x}$ and for a matrix $M\in\mathbb{R}^{d\times d}$, we denote
    $|M|:=\sup_{a\in\mathbb{R}^{d},|a|\leq 1}|Ma|.$
  • 7.
    $(\Omega,\mathcal{H},\mathbb{P})$ is a complete probability space.
  • 8.
    For $p\in[1,\infty]$, ${E}$ is a finite dimensional Euclidian vector space and for a $\sigma$-field $\mathcal{H}$, $\mathcal{L}^{p}(\mathcal{H},{E})$, denotes the set of $\mathcal{H}$-meassurable random variable $X$ with values in ${E}$ such that $\|X\|_{p}:=\left(\mathbb{E}\left[|X|^{p}\right]\right)^{\frac{1}{p}}<\infty$ for $p<\infty$ and for $p=\infty$, $\|X\|_{\infty}:=\mathrm{esssup}|X(\omega)|<\infty$.
  • 9.
    For a filtration $\mathbb{G}$, $p\in[1,+\infty]$ and $I\in\mathbb{N}^{*}$, +$\mathscr{L}^{p}_{+}(\mathbb{G},(0,\infty)^{I})$ is the set of continuous-time processes that are $\mathbb{G}$-adapted valued in $(0,\infty)^{I}$ and which satisfy
    $\lVert X_{t}\rVert_{p}<\infty\text{ for all }t\in\mathbb{R}_{+}.$
  • 10.
    If $X$ and $Y$ are two random variables $\mathbb{R}^{d}$-valued, for $x\in\mathbb{R}^{d}$, we note $Y|X=x$ the conditional distribution of $Y$ given $X=x$, and $Y|\mathcal{F}$ the conditional distribution of $Y$ given the filtration $\mathcal{F}$.
  • 11.
    If X is a log-Normal random variable, we note $X\sim\mathcal{LN}(\mu,\sigma^{2})$ with $\mu:=\mathbb{E}[\log(X)]$ (the mean of $\log(X)$) and $\sigma^{2}:=\mathbb{V}[\log(X)]$ (the variance of $\log(X)$).
  • 12.
    For $x,y\in\mathbb{R}$, we note $\Phi_{2}(x,y;\rho)$ is the cumulative distribution function of the bi-variate Gaussian vector $(X,Y)$ with correlation $\rho$ on the space $[-\infty,x]\times[-\infty,y]$.
  • 13.
    For $A\subset\mathbb{R}$ and $B\subset\mathbb{R}^{I}$, if $f:A\to B,t\mapsto f(t)$ is a differentiable function, we note $\dot{f}$ its first derivative. We also note $\mathfrak{D}(A,B)$ the set of derivative functions from $A$ to $B$.
1 Introduction
There is now evidence that climate change has adverse effects on the environment, on global warming, and on human societies. Climate risk has two components: physical risk which arises from alterations in climate variables, like increasing temperatures and melting ice, or from extreme weather events, such as droughts or typhoons that can damage infrastructure and endanger populations; and transition risk which comes from the necessity of transitioning to a low-carbon economy, leading to the implementation of regulatory/political measures, potential technological changes, and the evolution of consumer preferences. We focus in this work on transition risk and particularly on political changes. The most well-known way to that is the carbon price. This means that the greenhouse gases (GHG) emissions of economic agents are charged. For instance, the carbon price can modify the three main components of credit risk, which are the borrower’s cash flows, the value of his assets, and the value of his guarantees. The aim of this work is to compute the losses of a credit portfolio when the bank, its borrowers, and their guarantees operate in an economy that undergoes the climate transition.
When an obligor (firm, government, or individual) defaults, the creditor (bank) stands to lose its money. One way to ensure the stability of the banks’ business and more generally the soundness of the whole financial system, is ideally, to anticipate when the default will happen and the amount a bank could lose. In order to achieve that, the Basel Committee on Banking Supervision, (2017) introduces four parameters: the probability of default (PD) which measures the default risk associated with each borrower, the exposure at default (EAD) which quantifies the outstanding debt at the time of default, the loss given default (LGD) which captures the expected percentage of EAD that is lost if the debtor defaults, and the effective maturity T which represents the duration of the credit. By using these parameters, banks can compute various risk measures (such as expected, unexpected, and stressed losses) which help later on to determine provisions, as well as economic and regulatory capital. Using Gordy, (2003), we can obtain the losses as a function of PD, LGD, EAD, and T. An essential part of a bank risk division is to estimate how the risk measures change with various factors such as time and economic conditions.
In Bouveret et al., (2025), we propose a methodology in discrete time without collateral (i.e., with LGD independent of the economy and climate transition). However, when a debtor defaults, the bank can lose all or part of its exposure. The fraction of the loss relative to EAD is LGD while the recovery rate is the fraction of EAD recovered so that $LGD=1-Recovery$. So modeling LGD or recovery is equivalent. According to Chalupka and Kopecsni, (2008), there are three types of LGD: ”market LGD is observed from market prices of defaulted bonds or marketable loans soon after the actual default event. Workout LGD is derived from a set of estimated cash flows resulting from a workout and collection process, properly discounted to a date of default. Thirdly, implied market LGD is derived from risky but not defaulted bond prices using a theoretical asset pricing model”. When consider the IRB approach [^1] Internal Rating Based (see Basel Committee on Banking Supervision, (2017)), LGD refers to Workout LGD and there are several techniques to model it. In economic modeling, as detailed by Bastos, (2010); Roncalli, (2020), LGD is a function of different factors which can be external to the issuer, specific to the issuer, or specific to the debt issuance. That function can be obtained through logistic regression, regression trees, or neural networks. In stochastic modeling, it is assumed that LGD follows a given (parametric or non-parametric) distribution. For example, LGD is commonly modeled by a Beta distribution as Roncalli, (2020)[Page 193] and Chalupka and Kopecsni, (2008).Fermanian, (2020), for his part, proposes a joint modeling of PD and LGD by writing the potential loss at default as the difference between the debt amount (EAD) and the assets at the default date.
In a credit portfolio, there are secured (or collateralized ) and unsecured loans. We call secured loans when there is collateral (also called guarantee) and unsecured loans when there is not. Of course, not all borrowers put up collateral when taking out loans, and there is evidence that loans with collateral may be riskier for lenders. Berger and Udell, (1990) note that banks require more collateral from riskier borrowers, and although seizing a guarantee when a borrower defaults reduces the bank’s loss, this does not always fully compensate the bank for the fact that the loan was initially riskier. For a secured loan, when the counterpart defaults, the bank liquidates the collateral, and if the EAD is not reached, it can recover the remaining amount by liquidating other assets (called residual recovery). These guarantees can be tangible assets (buildings, business equipments, inventories, etc.) or intangible assets (cash deposits, public bonds, securities, etc.) as noted by Berger and Udell, (1990), Blazy and Weill, (2013). For secured loans, the recovery (that is, $1-\mathrm{LGD}$) is therefore made up of both the value of the collateral at the date of default and the value of the residual recovery Frontczak and Rostek, (2015), and Pelizza and Schenk-Hoppé, (2020). However, in most LGD modeling approaches, the presence or absence of the guarantees is often overlooked. We will tackle these limits here by considering two examples of guarantees: either a security or a (commercial or residential) building, which both will be affected by the climate transition.
A security can represent ownership in a corporation in the form of stock, a creditor relationship with a governmental body, or a corporation represented by owning that entity’s bond; or rights to ownership as represented by an option. A security generates a stream of cash flows. The proxy of the security value is the (in)finite sum of the present value of the future cash flows. As we already know from what we propose in Bouveret et al., (2025), the value of the security (particularly if it is a firm) will be affected by the transition risk. We will therefore revisit the results of Bouveret et al., (2025) where carbon price dynamics affects the firm value and credit risk measures such as probability of default, expected and unexpected losses. In particular, we redesign the multisectoral model with carbon price, the firm valuation model, and the credit risk model proposed in continuous time.
In the same way, a commercial or residential building price will be affected by the climate transition, for example through Energy Performance (or Energy Efficiency) as mentioned in Aydin et al., (2020); Franke and Nadler, (2019). Ter Steege and Vogel, (2021) quantify the depreciation by writing the price difference per square meter between two properties with different energy efficiency as the sum of the discounted value of the (expected) energy cost differences. Sopgoui, (2024) enhances this work by assuming that initially the owner of the building incurs additional energy costs (written as a function of the carbon price) due to the inefficiency of his property. Then, he may decide to spend money on renovations to make his building energy efficient. After the renovation, he no longer incurs additional energy costs. Therefore, he obtains the value of his building as the difference between the price of an equivalent efficient building following an exponential Ornstein-Uhlenbeck as well as the actualized renovation costs and the actualized sum of the future additional energy costs.
Thanks to our micro-founded approach to a credit portfolio, we obtain explicit expressions of risk measures (especially probability of default, loss given default, expected and unexpected losses) depending on variables and parameters traditionally found in the literature, such as the the borrower’s cash flows, its exposure, the calculation time, the nature of the collateral, the liquidation time, date, and costs, the interest rate, the productivity, etc. Additionally, they now depend on climate transition-specific parameters, such as carbon price, carbon intensities, energy efficiency, or renovation costs. In summary, our framework enables the calculation of the potential loss of a collateralized loan and also measures the variation in this loss depending on the borrower’s and the collateral’s virtuousness in terms of GHG emissions. Precisely, we will show that (1) expected and unexpected losses increase when the price of carbon increases, (2) the presence of collateral significantly reduces expected and unexpected losses, and (3) the positive effect of collateral on losses is reduced if the collateral is energy inefficient (for a building) or depends on a polluting sector (financial asset). And at each stage, variations can be precisely quantified.
The rest of the present work is organized as follows. In section 2, we clarify the problem that this paper aims to solve. We revisit in  section 3 the results of Bouveret et al., (2025) in a continuous time setting, namely a multisectoral economic model with carbon price, a firm valuation model, and a credit risk model. In section 4, we define the loss at default as the difference between EAD and the liquidated collateral, which will help us to define LGD. If the collateral is a financial asset, we model it in section 4.2 by the continuous time version of the discounted cash flows, where the cash flows growth is driven by the instantaneous consumption growth, the instantaneous growth of a carbon price’s function and a Brownian motion. If the collateral is a building, we will use the housing valuation under the climate transition proposed in Sopgoui, (2024) to compute the loss of the portfolio. The last sections (5 and 6) are dedicated to estimations, simulations, and discussion.
2 The problem
We consider a bank credit portfolio composed of $N\in\mathbb{N}^{*}$ firms in a closed economy (in other words, no import and no export). In credit risk assessment, one of the first steps is to create homogeneous sub-portfolios of firms. As we are dealing here with climate transition risk, we would like to classify firms by carbon intensity – the quantity of GHG in tons emits for each unit of production (in ton per euro) [^2] To obtain this, one could simply divide each company’s greenhouse gas emissions by its revenue. – so that firms with similar carbon intensities belong to a same homogeneous sub-portfolio. It should be noted that in the absence of a climate transition, firms are traditionally clustered in terms of industry, geography, size, and credit rating, for example.
We thus assume $I\in\mathbb{N}^{*}$ ($I\leq N$) homogeneous carbon emission sectors in the economy. Nevertheless, as we rarely have the firm individual carbon emissions/intensities (especially for small and medium enterprises), we assume that each company has the carbon intensity of its industry sector. This amounts to grouping ”industry sectors” into $I$ ”carbon intensity sectors”. From now on, sectors are to be interpreted as carbon emission sectors.
Definition 2.1.
We divide our portfolio into $I$ disjunct sub-portfolios $g_{1},\ldots,g_{I}$ so that each sub-portfolio represents a single risk class and the firms in each sub-portfolio belong to a single carbon emission sectors. From now on, we denote $\mathcal{I}$ the set of sectors with cardinal $I\in\mathbb{N}^{*}$. We also fix $n_{i}:=\min{\{n\in\{1,\ldots,N\}\text{ such that }n\in g_{i}\}}$ for each $i\in\mathcal{I}$. Therefore, firm $n_{i}$ is a representative of the group $i$.
We would like to know how the whole portfolio loss and sub-portfolios losses would be affected should the regulator introduce a carbon price in the economy, in order to mitigate the effects of climate change. This precisely amounts to quantifying the distortion over time of credit risk measures created by the introduction of a carbon price. For example, if the government decides to charge firms and households GHG emissions between 2025 and 2035, a bank would like to estimate today how the probability of a company to default in 2030 is impacted.
The bank’s potential loss caused by a firm depends essentially on the default date and on the liquidation of the guarantees if they exist. The firm as well as the guarantee belong to the same economy which is subject to the climate transition. Thus, we build in the first stage a dynamic, stochastic, and multisectoral economic model in which direct and indirect GHG emissions from companies as well as direct GHG emissions from households are charged. We choose a representative firm in each sector and a representative household for the whole economy. By observing that each firm belongs to a sector and its cash flows are a proportion of its sales. The latter are themselves a proportion of the sectoral output. We obtain the cash flows dynamics that we use to determine the value of firms in an environment where GHG emissions are charged. Then, starting from a default model in which a company defaults if its value falls below its debt, we calculate the probability of default of each firm. Finally, considering that the collateral is a GHG-emitting asset in the economy, we compute the distortion of the (associated statistics of) loss – defined as the difference between the exposure and the liquidated collateral – by the introduction of a carbon price.
3 Main assumptions and results of Bouveret et al., (2025) in continuous time
In Bouveret et al., (2025), we propose an end-to-end methodology, starting from a transition scenario modeled by the carbon price and carbon intensities, to the impact on different credit risk measures – especially the probability of default, the expected and unexpected losses – with both EAD and LGD deterministic and without collateral. In this section, we revisit that framework in continuous time. Precisely, we decline, in continuous time, the two standing assumptions as well as the three main results respectively on the dynamic stochastic multisectoral model with carbon emissions costs, on the firm valuation model, and on the structural credit risk model. Most of the proofs can be derived from the discrete time so we will skip them or detail them in appendix B.
3.1 A Multisectoral Model with Carbon price
The goal here is to derive the dynamics of macroeconomic variables (GHG emissions, output, consumption, labor, and intermediary inputs) per sector and per climate transition scenario.
Each sector $i\in\mathcal{I}$ has a representative firm which produces a single good, so that we can associate sector, firm, and good. We introduce the following standing assumption which describes the productivity, which is considered to have stationary Ornstein-Uhlenbeck dynamics.
Standing Assumption 3.1.
We define the $\mathbb{R}^{I}$-valued process $\mathcal{A}$ which evolves according to
$\left\{\begin{array}[]{rl}\mathrm{d}\mathcal{Z}_{t}&=-\Gamma\mathcal{Z}_{t}\mathrm{d}t+\Sigma\mathrm{d}B_{t}^{\mathcal{Z}}\\ \mathrm{d}\mathcal{A}_{t}&=\left(\mu+\varsigma\mathcal{Z}_{t}\right)\mathrm{d}t\end{array}\quad\textrm{for all }t\in\mathbb{R}_{+},\right.$ (3.1)
where $(B_{t}^{\mathcal{Z}})_{t\in\mathbb{R}_{+}}$ is a $I$-dimensional standard Brownian Motion, and where the constants $\mu,\mathcal{A}_{0}\in\mathbb{R}^{I}$, the matrices $\Gamma,\Sigma\in\mathbb{R}^{I\times I}$, $\mathcal{Z}_{0}\sim\mathcal{N}\left(0,\Sigma\Sigma^{\top}\right)$, and $0<\varsigma\leq 1$ is an intensity of noise parameter that is fixed: it will be used later to obtain a tractable proxy of the firm value. Moreover, $\Sigma$ is a positive definite matrix and $-\Gamma$ is a Hurwitz matrix i.e. its eigenvalues have strictly negative real parts.
The processes $\mathcal{Z}^{i}$ and $\mathcal{A}^{i}$ play a major role in our factor productivity model since, for any $i\in\mathcal{I}$, the total factor productivity of sector $i$ is defined as $A^{i}:=\exp{(\mathcal{A}^{i})}$, so that $\mathcal{Z}^{i}$ is the log-productivity growth and $\mathcal{A}^{i}$ is the cumulative log-productivity growth. In the rest of the paper, the terminology ”productivity” will be used within a context that will allow the reader to understand if the term refers to $\mathcal{Z}^{i}$, $\mathcal{A}^{i}$, or $A^{i}$.
We also introduce the following filtration $\mathbb{G}:=(\mathcal{G}_{t})_{t\in\mathbb{R}_{+}}$ with $\mathcal{G}_{0}:=\sigma(\mathcal{Z}_{0})$ and for $t>0$, $\mathcal{G}_{t}:=\sigma\left(\left\{\mathcal{Z}_{0},B_{s}^{\mathcal{Z}}:s\leq t\right\}\right)$.
Remark 3.2 (OU process).
We have the following results on OU that we will use later:
  1. 1.
    According to Gobet and She, (2016)[Proposition 1], if one assumes that $\mathcal{Z}_{0}$ and $B^{\mathcal{Z}}$ are independent and $\mathcal{Z}_{0}$ is square integrable, then, there exists a unique square integrable solution to the $I$-dimentional Ornstein-Uhlenbeck process $\mathcal{Z}$ satisfying $\mathrm{d}\mathcal{Z}_{t}=-\Gamma\mathcal{Z}_{t}\mathrm{d}t+\Sigma\mathrm{d}B_{t}^{\mathcal{Z}}$, represented as
    $\mathcal{Z}_{t}=e^{-\Gamma t}\left(\mathcal{Z}_{0}+\int_{0}^{t}e^{\Gamma u}\Sigma\mathrm{d}B_{u}^{\mathcal{Z}}\right),\quad\textrm{for all }t\in\mathbb{R}_{+}.$
    Additionally, for any $t,h\geq 0$, the distribution of $\mathcal{Z}_{t+h}$ conditional on $\mathcal{G}_{t}$ is Gaussian $\mathcal{N}\left(M^{\mathcal{Z},h}_{t},\Sigma^{\mathcal{Z},h}_{t}\right)$, with the mean vector
    $M^{\mathcal{Z},h}_{t}:=\mathbb{E}[\mathcal{Z}_{t+h}|\mathcal{G}_{t}]=e^{-\Gamma h}\mathcal{Z}_{t},$ (3.2)
    and the covariance matrix
    $\Sigma^{\mathcal{Z},h}_{t}:=\mathbb{V}[\mathcal{Z}_{t+h}|\mathcal{G}_{t}]=\int_{0}^{h}e^{-\Gamma u}\Sigma\Sigma^{\top}e^{-\Gamma^{\top}u}\mathrm{d}u.$ (3.3)
  2. 2.
    Since $-\Gamma$ is a Hurwitz matrix, then if we note $\lambda_{\Gamma}:=\max_{\lambda\in\lambda(\Gamma)}Re(\lambda)$, there exists $c_{\Gamma}>0$ so that $|e^{-\Gamma t}|<c_{\Gamma}e^{-\lambda_{\Gamma}t}$ for all $t\geq 0$. Therefore, according to Gobet and She, (2016)[Proposition 2], $\mathcal{Z}$ has a unique stationary distribution which is Gaussian with mean 00 and covariance $\int_{0}^{+\infty}e^{-\Gamma u}\Sigma\Sigma^{\top}e^{-\Gamma^{\top}u}\mathrm{d}u$.
  3. 3.
    We show in Sopgoui, (2024)[Appendix A] that for any $t,h\geq 0$, we have
    $\mathcal{A}_{t+h}=\mathcal{A}_{t}+\int_{t}^{t+h}(\mu+\varsigma\mathcal{Z}_{s})\mathrm{d}s=\mu h+\varsigma\int_{t}^{t+h}\mathcal{Z}_{s}\mathrm{d}s,$
    and conditionally on $\mathcal{G}_{t}$, $\mathcal{A}_{t+h}$ has an $I$-dimensional normal distribution with the mean vector $M^{\mathcal{A},h}_{t}:=\mu h+\varsigma\Upsilon_{h}\mathcal{Z}_{t}+\mathcal{A}_{t}$, with
    $\Upsilon_{h}:=\int_{0}^{h}e^{-\Gamma s}\mathrm{d}s=\Gamma^{-1}(\mathbf{I}_{I}-e^{-\Gamma h}),$ (3.4)
    and the covariance matrix
    $\Sigma^{\mathcal{A},h}_{t}:=\varsigma^{2}\Gamma^{-1}\left(\int_{0}^{h}\left(e^{-\Gamma u}-\mathbf{I}_{I}\right)\Sigma\Sigma^{\top}\left(e^{-\Gamma u}-\mathbf{I}_{I}\right)\mathrm{d}u\right)(\Gamma^{-1})^{\top}=\varsigma^{2}\int_{0}^{h}\Upsilon_{u}\Sigma\Sigma^{\top}\Upsilon_{u}^{\top}\mathrm{d}u.$ (3.5)
  4. 4.
    For later use, we define $\mathcal{A}^{\circ}_{t}:=\mathcal{A}_{t}-\mathcal{A}_{0}$, and observe that $(\mathcal{A}^{\circ}_{t},\mathcal{Z}_{t})_{t\geq 0}$ is a Markov process.
Firms emit GHG when they consume intermediary input from other sectors and when they produce output. Likewise, households emit GHG when they consume. All these emissions are charged through a carbon price dynamics. For the whole economy, we introduce a deterministic and exogenous carbon price in euro per ton of CO2-equivalent. It allows us to model the impact of the transition pathways on the whole economy. We shall then assume the following setting.
Standing Assumption 3.3.
We introduce the carbon price and the carbon intensities (the quantity of GHG in tons emits for each unit of production/consumption) processes:
  1. 1.
    Let $0\leq t_{\circ}<t_{\star}$ be given. The complete carbon price sequence noted $\delta$ satisfies
    • (a)
      for $t\in[0;t_{\circ}]$, $\delta_{t}=\delta_{0}\in\mathbb{R}_{+}$, namely the carbon price is constant;
    • (b)
      for $t\in(t_{\circ},t_{\star})$, $\delta_{t}\in\mathbb{R}_{+}$, the carbon price may evolve;
    • (c)
      for $t\geq t_{\star}$, $\delta_{t}=\delta_{t_{\star}}\in\mathbb{R}_{+}$, namely the carbon price is constant.
    We assume moreover that $t\mapsto\delta_{t}$ is $\mathfrak{D}(\mathbb{R}_{+},\mathbb{R}_{+})$.
  2. 2.
    We also introduce carbon intensities as the sequences $\tau$, $\zeta$, and $\kappa$ being respectively $\mathbb{R}_{+}^{I}$, $\mathbb{R}_{+}^{I\times I}$, and $\mathbb{R}_{+}^{I}$-processes, and representing respectively carbon intensities on firm’s output, on firm’s intermediary consumption, and on household’s consumption, and satisfying for all $t\in\mathbb{R}_{+}$ and
    $\mathfrak{y}\in\{\tau^{1},\ldots,\tau^{I},\zeta^{11},\zeta^{12},\ldots,\zeta^{II-1},\zeta^{II},\kappa^{1},\ldots,\kappa^{I}\}$,
    $\mathfrak{y}_{t}=\left\{\begin{array}[]{ll}&\mathfrak{y}_{0}\exp{\left(g_{\mathfrak{y},0}\frac{1-\exp{(-\theta_{\mathfrak{y}}t)}}{\theta_{\mathfrak{y}}}\right)}\qquad\text{if }0\leq t\leq t_{\star}\\ &\mathfrak{y}_{0}\exp{\left(g_{\mathfrak{y},0}\frac{1-\exp{(-\theta_{\mathfrak{y}}t_{\star})}}{\theta_{\mathfrak{y}}}\right)}\qquad\text{else},\end{array}\right.$ (3.6)
    with $\mathfrak{y}_{0},g_{\mathfrak{y},0},\theta_{\mathfrak{y}}>0$. For each $t\geq 0$, we call $\mathfrak{y}_{t}\delta_{t}$ the emissions cost rate at time $t$.
  3. 3.
    For each $i\in\mathcal{I}$ and for each $t\in\mathbb{R}_{+}$,
    $\delta_{t}\max_{i\in\mathcal{I}}{\tau^{i}_{0}}<1.$ (3.7)
Remark 3.4.
At each time $t\geq 0$ and for $i\in\mathcal{I}$,
  • 1.
    $\tau_{t}^{i}$ represents the carbon intensity of the firms production of sector $i$ at time $t$;
  • 2.
    $\kappa_{t}^{i}$, the carbon intensity of the household final consumption in sector $i$ at time $t$,
  • 3.
    and for $j\in\mathcal{I}$, $\zeta^{ji}_{t}$, the carbon intensity of the firms of sector $i$ intermediate consumption in sector $j$ at time $t$.
In the following, we will note for all $t\geq 0$,
$\mathfrak{d}_{t}:=(\delta_{t}\tau_{t},\delta_{t}\zeta_{t},\delta_{t}\kappa_{t}).$ (3.8)
An example of carbon price process
We assume the regulator fixes $t_{\circ}\geq 0$ when the transition starts and the transition horizon time $t_{\star}>t_{\circ}$, the carbon price at the beginning of the transition $P_{carbon}>0$, at the end of the transition $\delta_{t_{\star}}>P_{carbon}$, and the annual growth rate $\eta_{\delta}>0$. Then, for all $t\geq 0$,
$\delta_{t}=\left\{\begin{array}[]{ll}\displaystyle P_{carbon},&\mbox{if }t\leq t_{\circ},\\ \displaystyle P_{carbon}e^{\eta_{\delta}(t-t_{\circ})},&\mbox{if }t\in(t_{\circ},t_{\star}],\\ \displaystyle\delta_{t_{\star}}=P_{carbon}e^{\eta_{\delta}(t_{\star}-t_{\circ})},&\mbox{otherwise}.\end{array}\right.$ (3.9)
In the example above that will be used in the rest of this work, we assume that the carbon price increases. However, there are several scenarios that could be considered, including a carbon price that would increase until a certain year before leveling off or even decreasing. We also assume an unique carbon price for the entire economy whereas we could proceed differently. For example, the carbon price could increase for production when stabilize or disappear on households in order to avoid social movements and so on. The framework can be adapted to various sectors as well as scenarios.
In our framework, a representative firm in each sector which maximizes its profits by choosing, at each time and for a given productivity, the quantities of labor and intermediary inputs, while, a representative household solves a dynamic optimization problem to decide how to allocate its consumption expenditures among the different goods and hours worked and among the different sectors. We assume that the utility function $U:\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*}\to\mathbb{R},(c,h)\mapsto\log{c}-\frac{h^{1+\varphi}}{1+\varphi}$ with $\varphi\geq 0$; $c$ and $h$ represents respectively the household consumption and the labor. Moreover, $\boldsymbol{\lambda}$ (respectively, $\psi$) are matrix in $(\mathbb{R}_{+}^{*})^{I\times I}$ (respectively vector in $(\mathbb{R}_{+}^{*})^{I}$) of the elasticities of intermediary inputs (respectively labor). We also assume a constant return to scale, namely
$\psi^{i}+\sum_{j\in\mathcal{I}}\boldsymbol{\lambda}^{ji}=1,\qquad\text{for each }i\in\mathcal{I}.$ (3.10)
Since the productivity and the carbon price processes are continuous, the firms and households problems are well posed and their solutions exist. More details are given in appendix B. In the following proposition, we give an explicit expression of the output.
Proposition 3.5.
For $(\overline{\tau},\overline{\zeta},\overline{\kappa},\overline{\delta})\in\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}^{I\times I}\times\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}$, let us note
$\Psi(\overline{\mathfrak{d}}):=\left(\psi^{i}\frac{1-\overline{\tau}^{i}\overline{\delta}}{1+\overline{\kappa}^{i}\overline{\delta}}\right)_{i\in\mathcal{I}}\ \quad\text{and}\quad\Lambda(\overline{\mathfrak{d}}):=\left(\boldsymbol{\lambda}^{ji}\frac{1-\overline{\tau}^{i}\overline{\delta}}{1+\overline{\zeta}_{t}^{ji}\overline{\delta}}\frac{1+\overline{\kappa}^{j}\overline{\delta}}{1+\overline{\kappa}^{i}\overline{\delta}}\right)_{j,i\in\mathcal{I}},$ (3.11)
with $\overline{\mathfrak{d}}:=(\overline{\delta}\overline{\tau},\overline{\delta}\overline{\zeta},\overline{\delta}\overline{\kappa})$. Assume that
  1. 1.
    $\mathbf{I}_{I}-\boldsymbol{\lambda}$ is not singular,
  2. 2.
    $\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{t})^{\top}$ is not singular for all $t\in\mathbb{R}_{+}$.
Then, for all $t\in\mathbb{R}_{+}$, there exists an unique couple of consumption and output $({C}_{t},{Y}_{t})$ solving the (dynamic stochastic) multisectoral model. Moreover, for all $t\in\mathbb{R}_{+}$.
  1. 1.
    if $\mathfrak{e}_{t}^{i}:=\frac{Y^{i}_{t}}{C^{i}_{t}}$ for $i\in\mathcal{I}$, we have
    $\mathfrak{e}_{t}=\mathfrak{e}(\mathfrak{d}_{t}):=(\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{t})^{\top})^{-1}\mathbf{1}.$ (3.12)
  2. 2.
    For $(\overline{\tau},\overline{\zeta},\overline{\kappa},\overline{\delta})\in\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}^{I\times I}\times\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}$,
    $v^{i}(\overline{\mathfrak{d}}):=\log\left((\mathfrak{e}(\overline{\mathfrak{d}})^{i})^{-\frac{\varphi\psi^{i}}{1+\varphi}}\left(\Psi^{i}(\overline{\mathfrak{d}})\right)^{\frac{\psi^{i}}{1+\varphi}}\prod_{j\in\mathcal{I}}\left(\Lambda^{ji}(\overline{\mathfrak{d}})\right)^{\boldsymbol{\lambda}^{ji}}\right)_{i\in\mathcal{I}}+((\mathbf{I}_{I}-\boldsymbol{\lambda})\log{(\mathfrak{e}(\overline{\mathfrak{d}}))}),$ (3.13)
    We obtain
    ${Y}_{t}=\exp{\left((\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}{\left(\mathcal{A}_{t}+v(\mathfrak{d}_{t})\right)}\right)}.$ (3.14)
  3. 3.
    Furthermore, since $\mathfrak{d}\in\mathfrak{D}(\mathbb{R}_{+},[0,1)^{I}\times(\mathbb{R}_{+})^{I\times I}\times(\mathbb{R}_{+})^{I})$, we directly have $\Psi(\overline{\mathfrak{d}}_{\cdot}),\Lambda(\overline{\mathfrak{d}}_{\cdot})\in\mathfrak{D}(\mathbb{R}_{+},\mathbb{R})$. Moreover, $\bar{\mathfrak{d}}\mapsto(\mathbf{I}_{I}-{\Lambda}(\bar{\mathfrak{d}})^{\top})^{-1}$ on $\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}^{I\times I}\times\mathbb{R}_{+}^{I}$) is differentiable, then $(\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{\cdot})^{\top})^{-1})\in\mathfrak{D}(\mathbb{R}_{+},\mathbb{R})$.
The output ${Y}$ is also positive, we can then introduce, from the third item, for all $t\geq 0$, $\mathrm{d}\log{({Y}_{t})}$ representing the instantaneous consumption growth. This proposition is an equivalent of Bouveret et al., (2025)[Theorem 1.10.] in continuous time. In the same spirit, we can also define and determine
  • 1.
    consumption, intermediary inputs, labor, and GHG emissions at each time as well as their logarithmic growth between two instants.
  • 2.
    The sensitivities of the growth of variables to the carbon price. This allows us to answer the following types of questions: what is the impact on all variables if in the whole economy, we only charge GHG emissions coming from the production of companies in sector $1$? The main difficulty here comes from the fact that the coefficients of the inverse of the elasticity matrix $(\boldsymbol{\lambda}^{ji})_{i,j\in\mathcal{I}}$ are not known.
  • 3.
    The law of the logarithmic growth of macroeconomic variables that we obtain by using the Gaussianity and stationarity of logarithmic growth.
  • 4.
    the evolution of the contribution of each sector in production/consumption as well as the evolution of GHG emissions.
3.2 A Firm Valuation Model
The goal in this section is to describe the dynamics of the value of a firm operating in an economy that undergoes the climate transition described by both the carbon price and carbon intensities.
Consider a firm $n\in\{1,\ldots,N\}$. For any time $t\in\mathbb{R}_{+}$ and firm $n$, we note $F^{n}_{t}$ the free cash flows of $n$ at $t$, and $r>0$ the discount rate, we introduce the following assumption:
Assumption 3.6.
The $\mathbb{R}$-valued process on the instantaneous growth of the cash flows of firm $n$ denoted by $\mathrm{d}\log{F^{n}_{t,\mathfrak{d}}}$ is linear in the economic factors (output growth by sector), specifically we set for all $t\in\mathbb{R}_{+}$,
$\mathrm{d}\log{F^{n}_{t,\mathfrak{d}}}=\tilde{\mathfrak{a}}^{n\cdot}\mathrm{d}\log{{Y}_{t}}+\sigma_{n}\mathrm{d}\mathcal{W}^{n}_{t}=\mathfrak{a}^{n\cdot}(\mathrm{d}\mathcal{A}_{t}+\mathrm{d}v(\mathfrak{d}_{t}))+\sigma_{n}\mathrm{d}\mathcal{W}^{n}_{t},$ (3.15)
for $\tilde{\mathfrak{a}}^{n\cdot}\in\mathbb{R}^{I}$ and $\mathfrak{a}^{n\cdot}=\tilde{\mathfrak{a}}^{n\cdot}(\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}$, where $(\mathcal{W}^{n}_{t})_{t\in\mathbb{R}_{+}}$ is a $\mathbb{R}^{N}$-Brownian motion with $\sigma_{n}>0$. Moreover, $B^{\mathcal{Z}}$ and $\mathcal{W}^{n}$ are independent.
We call $\tilde{\mathfrak{a}}^{n\cdot}=(\tilde{\mathfrak{a}}^{ni})_{1\leq i\leq I}$ (respectively $\mathfrak{a}^{n\cdot}=(\mathfrak{a}^{ni})_{1\leq i\leq I}$) factor loadings, quantifying the extent to which $\mathrm{d}\log{F^{n}_{,\mathfrak{d}}}$ is related to $\mathrm{d}\log{{Y}}$ (respectively $\mathrm{d}\mathcal{A}$). We define the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ by $\mathcal{F}_{t}=\sigma\left(\mathcal{G}_{t}\cup\sigma\left\{\mathfrak{b}_{s}:s\in[0,t]\right\}\right)$ for $t\geq 0$, and we denote $\mathbb{E}_{t}[\cdot]:=\mathbb{E}[\cdot|\mathcal{F}_{t}]$.
Recall that the economic motivation behind (3.15) comes from the fact that if firm $n$ belongs to sector $i$, then its production is proportional to the sectoral output and its cash flows are proportional to its production (as in the Dechow-Kothari-Watts model in Barth et al., (2001)). Thus, we obtain a relation between the cash flows of firm $n$ and the total output of sector $i$. The assumption $\tilde{\mathfrak{a}}^{n\cdot}\in\mathbb{R}^{I}$ stems from the fact that a company is not restricted to one sector only in general. However, since we are considering the emission sector here, we expect that each firm $n$ only belongs to one sector ($i$ for example). Therefore $\mathfrak{a}^{nj}=0$ for all $i\neq j$ and hence $|\mathfrak{a}^{ni}|=\max_{j\in\mathcal{I}}|\mathfrak{a}^{nj}|$.
Let $r\geq 0$ representing the interest rate, by the continuous form of the discounted cash flows valuation, the value $V^{n}_{t,\mathfrak{d}}$ of the firm $n$, at time $t$, is
$V^{n}_{t,\mathfrak{d}}:=\mathbb{E}_{t}\left[\int_{t}^{+\infty}e^{-r(s-t)}F^{n}_{s,\mathfrak{d}}\mathrm{d}s\right].$ (3.16)
For $n\in\{1,\ldots,N\}$ and $t\geq 0$, describing $V^{n}_{t,\mathfrak{d}}$ as a function of the underlying processes driving the economy does not lead to an easily tractable formula. But this allows us to write it as a fixed-point problem which can be solved by numerical methods such as deep learning methods (see Hammad et al., (2022); Kobeissi and Bach, (2022)) or Picard iteration (see Berinde and Takens, (2007)). However, to facilitate the forthcoming credit risk analysis, when $\varsigma$ (introduced in 3.1) is close to 00, we approach $V^{n}_{t,\mathfrak{d}}$ by the quantity
$\mathcal{V}_{t,\mathfrak{d}}^{n}:=F^{n}_{t,\mathfrak{d}}\int_{t}^{+\infty}e^{-r(s-t)}\mathbb{E}_{t}\left[\exp\left((s-t)\mathfrak{a}^{n\cdot}\mu+\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)+\sigma_{n}(\mathcal{W}^{n}_{s}-\mathcal{W}^{n}_{t})\right)\mathrm{d}s\right],$ (3.17)
which is defined using the first order expansion of $\frac{V^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}$ and that we describe as a proxy the firm $n$ value at time $t$. We will work directly with $\mathcal{V}^{n}_{t,\mathfrak{d}}$ instead of $V^{n}_{t,\mathfrak{d}}$. We have the following proposition, whose proof is given in B.4.
Proposition 3.7.
For any $n\in\{1,\ldots,N\}$ and for all $t\in\mathbb{R}_{+}$.
  1. 1.
    Assume that $\varrho_{n}:=\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu-r<0$, then $\mathcal{V}_{t,\mathfrak{d}}^{n}$ is well defined and
    $\mathcal{V}^{n}_{t,\mathfrak{d}}=F^{n}_{0}\mathfrak{R}^{n}_{t}(\mathfrak{d})\exp{\left(\mathfrak{a}^{n\cdot}(\mathcal{A}_{t}-v(\mathfrak{d}_{0}))\right)}\exp\left(\sigma_{n}\mathcal{W}^{n}_{t}\right),$ (3.18)
    where
    $\mathfrak{R}^{n}_{t}(\mathfrak{d}):=\int_{0}^{\infty}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s.$ (3.19)
  2. 2.
    Moreover, with $t_{\circ}$ and $t_{\star}$ defined in 3.3, we obtain the following explicit form,
    $\mathfrak{R}^{n}_{t}(\mathfrak{d})=\left\{\begin{array}[]{ll}\displaystyle-\frac{e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})}}{\varrho_{n}},&\mbox{if }t\geq t_{\star},\\ \displaystyle\int_{0}^{t_{\star}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s-\frac{e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})+\varrho_{n}(t_{\star}-t)}}{\varrho_{n}},&\mbox{if }0\leq t<t_{\star},.\end{array}\right.$ (3.20)
  3. 3.
    Assume that
    $\rho_{n}:=\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu+\frac{1}{2}\varsigma^{2}\frac{c_{\Gamma}^{2}}{\lambda_{\Gamma}^{2}}|\mathfrak{a}^{n\cdot}|^{2}|\Sigma|^{2}<r,$ (3.21)
    therefore $V_{t,\mathfrak{d}}^{n}$ is well defined and there exists a constant $C$ such that $\mathbb{E}\left[\left|\frac{V^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}-\frac{\mathcal{V}^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}\right|\right]\leq C\varsigma$, for all $\varsigma>0$.
In $(iii)$ of proposition 3.7, $\mathbb{E}\left[\left|\frac{V^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}-\frac{\mathcal{V}^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}\right|\right]\leq C\varsigma$ means that when the noise term in the productivity process is small, the terms $\frac{\mathcal{V}^{n}_{t}}{F^{n}_{t}}$ and $\frac{V^{n}_{t}}{F^{n}_{t}}$ become closer. The following corollary gives (conditional) laws of the (proxy) of the firm value $\mathcal{V}_{\cdot,\mathfrak{d}}^{n}$.
Corollary 3.8.
For all $t,T\geq 0$.
  1. 1.
    We note $\mathfrak{m}^{n}(\mathfrak{d},t,\mathcal{A}_{t}^{\circ}):=\log{(F^{n}_{0}\mathfrak{R}^{n}_{t}(\mathfrak{d}))}+\left(\mathfrak{a}^{n\cdot}(\mathcal{A}_{t}^{\circ}-v(\mathfrak{d}_{0}))\right)$ and we have
    $\log{\mathcal{V}_{t,\mathfrak{d}}^{n}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\mathfrak{m}^{n}(\mathfrak{d},t,\mathcal{A}_{t}^{\circ}),t\sigma_{n}^{2}\right).$ (3.22)
  2. 2.
    We note $\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t}):=\log{(F^{n}_{0}\mathfrak{R}^{n}_{t+T}(\mathfrak{d}))}+\mathfrak{a}^{n\cdot}(\mu T+\varsigma\Upsilon_{T}\mathcal{Z}_{t}+\mathcal{A}_{t}^{\circ}-v(\mathfrak{d}_{0}))$ and $\mathcal{L}^{n}(t,T):=\mathfrak{a}^{n\cdot}\Sigma^{\mathcal{A},T}_{t}\mathfrak{a}^{n\cdot}+(t+T)\sigma_{n}^{2}$, we have
    $\log{\mathcal{V}_{t+T,\mathfrak{d}}^{n}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t}),\mathcal{L}^{n}(t,T)\right).$ (3.23)
3.3 A Credit Risk Model without collateral
In this section, we aim to calculate the probability of default of the previous firm (i.e. operating in an economy that undergoes the climate transition described by both the carbon price and carbon intensities).
As Basel Committee on Banking Supervision, (2017), we introduce four credit risk parameters: the probability of default (PD) measures the default risk associated with each borrower, the exposure at default (EAD) measures the outstanding debt at the time of default, the loss given default (LGD) denotes the expected percentage of EAD that is lost if the debtor defaults, and the effective maturity $T$ represents the duration of the credit. With these four parameters, we can compute the portfolio loss $L$, with some assumptions:
Assumption 3.9.
Consider a portfolio of $N\in\mathbb{N}^{*}$ credits. For $1\leq n\leq N$,
  1. (1)
    Firm $n$ has issued two classes of securities: equity and debt.
  2. (2)
    $(\mathrm{EAD}_{t}^{n})_{t\geq 0}$ is a $\mathbb{R}_{+}^{*}$-valued continuous and deterministic process, and for all $t\geq 0$, the family $(\mathrm{EAD}_{t}^{n})_{n=1,\ldots,N}$ is a sequence of positive constants such that
    1. (a)
      $\sum_{n\geq 1}\mathrm{EAD}_{t}^{n}=+\infty$;
    2. (b)
      there exists $\upsilon>0$ such that $\frac{\mathrm{EAD}_{t}^{n}}{\sum_{n=1}^{N}\mathrm{EAD}_{t}^{n}}=\mathcal{O}(N^{-(\frac{1}{2}+\upsilon)})$, as $N$ tends to infinity.
  3. (3)
    $(\mathrm{LGD}_{t}^{n})_{t\geq 0}$ is a $(0,1]$-valued continuous and deterministic process;
  4. (4)
    $(\mathcal{D}^{n}_{t})_{t\geq 0}$ is a $\mathbb{R}_{+}$-valued continuous and deterministic process, representing the debt of firm $n$ at time $t$. We will also denote $D^{n}_{t}:=\displaystyle\frac{\mathcal{D}_{t}^{n}}{\mathbb{E}[F_{t,0}^{n}]}$ representing the debt to cash flows ratio.
  5. (5)
    The value of the firm $n$ at time $t$ is assumed to be a tradable asset given by $V^{n}_{t,\mathfrak{d}}$ defined in (3.18).
According to Kruschwitz and Löffler, (2020), there are two ways to handle the default of a company: for a given financing policy, a levered firm is
  • 1.
    in danger of illiquidity if the cash flows do not suffice to fulfill the creditors’ payment claims (interest and net redemption) as contracted,
  • 2.
    over-indebted if the market value of debt exceeds the firm’s market value.
We follow in the present work the second definition of default proposed: a firm default when it is over-indebted, that is in fact the same approach used in the structural credit risk models. We retain this definition in this continuous setting in the same way as Bouveret et al., (2025). Therefore, the over-indebtedness of entity $n$ occurs at time $t$ when the firm value $\mathcal{V}^{n}_{t,\mathfrak{d}}$ falls below a given barrier $\mathcal{D}_{t}^{n}$, related to the net debt, namely on the event $\left\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\right\}$.
However, it should be noted that in a continuous time setting, it can be interesting to work in the Black and Cox, (1976) model. Here, the default event depends on the trajectory of the firm value process $\mathcal{V}$. Therefore, at a given time $t$, the firm defaults if it has been over-indebted at least one time during the period $[0,t]$, that is $\left\{\exists s\in[0,t]\text{ such that }\mathcal{V}^{n}_{s,\mathfrak{d}}<\mathcal{D}^{n}_{s}\right\}$. Thus, the default time is given by $\tau^{n}:=\inf\left\{t\geq 0,\quad\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\right\}$.
Then, if we are interested in the probability of the firm $n$ defaulting before $t$ conditionally to $\mathcal{G}_{t}$ that is noted $PD^{n}_{t,\mathfrak{d}}$, we have
$PD^{n}_{t,T,\mathfrak{d}}=\mathbb{P}\left(\tau^{n}\leq t|\mathcal{G}_{t}\right)=\mathbb{P}\left(\inf_{0\leq s\leq t}\mathcal{V}^{n}_{s,\mathfrak{d}}<\mathcal{D}^{n}_{s}\middle|\mathcal{G}_{t}\right)=\mathbb{P}\left(\inf_{0\leq s\leq t}\log{\mathcal{V}^{n}_{s,\mathfrak{d}}}<\log{\mathcal{D}^{n}_{s}}\middle|\mathcal{G}_{t}\right).$
But for $0\leq s\leq t$ and from (3.18),
$\log{\mathcal{V}_{s,\mathfrak{d}}^{n}}=\log{\left(F_{0,\mathfrak{d}}^{n}\mathfrak{R}^{n}_{s}(\mathfrak{d})\right)}+\mathfrak{a}^{n\cdot}(\mathcal{A}_{s}^{\circ}-v(\mathfrak{d}_{0}))+\sigma_{n}\mathcal{W}^{n}_{s},$
therefore $\log{\mathcal{V}^{n}_{\cdot,\mathfrak{d}}}$ is a Gaussian process. However, as Azaıs and Wschebor, (2000) summarizes, the computation of the distribution function of the random variable $\inf_{0\leq s\leq t}\log{\mathcal{V}^{n}_{s,\mathfrak{d}}}$ is by means of a closed formula is known only for a very restricted number of stochastic processes as the Brownian Motion, the Brownian Bridge, the Brownian Motion with a linear drift, and the stationary Gaussian processes with relatively simple with covariance. This is not the case here. Consequently, we extend Merton, (1974) model instead of Black and Cox, (1976) model.
At each time $t\geq 0$, we are interested in the probability that firm $n$ is over-indebted at a certain date $t+T$, we note $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ and we have
$\mathrm{PD}^{n}_{t,T,\mathfrak{d}}:=\mathbb{P}\left(\mathcal{V}^{n}_{t+T,\mathfrak{d}}\leq\mathcal{D}^{n}_{t+T}\middle|\mathcal{G}_{t}\right).$ (3.24)
We have the following proposition whose proof is a direct application of Corollary 3.8.
Proposition 3.10 (Probability of default).
For $t\geq 0$, $T\geq 0$, and $n\in\{1,\ldots,N\}$, the (conditional) probability of default of the entity $n$ at time $t$ over the horizon $T$ is
$\mathrm{PD}^{n}_{t,T,\mathfrak{d}}=\Phi\left(\frac{\log(\mathcal{D}_{t+T}^{n})-\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})}{\sqrt{\mathcal{L}^{n}(t,T)}}\right),$ (3.25)
where $\mathcal{K}^{n}(\mathfrak{d},t,T,a,\theta)$ as well as $\mathcal{L}^{n}(t,T)$ are defined in Corollary 3.8 and $\Phi$ the Gaussian cumulative distribution function.
The previous results tell us that the probability of the over-indebtedness depends on parameters specific to:
  1. 1.
    the climate transition
    • (a)
      the carbon price $\delta$,
    • (b)
      the carbon intensities $\tau,\zeta,\kappa$,
  2. 2.
    to the company (the contract),
    • (a)
      the factors loading $\mathfrak{a}^{n\cdot}$ and the standard deviation of the cash flows $\sigma_{n}$,
    • (b)
      the time $t$ when it is computed,
    • (c)
      the potential date of the over-indebtedness $t+T$,
    • (d)
      the over-indebtedness’s barrier $\mathcal{D}$,
  3. 3.
    the economy to which the company belongs to:
    • (a)
      the productivity $\mathcal{Z}$ and $\mathcal{A}$ (and their parameters) of the economy,
    • (b)
      the interest rate $r$.
Remark 3.11.
Motivated by the uncertainty of the climate transition, a natural extension is to introduce a stochastic carbon price (see, for example, Le Guenedal and Tankov, (2025)). Regardless of whether it is correlated with the productivity process, the expressions of the macroeconomic variables (for example, the output in proposition 3.5) remain unchanged; however, their dynamics do differ. Consequently, primarily due to the form of $v$ defined in (3.13) – whose law is a product and quotient of Gaussian random variables –, we could no longer obtain a closed-form expression for the firm value – because filtration $\mathbb{G}$ would not only be adapted to productivity $\Theta$, but also to the carbon price $\delta$ – and then for the probability of default. Numerical methods are therefore required.
In the next section, we will define and express LGD as a function of some guarantees which are affected by the climate transition. Therefore, by assuming that EAD are deterministic and independent of the carbon price, we will also introduce the expected and unexpected losses.
4 LGD with stochastic collaterals in continuous time
We are in the same framework as in the previous section, but item 3 of 3.9 is not satisfied anymore, therefore the bank could require from each counterpart $1\leq n\leq N$ a (single) collateral $C^{n}$ to secure its debt. Collateral can take the form of a physical asset such as real estate, business equipment, or inventory, or it can be a financial asset such as invoices, cash, or investments. If a firm is over-indebted at time $t$, we assume that the liquidation ends at $t+a$ with $a\in\mathbb{R}_{+}$ where $a$ is the liquidation delay. Moreover, $k\in[0,1)$ represents the fraction of the collateral used to cover liquidations auctions, as well as other legal and administrative procedures.
A firm is over-indebted at time $t\geq 0$ if the market value of its debt $\mathcal{D}^{n}_{t}$ exceed its market value $\mathcal{V}^{n}_{t,\mathfrak{d}}$, namely $\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\}$. At time $t$, if the company $n$ in the portfolio defaults i.e. $\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}~{}$, the bank recovers $(1-k)e^{-ra}C^{n}_{t+a}$ after the collateral liquidation. In general, the liquidations do not cover all the debt, i.e. $\mathrm{EAD}_{t}^{n}\geq(1-k)e^{-ra}C^{n}_{t+a}$, the bank deploys further actions to recover an additional fraction. We note that fraction $\gamma\in[0,1)$. Therefore, the bank recovers $\gamma(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})_{+}$ by other tools.
The potential loss that would be recorded due to the firm default event is the difference between the debt amount $\mathrm{EAD}_{t}^{n}$ and the amount received after the recovery processes on the time horizon. Consequently, if there is not default ($\mathcal{V}^{n}_{t,\mathfrak{d}}\geq\mathcal{D}^{n}_{t}$) or there is default and the collateral liquidated exceed the exposure ($(1-k)e^{-ra}C^{n}_{t+a}\geq\mathrm{EAD}_{t}^{n}$), the loss is zero, and if there is default and if the exposure exceed the collateral liquidated, the loss is
$\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a}-\gamma(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})=(1-\gamma)\left(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a}\right),$ (4.1)
where the constant $r$ is the discount rate. The loss, $L^{N}_{t}$, of the portfolio at time $t$, is in fact, defined as
$L^{N}_{t}:=\sum_{n=1}^{N}(1-\gamma)(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\}}.$ (4.2)
The following result is similar with the one introduced in Bouveret et al., (2025)[Theorem 3.5]. It gives a proxy of the loss of the portfolio.
Theorem 4.1 (Definition of LGD).
For all $t\in\mathbb{R}_{+}$, define
$\mathrm{L}^{\mathbb{G},N}_{t}:=\mathbb{E}\left[L^{N}_{t}\middle|\mathcal{G}_{t}\right]=\sum_{n=1}^{N}\mathrm{EAD}_{t}^{n}\cdot\mathrm{LGD}_{t,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t,\mathfrak{d}}^{n},$ (4.3)
where
$\displaystyle\mathrm{PD}_{t,\mathfrak{d}}^{n}$ $:=\mathbb{P}\left(\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}|\mathcal{G}_{t}\right)=\Phi\left(\frac{\log(\mathcal{D}_{t}^{n})-\mathfrak{m}^{n}(\mathfrak{d},t,\mathcal{A}_{t})}{\sigma^{n}\sqrt{t}}\right),$ (4.4a)
$\displaystyle\mathrm{LGD}_{t,\mathfrak{d}}^{n}$ $:=(1-\gamma)\mathbb{E}\left[\left(1-(1-k)e^{-ra}\frac{C^{n}_{t+a}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right].$ (4.4b)
Under Assumptions 3.9, we have $L^{N}_{t}-\mathrm{L}^{\mathbb{G},N}_{t}$ converges to zero almost surely as $N$ tends to infinity, for all $t\in\mathbb{R}_{+}$.
Proof.
Let $t\in\mathbb{R}_{+}$,
$\begin{split}\mathrm{L}^{\mathbb{G},N}_{t}&=\mathbb{E}\left[L^{N}_{t}\middle|\mathcal{G}_{t}\right]=\mathbb{E}\left[\sum_{n=1}^{N}(1-\gamma)(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\}}\middle|\mathcal{G}_{t}\right].\end{split}$
$(\mathrm{EAD}_{t}^{n})_{n\in\{1,\ldots,N\}}$ is deterministic, we have:
$\begin{split}\mathrm{L}^{\mathbb{G},N}_{t}&=\sum_{n=1}^{N}\mathrm{EAD}_{t}^{n}\cdot\mathbb{E}\left[(1-\gamma)\left(1-(1-k)e^{-ra}\frac{C^{n}_{t+a}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\}}\middle|\mathcal{G}_{t}\right]\\ &=\sum_{n=1}^{N}\mathrm{EAD}_{t}^{n}\cdot\mathbb{E}\left[(1-\gamma)\left(1-(1-k)e^{-ra}\frac{C^{n}_{t+a}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right]\cdot\mathbb{P}\left[\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\middle|\mathcal{G}_{t}\right].\\ \end{split}$
The rest of the proof requires a version of the strong law of large numbers (Appendix of (Gordy, , 2003, Propositions 1, 2)), where the systematic risk factor is $\mathcal{G}_{t}$. ∎
Explicitly, in the above theorem, we assume that our portfolio is perfectly fine grained, so that we can approximate $L_{t}^{N}$ – the portfolio loss – by $L_{t}^{\mathbb{G},N}$ – the conditional expectation of loss given the systemic factor. By construction, the loss given default noted $\mathrm{LGD}$ is the percentage of the total exposure that the bank loses when a over-indebtedness occurs. The literature on LGD modeling is fairly extensive. We can distinguish namely, economic modeling Bastos, (2010); Roncalli, (2020) and stochastic modeling Roncalli, (2020)[Page 193], Chalupka and Kopecsni, (2008). As the definition of $\mathrm{PD}$ does not change compare to what we did in section 3 and as $\mathrm{EAD}$ is given, we will focus on $\mathrm{LGD}$ modeling. We can first remark that $0\leq\mathrm{LGD}_{t,\mathfrak{d}}^{n}\leq 1-\gamma$, then the presence of a collateral necessarily reduces $\mathrm{LGD}$.
Other key quantities for the bank to understand the (dynamics of the) risk in the portfolio are the (expected and unexpected) losses and probability of default conditionally to the (information generated by the) risk factors. Precisely, for a date $t$ and a horizon $T$, a bank computes some risk measures at $t$ of its portfolio maturing at horizon $T$.
Definition 4.2 (Projected losses).
Let $t\geq 0$ be the time when the risk measure is computed for a period $T\geq 0$. As classically done, the potential loss is separated into three components:
  • 1.
    The (conditional) expected loss (EL) – the amount that a bank expects to lose on a credit exposure – reads
    $\mathrm{EL}^{N,T}_{t}:=\mathbb{E}\left[\mathrm{L}^{\mathbb{G},N}_{t+T}\middle|\mathcal{G}_{t}\right].$ (4.5)
  • 2.
    The unexpected loss (UL) – the amount by which potential credit losses might exceed the EL – reads for $\alpha\in(0,1)$,
    $\mathrm{UL}^{N,T}_{t}\!(\alpha):=\mathrm{VaR}^{\alpha,N,T}_{t}-\mathrm{EL}^{N,T}_{t},\quad\text{where}\quad 1-\alpha=\mathbb{P}\left[L_{t+T}^{\mathbb{G},N}\leq\mathrm{VaR}^{\alpha,N,T}_{t}\middle|\mathcal{G}_{t}\right],$ (4.6)
    and the scalar $\mathrm{VaR}^{\alpha,N,T}_{t}$ called value-at-risk.
  • 3.
    The stressed loss (or expected shortfall or ES) – the amount by which potential credit losses might exceed the capital requirement – reads:
    $\mathrm{ES}^{N,T}_{t}(\alpha):=\mathbb{E}\left[L^{N}_{t+T}\middle|L^{N}_{t+T}\geq\mathrm{VaR}^{\alpha,N,T}_{t},\mathcal{G}_{t}\right],\qquad\text{for }\alpha\in(0,1).$ (4.7)
Knowing PD, LGD, and EAD is enough to determine the losses. We assume EAD is given (item 2 in 3.9). PD is given by (3.25) in section 3.3. We will now determine LGD when a collateral exists. We focus on two types: a financial asset and a property in housing market. We introduce the two following lemmas which will help later on to set up explicit formulas for LGD. Let $n\in\{1,\ldots,N\}$.
Lemma 4.3.
Assume that a stochastic process $K^{n}$ satisfies for all $t\in\mathbb{R}_{+}$,
  1. 1.
    $\log{K^{n}}_{t}|\mathcal{G}_{t}\sim\mathcal{N}(m_{t}^{n},(\sigma^{n}_{t})^{2})$ with $m^{n}_{t}\in\mathbb{R}$ and $\sigma^{n}_{t}>0$,
  2. 2.
    and $K^{n}_{t}|\mathcal{G}_{t}$ and $\mathcal{V}^{n}_{t,\mathfrak{d}}|\mathcal{G}_{t}$ are independent.
Therefore, for $t,u\in\mathbb{R}_{+}$,
$\mathbb{E}\left[\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right]=u\Phi\left(\frac{w_{t}^{n}}{\sigma^{n}_{t}}\right)-\exp{\left(-w_{t}^{n}+\frac{1}{2}(\sigma^{n}_{t})^{2}\right)}\Phi\left(\frac{w_{t}^{n}}{\sigma^{n}_{t}}-\sigma^{n}_{t}\right),$ (4.8)
where
$w_{t}^{n}:=\log{\left(u\frac{\mathrm{EAD}_{t}^{n}}{1-k}\right)}-m_{t}^{n}.$ (4.9)
Proof.
Let $t,u\in\mathbb{R}_{+}$, we have
$\mathbb{E}\left[\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right]$
$\qquad=\mathbb{E}\left[\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{G}_{t}\right]\quad\text{because}\quad K^{n}_{t}|\mathcal{G}_{t}\text{ and }\mathcal{V}_{t,\mathfrak{d}}^{n}|\mathcal{G}_{t}\text{ are independent}$
$\qquad=\mathbb{E}\left[\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\right)\mathbf{1}_{\{u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\geq 0\}}\middle|\mathcal{G}_{t}\right]$
$\qquad=u\mathbb{P}\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\geq 0\middle|\mathcal{G}_{t}\right)-\frac{(1-k)}{\mathrm{EAD}_{t}^{n}}\mathbb{E}\left[K^{n}_{t}\mathbf{1}_{\{u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\geq 0\}}\middle|\mathcal{G}_{t}\right].$
However, $\log{K^{n}_{t}}|\mathcal{G}_{t}\sim\mathcal{N}(m_{t}^{n},(\sigma^{n}_{t})^{2})$. We also consider $w_{t}^{n}$ defined in (4.9), therefore
$\mathbb{P}\left(u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\geq 0\middle|\mathcal{G}_{t}\right)=\Phi\left(\frac{w_{t}^{n}}{\sigma^{n}_{t}}\right).$
We also have
$\mathbb{E}\left[K^{n}_{t}\mathbf{1}_{\{u-(1-k)\frac{K^{n}_{t}}{\mathrm{EAD}_{t}^{n}}\geq 0\}}\middle|\mathcal{G}_{t}\right]=\exp{\left(-w_{t}^{n}+\frac{1}{2}(\sigma^{n}_{t})^{2}\right)}\Phi\left(\frac{w_{t}^{n}}{\sigma^{n}_{t}}-\sigma^{n}_{t}\right).$
The conclusion follows. ∎
Lemma 4.4.
Assume that a stochastic process $K^{n}$ satisfies, for each $t,T\in\mathbb{R}_{+}$,
  1. 1.
    $\log{K^{n}}_{t+T}|\mathcal{G}_{t}\sim\mathcal{N}(m_{t,T}^{n},(\sigma^{n}_{t,T})^{2})$ with $m^{n}_{t,T}\in\mathbb{R}$ and $\sigma^{n}_{t,T}>0$,
  2. 2.
    and $\begin{bmatrix}\log{\mathcal{V}_{t+T}^{n}}\\ \log{K_{t+T+a}^{n}}\end{bmatrix}|\mathcal{G}_{t}\sim\mathcal{N}\left(\begin{bmatrix}\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})\\ \overline{\mathcal{K}}^{n}_{t,T+a}\end{bmatrix},\begin{bmatrix}\mathcal{L}^{n}(t,T)&cv_{t,T,a}^{n}\\ cv_{t,T,a}^{n}&\overline{\mathcal{L}}_{t,T+a}^{n}\end{bmatrix}\right)$, where ¯$\overline{\mathcal{L}}_{t,T+a}^{n}>0$, $cv_{t,T,a}^{n},\overline{\mathcal{K}}^{n}_{t,T+a}\in\mathbb{R}$, and $\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})$ as well as $\mathcal{L}^{n}(t,T)$ are defined in Corollary 3.8.
Therefore, for $t,T,u\in\mathbb{R}_{+}$, we have
$\begin{split}&\mathbb{E}\left[\left(1-(1-k)e^{-ra}\frac{\mathcal{C}^{n}_{t+T+a,\mathfrak{d}}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right]\\ &\qquad=u\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}});\rho_{t,T,a}^{n}\right)-\exp{\left(\frac{1}{2}\overline{\mathcal{L}}_{t,T+a}^{n}-\sqrt{\overline{\mathcal{L}}_{t,T+a}^{n}}\overline{\omega}_{t,T,a}^{n}\right)}\times\\ &\qquad\qquad\qquad\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n}-\sqrt{\overline{\mathcal{L}}_{t,T+a}^{n}},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}})-\rho_{t,T,a}^{n}\sqrt{\overline{\mathcal{L}}_{t,T+a}^{n}};\rho_{t,T,a}^{n}\right),\end{split}$ (4.10)
where $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ is defined in (3.25) and where
$\rho_{t,T,a}^{n}:=\frac{cv_{t,T,a}^{n}}{\sqrt{\mathcal{L}^{n}(t,T)\overline{\mathcal{L}}^{n}_{t,T+a}}},\quad\text{and}\quad\overline{\omega}_{t,T,a}^{n}:=\frac{\log{\left(u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\right)}-\overline{\mathcal{K}}^{n}_{\mathfrak{d},t,T+a}}{\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}}}.$
Proof.
Let $t,T,u\in\mathbb{R}_{+}$, we have
$\begin{split}&\mathbb{E}\left[\left(u-(1-k)e^{-ra}\frac{K_{t+T+a}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right]\\ &\qquad=\mathbb{E}\left[\left(u-(1-k)e^{-ra}\frac{K_{t+T+a}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)\cdot\mathbf{1}_{u-(1-k)e^{-ra}\frac{K_{t+T+a}^{n}}{\mathrm{EAD}_{t+T}^{n}}\geq 0}\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right]\\ &\qquad=\mathbb{E}\left[\left(u-(1-k)e^{-ra}\frac{K_{t+T+a}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},\quad K_{t+T+a}^{n}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}}\middle|\mathcal{G}_{t}\right]\\ &\qquad=u\mathbb{P}\left[\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},K_{t+T+a}^{n}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}\middle|\mathcal{G}_{t}\right]\\ &\qquad\qquad-\frac{(1-k)e^{-ra}}{\mathrm{EAD}_{t+T}^{n}}\mathbb{E}\left[K_{t+T+a}^{n}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},K_{t+T+a}^{n}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}}\middle|\mathcal{G}_{t}\right].\end{split}$
However $\begin{bmatrix}\log{\mathcal{V}_{t+T}^{n}}\\ \log{K_{t+T+a}^{n}}\end{bmatrix}|\mathcal{G}_{t}\sim\mathcal{N}\left(\begin{bmatrix}\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})\\ \overline{\mathcal{K}}^{n}_{t,T+a}\end{bmatrix},\begin{bmatrix}\mathcal{L}^{n}(t,T)&cv_{t,T,a}^{n}\\ cv_{t,T,a}^{n}&\overline{\mathcal{L}}_{t,T+a}^{n}\end{bmatrix}\right)$, therefore we have
$\mathbb{P}\left[\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},K^{n}_{t+T+a}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}\middle|\mathcal{G}_{t}\right]$
$\qquad=\Phi_{2}\left(\frac{\log{u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}}-\overline{\mathcal{K}}^{n}_{t,T+a}}{\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}}},\frac{\log{\mathcal{D}^{n}_{t+T}}-\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})}{\sqrt{\mathcal{L}^{n}(t,T)}};\frac{cv_{t,T,a}}{\sqrt{\mathcal{L}^{n}(t,T)\overline{\mathcal{L}}^{n}_{t,T+a}}}\right),$
and
$\begin{split}&\frac{(1-k)e^{-ra}}{\mathrm{EAD}_{t+T}^{n}}\mathbb{E}\left[K^{n}_{t+T+a}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},K^{n}_{t+T+a}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}}\middle|\mathcal{G}_{t}\right]\\ &\qquad=\mathbb{E}\left[e^{\log{K^{n}_{t+T+a}}}\cdot\mathbf{1}_{\left\{\frac{\log{K^{n}_{t+T+a}}-\overline{\mathcal{K}}^{n}_{t,T+a}}{\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}}}\leq\overline{\omega}_{t,T,a}^{n},\frac{\log{\mathcal{V}^{n}_{t+T,\mathfrak{d}}}-\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})}{\sqrt{\mathcal{L}^{n}(t,T)}}<\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}})\right\}}\middle|\mathcal{G}_{t}\right]\end{split}$
However, according to (A.3) in section A.2, $\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]=e^{\frac{1}{2}\sigma^{2}}\Phi_{2}\left(x-\sigma,y-\rho\sigma;\rho\right)$, therefore,
$\frac{(1-k)e^{-ra}}{\mathrm{EAD}_{t+T}^{n}}\mathbb{E}\left[e^{\log{K^{n}_{t+T+a}}}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},K^{n}_{t+T+a}\leq u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\}}\middle|\mathcal{G}_{t}\right]$
$\qquad=\exp{\left(\frac{1}{2}\overline{\mathcal{L}}^{n}_{t,T+a}-\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}}\overline{\omega}_{t,T,a}^{n}\right)}\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n}-\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}})-\frac{cv_{t,T,a}}{\sqrt{\mathcal{L}^{n}(t,T)}};\frac{cv_{t,T,a}}{\sqrt{\mathcal{L}^{n}(t,T)\overline{\mathcal{L}}^{n}_{t,T+a}}}\right);$
Moreover, from proposition 3.10,
$\mathbb{P}\left[\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\middle|\mathcal{G}_{t}\right]=\Phi\left(\frac{\log(\mathcal{D}_{t+T}^{n})-\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})}{\sqrt{\mathcal{L}^{n}(t,T)}}\right).$
This concludes the proof. ∎
4.1 When there is not collateral
When firm $n$ does not have a collateral, therefore $C^{n}=0$ and from (4.4b), LGD is
$\mathrm{LGD}_{t,\mathfrak{d}}^{n}=1-\gamma.$ (4.11)
4.2 When collateral is a financial asset
Here we assume that the collateral of the firm $n$ is an investment in a financial asset. Precisely, we assume that that investment is a proportion $\alpha^{n}\in(0,1]$ of a given firm located in the economy described in section 3. Consequently, it is subjected to the same constraints in terms of productivity and of carbon transition scenarios as firm $n$. As any investment, it should generate a stream of cash flows so that at each time, we can compute its value by using the discounted cash flows model introduced in (3.16).
Let note the collateral cash flows $(\overline{F}_{t}^{n})_{t\in\mathbb{R}_{+}}$, its dynamics is similar to the firm cash flows introduced in 3.6. We have for all $t\in\mathbb{R}_{+}$,
$\mathrm{d}\overline{F}^{n}_{t,\mathfrak{d}}=\mathfrak{\overline{a}}^{n\cdot}((\mu+\varsigma\mathcal{Z}_{t})\mathrm{d}t+\mathrm{d}v(\mathfrak{d}_{t}))+\overline{\sigma}_{n}\mathrm{d}\overline{\mathcal{W}}^{n}_{t},$ (4.12)
where $\mathfrak{\overline{a}}^{n\cdot}\in\mathbb{R}^{I}$ and where $(\overline{\mathcal{W}}_{t})_{t\in\mathbb{R}_{+}}$ is a $\mathbb{R}^{N}$-Brownian motion with $\overline{\sigma}_{n}>0$. Moreover, $B^{\mathcal{Z}}$ (noise of productivity), $\overline{\mathcal{W}}^{n}$ (noise of collateral), and $(\mathcal{W}^{n})_{n\in\{1,\ldots,N\}}$ (noise of debtors) are independent. We also note $\overline{\tilde{\mathfrak{a}}}^{n\cdot}=\overline{\mathfrak{a}}^{n\cdot}(\mathbf{I}_{I}-\boldsymbol{\lambda})$.
Remark 4.5.
We have assumed that $\overline{\mathcal{W}}^{n}$ and $\mathcal{W}^{n}$ are not correlated, but this is not always the case. For example, if the depreciation of the firm value heading to its over-indebtedness implies the depreciation of the collateral value, then we should have a positive correlation.
Inspired by (3.16), the collateral value at time $t$ is
$C^{n}_{t,\mathfrak{d}}:=\alpha^{n}\mathbb{E}_{t}\left[\int_{s=t}^{+\infty}e^{-rs}\overline{F}^{n}_{s,\mathfrak{d}}\mathrm{d}s\right],$
and by (3.17), a proxy collateral value as
$\mathcal{C}_{t,\mathfrak{d}}^{n}:=\alpha^{n}\overline{F}^{n}_{t,\mathfrak{d}}\int_{t}^{+\infty}e^{-r(s-t)}\mathbb{E}_{t}\left[\exp\left((s-t)\overline{\mathfrak{a}}^{n\cdot}\mu+\overline{\mathfrak{a}}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)+\sigma_{n}(\overline{\mathcal{W}}^{n}_{s}-\overline{\mathcal{W}}^{n}_{t})\right)\mathrm{d}s\right].$ (4.13)
Therefore, the following proposition (whose the proof is inspired by Lemma 3.7 and corollary 3.8) gives a proxy of the collateral value.
Proposition 4.6.
For any $n\in\{1,\ldots,N\}$ and for all $t\in\mathbb{R}_{+}$
  1. 1.
    Assume that $\overline{\varrho}_{n}:=\frac{1}{2}\overline{\sigma}_{\mathfrak{b}_{n}}^{2}+\mathfrak{\overline{a}}^{n\cdot}\mu-r<0$. Given the carbon emissions costs sequence $\mathfrak{d}$, the proxy of collateral value defined in (4.13), is well defined and
[…]
    where
[…]
  2. 2.
[…]
    $\log{\mathcal{C}^{n}_{t,\mathfrak{d}}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\mathfrak{\overline{m}}^{n}(\mathfrak{d},t,\mathcal{A}_{t}^{\circ}),t\overline{\sigma}_{n}^{2}\right),$ (4.16)
    and we note $\overline{\mathcal{K}}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t}):=\log{(\alpha^{n}\overline{F}^{n}_{0}\overline{\mathfrak{R}}^{n}_{t+T}(\mathfrak{d}))}+\overline{\mathfrak{a}}^{n\cdot}(\mu T+\varsigma\Upsilon_{T}\mathcal{Z}_{t}+\mathcal{A}_{t}^{\circ}-v(\mathfrak{d}_{0}))$ and ¯$\overline{\mathcal{L}}^{n}(t,T):=\overline{\mathfrak{a}}^{n\cdot}\Sigma^{\mathcal{A}}_{t,t+T}\overline{\mathfrak{a}}^{n\cdot}+(t+T)\overline{\sigma}_{n}^{2}$, and we have
    $\log{\mathcal{C}_{t+T}^{n}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\overline{\mathcal{K}}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t}),\overline{\mathcal{L}}^{n}(t,T)\right).$ (4.17)
  3. 3.
    Assume that
    $\overline{\rho}_{n}:=\frac{1}{2}\overline{\sigma}_{n}^{2}+\overline{\mathfrak{a}}^{n\cdot}\mu+\frac{1}{2}\varsigma^{2}\frac{c_{\Gamma}^{2}}{\lambda_{\Gamma}^{2}}\lVert\overline{\mathfrak{a}}^{n\cdot}\rVert^{2}\lVert\Sigma\rVert^{2}<r,$ (4.18)
    therefore $C_{t,\mathfrak{d}}^{n}$ is well defined and there exists a constant $\overline{C}$ such that $\mathbb{E}\left[\left|\frac{C^{n}_{t,\mathfrak{d}}}{\overline{F}^{n}_{t,\mathfrak{d}}}-\frac{\mathcal{C}^{n}_{t,\mathfrak{d}}}{\overline{F}^{n}_{t,\mathfrak{d}}}\right|\right]\leq\overline{C}\varsigma$, for all $\varsigma>0$.
Proof.
Let $n\in\{1,\ldots,N\}$ and $t\in\mathbb{R}_{+}$, (4.14) directly comes from proposition 3.7. The proofs of the three points are equivalent to B.4. Let us develop the conditional laws. From (4.14), we have
$\log{\mathcal{C}^{n}_{t}}=\log{\alpha^{n}\overline{F}^{n}_{0}\overline{\mathfrak{R}}^{n}_{t}(\mathfrak{d})}+\overline{\mathfrak{a}}^{n\cdot}(\mathcal{A}_{t}^{\circ}-v(\mathfrak{d}_{0}))+\overline{\sigma}_{n}\overline{\mathcal{W}}^{n}_{t}.$
Because $\overline{\mathcal{W}}^{n}$ is a Brownian motion, $\overline{\mathcal{W}}^{n}_{t}\sim\mathcal{N}\left(0,t\right)$ and, $\overline{\mathcal{W}}^{n}$ and $B^{\mathcal{Z}}$ are independent, we obtain $\log{\mathcal{C}_{t}^{n}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\overline{\mathfrak{m}}^{n}(\mathfrak{d},t,\mathcal{A}_{t}^{\circ}),t\overline{\sigma}_{n}^{2}\right)$. Let also $T\in\mathbb{R}_{+}$, we have
$\begin{split}\log{\mathcal{C}^{n}_{t+T}}&=\log{\overline{F}^{n}_{0}\overline{\mathfrak{R}}^{n}_{t+T}(\mathfrak{d})}+\overline{\mathfrak{a}}^{n\cdot}(\mathcal{A}_{t+T}^{\circ}-v(\mathfrak{d}_{0}))+\overline{\mathcal{W}}^{n}_{t+T}.\end{split}$
From remark 3.2, $\mathcal{A}_{t+T}|\mathcal{G}_{t}\sim\mathcal{N}\left(M^{\mathcal{A},T}_{t},\Sigma^{\mathcal{A},T}_{t}\right)$ and because $\overline{\mathcal{W}}^{n}$ is a Brownian motion, $\overline{\mathcal{W}}^{n}_{t+T}\sim\mathcal{N}\left(0,t+T\right)$. Moreover, $\overline{\mathcal{W}}^{n}$ and $B^{\mathcal{Z}}$ are independent. We have
$\log{\mathcal{C}^{n}_{t+T}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\log{\alpha^{n}\overline{F}^{n}_{0}\overline{\mathfrak{R}}^{n}_{t+T}(\mathfrak{d})}+\overline{\mathfrak{a}}^{n\cdot}(M^{\mathcal{A},T}_{t}-v(\mathfrak{d}_{0})),\overline{\mathfrak{a}}^{n\cdot}\Sigma^{\mathcal{A},T}_{t}\overline{\mathfrak{a}}^{n\cdot}+(t+T)\overline{\sigma}_{n}^{2}\right).$
The conclusion follows. ∎
From the (proxy of the) collateral value $\mathcal{C}$, we can then derive a precised expression of $\mathrm{LGD}$ based on theorem 4.1. We have:
Theorem 4.7.
When $a=0$ (no liquidation delay), the Loss Given Default of the obligor $n$ over-indebted at time $t\in\mathbb{R}_{+}$, conditional on $\mathcal{G}_{t}$ is
$\mathrm{LGD}_{t,\mathfrak{d}}^{n}=(1-\gamma)\left[\Phi\left(\frac{w_{t}^{n}}{\overline{\sigma}_{\mathfrak{b}_{n}}\sqrt{t}}\right)-\exp{\left(-w_{t}^{n}+\frac{1}{2}t\overline{\sigma}_{n}^{2}\right)}\Phi\left(\frac{w_{t}^{n}}{\overline{\sigma}_{n}\sqrt{t}}-\overline{\sigma}_{n}\sqrt{t}\right)\right],$ (4.19)
where
$w_{t}^{n}:=\log{\left(\frac{\mathrm{EAD}_{t}^{n}}{1-k}\right)}-\mathfrak{\overline{m}}^{n}(\mathfrak{d},t,\mathcal{A}_{t}).$ (4.20)
Proof.
Let $t\in\mathbb{R}_{+}$. By remarking that
  1. 1.
    $\log{\mathcal{C}^{n}_{t,\mathfrak{d}}}|\mathcal{G}_{t}\sim\mathcal{N}\left(\mathfrak{\overline{m}}^{n}(\mathfrak{d},t,\mathcal{A}_{t}),t\overline{\sigma}_{n}^{2}\right)$,
  2. 2.
    and $\mathcal{C}_{t,\mathfrak{d}}^{n}|\mathcal{G}_{t}$ and $\mathcal{V}_{t,\mathfrak{d}}^{n}|\mathcal{G}_{t}$ are independent,
we can simply apply Lemma 4.3 with $u=1$. ∎
We can also remark that the situation where there is no collateral corresponds to $\overline{F}^{n}_{0}=0$. We then have
$\begin{split}\overline{F}^{n}_{0}\to 0&\implies\log{(\overline{F}^{n}_{0})}\to-\infty\implies\mathfrak{\overline{m}}^{n}(\mathfrak{d},t,\mathcal{A}_{t}^{\circ})\to-\infty\implies w_{t}^{n}\to+\infty\implies\mathrm{LGD}_{t,\mathfrak{d}}^{n}\to 1-\gamma.\end{split}$
For each $t,T\in\mathbb{R}_{+}$, we introduce now the (conditional) $\mathrm{LGD}$ of the entity $n$ at time $t$ on the horizon $T$, namely
$\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}:=(1-\gamma)\mathbb{E}\left[\left(1-(1-k)e^{-ra}\frac{\mathcal{C}^{n}_{t+T+a,\mathfrak{d}}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T},\mathcal{G}_{t}\right].$
It is precisely about calculating at date $t$ the proportion of the exposure that the bank would lose if the counterpart $n$ is over-indebted at date $t+T$.
Proposition 4.8 (Projected PD and LGD).
For each $t,T\in\mathbb{R}_{+}$, the (conditional) $\mathrm{LGD}$ of the entity $n$ at time $t$ on the horizon $T$, reads
$\begin{split}\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}&=\frac{1-\gamma}{\mathrm{PD}^{n}_{t,T,\mathfrak{d}}}\left[\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}});\rho_{t,T,a}^{n}\right)-\exp{\left(\frac{1}{2}\overline{\mathcal{L}}^{n}(t,T+a)-\sqrt{\overline{\mathcal{L}}^{n}(t,T+a)}\overline{\omega}_{t,T,a}^{n}\right)}\times\right.\\ &\qquad\qquad\qquad\left.\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n}-\sqrt{\overline{\mathcal{L}}^{n}(t,T+a)},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}})-\rho_{t,T,a}^{n}\sqrt{\overline{\mathcal{L}}^{n}(t,T+a)};\rho_{t,T,a}^{n}\right)\right],\end{split}$ (4.21)
where
$\rho_{t,T,a}^{n}:=\frac{\varsigma^{2}\mathfrak{a}^{n\cdot}\Gamma^{-1}\left(\int_{0}^{T}\left(e^{-\Gamma u}-\mathbf{I}_{I}\right)\Sigma\Sigma^{\top}\left(e^{-\Gamma(u+a)}-\mathbf{I}_{I}\right)\mathrm{d}u\right)(\overline{\mathfrak{a}}^{n\cdot}\Gamma^{-1})^{\top}}{\sqrt{\mathcal{L}^{n}(t,T)\overline{\mathcal{L}}^{n}(t,T+a)}},$
and
$\overline{\omega}_{t,T,a}^{n}:=\frac{\log{\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}}-\overline{\mathcal{K}}^{n}(\mathfrak{d},t,T+a,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})}{\sqrt{\overline{\mathcal{L}}^{n}(t,T+a)}},$
and where $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ defined in proposition 3.10.
Proof.
Let $t,T\in\mathbb{R}_{+}$, from (4.3) and (4.5),
$\mathrm{EL}^{N,T}_{t}:=\mathbb{E}\left[\mathrm{L}^{\mathbb{G},N}_{t+T}\middle|\mathcal{G}_{t}\right]=\mathbb{E}\left[\sum_{n=1}^{N}L_{n,t+T}^{\mathbb{G}}\middle|\mathcal{G}_{t}\right]=\sum_{n=1}^{N}\mathbb{E}\left[L_{n,t+T}^{\mathbb{G}}\middle|\mathcal{G}_{t}\right].$
But for $1\leq n\leq N$, we have
$\begin{split}\mathbb{E}\left[L_{n,t+T}^{\mathbb{G}}\middle|\mathcal{G}_{t}\right]&=\mathbb{E}\left[\mathrm{EAD}_{t+T}^{n}\mathrm{LGD}^{n}_{t+T,\mathfrak{d}}\mathrm{PD}^{n}_{t+T,\mathfrak{d}}\middle|\mathcal{G}_{t}\right]\\ &=\mathrm{EAD}_{t+T}^{n}\mathbb{E}\left[\mathrm{LGD}^{n}_{t+T,\mathfrak{d}}\mathrm{PD}_{t+T,\mathfrak{d}}^{n}\middle|\mathcal{G}_{t}\right]\quad\text{as }\mathrm{EAD}_{t+T}^{n}\text{ is deterministic}\\ &=\mathrm{EAD}_{t+T}^{n}\mathbb{E}\left[\mathbb{E}\left[(1-\gamma)\left(1-(1-k)e^{-ra}\frac{\mathcal{C}^{n}_{t+T+a,\mathfrak{d}}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t+T}\right]\middle|\mathcal{G}_{t}\right]\\ &=(1-\gamma)\mathrm{EAD}_{t+T}^{n}\mathbb{E}\left[\left(1-(1-k)e^{-ra}\frac{\mathcal{C}^{n}_{t+T+a,\mathfrak{d}}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right].\end{split}$
[…]
$cov(\log{F_{t+T}^{n}},\log{\mathcal{C}_{t+T+a}^{n}}|\mathcal{G}_{t})=\mathbb{E}[\log{F_{t+T}^{n}}\log{\mathcal{C}_{t+T+a}^{n}}|\mathcal{G}_{t}]-\mathbb{E}[\log{F_{t+T}^{n}}|\mathcal{G}_{t}]\mathbb{E}[\log{\mathcal{C}_{t+T+a}^{n}}|\mathcal{G}_{t}].$
However,
$\mathbb{E}[\log{F_{t+T}^{n}}\log{\mathcal{C}_{t+T+a}^{n}}|\mathcal{G}_{t}]$
$\qquad=\mathbb{E}\left[\left(\log{\left(F^{n}_{0}\mathfrak{R}^{n}_{t+T}(\mathfrak{d})e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{0})}\right)}+\mathfrak{a}^{n\cdot}\mathcal{A}_{t+T}^{\circ}+\sigma_{n}\mathcal{W}^{n}_{t+T}\right)\right.$
[…]
$\qquad\qquad\left.+\log{\left(F^{n}_{0}\mathfrak{R}^{n}_{t+T}(\mathfrak{d})e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{0})}\right)}(\overline{\mathfrak{a}}^{n\cdot}\mathcal{A}_{t+T+a}^{\circ}+\overline{\sigma}_{n}\overline{\mathcal{W}}_{t+T+a}^{n})+\overline{\sigma}_{n}\overline{\mathcal{W}}_{t+T+a}^{n}\sigma_{n}\mathcal{W}^{n}_{t+T}\right.$
[…]
$\qquad\qquad+\log{\left(F^{n}_{0}\mathfrak{R}^{n}_{t+T}(\mathfrak{d})e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{0})}\right)}\overline{\mathfrak{a}}^{n\cdot}\mathbb{E}[\mathcal{A}_{t+T+a}^{\circ}|\mathcal{G}_{t}]$
[…]
By also developing $\mathbb{E}[\log{F_{t+T}^{n}}|\mathcal{G}_{t}]\mathbb{E}[\log{\mathcal{C}_{t+T}^{n}}|\mathcal{G}_{t}]$, we obtain
$\begin{split}&cov(\log{F_{t+T}^{n}},\log{\mathcal{C}_{t+T}^{n}}|\mathcal{G}_{t})=cov(\mathfrak{a}^{n\cdot}\mathcal{A}_{t+T}^{\circ},\overline{\mathfrak{a}}^{n\cdot}\mathcal{A}_{t+T}^{\circ}|\mathcal{G}_{t})\\ &\qquad=\varsigma^{2}\mathfrak{a}^{n\cdot}\Gamma^{-1}\left(\int_{0}^{T}\left(e^{-\Gamma u}-\mathbf{I}_{I}\right)\Sigma\Sigma^{\top}\left(e^{-\Gamma(u+a)}-\mathbf{I}_{I}\right)\mathrm{d}u\right)(\overline{\mathfrak{a}}^{n\cdot}\Gamma^{-1})^{\top}:=cv_{t,T,a}.\end{split}$
We obtain
$\begin{bmatrix}\log{\mathcal{V}_{t+T}^{n}}\\ \log{\mathcal{C}_{t+T}^{n}}\end{bmatrix}|\mathcal{G}_{t}\sim\mathcal{N}\left(\begin{bmatrix}\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})\\ \overline{\mathcal{K}}^{n}(\mathfrak{d},t,T+a,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})\end{bmatrix},\begin{bmatrix}\mathcal{L}^{n}(t,T)&cv_{t,T,a}\\ cv_{t,T,a}&\overline{\mathcal{L}}^{n}(t,T+a)\end{bmatrix}\right).$ (4.22)
Then, we use the Lemma 4.4 with $u=1$ to conclude the proof. ∎
We remark that the carbon price introduced in our economy affect both PD through the obligor cash flows and LGD through the collateral cash flows. See more remarks in section 4.5.
4.3 When collateral is commercial or residential property
In this section, we assume that loans are backed by either residential or commercial building. The problem here is then to model the real estate market in the presence of the climate transition risk. The latter is represented by energy efficiency as well as the carbon price. We would like to compute the value of a dwelling at time time $t$. We use exactly as in Sopgoui, (2024) the actualized sum of the cash flows before the renovation date (taking into account the additional energy costs due to inefficiency of the building), at the renovation date, and after the renovation date (when the building becomes efficient). Moreover, the agent chooses rationally the date of renovation which maximizes the value of his property. Therefore, according to Sopgoui, (2024)[Theorem 2.4], we have the following proposition.
Theorem 4.9.
Assume that the following conditions are satisfied:
  1. 1.
    the carbon price function $\delta:t\mapsto\delta_{t}$ is non decreasing on $\mathbb{R}_{+}$ and deterministic;
  2. 2.
    the energy price $\mathfrak{f}(\cdot,\mathfrak{p})$ is non decreasing on $\mathbb{R}_{+}$ for each type of energy $\mathfrak{p}$.
Then, the market value of the building serving as the collateral to firm $n$ at $t\geq 0$, given the carbon price sequence $\delta$, is given by
$\mathcal{C}^{n}_{t,\delta}=C^{n}_{t}-R_{n}X_{t,\delta}^{n},$ (4.23)
where
$X_{t,\delta}^{n}:=\mathfrak{c}(\alpha^{n},\alpha^{\star})e^{-\bar{r}(\mathfrak{t}_{n}-t)}+(\alpha^{n}-\alpha^{\star})\int_{t}^{\mathfrak{t}_{n}}\mathfrak{f}(\delta_{u},\mathfrak{p})e^{-\bar{r}(u-t)}\mathrm{d}u,$ (4.24)
and where the optimal date of renovations $\mathfrak{t}_{n}\in[t,+\infty]$ is given by
$\mathfrak{t}_{n}=$ $t$$ if $\mathfrak{f}(\delta_{\theta},\mathfrak{p})-\bar{r}\frac{\mathfrak{c}(\alpha^{n},\alpha^{\star})}{\alpha^{n}-\alpha^{\star}}>0$ for all $\theta\in[t,\infty)$ (4.25)
$\mathfrak{t}_{n}=$ +$\infty$+ ∞ if $\mathfrak{f}(\delta_{\theta},\mathfrak{p})-\bar{r}\frac{\mathfrak{c}(\alpha^{n},\alpha^{\star})}{\alpha^{n}-\alpha^{\star}}<0$ for all $\theta\in[t,\infty)$ (4.26)
$\mathfrak{t}_{n}=$ $\theta^{\star}$ the unique solution of $\mathfrak{f}(\delta_{\theta},\mathfrak{p})=\bar{r}\frac{\mathfrak{c}(\alpha^{n},\alpha^{\star})}{\alpha^{n}-\alpha^{\star}}$ on $\theta\in[t,\infty)$. (4.27)
Moreover,
$C^{n}_{t}$ $:=R_{n}C^{n}_{0}e^{K_{t}},$ (4.28)
where
$\displaystyle\mathrm{d}K_{t}$ $=\left(\dot{\chi}_{t}+\nu(\chi_{t}-K_{t})\right)\mathrm{d}t+\overline{\sigma}\mathrm{d}\overline{B}_{t},$ (4.29a)
$\mathrm{d}\overline{B}_{t}$ $=\rho^{\top}\mathrm{d}B^{\mathcal{Z}}_{t}+\sqrt{1-\lVert\rho\rVert^{2}}\mathrm{d}\overline{\mathcal{W}}_{t},$ (4.29b)
with $(\overline{\mathcal{W}}_{t})_{t\in\mathbb{R}_{+}}$ is a standard Brownian motion independent to $B^{\mathcal{Z}}$ introduced in 3.1 and driving the productivity of the economy. Moreover, $C^{n}_{0}$, $r$, $R_{n},\overline{\sigma}>0$, $\rho\in\mathbb{R}_{+}^{I}$, and $\chi\in\mathfrak{D}(\mathbb{R}_{+},\mathbb{R}_{+})$. We introduce the following filtration $\mathbb{U}:=(\mathcal{U}_{t})_{t\in\mathbb{R}_{+}}$ with respect to $t\geq 0$, $\mathcal{U}_{t}:=\sigma\left(\left\{\overline{W}_{s},B^{\mathcal{Z}}_{s}:s\leq t\right\}\right)$.
The following corollary gives the conditional distribution of the collateral. Its proof is straightforward and is detailed in Sopgoui, (2024).
Corollary 4.10.
For $0\leq t\leq T$, the law of $C^{n}_{t+T}=R_{n}C_{0}^{n}\exp{(K_{t+T})}$ conditional on $\mathcal{G}_{t}$ is log-Normal $\mathcal{LN}(m_{t,T}^{n},v_{t,T}^{n})$ with
$m_{t,T}^{n}:=\log{(R_{n}C_{0}^{n})}+\chi_{t+T}-\left(\chi_{0}-K_{0}\right)e^{-\nu(t+T)}+\overline{\sigma}\rho^{\top}\int_{0}^{t}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s},$ (4.30)
and
$v_{t,T}^{n}:=\frac{(\overline{\sigma}\lVert\rho\rVert)^{2})}{2\nu}\left(1-e^{-2\nu T}\right)+\frac{(\overline{\sigma})^{2}(1-\lVert\rho\rVert^{2})}{2\nu}\left(1-e^{-2\nu(t+T)}\right).$ (4.31)
An example of the energy price function
We can assume that the price of each type of energy $\mathfrak{p}$ is a linear function of the carbon price, therefore
$\mathfrak{f}:(\delta_{t},\mathfrak{p})\mapsto\mathfrak{f}^{\mathfrak{p}}_{1}\delta_{t}+\mathfrak{f}^{\mathfrak{p}}_{0}\qquad t\geq 0,$ (4.32)
with $\mathfrak{f}^{\mathfrak{p}}_{1},\mathfrak{f}^{\mathfrak{p}}_{0}>0$ and $\delta$ is the carbon price defined in the 3.3 or an example given in (3.9).
An example of the renovation costs function
We can consider that the costs of renovation of a dwelling $\mathfrak{c}$, to move its energy efficiency from $x$ to $y$, is
$\mathfrak{c}:(x,y)\mapsto c_{0}|x-y|^{1+c_{1}},$ (4.33)
with $c_{0}>0$ and $c_{1}\geq-1$. This choice of $\mathfrak{c}$ allows us to model that when a building has a bad energy efficiency, its renovation is costly. This is similar to that of Sopgoui, (2024) and is inspired on the data given in Caumont et al., (2016), which summarizes the results of the survey on energy renovation of single-family homes in France.
An example of the optimal renovation time
With the example of the carbon price in (3.9), the example of the energy price in (4.32), and the example of the renovation costs in (4.33), the optimal renovation time, solution of (4.27) is given by
$\mathfrak{t}_{n}=t_{\circ}+\frac{1}{\eta_{\delta}}\log{\left(\frac{c_{0}r|\alpha^{n}-\alpha^{\star}|^{c_{1}}-\mathfrak{f}^{\mathfrak{p}}_{0}}{\mathfrak{f}^{\mathfrak{p}}_{1}P_{carbon}}\right)}.$ (4.34)
We can clearly remark that the optimal renovation date depends on the climate transition policy ($P_{carbon}$ and $\eta_{\delta}$), on the energy prices ($\mathfrak{f}^{\mathfrak{p}}_{0}$ and $\mathfrak{f}^{\mathfrak{p}}_{1}$), on the renovation costs ($c_{0}$ and $c_{1}$), and on the energy efficiencies ($\alpha^{n}$ and $\alpha^{\star}$).
By using the housing price under the climate transition as given in theorem 4.9, we can then derive a precised expression of $\mathrm{LGD}$ when the collateral exists and is a building. We have:
Theorem 4.11.
Let $1\leq n\leq N$. When $a=0$ (no liquidation delay), the Loss Given Default of the obligor $n$ is over-indebted at time $t\in\mathbb{R}_{+}$, conditional on $\mathcal{G}_{t}$, is
$\mathrm{LGD}_{t,\delta}^{n}=(1-\gamma)\left[\left(1+(1-k)\frac{R_{n}X_{t,\delta}^{n}}{\mathrm{EAD}_{t}^{n}}\right)\Phi\left(\frac{w_{t}^{n}}{\sqrt{v_{t,0}^{n}}}\right)-\exp{\left(-w_{t}^{n}+\frac{1}{2}v_{t,t}^{n}\right)}\Phi\left(\frac{w_{t}^{n}}{\sqrt{v_{t,0}^{n}}}-\sqrt{v_{t,0}^{n}}\right)\right],$ (4.35)
where
$w_{t}^{n}:=\log{\left(\frac{\mathrm{EAD}_{t}^{n}}{(1-k)R_{n}}+X_{t,\delta}^{n}\right)}-m_{t,0}^{n},$ (4.36)
and with $m_{t,0}^{n}$ and $v_{t,t}^{n}$ defined in Corollary 4.10, and $X_{t,\delta}^{n}$ defined in (4.24).
We can also verify that when there is not collateral corresponding to $C^{n}_{0}=0$. We then have
$\begin{split}C^{n}_{0}\to 0&\implies\log{(R_{n}C^{n}_{0})}\to-\infty\implies m_{t,0}^{n}\to-\infty\implies w_{t}^{n}\to+\infty\implies\mathrm{LGD}_{t,\delta}^{n}\to 1-\gamma.\end{split}$
It is even worse when the costs associated with the transition explode, LGD also explodes as
$X_{t,\delta}^{n}\to+\infty\implies w_{t}^{n}\to+\infty\implies\mathrm{LGD}_{t,\delta}^{n}\to+\infty.$
Proof.
Let $t\geq 0$ and $1\leq n\leq N$. By remarking that
  1. 1.
    $\log{C^{n}_{t,\mathfrak{d}}}|\mathcal{G}_{t}\sim\mathcal{N}\left(m_{t,0}^{n},v_{t,0}^{n}\right)$, and
  2. 2.
    $C_{t,\mathfrak{d}}^{n}|\mathcal{G}_{t}$ and $\mathcal{V}_{t,\mathfrak{d}}^{n}|\mathcal{G}_{t}$ are independent,
  3. 3.
    and from (4.4b) when $a=0$, we have
    $\mathrm{LGD}_{t,\delta}^{n}$ $=(1-\gamma)\mathbb{E}\left[\left(1-(1-k)\frac{\mathcal{C}^{n}_{t,\delta}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right]$
    $=(1-\gamma)\mathbb{E}\left[\left(1+(1-k)\frac{R_{n}X^{n}_{t,\delta}}{\mathrm{EAD}_{t}^{n}}-(1-k)\frac{C^{n}_{t,\delta}}{\mathrm{EAD}_{t}^{n}}\right)_{+}\middle|\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t},\mathcal{G}_{t}\right],$
we can simply apply Lemma 4.3 with $u=1+(1-k)\frac{R_{n}X^{n}_{t,\delta}}{\mathrm{EAD}_{t}^{n}}$. ∎
Once again, we want to compute the (conditional) Loss Given Default of the entity $n$ at time $t$ on the horizon $T$. We can formalize that in the following proposition:
Proposition 4.12 (Projected LGD).
For each $t,T\geq 0$ and $1\leq n\leq N$, the (conditional) Loss Given Default of the entity $n$ at time $t$ on the horizon $T$, reads
$\begin{split}\mathrm{LGD}^{n}_{t,T,\delta}&=\frac{1-\gamma}{\mathrm{PD}^{n}_{t,T,\mathfrak{d}}}\left[\left(1+(1-k)e^{-ra}\frac{R_{n}X_{t+T+a,\delta}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}});\rho_{t,T,a}^{n}\right)\right.\\ &\qquad\qquad\quad\left.-\exp{\left(\frac{1}{2}v_{t,T+a}^{n}-\sqrt{v_{t,T+a}^{n}}\overline{\omega}_{t,T,a}^{n}\right)}\Phi_{2}\left(\overline{\omega}_{t,T,a}^{n}-\sqrt{v_{t,T+a}^{n}},\Phi^{-1}(\mathrm{PD}^{n}_{t,T,\mathfrak{d}})-\rho_{t,T,a}^{n}\sqrt{v_{t,T+a}^{n}};\rho_{t,T,a}^{n}\right)\right],\end{split}$ (4.37)
where
$\rho_{t,T,a}^{n}:=\overline{\sigma}\varsigma\frac{\mathfrak{a}^{n\cdot}\Gamma^{-1}\left(\int_{0}^{T}e^{-\nu(u+a)}\left(\mathbf{I}_{I}-e^{-\Gamma u}\right)\Sigma\mathrm{d}u\right)\rho}{\sqrt{\mathcal{L}^{n}(t,T)\times v_{t,T+a}^{n}}},$
and
$\overline{w}_{t,T,a}^{n}:=\frac{\log{\left(\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)R_{n}e^{-ra}}+X_{t+T+a,\delta}^{n}\right)}-m_{t,T+a}^{n}}{\sqrt{v_{t,T+a}^{n}}},$
and with $m_{t,t+T+a}^{n}$ and $v_{t,t+T+a}^{n}$ defined in Corollary 4.10 and $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ defined in proposition 3.10.
Proof.
Let $t,T\geq 0$ and $1\leq n\leq N$. From the beginning of the proof of proposition 4.8, we have
$\mathbb{E}\left[L_{n,t+T}^{\mathbb{G}}\middle|\mathcal{G}_{t}\right]$ $=(1-\gamma)\mathrm{EAD}_{t+T}^{n}\mathbb{E}\left[\left(1-(1-k)e^{-ra}\frac{\mathcal{C}^{n}_{t+T+a,\mathfrak{d}}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right],$
then
$\mathbb{E}\left[L_{n,t+T}^{\mathbb{G}}\middle|\mathcal{G}_{t}\right]$ $=(1-\gamma)\mathrm{EAD}_{t+T}^{n}\mathbb{E}\left[\left(1+(1-k)e^{-ra}\frac{R_{n}X_{t+T+a,\delta}^{n}}{\mathrm{EAD}_{t+T}^{n}}-(1-k)e^{-ra}\frac{\mathcal{C}_{t+T+a,\delta}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t+T,\mathfrak{d}}<\mathcal{D}^{n}_{t+T}\}}\middle|\mathcal{G}_{t}\right].$
Remark that
$\begin{split}\log{C_{0}^{n}e^{K_{t+T+a}}}&=\log{(C_{0}^{n})}+\chi_{t+T+a}-(\chi_{0}-K_{0})e^{-\nu(t+T+a)}\\ &\qquad+\overline{\sigma}\rho^{\top}\int_{0}^{t+T+a}e^{-\nu(t+T+a-s)}\mathrm{d}B_{s}^{\mathcal{Z}}+\overline{\sigma}\sqrt{1-\lVert\rho\rVert^{2}}\int_{0}^{t+T+a}e^{-\nu(t+T+a-s)}\mathrm{d}\overline{\mathcal{W}}_{s},\end{split}$
and recall that $\log{(\mathcal{V}^{n}_{t+T,\mathfrak{d}})}=\log{(F^{n}_{0}\mathfrak{R}^{n}_{t+T}(\mathfrak{d}))}+\mathfrak{a}^{n\cdot}(\mathcal{A}_{t+T}^{\circ}-v(\mathfrak{d}_{0}))+\sigma_{n}\mathcal{W}^{n}_{t+T}$. Therefore,
$\begin{split}cov\left(\log{C_{0}^{n}e^{K_{t+T+a}}},\log{F^{n}_{t+T}}\middle|\mathcal{G}_{t}\right)&=cov\left(\mathfrak{a}^{n\cdot}\mathcal{A}_{t+T}^{\circ},\overline{\sigma}\rho^{\top}\int_{0}^{t+T+a}e^{-\nu(t+T+a-s)}\mathrm{d}B_{s}^{\mathcal{Z}}\middle|\mathcal{G}_{t}\right)\\ &=\overline{\sigma}\mathfrak{a}^{n\cdot}cov\left(\mathcal{A}_{t+T}^{\circ},\int_{0}^{t+T+a}e^{-\nu(t+T+a-s)}\mathrm{d}B_{s}^{\mathcal{Z}}\middle|\mathcal{G}_{t}\right)\rho\\ &=\overline{\sigma}\varsigma\mathfrak{a}^{n\cdot}\Gamma^{-1}\left(\int_{0}^{T}e^{-\nu(u+a)}\left(\mathbf{I}_{I}-e^{-\Gamma u}\right)\Sigma\mathrm{d}u\right)\rho:=cv_{t,T,a}.\end{split}$
Consequently, we can write
$\begin{bmatrix}\log{F_{t+T}^{n}}\\ \log{C_{0}^{n}e^{K_{t+T+a}}}\end{bmatrix}|\mathcal{G}_{t}\sim\mathcal{N}\left(\begin{bmatrix}\mathcal{K}^{n}(\mathfrak{d},t,T,\mathcal{A}_{t}^{\circ},\mathcal{Z}_{t})\\ m_{t,T+a}^{n}\end{bmatrix},\begin{bmatrix}\mathcal{L}^{n}(t,T)&cv_{t,T,a}\\ cv_{t,T,a}&v_{t,T+a}^{n}\end{bmatrix}\right).$ (4.38)
Then, we use the Lemma 4.4 with $u=1+(1-k)e^{-ra}\frac{R_{n}X_{t+T+a,\delta}^{n}}{\mathrm{EAD}_{t+T}^{n}}$ to conclude the proof. ∎
We can remark that $\mathrm{LGD}^{n}_{t,\delta}$ as well as $\mathrm{LGD}^{n}_{t,T,\delta}$ are also functions of the optimal renovation time $\mathfrak{t}_{n}$. Furthermore, if both the financial asset and the housing price are affected by the climate transition through their dependence on the carbon price sequence $\delta$, the financial asset depends also on the carbon price intensities (of firms production/consumption and of households consumption) $(\tau,\zeta,\kappa)$ which are not specific to a given company but to the economy as a whole. The housing price is clearly affected by specific climate factors, namely the energy efficiency $\alpha^{n}$ and the renovation date $\mathfrak{t}_{n}$.
4.4 Expected and Unexpected losses
Let us recall that we have a portfolio with N loans. We assume that loans from 1 to $N_{1}$ are unsecured, loans from $N_{1}+1$ to $N_{2}$ are secured by a financial asset as collateral, and loans from $N_{2}+1$ to $N$ are secured by a commercial or residential property as collateral.
We write the expression of the portfolio EL and UL as functions of the parameters and of the processes introduced above, and introduce the entity’s probability of default.
We can therefore give expressions of EL and UL. Let $t,T\geq 0$, the (conditional) Expected Loss of the portfolio at time $t$ on the horizon $T$ defined in (4.5), reads
$\begin{split}\mathrm{EL}^{N,T}_{t}&=\mathbb{E}\left[\mathrm{L}^{\mathbb{G},N}_{t+T}\middle|\mathcal{G}_{t}\right]=\sum_{n=1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}\\ &=\sum_{n=1}^{N_{1}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}+\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}\\ &\qquad\qquad+\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\delta}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}.\end{split}$ (4.39)
We can then compute each term conditionally to $\mathcal{G}_{t}$.
  1. 1.
    Given that $\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}=1-\gamma$ for $1\leq n\leq N_{1}$, to compute $\sum_{n=1}^{N_{1}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$, all you have to do is calculate $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$.
  2. 2.
    Given that $N_{1}+1\leq n\leq N_{2}$, the collaterals are financial assets, therefore, to compute $\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$, we compute first $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$. Then we compute $\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}$ directly through (4.21).
  3. 3.
    Given that $N_{2}+1\leq n\leq N$, the collaterals are properties, therefore, to compute $\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\delta}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$, we compute first $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$. Then we compute $\mathrm{LGD}^{n}_{t,T,\delta}$ directly through (4.37).
For $\alpha\in(0,1)$, the (conditional) Unexpected Loss of the portfolio at time $t$ on the horizon $T$, cannot be obtained in closed-form as EL. Precisely, there is not a closed-form expression neither of $\mathrm{UL}^{\alpha,N,T}_{t}$ nor of $\mathrm{VaR}^{\alpha,N,T}_{t}$. But we can describe how to compute $\mathrm{VaR}^{\alpha,N,T}_{t}$ given that $\mathbb{P}\left[L_{t+T}^{\mathbb{G},N}\leq\mathrm{VaR}^{\alpha,N,T}_{t}\middle|\mathcal{G}_{t}\right]$ as introduced in (4.6). First, let us note that from theorem 4.1, we have
$\begin{split}L_{t+T}^{\mathbb{G},N}&=\sum_{n=1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}\\ &=\sum_{n=1}^{N_{1}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}+\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}\\ &\qquad\qquad+\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\delta}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}.\end{split}$
We can then describe each term’s law conditionally to $\mathcal{G}_{t}$.
  1. 1.
    $\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}=1-\gamma$ and from (3.24), we have $\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ which depends on $\mathcal{A}_{t+T}$. Then to simulate law of $\sum_{n=1}^{N_{1}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ conditional on $\mathcal{G}_{t}$, just simulate $\mathcal{A}_{t+T}|\mathcal{G}_{t}$.
  2. 2.
    From (4.19), we have $\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}$ which depends on $\mathcal{A}_{t+T}$ through $w_{t+T}^{n}$ defined in (4.20). We said in the previous item that $\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ depends on $\mathcal{A}_{t+T}$. Therefore, to simulate law of $\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ conditional on $\mathcal{G}_{t}$, just simulate $\mathcal{A}_{t+T}|\mathcal{G}_{t}$.
  3. 3.
    From (4.35), we have $\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}$ which depends on $\int_{0}^{t+T}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s}$ through $w_{t+T}^{n}$ defined in (4.36). We said in the previous item that $\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ depends on $\mathcal{A}_{t+T}$. Therefore, to simulate law of $\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}_{t+T,\delta}^{n}\cdot\mathrm{PD}_{t+T,\mathfrak{d}}^{n}$ conditional on $\mathcal{G}_{t}$, just simulate $\mathcal{A}_{t+T}|\mathcal{G}_{t}$ and $\int_{0}^{t+T}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s}|\mathcal{G}_{t}$ (which are in fact the same because both $\mathcal{A}_{t+T}$ and $\int_{0}^{t+T}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s}$ are $\mathcal{G}_{t+T}$-measurable).
4.5 Remarks on the determinants of LGD
The results  (4.19) and (4.37) tell us that, in the case the collateral is an investment, Loss Given Default depends on:
  1. 1.
    the carbon price $\delta$ for both (4.19) and (4.37),
  2. 2.
    parameters specific to the company (the contract),
    • (a)
      the time $t$ when it is computed,
    • (b)
      the date of default $t+T$,
    • (c)
      the Exposure at Default $EAD$,
  3. 3.
    parameters specific to the collateral,
    • (a)
      its liquidation time $t+T+a$,
    • (b)
      the liquidation costs $k$,
    • (c)
      the correlation of its cash flows with the environment $\overline{\mathfrak{a}}$,
    • (d)
      the standard deviation of its cash flows $\overline{\sigma}_{\mathfrak{b}}$,
    • (e)
      the fraction of recovery from other means $\gamma$,
  4. 4.
    the nature of the collateral:
    • (a)
      if it is a financial asset, then parameters related to the carbon intensities $\tau,\zeta,\kappa$,
    • (b)
      if it is a building, then parameters related to the energy efficiency $\alpha$, type of energy $\mathfrak{p}$, and renovation costs $\mathfrak{c}$,
  5. 5.
    parameters specific to the economy to which the colateral belongs to:
    • (a)
      the (cumulative) productivity $\mathcal{A}$ (and its parameters) of the economy,
    • (b)
      the interest rates $r$ and $\bar{r}$.
Some of these typical risk drivers are reported by Chalupka and Kopecsni, (2008).
We could also look at the sensitivities of the LGD to each of these variables and parameters. However, the expressions of LGD we obtained are not very tractable so that it would be difficult to get detailed expressions of theses sensitivities. If necessary, they can be calculated using numerical methods (see Caprioli et al., (2023); Fermanian and Scaillet, (2005)).
5 Numerical experiments, estimation and calibration
In this section, we describe how the parameters of multisectoral model, of the firm valuation model, and of the credit risk model are estimated given the historical macroeconomic variables (consumption, labour, output, GHG emissions, housing prices, etc.) as well as the historical credit portfolio data (firms rated and defaulted, collateral, etc.) In a second step, we give the expression of the risk measures (PD, LGD, EL, and UL) introduced in the previous sections, that we compute using Monte Carlo simulations.
5.1 Calibration and estimation
We will calibrate the model parameters on a set of data ranging from time $\mathfrak{t}_{0}$ to $\mathfrak{t}_{1}$. In practice, $\mathfrak{t}_{0}=1978$ and $\mathfrak{t}_{1}=t_{\circ}=2021$. From now on, we will discretize the observation interval into $M\in\mathbb{N}^{*}$ steps $t_{m}=\mathfrak{t}_{0}+\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}m$ for $0\leq m\leq M$. We note $\mathcal{T}^{M}:=\{t_{0},t_{1},\ldots,t_{M}\}$. We will not be interested in convergence results here. For the sake of clarity, we will omit the dependence of each estimated parameter/function on $\mathfrak{t}_{0}$ and $\mathfrak{t}_{1}$. We then note $\widehat{x}$ the estimation of $x$.
5.1.1 Estimation of carbon intensities
For each sector $i\in\mathcal{I}$ and for $0\leq m\leq M$, we observe the output $Y_{t_{m}}^{i}$, the labor $N_{t_{m}}^{i}$, the intermediary input $(Z_{t_{m}}^{ji})_{j\in\mathcal{I}}$, and the consumption $C_{t_{m}}^{i}$ (recall that the transition starts at year $t_{\circ}$). For the sake of clarity, we will omit the dependence of each estimated parameter on $M$.
To calibrate each carbon intensity $\mathfrak{y}\in\{\tau^{1},\ldots,\tau^{I},\zeta^{11},\zeta^{12},\ldots,\zeta^{II-1},\zeta^{II},\kappa^{1},\ldots,\kappa^{I}\}$, we follow exactly the same process already presented in Bouveret et al., (2025). The main difference is that after calibration, we can compute $\mathfrak{y}$ for each $t\in\mathbb{R}_{+}$. Afterwards, if we consider the example of the carbon price introduces in (3.9), we can compute the emissions cost rate $\widehat{\mathfrak{d}}_{t}$.
5.1.2 Estimation of economic parameters
As in Galí, (2015), we assume a unitary Frisch elasticity of labor supply so $\varphi=1$ and the utility of consumption is logarithmic so $\sigma=1$, while we calibrate $({\boldsymbol{\lambda}}^{ij})_{i,j\in\mathcal{I}}$ and  $({\psi}^{i})_{i\in\mathcal{I}}$ in the same way as in Bouveret et al., (2025). We can then compute the functions $\Psi$ and $\Lambda$ defined in proposition 3.5, followed by the function $\widehat{v}^{i}$ as defined in (3.13). We can also compute the output growth $\left(\Delta^{{Y}}_{t_{m}}=(\log({Y}_{t_{m}}^{i})-\log({Y}_{t_{m-1}}^{i}))_{j\in\mathcal{I}}\right)_{1\leq m\leq M}$ directly from data.
Without carbon tax in any sector, it follows from (3.14) in Corollary 3.5 that, for each $1\leq m\leq M$, the computed consumption growth $\Delta^{{Y}}_{t_{m}}$ is equal to $\Delta^{{Y}}_{t_{m}}=\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}(\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}})^{-1}\widehat{\Theta}_{t_{m}}$ when $\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}}$ is not singular; hence $\widehat{\Theta}_{t_{m}}=\frac{M}{\mathfrak{t}_{1}-\mathfrak{t}_{0}}(\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}})\Delta^{{Y}}_{t_{m}}$. We can then compute the estimations $\widehat{\mu}$, $\widehat{\Gamma}$, $\widehat{\Sigma}$, and $\widehat{\varsigma}$, of parameters $\mu$, $\Gamma$, $\Sigma$, and $\varsigma$ (all defined in 3.1), as detailed in Sopgoui, (2024)[Section 3.1.1.].
5.1.3 Estimation of firm and of the credit model parameters
Recall that we have a portfolio with $N\in\mathbb{N}^{*}$ firms (or credit) at time $t_{\circ}$. For each firm $n\in\{1,\ldots,N\}$, we have its historical cash flows $(F_{t_{m}}^{n})_{1\leq m\leq M}$, hence its log-cash flow growths. For any $t\in\mathcal{T}^{M}$ and $1\leq i\leq I$, we denote by $r_{t}^{i}$ (resp. $d_{t}^{i}$) the number of firms in $g_{i}$ rated at the beginning of the year $t$ (resp. defaulted during the year $t$). In particular, $r_{\mathfrak{t}_{0}}=\#g_{i}$. Within each group $g_{i}$, all the firms behave in the same way as there is only one risk class. Since each sub-portfolio constitutes a single risk class, we have for each $n\in g_{m}$, $\mathfrak{a}^{n}=\mathfrak{a}^{n_{i}}$, $\sigma_{\mathfrak{b}^{n}}=\sigma_{\mathfrak{b}^{n_{i}}}$, and $B^{n}=B^{n_{i}}$. We then proceed as follows:
  1. 1.
    Knowing the output growth $\left(\Delta^{{Y}}_{t}\right)_{t\in\mathcal{T}^{M}}$, we calibrate the factor loading $\mathfrak{a}_{n_{i}}$ and the standard deviation $\sigma_{n_{i}}$, according to 3.6, appealing to the regression. For all $1\leq m\leq M$,
    $\sum_{n\in g_{i}}\log{F_{t_{m}}^{n}}-\log{F_{t_{m-1}}^{n}}=(\#g_{i})\mathfrak{a}^{n_{i}}\Delta^{{Y}}_{t_{k}}+\sqrt{\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}\#g_{i}}\sigma_{\mathfrak{b}^{n_{i}}}\mathfrak{u}_{t_{m}}\quad\text{where}\quad\mathfrak{u}_{t_{m}}\sim\mathcal{N}(0,1).$ (5.1)
  2. 2.
    We then obtain the barrier $B^{n_{i}}$ by maximum likelihood estimation (MLE) as detailed in (Gordy and Heitfield, , 2002, Section 3). We have
    $\widehat{B}^{n_{i}}:=\operatorname*{arg\,max}_{B^{n_{i}}\in\mathbb{R}^{+}}\mathcal{L}(B^{n_{i}}),$
    where $\mathcal{L}(B^{n_{i}})$ is the log-likelihood function defined by
[…]
    and where
    $\mathbb{P}[D^{n_{i}}=d_{t_{m}}^{i}|(\mathcal{A}_{t_{m}}^{\circ},\mathcal{Z}_{t_{m}})]={\binom{r_{t_{m}}^{i}}{d_{t_{m}}^{i}}}(\mathrm{PD}^{n_{i}}_{t_{m},1,0})^{d_{t_{m}}^{i}}\Big{(}1-\mathrm{PD}^{n_{i}}_{t_{m},1,0}\Big{)}^{r_{t_{m}}^{i}-d_{t_{m}}^{i}},$
    with $D^{n_{i}}$ the Binomial random variable standing for the conditional number of defaults, and $\mathrm{PD}^{n_{i}}_{t_{m},1,0}$ in eq. 3.25, depending on $\sigma_{\mathfrak{b}^{n_{i}}}=\widehat{\sigma}_{\mathfrak{b}^{n_{i}}}$, $\mathfrak{a}^{n_{i}}=\widehat{\mathfrak{a}}^{n_{i}}$, for $1\leq m\leq M$, $\delta_{t_{m}}=0$ and on $B^{n_{i}}$.
5.1.4 Calibration of collateral
Recall that we have a portfolio with $N\in\mathbb{N}^{*}$ firms (or credit) at time $t_{\circ}$. For each firm $n\in\{1,\ldots,N\}$, if the collateral is
A financial asset
We have its historical cash flows $(\overline{F}_{t_{m}}^{n})_{0\leq m\leq M}$, hence its log-cash flow growths. Recall that, even if two firms belong to the same sub-portfolio, there is no reason that their collaterals behave in the same way. We also know the output growth $\left(\Delta^{{Y}}_{t_{m}}\right)_{1\leq m\leq M}$. We then have,
  1. 1.
    the proportion $\alpha^{n}$ of the investment representing the collateral is known.
  2. 2.
    We calibrate the factor loading $\widehat{\overline{\mathfrak{a}}}_{n}$ and the standard deviation $\widehat{\overline{\sigma}}_{n}$, according to (4.12), appealing to the regression. For all $1\leq m\leq M$
    $\log{\overline{F}_{t_{m}}^{n}}-\log{\overline{F}_{t_{m-1}}^{n}}={\overline{\mathfrak{a}}}^{n}\Delta^{{Y}}_{t_{m}}+\sqrt{\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}}\overline{\sigma}_{n}\mathfrak{u}_{t_{m}}\quad\text{where}\quad\mathfrak{u}_{t_{m}}\sim\mathcal{N}(0,1).$ (5.2)
A commercial or residential property
We assume that in the past, carbon price did not have impact on the dwelling price so that for all $t\in\mathcal{T}^{M}$, $X_{t,\delta}^{n}$ defined in (4.24) is zero. Moreover, $C^{n}_{0}$ and $R_{n}$ defined in (4.28), the value of the collateral at 00 and the surface, are known. All that remains is to calibrate the parameters of the process $K$ defined in (4.29a) and (4.29b). Let us consider a real estate index $(REI_{t_{m}})_{0\leq m\leq M}$. We assume that the long-term average of the real estate index $\chi$, introduced in (4.29a) is linear and for $t\in\mathbb{R}_{+}$, $\chi_{t}=\varrho t+\vartheta$. We then estimate $\varrho$, $\vartheta$, $\nu$, $\overline{\sigma}$, and $\rho$ as Sopgoui, (2024)[Section 3.1.2].
5.2 Simulations
In this section as well, the idea here is not to (re)demonstrate or improve convergence results. Recall the climate transition should start at $t_{\circ}$ and end at $t_{\star}$. Therefore, we would like to compute the risk measures for $t\in[t_{\circ},t_{\star}]$.
5.2.1 Of the productivities $\mathcal{Z}$ and $\mathcal{A}$
Let $K~{}\in\mathbb{N}^{*}$, for $0\leq k\leq K$, we note $u_{k}=t_{\circ}+\frac{t_{\star}-t_{\circ}}{K}k$ for $0\leq k\leq K$. We would like to simulate $\mathcal{Z}_{u_{k}}$ and $\mathcal{A}_{u_{k}}$. For $\mathcal{Z}$, we adopt the Euler-Maruyana (see Maruyama, (1955); Kanagawa, (1988)) scheme as in Sopgoui, (2024)[Section 3.2.1].
5.2.2 Of the probability of over-indebtedness PD and of LGD
For $n\in\{1,\ldots,N\}$ and $t_{\circ}\leq t\leq t_{\star}$, We would like to compute $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ as defined in (3.25) as well as $\mathrm{LGD}_{t,T,\mathfrak{d}}^{n}$ defined in (4.21) and $\mathrm{LGD}_{t,T,\delta}^{n}$ in (4.37). After simulating $\mathcal{Z}_{t}$ and $\mathcal{A}_{t}$ as described in 5.2.1, we get $\widehat{\mathcal{Z}}_{t}$ and $\widehat{\mathcal{A}}_{t}$. Then, for each $1\leq i\leq I$ and for each $n\in g_{i}$, we have
  1. 1.
    from (3.25), the estimated probability of default of firm $n$ is
    $\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}=\Phi\left(\frac{\log(\mathcal{D}_{t+T}^{n})-\widehat{\mathcal{K}}^{n}(\widehat{\mathfrak{d}},t,T,\widehat{\mathcal{A}}_{t},\widehat{\mathcal{Z}}_{t})}{\widehat{\mathcal{L}}^{n}(t,T)}\right),$ (5.3)
    with
    $\widehat{\mathcal{K}}^{n}(\widehat{\mathfrak{d}},t,T,\widehat{\mathcal{A}}_{t},\widehat{\mathcal{Z}}_{t})=\log{(F^{n}_{0}\widehat{\mathfrak{R}}^{n}_{t+T}(\widehat{\mathfrak{d}}))}+\widehat{\mathfrak{a}}^{n\cdot}(\widehat{\mu}T+\widehat{\varsigma}\Upsilon_{T}\widehat{\mathcal{Z}}_{t}+\widehat{\mathcal{A}}_{t}-\widehat{v}(\widehat{\mathfrak{d}}_{0})),$ (5.4)
    and
    ^$\widehat{\mathcal{L}}^{n}(t,T):=\widehat{\varsigma}^{2}\widehat{\mathfrak{a}}^{n\cdot}\widehat{\Gamma}^{-1}\left(\frac{T}{L}\sum_{l=0}^{L}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\widehat{\Sigma}\widehat{\Sigma}^{\top}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\right)(\widehat{\mathfrak{a}}^{n\cdot}\widehat{\Gamma}^{-1})^{\top}+(t+T)\widehat{\sigma}_{n}^{2},$ (5.5)
    where $F^{n}_{0}$ and $\mathcal{D}_{t+T}^{n}$ are know, $\widehat{\mathfrak{d}}$ defined in section 5.1.1, ^$\widehat{\mathfrak{R}}^{n}_{t+T}(\widehat{\mathfrak{d}}))$ in (3.19) in proposition 3.7, $\widehat{\Gamma},\widehat{\varsigma},\widehat{v}$ in section 5.1.2, $\widehat{\mathfrak{a}}^{n\cdot},\widehat{\sigma}^{n}$ in section 5.1.3, $\widehat{\Upsilon}_{T}:=\widehat{\Gamma}^{-1}(\mathbf{I}_{I}-e^{-\widehat{\Gamma}T})$ and with $u_{l}:=\frac{Tl}{L},l=0,\ldots,L$.
  2. 2.
    If the collateral of loan $n$ is a financial asset, from (4.21),
    $\begin{split}\widehat{\mathrm{LGD}}^{n}_{t,T,\mathfrak{d}}&=\frac{1-\gamma}{\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}}\left[\Phi_{2}\left(\widehat{\overline{\omega}}_{t,T,a}^{n},\Phi^{-1}\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}});\widehat{\rho}_{t,T,a}^{n}\right)-\exp{\left(\frac{1}{2}\widehat{\overline{\mathcal{L}}}^{n}(t,T+a)\right)}\times\right.\\ &\qquad\qquad\qquad\left.\Phi_{2}\left(\widehat{\overline{\omega}}_{t,T,a}^{n}-\sqrt{\widehat{\overline{\mathcal{L}}}^{n}(t,T+a)},\Phi^{-1}(\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}})-\widehat{\rho}_{t,T,a}^{n}\sqrt{\widehat{\overline{\mathcal{L}}}^{n}(t,T+a)};\widehat{\rho}_{t,T,a}^{n}\right)\right],\end{split}$ (5.6)
    where $\widehat{\overline{\mathcal{K}}}$ and $\widehat{\overline{\mathcal{L}}}$ are computed in the same way $\widehat{{\mathcal{K}}}$ and $\widehat{{\mathcal{L}}}$ were in (5.4) and (5.5). Moreover,
    $\widehat{\rho}_{t,T,a}^{n}:=\frac{\widehat{\varsigma}^{2}\widehat{\mathfrak{a}}^{n\cdot}\widehat{\Gamma}^{-1}\left(\frac{T}{L}\sum_{l=0}^{L}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\widehat{\Sigma}\widehat{\Sigma}^{\top}\left(e^{-\widehat{\Gamma}(u_{l}+a)}-\mathbf{I}_{I}\right)\right)(\widehat{\overline{\mathfrak{a}}}^{n\cdot}\widehat{\Gamma}^{-1})^{\top}}{\sqrt{\widehat{\mathcal{L}}^{n}(t,T)\widehat{\overline{\mathcal{L}}}^{n}(t,T+a)}},$
    and
    $\widehat{\overline{\omega}}_{t,T,a}^{n}:=\frac{\log{\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}}-\widehat{\overline{\mathcal{K}}}^{n}(\widehat{\mathfrak{d}},t,T+a,\widehat{\mathcal{A}}_{t}^{\circ},\widehat{\mathcal{Z}}_{t})}{\sqrt{\widehat{\overline{\mathcal{L}}}^{n}(t,T+a)}},$
[…]
  3. 3.
    If the collateral of loan $n$ is a commercial or residential property, we compute in order (4.30), (4.31), (4.36), and (4.37). Since $\chi_{t}=\varrho t+\vartheta$ and $C_{0}^{n}$ are known
    $\widehat{m}_{t,T+a}^{n}:=\log{(R_{n}C_{0}^{n})}+(\widehat{\varrho}t+\widehat{\vartheta})-(\widehat{\vartheta}-K_{0})e^{-\widehat{\nu}(t+T+a)}+\widehat{\overline{\sigma}}\widehat{\rho}^{\top}\sum_{k=0}^{L}e^{-\widehat{\nu}((t+T+a)-\frac{kt}{L})}\eta_{u_{\frac{kT}{L}}},$
    where $\eta_{\frac{kt}{L}}\sim\mathcal{N}\left(0,\frac{t}{L}\mathbf{I}_{I}\right),k=0,\ldots,L$ with $L\in\mathbb{N}^{*}$, and
    $\widehat{v}_{t,T+a}^{n}:=\frac{(\widehat{\overline{\sigma}}\widehat{\rho})^{2})}{\widehat{\nu}}\left(1-e^{-2\widehat{\nu}(T+a)}\right)+\frac{(\widehat{\overline{\sigma}})^{2}(1-(\widehat{\rho})^{2})}{\widehat{\nu}}\left(1-e^{-2\widehat{\nu}(t+T+a)}\right).$
    Therefore, we have
    $\begin{split}\widehat{\mathrm{LGD}}^{n}_{t,T,\delta}&=\frac{1-\gamma}{\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}}\left[\left(1+(1-k)e^{-ra}\frac{R_{n}\widehat{X}_{t+T+a,\delta}^{n}}{\mathrm{EAD}_{t+T}^{n}}\right)\Phi_{2}\left(\widehat{\overline{w}}_{t,T,a}^{n},\Phi^{-1}(\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}});\widehat{\rho}_{t,T,a}^{n}\right)\right.\\ &\qquad\qquad\quad\left.-\exp{\left(\frac{1}{2}\widehat{v}_{t,T+a}^{n}-\sqrt{\widehat{v}_{t,T+a}^{n}}\widehat{\overline{w}}_{t,T,a}^{n}\right)}\Phi_{2}\left(\widehat{\overline{w}}_{t,T,a}^{n}-\sqrt{\widehat{v}_{t,T+a}^{n}},\Phi^{-1}(\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}})-\widehat{\rho}_{t,T,a}^{n}\sqrt{\widehat{v}_{t,T+a}^{n}};\widehat{\rho}_{t,T,a}^{n}\right)\right],\end{split}$ (5.7)
    where
    $\widehat{\rho}_{t,T,a}^{n}:=\widehat{\overline{\sigma}}\widehat{\varsigma}\frac{\widehat{\mathfrak{a}}^{n\cdot}\widehat{\Gamma}^{-1}\left(\frac{T}{L}\sum_{l=0}^{L}e^{-\widehat{\nu}(u_{l}+a)}\left(\mathbf{I}_{I}-e^{-\widehat{\Gamma}u_{l}}\right)\widehat{\Sigma}\right)\widehat{\rho}}{\sqrt{\widehat{\mathcal{L}}^{n}(t,T)\times\widehat{v}_{t,t+T+a}^{n}}},$
    and
    $\widehat{\overline{w}}_{t,T,a}^{n}:=\log{\left(\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)R_{n}e^{-ra}}+\widehat{X}_{t+T+a,\delta}^{n}\right)}-\widehat{m}_{t,T+a}^{n},$
    and $\widehat{X}$ is obtained by considering that from (4.24),
    $\widehat{X}_{t,\delta}^{n}=\mathfrak{c}(\alpha^{n},\alpha^{\star})e^{-r(\mathfrak{t}_{n}-t)}+(\alpha^{n}-\alpha^{\star})\frac{(\mathfrak{t}_{n}-t)}{P}\sum_{p=1}^{P}\mathfrak{f}(\delta_{v_{p}},\mathfrak{p})e^{-r(v_{p}-t)},$
    and where $\gamma$, $k$, $r$, $R_{n}$, and $\mathrm{EAD}_{t}^{n}$ are known, $\mathfrak{t}_{n}$ given by (4.34), $u_{l}:=\frac{(t_{\star}-t)l}{L},l=0,\ldots,L$, and $v_{p}:=\frac{(\mathfrak{t}_{n}-t)p}{P},l=0,\ldots,P$.
5.2.3 Of the (un)expected losses EL and UL
For EL, the result is direct by using $\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}$ in (5.3), $\widehat{\mathrm{LGD}}^{n}_{t,T,\mathfrak{d}}$ in (5.6), and $\widehat{\mathrm{LGD}}^{n}_{t,T,\delta}$ in (5.7), we have from (4.39),
$\begin{split}\widehat{\mathrm{EL}}^{N,T}_{t,\mathfrak{d}}&:=\sum_{n=1}^{N_{1}}(1-\gamma)\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}+\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}^{n}_{t,T,\mathfrak{d}}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}\\ &\qquad\qquad+\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}^{n}_{t,T,\delta}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}.\end{split}$ (5.8)
For UL, we use
$\begin{split}\widehat{L}_{t+T}^{\mathbb{G},N}&=\sum_{n=1}^{N_{1}}(1-\gamma)\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{PD}}_{t+T,\mathfrak{d}}^{n}+\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}_{t+T,\mathfrak{d}}^{n}\cdot\widehat{\mathrm{PD}}_{t+T,\mathfrak{d}}^{n}\\ &\qquad\qquad+\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}_{t+T,\delta}^{n}\cdot\widehat{\mathrm{PD}}_{t+T,\mathfrak{d}}^{n},\end{split}$ (5.9)
by noting that $\mathrm{PD}_{t+T,\mathfrak{d}}^{n}=\mathrm{PD}^{n}_{t+T,0,\mathfrak{d}}$, $\mathrm{LGD}_{t+T,\mathfrak{d}}^{n}=\mathrm{LGD}^{n}_{t+T,0,\mathfrak{d}}$, and $\mathrm{LGD}_{t+T,\delta}^{n}=\mathrm{LGD}^{n}_{t+T,0,\delta}$. Therefore, as $\widehat{L}_{t+T}^{\mathbb{G},N}$ depends on $(\widehat{\mathrm{PD}}_{t+T,\mathfrak{d}}^{n},\widehat{\mathrm{LGD}}_{t+T,\mathfrak{d}}^{n},\widehat{\mathrm{LGD}}_{t+T,\delta}^{n})$ which depends on $(\widehat{\mathcal{A}}_{t+T},\widehat{\mathcal{Z}}_{t+T})$. However, we want to compute $\mathrm{VaR}^{\alpha,N,T}_{t}$ so that $\mathbb{P}\left[L_{t+T}^{\mathbb{G},N}\leq\mathrm{VaR}^{\alpha,N,T}_{t}\middle|\mathcal{G}_{t}\right]$. Then, we simulate $D\in\mathbb{N}^{*}$ couples noted $(\widehat{\mathcal{A}}_{t+T|t}^{p},\widehat{\mathcal{Z}}_{t+T|t}^{p})_{1\leq p\leq D}$ so that $\widehat{\mathcal{Z}}_{t+T|t}^{p}=^{d}\mathcal{Z}_{t+T}|\mathcal{G}_{t}$ and $\widehat{\mathcal{A}}_{t+T|t}^{p}=^{d}\mathcal{A}_{t+T}|\mathcal{G}_{t}$. That is straightforward and
$\widehat{\mathcal{Z}}_{t+T|t}^{p}|\mathcal{G}_{t}\sim\mathcal{N}\left(e^{-\widehat{\Gamma}T}\widehat{\mathcal{Z}}_{t},\frac{T}{L}\sum_{l=0}^{L}e^{-\widehat{\Gamma}u_{l}}\widehat{\Sigma}\widehat{\Sigma}^{\top}e^{-\widehat{\Gamma}^{\top}u_{l}}\right),$ (5.10)
and
$\widehat{\mathcal{A}}_{t+T|t}^{p}|\mathcal{G}_{t}\sim\mathcal{N}\left(\widehat{\mu}T+\widehat{\varsigma}\widehat{\Upsilon}_{T}\widehat{\mathcal{Z}}_{t}+\widehat{\mathcal{A}}_{t},\widehat{\varsigma}^{2}\widehat{\Gamma}^{-1}\left[\frac{T}{L}\sum_{l=0}^{L}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\widehat{\Sigma}\widehat{\Sigma}^{\top}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\right](\widehat{\Gamma}^{-1})^{\top}\right),$ (5.11)
with $u_{l}:=\frac{Tl}{L},l=0,\ldots,L$. We also need to simulate $\mathfrak{h}_{t+T}|\mathcal{G}_{t}:=\int_{0}^{t+T}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s}|\mathcal{G}_{t}$ (which comes from $m_{t+T,0}^{n}$ in (4.30)). As $\mathfrak{h}_{t+T}|\mathcal{G}_{t}\sim\mathcal{N}\left(\int_{0}^{t}e^{-\nu(t+T-s)}\mathrm{d}B^{\mathcal{Z}}_{s},\frac{1-e^{-2\nu T}}{2\nu}\mathbf{I}_{I}\right)$, we have
$\widehat{\mathfrak{h}}_{t+T}^{p}|\mathcal{G}_{t}\sim\mathcal{N}\left(\sum_{k=0}^{L}e^{-\widehat{\nu}((t+T)-\frac{kt}{L})}\eta_{u_{\frac{kT}{L}}},\frac{1-e^{-2\widehat{\nu}T}}{2\widehat{\nu}}\mathbf{I}_{I}\right),\quad\eta_{\frac{kt}{L}}\sim\mathcal{N}\left(0,\frac{t}{L}\mathbf{I}_{I}\right).$
Then, the unexpected loss is
$\widehat{\mathrm{UL}}^{N,T}_{t,\delta,\alpha}:=q_{\alpha,D}\left(\left\{(\widehat{L}_{t+T}^{\mathbb{G},N})^{1},(\widehat{L}_{t+T}^{\mathbb{G},N})^{2},\ldots,(\widehat{L}_{t+T}^{\mathbb{G},N})^{D}\right\}\right)-\widehat{\mathrm{EL}}^{N,T}_{t,\mathfrak{d}},$ (5.12)
where $(\widehat{L}_{t+T}^{\mathbb{G},N})^{p}$ is obtained by replacing $(\widehat{\mathcal{A}}_{t+T},\widehat{\mathcal{Z}}_{t+T})$ in (5.9) by $(\widehat{\mathcal{A}}_{t+T|t}^{p},\widehat{\mathcal{Z}}_{t+T|t}^{p})$, and where $q_{\alpha,M}(\{Y^{1},\ldots,Y^{D}\})$ denotes the empirical $\alpha$-quantile of the distribution of $Y$.
5.3 Summary of the process
More concretely, the goal is to calculate the expected and unexpected losses at each $t\in[\mathfrak{t}_{0},\mathfrak{t}_{1}]$. To achieve that, we use (1) the number of firms rated $r_{t}$ and defaulted $d_{t}$ in each $t\in\mathcal{T}^{M}$, (2) all the firms’ cash flows $(F_{t}^{n})_{1\leq n\leq N}$ in each $t\in\mathcal{T}^{M}$, (3) for each borrower $n$, if a collateral exists and is a financial asset, its cash flows $(\overline{F}_{t}^{n})_{t\in\mathcal{T}^{M}}$ and if it is a building, its surface $R_{n}$ and its price at $\mathfrak{t}_{1}$, $C_{0}^{n}$, (4) the macroeconomic variables as well as the carbon intensities by sector, and the real estate index observed $REI_{t}$ in each $t\in\mathcal{T}^{M}$, and (5) the carbon price dynamics $(\delta_{t})_{t\in\{t_{\circ},\ldots,t_{\star}\}}$ given by the regulator. We proceed as follows.
  1. 1.
    From the macroeconomic historical data, we estimate the productivity parameters $\widehat{\Gamma}$, $\widehat{\mu}$$\widehat{\Sigma}$ and $\widehat{\varsigma}$, as well as the elasticities $\widehat{\psi}$ and $\widehat{\boldsymbol{\lambda}}$ as described in section 5.1.2. We also estimate the real estate market parameters $\widehat{\varrho}$, $\widehat{\vartheta}$, $\widehat{\nu}$, $\widehat{\overline{\sigma}}$, and $\widehat{\rho}$ as described in section 5.1.4.
  2. 2.
    For each $i\in\{1,\ldots,I\}$, we estimate the parameters $B^{n_{i}}$, $\sigma_{\mathfrak{b}^{n_{i}}}$, $\mathbf{a}^{n_{i}}$ using section 5.1.3, yielding $\widehat{B}^{n_{i}}$, $\widehat{\sigma}_{\mathfrak{b}^{n_{i}}}$, $\widehat{\mathbf{a}}^{n_{i}}$.
  3. 3.
    We compute the carbon price dynamics $(\delta_{t})_{t_{\circ}\leq t\leq t_{\star}}$ and the carbon intensities $(\tau_{t})_{t_{\circ}\leq t\leq t_{\star}}$, $(\zeta_{t})_{t_{\circ}\leq t\leq t_{\star}}$, and $(\kappa_{t})_{t_{\circ}\leq t\leq t_{\star}}$ as defined in 3.3 as well as the emissions cost rate $(\mathfrak{d}_{t})_{t_{\circ}\leq t\leq t_{\star}}$ defined in (3.8) and the output carbon cost function $v$ defined in (3.13).
  4. 4.
    For $t_{\circ}\leq t\leq t_{\star}$, we fix a large enough integer $D$, and simulate $D$ paths of the productivity process $(\widehat{\mathcal{Z}}_{t}^{d})_{1\leq d\leq D}$, then we derive $((\mathcal{A}_{t}^{\circ})^{d})_{1\leq d\leq D}$ as defined in section 5.2.1. Then, for each $p$, we simulate $D$ paths $(\widehat{\mathcal{Z}}_{t+T|t}^{d^{\prime}}|\mathcal{G}_{t}^{d})_{1\leq d\leq D}$ and $(\widehat{\mathcal{Z}}_{t+T|t}^{d^{\prime}}|\mathcal{G}_{t}^{d})_{1\leq d^{\prime}\leq D}$ using (5.10) and (5.11) respectively.
  5. 5.
    For each $n\in\{1,\ldots,N\}$ and for each $1\leq d,d^{\prime}\leq D$,
    • (a)
      we compute the probability of default $\widehat{\mathrm{PD}}_{t+1,\mathfrak{d}}^{n}$ using (5.3).
    • (b)
      Depending on the nature of the collateral, we compute $\widehat{\mathrm{LGD}}_{t+1,\delta}^{n}$ , using either (4.11), (5.6), or (5.7)
  6. 6.
    We finally compute the expected (resp. unexpected) losses $\widehat{\mathrm{EL}}^{N}_{t,\mathfrak{d}}$ (resp. $\widehat{\mathrm{UL}}^{N,T}_{t,\delta,\alpha}$), for each $t_{\circ}\leq t\leq t_{\star}$, using (5.8) (resp. (5.12)).
6 Discussion
In this section, we describe the data used to calibrate the different parameters, we perform some simulations, and we comment the results.
6.1 Data
As in Bouveret et al., (2025) and Sopgoui, (2024), we work on data related to the French economy.
  1. 1.
    Due to data availability (precisely, we do not find public monthly/quaterly data for the intermediary inputs), we consider an annual frequency.
  2. 2.
    Annual consumption, labor, output, and intermediary inputs come from INSEE [^3] The French National Institute of Statistics and Economic Studies from 1978 to 2021 (see INSEE, (2023) for details) and are expressed in billion euros, therefore $\mathfrak{t}_{0}=1978$, $\mathfrak{t}_{1}=2021$, and $M=44$.
  3. 3.
    For the climate transition, we consider a time horizon of ten years with $t_{\circ}=2021$ as starting point, a time step of one year and $t_{\star}=2030$ as ending point. In addition, we will be extending the curves to 2034 to see what happens after the transition, even though the results will be calculated and analyzed during the transition.
  4. 4.
    The 38 INSEE sectors are grouped into four categories: Very High Emitting, Very Low Emitting, Low Emitting, and High Emitting, based on their carbon intensities.
  5. 5.
    The carbon intensities are calibrated on the realized emissions from Eurostat, (2023) (expressed in tonnes of CO2-equivalent) between 2008 and 2021.
  6. 6.
    Metropolitan France housing price index comes from OECD [^4] The Organisation for Economic Co-operation and Development data and are from 1980 to 2021 (see OECD Stat, (2024) for details) in Base 2015. We renormalize in Base 2021 (i.e. in 2021, the housing price index is equal to $100$).
6.2 Definition of the climate transition
We consider four deterministic transition scenarios giving four deterministic carbon price trajectories. The scenarios used come from the NGFS [^5] Network for Greening the Financial System. simulations, whose descriptions are given by NGFS, (2022) as follows:
  • 1.
    Net Zero 2050 is an ambitious scenario that limits global warming to $1.5^{\circ}C$ through stringent climate policies and innovation, reaching net zero $\mathrm{CO}_{2}$ [^6] Carbon dioxide. emissions around 2050. Some jurisdictions such as the United States of America, the European Union, and Japan reach net zero for all GHG by this point.
  • 2.
    Divergent Net Zero reaches net-zero by 2050 but with higher costs due to divergent policies introduced across sectors and a quicker phase out of fossil fuels.
  • 3.
    Nationally Determined Contributions (NDCs) includes all pledged policies even if not yet implemented.
  • 4.
    Current Policies assumes that only currently implemented policies are preserved, leading to high physical risks.
For each scenario, the regulator decides the carbon price $P_{carbon,0}$ (in euro per ton of CO2-equivalent) at $t_{0}$ and the speed rate of the transition $\eta_{\delta}$. We summarize them in table 1 below.
Current Policies NDCs Divergent Net Zero Net Zero 2050
$P_{carbon,0}$ (in euro/ton) 30.957 33.321 32.963 34.315
$\eta_{\delta}$ (in %) 1.693 7.994 12.893 17.935
Table 1: Carbon price parameters
We can then compute the carbon price, whose evolution is plotted in Figure 1(a), at each date using (3.9).
For the energy price, we consider electricity as the unique source of energy. Then, we assume a linear relation between the electricity and the carbon price inspired by Abrell et al., (2023), where a variation of the carbon price is linked withe the variation of the electricity by a the pass-through rate noted $k$  [^7] We can definitively consider other types of energy. Moreover, instead of working with a linear relation (with a pass-through rate) between the energy price and the carbon price, we could consider a more complex model such as NiGEM (the National Institute Global Econometric Model in Hantzsche et al., (2018)).. This means that $\mathfrak{f}^{elec}_{1}$ and $\mathfrak{f}^{elec}_{2}$ define in (4.32) are respectively $k$ and $P_{elec,0}-k\times P_{carbon,0}$. For France, we take the electricity price $P_{elec,0}=0.2161$ euro per Kilowatt-hour and $k=0.55$ (see Abrell et al., (2023)) ton per Kilowatt-hour. Its evolution is plotted in Figure 1(b).
(a) Carbon price
(b) Energy price
Figure 1: per scenario and per year
For the renovation costs to improve a building for the energy efficiency $\alpha$ to $\alpha^{\star}$ as defined in (4.33), we take $c_{0}=0.01$ euro per kilowatt-hour and per square meter (€/kWh.m2) and $c_{1}=0.1$.
6.3 Estimations
6.3.1 The carbon intensities
We use the realized GHG emissions as well as the macroeconomic variables and their frequency being the same as in Bouveret et al., (2025), we use the same estimations. But after that, we can compute the carbon intensities at each date in $\mathbb{R}_{+}$ using (3.6).
6.3.2 Economic and housing pricing index (HPI) parameters
We keep the values of $\phi$, $\sigma$, $(\chi^{i})_{i\in\mathcal{I}}$, and $(\boldsymbol{\lambda}^{ji})_{i,j\in\mathcal{I}}$ already estimated. For the productivity process, we switched from a vector autoregressive process to an Ornstein-Uhlenbeck process. We therefore calibrate ${\mu}$, ${\varsigma}$, ${\Sigma}$, and ${\Gamma}$ as detailed in section 5.1.2 and we obtain the same results as in Sopgoui, (2024)[Section 4.3.1.].
We write the housing price index $K$ in Base 2021 and we apply the logarithm function. This means that $K_{t_{0}}=0$. We can therefore calibrate $\varrho,\vartheta,\nu,\overline{\sigma}$, and $\rho$. The values are presented in Sopgoui, (2024)[Table 5].
6.4 Simulations and discussions
In the previous work in discrete time, we simulate for different climate transition scenario between $t_{\circ}=2021$ and $t_{\star}=2030$, the annual evolution of (1) the output growth per sector (2) the output share per sector in the total output, (3) the firms direct GHG emissions per sector, (4) a given firm value and distribution, (5) the probabilities of default of fictive sub-portfolio of 4 firms each and of the resulting portfolio, (6) the expected and the unexpected losses of the previous (sub-)portfolios when the LGD are constant and deterministic, (7) the sensitivities of the losses to the carbon price.
In the current simulations, since we are keeping the same data at the same frequency (annual), the main change is then to replace the Vector Autoregressive process by the O.-U. process. Therefore, the comments already made for (1) to (5) concerning the trends, the impact of the carbon price, the difference of scenarios, the relation between sectors, etc. do not change. We will focus here on the LGD and on the losses, with different type of collateral.
We will assume $D=500$. This may seem low but it is actually $500\times 500$ simulations. This is computationally-intensive. In fact, for each year $t$, we firstly simulate $500$ times $(\mathcal{Z}_{t},\mathcal{A}_{t})$. However, we are interested in one year risk measures  [^8] Recall for example that we are in 2025, we would like to compute the probability that a firm defaults between 1 January 2027 and 31 December 2027., therefore we also simulate $500$ times $(\mathcal{Z}_{t+1},\mathcal{A}_{t+1})$ given $(\mathcal{Z}_{t},\mathcal{A}_{t})$. Moreover, we considered 4 sectors and for each step, there are numerical integrals, matrices products, etc. to compute.
6.4.1 Impact of the carbon price on Loss Given Default
When there is no guarantee, we assume as in the previous work that $\mathrm{LGD}$ is equal to $45\%$ so that $\gamma=0.55$. To illustrate the case where there is guarantee, we consider, both if the collateral is a financial asset and a building, $\mathrm{EAD}$ starts at $200$ and growths annually as the economic total output growth in the Current Policies scenario (see Table 2 below).
Year 2021 2022 2023 2024 2025 2026 2027 2028 2029 2030 2031 2032 2033 2034
EAD 200. 202.8 206.6 209.9 213.1 216.2 219.5 222.6 226.2 229.6 233.1 236.3 239.9 243.9
Table 2: EAD per year
If the collateral is a financial asset
We consider 4 firms so that firm $1$, $2$, $3$, and $4$ respectively belong to the Very High Emitting, High Emitting, Low Emitting, and Very Low Emitting groups. Each firm is characterized by its cash flows $F_{t_{\circ}-1}$ at $t_{\circ}-1$, the standard deviation of its cash flows $\sigma_{\mathfrak{b}}$, and the contribution $\mathfrak{a}$ of sectoral output growth to its cash flows growth as detailed in table 3. The chosen interest rate $r=5\%$.
Firm $1$ $2$ $3$ $4$
$\sigma_{\mathfrak{b}^{n}}$ 0.05 0.05 0.05 0.05
$F_{0}^{n}$ 1.0 1.0 1.0 1.0
$\mathfrak{a}^{n}$(Very High) 1.0 0.0 0.0 0.0
$\mathfrak{a}^{n}$(High) 0.0 1.0 0.0 0.0
$\mathfrak{a}^{n}$(Low) 0.0 0.0 1.0 0.0
$\mathfrak{a}^{n}$(Very Low) 0.0 0.0 0.0 1.0
Table 3: Characteristics of the firms
We compute here for $D=500$ simulations of the productivity processes $(\mathcal{Z},\mathcal{A})$, the loss given default of 4 loans with the same exposure but with 4 different financial assets collateral described in Table 3.
Emissions level No collateral Firm 1 Firm 2 Firm 3 Firm 4
Current Policies 45. 32.934 31.960 34.561 29.281
NDCs 45. 33.177 32.184 34.609 29.357
Divergent Net Zero 45. 33.485 32.471 34.673 29.459
Net Zero 2050 45. 33.995 32.940 34.784 29.640
Table 4: Average annual LGD per scenario between 2021 and 2030 (in %)
Both in Table 4 and in Figure 2, we can first see that the presence of guarantees reduce $\mathrm{LGD}$. Without collateral, we assume 45%, and with collateral, for all scenarios and for different characteristics of firms, $\mathrm{LGD}$ is less than 45%.
Figure 2: LGD with a financial asset as collateral
However, the decreasing of LGD depends on the scenarios. When the scenario becomes tougher, the impact of the presence of the collateral on LGD is lessened. This is logical and due to the fact that the value of the liquidated asset loses value when the price of carbon rises. The decreasing of the LGD also depends on the distinctive characteristics of the guarantees. Precisely, each firm in Table 3, serving as collateral, belongs to a unique and distinct sector (through $\mathfrak{a}$), which go from the more to the less polluting. Therefore the more the collateral is in a polluting sector, the less it reduces $\mathrm{LGD}$.
If the collateral is a building
We consider 5 apartments of 25 square meters whose price of the square meter fixed to 4000 euros in $t_{\circ}=2021$ is the same for all, but whose the energy efficiency are different. Moreover, we assume that the optimal energy efficiency equals to $\alpha^{\star}=70$ kilowatt hour per square meter per year (see Total Energies, (2024)) is reached.
Building $1$ $2$ $3$ $4$ $5$
$C_{n}^{0}$ 4000 4000 4000 4000 4000
$\alpha^{n}$ 320. 253. 187. 120. 70.
$R_{n}$ 25.0 25.0 25.0 25.0 25.
Table 5: Characteristics of the building
We use the $D=500$ trajectories of the productivity processes $(\mathcal{Z},\mathcal{A})$ simulated above. We compute the loss given default of 4 loans with the same exposure but with the 4 buildings described in Table 5 as collateral.
Emissions level No collateral Building 1 Building 2 Building 3 Building 4
Current Policies 45. 36.020 36.152 35.928 36.095
NDCs 45. 38.383 37.752 36.922 36.499
Divergent Net Zero 45. 38.939 38.303 37.377 36.751
Net Zero 2050 45. 39.102 38.524 37.615 36.908
Table 6: Average annual LGD per scenario between 2021 and 2030 (in %)
As we see in tables 6 and 3, all the comments made for a financial asset as collateral are valid for a building: the presence of a collateral reduces LGD, but this reduction is smaller when the climate transition scenario becomes tougher.
Figure 3: LGD with a building as collateral
There are two main differences. First, the more the building is energetically inefficient, the more LGD increases (it is the same above when the financial asset belongs to a very polluting sector). Secondly, LGD decreases when time increases. This is a consequence of the dynamics of the impact of the carbon price of the housing market (as described in Sopgoui, (2024)): as we approach the optimal renovation date, the prices of energy-inefficient buildings rise and converge progressively towards the prices of energy-efficient buildings (we can see on Sopgoui, (2024)[Figure 3]). LGD follows the same behaviour logically but with an inverse monotony.
6.4.2 Expected and unexpected loss
To this aim, to keep things simple, we will consider a credit portfolio of $N=12$ loans contracted by the firms described in Table 7 below.
Loans 1 2 3 4 5 6 7 8 9 10 11 12
$\mathrm{EAD}_{n}$ 200. 200. 200. 200. 200. 200. 200. 200. 200. 200. 200. 200.
$F_{0}^{n}$ 1. 1. 1. 1. 1. 1. 1. 1. 1. 1. 1. 1.
$B^{n}$ 3.76 3.98 3.75 4.41 3.76 3.98 3.75 4.41 3.76 3.98 3.75 4.41
$\sigma_{\mathfrak{b}^{n}}$ 0.05 0.05 0.05 0.05 0.05 0.05 0.05 0.05 0.05 0.05 0.05 0.05
$\mathfrak{a}^{n}$(Very High) 1. 0. 0. 0. 1. 0. 0. 0. 1. 0. 0. 0.
$\mathfrak{a}^{n}$(High) 0. 1. 0. 0. 0. 1. 0. 0. 0. 1. 0. 0.
$\mathfrak{a}^{n}$(Low) 0. 0. 1. 0. 0. 0. 1. 0. 0. 0. 0. 0.
$\mathfrak{a}^{n}$(Very Low) 0. 0. 0. 1. 0. 0. 0. 1. 0. 0. 0. 0.
Collateral type No No No No Fa Fa Fa Fa Ho Ho Ho Ho
$\overline{F}_{0}^{n}$ 1. 1. 1. 1.
$\overline{\sigma}_{\mathfrak{b}^{n}}$ 0.05 0.05 0.05 0.05
$\overline{\mathfrak{a}}^{n}$(Very High) 1. 0. 0. 0.
$\overline{\mathfrak{a}}^{n}$(High) 0. 1. 0. 0.
$\overline{\mathfrak{a}}^{n}$(Low) 0. 0. 1. 0.
$\overline{\mathfrak{a}}^{n}$(Very Low) 0. 0. 0. 1.
$C_{n}^{0}$ 4000. 4000. 4000. 4000.
$R_{n}$ 25. 25. 25. 25.
$\alpha^{n}$ 320. 253. 187. 120.
Table 7: Characteristics of the portfolio (No = no collateral, Fa = Financial asset collateral, Ho = housing collateral)
We can remark that, for each $k=0,1,2$, firms $14k+$ 1, $24k+$ 2, $34k+$ 3, and $44k+$ 4 respectively belong to the Very High Emitting, High Emitting, Low Emitting, and Very Low Emitting groups. Moreover, we assume that
  • 1.
    the loans of the firms $1$, $2$, $3$, and $4$ are not collateralized;
  • 2.
    the loans of the firms $5$, $6$, $7$, and $8$ are collateralized by financial assets described in Table 3;
  • 3.
    the loans of the firms $9$, $10$, $11$, and $12$ are collateralized by a building described in Table 5.
We want to calculate the expected (respectively unexpected) loss noted $\mathrm{EL}$ (respectively $\mathrm{UL}$) for each loan $n=1,\ldots,12$, by using (5.8) (respectively (5.12)).
Emissions level 1 2 3 4 5 6 7 8 9 10 11 12
Current Policies 1.00 1.00 1.00 1.00 0.79 0.77 0.81 0.71 0.58 0.59 0.58 0.58
NDCs 1.28 1.17 1.04 1.02 1.01 0.90 0.85 0.72 0.91 0.79 0.66 0.62
Divergent Net Zero 1.74 1.41 1.10 1.05 1.38 1.10 0.90 0.75 1.29 0.99 0.72 0.65
Net Zero 2050 2.85 1.91 1.21 1.11 2.27 1.47 0.98 0.79 2.13 1.37 0.81 0.70
Table 8: Average annual EL per scenario between 2021 and 2030 (in %)
Table 8 (respectively Table 9) shows average annual EL (respectively UL) normalized to the EL without collateral observed in the scenario Current Policies. We can make two key observations that were to be expected from the PD and LGD calculations:
  1. 1.
    Whether collateral is involved or not, we can see that EL and UL increase as the transition hardens. This is to be expected, since PD and LGD behave in the same way.
  2. 2.
    When a loan is collateralized, it significantly reduces the bank’s expected and unexpected losses. And for collateralized loans, these losses increase if the collateral has a high carbon footprint: in particular, if the collateral is a financial asset whose value growth is driven by a polluting sector or if it is a building that is not energy efficient.
Emissions level 1 2 3 4 5 6 7 8 9 10 11 12
Current Policies 1.00 1.00 1.00 1.00 0.81 0.81 0.84 0.82 0.59 0.57 0.56 0.62
NDCs 1.18 1.01 1.00 1.00 0.95 0.82 0.84 0.82 0.85 0.67 0.62 0.64
Divergent Net Zero 1.42 1.02 1.00 1.00 1.16 0.83 0.84 0.82 1.05 0.71 0.64 0.65
Net Zero 2050 1.80 1.02 1.0 0.99 1.49 0.85 0.85 0.82 1.34 0.73 0.65 0.66
Table 9: Average annual UL per scenario between 2021 and 2030 (in %)
We can finally say that introducing and increasing the carbon price in the economy will
  1. 1.
    increase bank charges (materialized by the level of provisions calculated from the expected loss $EL$) charged to clients or by the operating income of banks;
  2. 2.
    and reduce the solvency and profitability of banks (translated by the economic capital calculated from the unexpected loss $UL$).
For example, consider the Net Zero 2050 scenario, for a borrower belonging to a very high GHG-emitting sector and its collateral too, the bank will see its associated EL increases by 187% and its associated UL by 83%. Conversely, if the borrower is in a very low GHG emitting sector and its collateral too, the bank will obtain its associated EL increases by 11% and its UL unchanged.
Conclusion
Following Bouveret et al., (2025), we developed here a framework to quantify the impacts of the carbon price on a credit portfolio (expected and unexpected) losses, when the obligor companies as well as their guarantees belong to an economy subject to the climate transition declined by carbon price. We start by describing a closed economy, driven by a productivity following a multidimensional Ornstein-Uhlenbeck and subject to a climate transition modeled through a dynamic and deterministic carbon price, by a dynamic stochastic multisectoral. Then, by using the discounted cash flow methodology with the cash flows, following a stochastic differential equation, depending on the productivity as well as the carbon price, we evaluate the obligor value that helps us later on to compute its probability of over-indebtedness. We then turn to the bank’s loss in the event of a borrower’s over-indebtedness and if its loan is collateralized. When that is the case, the potential loss of the bank is written as the difference between the debt amount (EAD) and the collateral liquidated. We finally distinguish two types of collateral: either a financial asset or a building, both belonging to the economy so affected by the productivity and the carbon price. This work opens the door to many extensions as a finer modeling of the real estate market, taking into account other types of guarantees, modeling the unsecured loans that we assumed constant, modelling the impact of the carbon price on the exposure. This work can be extended to include physical risk. In order to implement that, we could for example introduce the damage multiplier defined through the temperature, and which will affected the output of each sector. Therefore, for each sector $i\in\mathcal{I}$, we will have
$\exp{\left(-\frac{\gamma_{i}}{2}T_{t}^{2}\right)},$ (6.1)
where $\gamma_{i}\geq 0$ is the damage function coefficient and $T_{t}$ is the mean temperature of the earth relative to its pre-industrial level.
Appendix A Proofs
A.1 Hurwitz matrix
Assume that $-\Gamma$ is a Hurwitz matrix, then
  1. 1.
    if we note $\lambda_{\Gamma}:=\max_{\lambda\in\lambda(\Gamma)}Re(\lambda)\geq 0$, there exists $c_{\Gamma}>0$ so that $\lVert e^{-\Gamma t}\rVert<c_{\Gamma}e^{-\lambda_{\Gamma}t}$ for all $t\geq 0$.
  2. 2.
    Moreover, for $t\geq 0$n $\Upsilon_{t}$ defined in (3.4) is such that
    $\lVert\Upsilon_{t}\rVert=\left\lVert\int_{0}^{t}e^{-\Gamma s}\mathrm{d}s\right\rVert\leq\int_{0}^{t}\left\lVert e^{-\Gamma s}\right\rVert\mathrm{d}s\leq c_{\Gamma}\int_{0}^{t}e^{-\lambda_{\Gamma}s}\mathrm{d}s\leq c_{\Gamma}\min{\left\{\frac{1}{\lambda_{\Gamma}},t\right\}}.$ (A.1)
A.2 Bivariate Gaussian
Assume that $X$ and $Y$ are two standard Gaussian with correlation coefficient $\rho$. We then have for $(x,y)\in\mathbb{R}^{2}$, the cdf,
$\Phi_{2}(x,y):=\mathbb{P}[X\leq x,Y\leq y]=\frac{1}{2\pi(1-\rho^{2})}\int_{-\infty}^{x}\int_{-\infty}^{y}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left(u^{2}+v^{2}-2\rho uv\right)\right)}\mathrm{d}u\mathrm{d}v.$ (A.2)
Let $\sigma>0$, we want to compute $\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]$. We have
$\begin{split}\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]&=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\int_{-\infty}^{x}\int_{-\infty}^{y}e^{\sigma u}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left(u^{2}+v^{2}-2\rho uv\right)\right)}\mathrm{d}u\mathrm{d}v\\ &=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\int_{-\infty}^{x}e^{\sigma u-\frac{1}{2(1-\rho^{2})}u^{2}}\int_{-\infty}^{y}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left(v^{2}-2\rho uv\right)\right)}\mathrm{d}v\quad\mathrm{d}u\\ &=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\int_{-\infty}^{x}e^{\sigma u-\frac{1}{2(1-\rho^{2})}u^{2}}\int_{-\infty}^{y}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left((v-\rho u)^{2}-\rho^{2}u^{2}\right)\right)}\mathrm{d}v\quad\mathrm{d}u\\ &=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\int_{-\infty}^{x}e^{\sigma u-\frac{1}{2}u^{2}}\int_{-\infty}^{y}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left((v-\rho u)^{2}\right)\right)}\mathrm{d}v\quad\mathrm{d}u\\ \end{split}$
But
$\int_{-\infty}^{y}\exp{\left(-\frac{1}{2(1-\rho^{2})}\left((v-\rho u)^{2}\right)\right)}\mathrm{d}v=\sqrt{2\pi(1-\rho^{2})}\Phi\left(\frac{y-\rho u}{\sqrt{1-\rho^{2}}}\right),$
therefore,
$\begin{split}\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]&=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{x}e^{\sigma u-\frac{1}{2}u^{2}}\Phi\left(\frac{y-\rho u}{\sqrt{1-\rho^{2}}}\right)\mathrm{d}u\\ &=\frac{e^{\frac{1}{2}\sigma^{2}}}{\sqrt{2\pi}}\int_{-\infty}^{x}e^{-\frac{1}{2}(u-\sigma)^{2}}\Phi\left(\frac{y-\rho u}{\sqrt{1-\rho^{2}}}\right)\mathrm{d}u\\ &=e^{\frac{1}{2}\sigma^{2}}\int_{-\infty}^{x-\sigma}\phi(u)\Phi\left(\frac{y-\rho\sigma}{\sqrt{1-\rho^{2}}}+\frac{-\rho}{\sqrt{1-\rho^{2}}}u\right)\mathrm{d}u.\end{split}$
However,
$\int_{-\infty}^{c}\Phi(a+bx)\phi(x)\mathrm{d}x=\Phi_{2}\left(c,\frac{a}{\sqrt{1+b^{2}}};\frac{-b}{\sqrt{1+b^{2}}}\right),$
we can then conclude that
$\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]=e^{\frac{1}{2}\sigma^{2}}\Phi_{2}\left(x-\sigma,y-\rho\sigma;\rho\right).$ (A.3)
Appendix B The multisectoral model in continuous time
For all $i\in\mathcal{I}$, let us consider the following $\mathbb{G}$-measurable and positive processes: $Y^{i}$ the production of sector $i$, $N^{i}$ the labor demand in sector $i$, and for all $j\in\mathcal{I}$, $Z^{ji}$ the consumption by sector $i$ of intermediate inputs produced by sector $j$.
B.1 The firm’s point of view
Aiming to work with a simple model, we follow (Galí, , 2015, Chapter 2). It then appears that the firm’s problem corresponds to an optimization performed at each period, depending on the state of the world. This problem will depend, in particular, on the productivity and the price processes introduced above. Moreover, it will also depend on $P^{i}$ and $W^{i}$, two $\mathbb{G}$-adapted positive stochastic processes representing respectively the price of good $i$ and the wage paid in sector $i\in\mathcal{I}$. We start by considering the associated deterministic problem below, when time and randomness are fixed.
Solution for the deterministic problem
We denote $\overline{a}\in(0,+\infty)^{I}$ the level of technology in each sector, $\overline{p}\in(0,\infty)^{I}$ the price of the goods produced by each sector, $\overline{w}\in(0,\infty)^{I}$ the nominal wage in each sector, $\overline{\tau}\in[0,1)^{I}$ and $\overline{\zeta}\in[0,1)^{I\times I}$ the price on production and consumption of goods. For $i\in\mathcal{I}$, we consider a representative firm of sector $i$, with technology described by the production function
+×$\mathbb{R}_{+}\times\mathbb{R}_{+}^{I}\ni(n,z)\mapsto F^{i}_{\overline{a}}(n,z)=\overline{a}^{i}n^{\psi^{i}}\prod_{j\in\mathcal{I}}(z^{j})^{\boldsymbol{\lambda}^{ji}}\in\mathbb{R}_{+},$ (B.1)
where $n$ represents the number of hours of work in the sector, and $z^{j}$ the firm’s consumption of intermediary input produced by sector $j$. The coefficients $\psi\in(\mathbb{R}_{+}^{*})^{I}$ and $\boldsymbol{\lambda}\in(\mathbb{R}_{+}^{*})^{I\times I}$ are elasticities satisfying (3.10). The management of firm $i$ then solves the classical problem of profit maximization
$\widehat{\Pi}^{i}_{(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta},\overline{\delta})}:=\sup_{(n,z)\in\mathbb{R}_{+}\times\mathbb{R}_{+}^{I}}\Pi^{i}(n,z),$ (B.2)
where, omitting the dependency in $(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})$,
$\Pi^{i}(n,z):=F^{i}_{\overline{a}}(n,z)\overline{p}^{i}-\overline{\tau}^{i}F^{i}_{\overline{a}}(n,z)\overline{p}^{i}\overline{\delta}-\overline{w}^{i}n-\sum_{j\in\mathcal{I}}z^{j}\overline{p}^{j}+z^{j}\overline{\zeta}^{ji}\overline{p}^{j}\overline{\delta}.$ (B.3)
Note that $F^{i}_{\overline{a}}(n,z)(1-\overline{\tau}^{i})\overline{p}^{i}$ represents the firm’s revenues after carbon price, that $\overline{w}^{i}n$ stands for the firm’s total compensations, and that $\sum_{j\in\mathcal{I}}z^{j}(1+\overline{\zeta}^{ji})\overline{p}^{j}$ is the firm’s total intermediary inputs. Now, we would like to solve the optimization problem for the firms, namely determine the optimal demands $\mathfrak{n}$ and $\mathfrak{z}$ as functions of $(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})$. Because we will lift these optimal quantities in a dynamical stochastic setting, we impose that they are expressed as measurable functions. We thus introduce:
Definition B.1.
An admissible solution to problem (B.2) is a pair of measurable functions
$(\mathfrak{n},\mathfrak{z}):(0,+\infty)^{I}\times(0,+\infty)^{I}\times(0,+\infty)^{I}\times[0,1)^{I}\times[0,1)^{I\times I}\rightarrow[0,+\infty)^{I}\times[0,+\infty)^{I\times I},$
such that, for each sector $i$, denoting $\overline{n}:=\mathfrak{n}^{i}(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})$ and $\overline{z}:=\mathfrak{z}^{\cdot i}(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})$,
$F^{i}_{\overline{a}}(\overline{n},\overline{z})(1-\overline{\tau}^{i}\overline{\delta})\overline{p}^{i}-\overline{w}^{i}\overline{n}-\sum_{j\in\mathcal{I}}\overline{z}^{j}(1+\overline{\zeta}^{ji}\overline{\delta})\overline{p}^{j}=\widehat{\Pi}^{i}_{(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta},\overline{\delta})},$
and $F^{i}_{\overline{a}}(\overline{n},\overline{z})>0$ (non-zero production), according to (B.2).
Remark B.2.
The solution obviously depends also on the coefficients $\psi$ and $\boldsymbol{\lambda}$. But these are fixed once and we will not study the dependence of the solution with respect to them.
Proposition B.3.
There exists admissible solutions in the sense of Definition B.1. Any admissible solution is given by for all $i\in\mathcal{I}$, $\mathfrak{n}^{i}>0$ and for all $(i,j)\in\mathcal{I}^{2}$,
$\mathfrak{z}^{ji}=\frac{\boldsymbol{\lambda}^{ji}}{\psi^{i}}\frac{\overline{w}^{i}}{(1+\overline{\zeta}^{ji}\overline{\delta})\overline{p}^{j}}\mathfrak{n}^{i}>0.$ (B.4)
Moreover, it holds that $\widehat{\Pi}^{i}_{(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta},\overline{\delta})}=0$ (according to (B.2)) and
$\displaystyle\mathfrak{n}^{i}$ $=\psi^{i}F^{i}_{\overline{a}}(\mathfrak{n}^{i},\mathfrak{z}^{\cdot i})\frac{(1-\overline{\tau}^{i}\overline{\delta})\overline{p}^{i}}{\overline{w}^{i}},$ (B.5a)
$\displaystyle\mathfrak{z}^{ji}$ $=\boldsymbol{\lambda}^{ji}F^{i}_{\overline{a}}(\mathfrak{n}^{i},\mathfrak{z}^{\cdot i})\frac{(1-\overline{\tau}^{i}\overline{\delta})\overline{p}^{i}}{(1+\overline{\zeta}^{ji}\overline{\delta})\overline{p}^{j}}.$ (B.5b)
Proof.
We study the optimization problem for the representative firm $i\in\mathcal{I}$. Since $\psi^{i}>0$ and $\boldsymbol{\lambda}^{ji}>0$, for all $j\in\mathcal{I}$, as soon as $n=0$ or $z^{j}=0$, for some $j\in\mathcal{I}$, the production is equal to 00. From problem (B.2), we obtain that necessarily $n\neq 0$ and $z^{j}\neq 0$ for all $j$ in this case. So an admissible solution, which has non-zero production, has positive components.
Setting $\overline{n}=\mathfrak{n}^{i}(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})>0$ and $\overline{z}=\mathfrak{z}^{\cdot i}(\overline{a},\overline{w},\overline{p},\overline{\tau},\overline{\zeta})>0$, the optimality of $(\overline{n},\overline{z})$ yields
$\partial_{n}\Pi^{i}(\overline{n},\overline{z})=0\text{ and for any }j\in\mathcal{I},\quad\partial_{z^{j}}\Pi^{i}(\overline{n},\overline{z})=0.$
We then compute
$\psi^{i}\frac{F^{i}_{\overline{a}}(\overline{n},\overline{z})}{\overline{n}}(1-\overline{\tau}^{i}\overline{\delta})\overline{p}^{i}-\overline{w}^{i}=0\text{ and for any }j\in\mathcal{I},\quad\boldsymbol{\lambda}^{ji}\frac{F^{i}_{\overline{a}}(\overline{n},\overline{z})}{\overline{z}^{j}}(1-\overline{\tau}^{i}\overline{\delta})\overline{p}^{i}-(1+\overline{\zeta}^{ji}\overline{\delta})\overline{p}^{j}=0,$
which leads to (B.4), (B.5a), and (B.5b). ∎
Dynamic setting
In B.3 below, we characterize the dynamics of the output and consumption processes using market equilibrium arguments. There, the optimal demand by the firm for intermediary inputs and labor is lifted to the stochastic setting where the admissible solutions then write as functions of the productivity, carbon price, price of goods/services; and wage processes, see Definition B.5. For all $i\in\mathcal{I}$, $Y^{i}$ representing the production of sector $i$, $N^{i}$ representing the labor demand in sector $i$, and for all $j\in\mathcal{I}$, $Z^{ji}$ representing the consumption by sector $i$ of intermediate inputs produced by sector $j$ are therefore positive and $\mathbb{G}$-adapted processes.
B.2 The household’s point of view
Let $(r_{t})_{t\geq 0}$ be the (exogenous) deterministic interest rate, valued in $\mathbb{R}_{+}$. At each time $t\geq 0$ and for each sector $i\in\mathcal{I}$, we denote
  • 1.
    $C_{t}^{i}$ the quantity consumed of the single good in the sector $i$, valued in $\mathbb{R}_{+}^{*}$;
  • 2.
    $H_{t}^{i}$ the number of hours of work in sector $i$, valued in $\mathbb{R}_{+}^{*}$.
We also introduce a time preference parameter $\beta\in[0,1)$ and a utility function $U:(0,\infty)^{2}\to\mathbb{R}$ given, for $\varphi\geq 0$, by $U(x,y):=\frac{x^{1-\sigma}}{1-\sigma}-\frac{y^{1+\varphi}}{1+\varphi}$ if $\sigma\in[0,1)\cup(1,+\infty)$ and by $U(x,y):=\log(x)-\frac{y^{1+\varphi}}{1+\varphi}$, if $\sigma=1$. We also suppose that
$\mathfrak{P}:=\sup_{t\geq 0,i\in\mathcal{I}}\mathbb{E}\left[\left(\frac{P_{t}^{i}}{W_{t}^{i}}\right)^{1+\varphi}\right]<+\infty.$ (B.6)
For any ${C},{H}\in\mathscr{L}^{1}_{+}(\mathbb{G},(0,\infty)^{I})$, we introduce the wealth process
$\mathrm{d}Q_{t}=r_{t}Q_{t}\mathrm{d}t+\sum_{i\in\mathcal{I}}W_{t}^{i}H_{t}^{i}-\sum_{i\in\mathcal{I}}P_{t}^{i}C_{t}^{i}-\sum_{i\in\mathcal{I}}\kappa_{t}^{i}P_{t}^{i}C_{t}^{i}\delta_{t},\qquad\text{for any }t\geq 0,$ (B.7)
with the convention $Q_{0}:=0$ and $r_{0}:=0$. Note that we do not indicate the dependence of $Q$ upon $C$ and $H$ to alleviate the notations.
For $t\geq 0$ and $i\in\mathcal{I}$, $P_{t}^{i}C_{t}^{i}$ represents the household’s consumption in the sector $i$ and $\kappa_{t}^{i}P_{t}^{i}C_{t}^{i}\delta_{t}$ is the cost paid by households due to their emissions when they consume goods $i$, so $\sum_{i\in\mathcal{I}}P_{t}^{i}C_{t}^{i}(1+\kappa_{t}^{i}\delta_{t})$ is the household’s total expenses. Moreover, $W_{t}^{i}H_{t}^{i}$ is the household’s labor income in the sector $i$, $(1+r_{t-1})Q_{t-1}$ the household’s capital income, and $(1+r_{t-1})Q_{t-1}+\sum_{i\in\mathcal{I}}W_{t}^{i}H_{t}^{i}$ the household’s total revenue.
We define $\mathscr{A}$ as the set of all couples $({C},{H})$ with ${C},{H}\in\mathscr{L}^{1}_{+}(\mathbb{G},(0,\infty)^{I})$ such that
$\left\{\begin{array}[]{ll}&\displaystyle\mathbb{E}\left[\sum_{i\in\mathcal{I}}\int_{t=0}^{\infty}\beta^{t}|U(C_{t}^{i},H_{t}^{i})|\mathrm{d}t\right]<\infty,\\ &\lim_{T\uparrow\infty}\mathbb{E}[Q_{T}|\mathcal{G}_{t}]\geq 0,\qquad\text{forall }t\geq 0.\end{array}\right.$
The representative household consumes the $I$ goods of the economy and provides labor to all the sectors. For any $({C},{H})\in\mathscr{A}$, let
$\mathcal{J}({C},{H}):=\sum_{i\in\mathcal{I}}\mathcal{J}_{i}({C}^{i},{H}^{i}),\qquad\textrm{with}\qquad\mathcal{J}_{i}({C}^{i},{H}^{i}):=\mathbb{E}\left[\int_{t=0}^{\infty}\beta^{t}U(C_{t}^{i},H_{t}^{i})\mathrm{d}t\right],\quad\textrm{for all }i\in\mathcal{I}.$
The representative household seeks to maximize its objective function by solving
$\max_{({C},{H})\in\mathscr{A}}\quad\mathcal{J}({C},{H}).$ (B.8)
We choose above a separable utility function as Miranda-Pinto and Young, (2019) does, meaning that the representative household optimizes its consumption and hours of work for each sector independently but under a global budget constraint. The following proposition provides an explicit solution to (B.8).
Proposition B.4.
Assume that (B.8) has a solution $(C,H)\in\mathscr{A}$. Then, for all $i,j\in\mathcal{I}$, the household’s optimality condition reads, for any $t\geq 0$,
$\displaystyle\frac{P_{t}^{i}}{W_{t}^{i}}$ $=\displaystyle\frac{1}{1+\kappa_{t}^{i}\delta_{t}}(H_{t}^{i})^{-\varphi}(C_{t}^{i})^{-\sigma},$ (B.9a)
$\displaystyle\frac{P_{t}^{i}}{P_{t}^{j}}$ $=\displaystyle\frac{1+\kappa_{t}^{j}\delta_{t}}{1+\kappa_{t}^{i}\delta_{t}}\left(\frac{C_{t}^{i}}{C_{t}^{j}}\right)^{-\sigma}.$ (B.9b)
Note that the discrete-time processes $C$ and $H$ cannot hit zero by definition of $\mathscr{A}$, so that the quantities above are well defined.
Proof.
Suppose that $\sigma\neq 1$. We first check that $\mathscr{A}$ is non empty. Assume that, for all $t\geq 0$ and $i\in\mathcal{I}$, $\tilde{C}_{t}^{i}=1$ and $\tilde{H}_{t}^{i}=\frac{P_{t}^{i}(1+\kappa_{t}^{i})}{W_{t}^{i}}$, then
$\mathbb{E}\left[\sum_{i\in\mathcal{I}}\int_{t=0}^{\infty}\beta^{t}|U(\tilde{C}_{t}^{i},\tilde{H}_{t}^{i})|\mathrm{d}t\right]$ $\leq\sum_{i\in\mathcal{I}}\int_{t=0}^{\infty}\beta^{t}\left(\frac{1}{1-\sigma}+\frac{1}{1+\varphi}\mathbb{E}\left[\left(\frac{P_{t}^{i}(1+\kappa_{t}^{i}\delta_{t})}{W_{t}^{i}}\right)^{1+\varphi}\right]\right)\mathrm{d}t.$
$\leq\sum_{i\in\mathcal{I}}\int_{t=0}^{\infty}\beta^{t}\left(\frac{1}{1-\sigma}+\frac{\mathfrak{P}(1+\kappa_{t}^{i}\delta_{t})^{1+\varphi}}{1+\varphi}\right)\mathrm{d}t<+\infty,$
using (B.6). We also observe that $Q$ built from $\tilde{H},\tilde{C}$ satisfies $Q_{t}=0$, for $t\geq 0$. Thus $(\tilde{H},\tilde{C})\in\mathscr{A}$.
Let now $(\widehat{{C}},\widehat{{H}})\in\mathscr{A}$ be such that $\mathcal{J}(\widehat{{C}},\widehat{{H}})=\max_{({C},{H})\in\mathscr{A}}\mathcal{J}({C},{H})$.
[…]
$\overline{{C}}:=\widehat{{C}}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}}\Delta^{(i,s)}\text{ and }\overline{{H}}:=\widehat{{H}}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}}\Delta^{(i,s)}\frac{{P}^{i}(1+\kappa^{i}\delta_{s})}{{W}^{i}}.$ (B.10)
We observe that for $(j,t)\neq(i,s)$, $\overline{{C}}^{j}_{t}=\widehat{{C}}^{j}_{t}$ and $\overline{H}^{j}_{t}=\widehat{H}^{j}_{t}$ and we compute
$\overline{{C}}^{i}_{s}\geq\widehat{{C}}^{i}_{s}-\theta^{(i,s)}\geq\frac{1}{2}\widehat{{C}}^{i}_{s}>0.$
Similarly, we obtain $\overline{H}^{i}_{s}>0$. We also observe that $\overline{C}\leq\frac{3}{2}\widehat{C}$ and $\overline{H}\leq\frac{3}{2}\widehat{H}$. Finally, we have that
$\sum_{j\in\mathcal{I}}W_{t}^{j}\overline{H}_{t}^{j}-\sum_{j\in\mathcal{I}}P_{t}^{j}(1+\kappa_{t}^{j}\delta_{t})\overline{C}_{t}^{j}=\sum_{j\in\mathcal{I}}W_{t}^{j}\widehat{H}_{t}^{j}-\sum_{j\in\mathcal{I}}P_{t}^{j}(1+\kappa_{t}^{j}\delta_{t})\widehat{C}_{t}^{j}.$
This allows us to conclude that $(\overline{{C}},\overline{{H}})\in\mathscr{A}$.
We have, by optimality of $(\widehat{{C}},\widehat{{H}})$,
$\mathcal{J}(\widehat{{C}},\widehat{{H}})-\mathcal{J}(\overline{{C}},\overline{{H}})=\sum_{j\in\mathcal{I}}\mathcal{J}_{j}(\widehat{{C}}^{j},\widehat{{H}}^{j})-\sum_{j\in\mathcal{I}}\mathcal{J}_{j}(\overline{{C}}^{j},\overline{{H}}^{j})\geq 0.$
However, for all $(t,j)\neq(s,i)$, $\overline{{C}}^{j}_{t}=\widehat{{C}}^{j}_{t}$ and $\overline{{H}}^{j}_{t}=\widehat{{H}}^{j}_{t}$, then
$\mathbb{E}\left[\beta^{s}U(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})\right]-\mathbb{E}\left[\beta^{s}U\left(\widehat{{C}}_{s}^{i}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}},\widehat{{H}}_{s}^{i}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}}\frac{P_{s}^{i}(1+\kappa_{s}^{i}\delta_{s})}{W_{s}^{i}}\right)\right]\geq 0,$
i.e.
$\frac{1}{h}\mathbb{E}\left[U(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})-U\left(\widehat{{C}}_{s}^{i}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}},\widehat{{H}}_{s}^{i}+\eta h\theta^{(i,s)}\mathbf{1}_{A^{s}}\frac{P_{s}^{i}(1+\kappa_{s}^{i}\delta_{s})}{W_{s}^{i}}\right)\right]\geq 0.$
Letting $h$ tend to 00, we obtain
$\mathbb{E}\left[\eta\theta^{(i,s)}\mathbf{1}_{A^{s}}\frac{\partial U}{\partial x}(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})+\eta\theta^{(i,s)}\mathbf{1}_{A^{s}}\frac{P_{s}^{i}(1+\kappa_{s}^{i}\delta_{s})}{W_{s}^{i}}\frac{\partial U}{\partial y}(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})\right]\geq 0.$
Since the above holds for all $A^{s}\in\mathcal{G}_{s}$, $\eta=\pm 1$ and since $\theta^{(i,s)}>0$, then
$\frac{\partial U}{\partial x}(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})+\frac{P_{s}^{i}(1+\kappa_{s}^{i}\delta_{s})}{W_{s}^{i}}\frac{\partial U}{\partial y}(\widehat{{C}}_{s}^{i},\widehat{{H}}_{s}^{i})=0,$
leading to (B.9a).
For $j\in\mathcal{I}\setminus\{i\}$ and $\theta^{(i,j,s)}:=\frac{1}{2}\left(1\wedge\frac{{P}^{j}_{s}(1+\kappa_{s}^{j}\delta_{s})}{{P}^{i}_{s}(1+\kappa_{s}^{i}\delta_{s})}\right)(1\wedge\widehat{C}^{i}_{s}\wedge\widehat{C}^{j}_{s})>0$, setting now
$\overline{{C}}:=\widehat{{C}}+\eta h\mathbf{1}_{A^{s}}\theta^{(i,j,s)}\left(\Delta^{(i,s)}-\Delta^{(j,s)}\frac{{P}^{i}(1+\kappa^{i}\delta_{s})}{{P}^{j}(1+\kappa^{j}\delta_{s})}\right)\quad\text{and}\quad\overline{{H}}:=\widehat{{H}},$
and using similar arguments as above, we obtain (B.9b).
When $\sigma=1$, we carry out an analogous proof. ∎
B.3 Markets equilibrium
We now consider that firms and households interact on the labor and goods markets.
Definition B.5.
A market equilibrium is a $\mathbb{G}$-adapted positive random process $(\overline{W},\overline{P})$ such that
  1. 1.
    Condition (B.6) holds true for $(\overline{W},\overline{P})$.
  2. 2.
    The goods’ and labor’s market clearing conditions are met, namely, for each sector $i\in\mathcal{I}$, and for all $t\geq 0$,
    $Y_{t}^{i}=C_{t}^{i}+\sum_{j\in\mathcal{I}}Z_{t}^{ij}\qquad\text{and}\qquad H_{t}^{i}=N_{t}^{i},$ (B.11)
    where $N_{t}=\overline{n}(A_{t},\overline{W}_{t},\overline{P}_{t},\kappa_{t},\zeta_{t})$, $Z_{t}=\overline{z}(A_{t},\overline{W}_{t},\overline{P}_{t},\kappa_{t},\zeta_{t})$, $Y=F_{A}(N,Z)$ with $(\overline{n},\overline{z})$ an admissible solution (B.5a)-(B.5b) to (B.2), from Proposition B.3 while $C$ and $H$ satisfy (B.9a)-(B.9b) for $(\overline{W},\overline{P})$.
In the case of the existence of a market equilibrium, we can derive equations that must be satisfied by the output production process $Y$ and the consumption process $C$.
Proposition B.6.
Assume that there exists a market equilibrium as in Definition B.5. Then, for $t\geq 0$, $i\in\mathcal{I}$, it must hold that
$\left\{\begin{array}[]{rl}Y_{t}^{i}&=\displaystyle C_{t}^{i}+\sum_{j\in\mathcal{I}}\Lambda^{ij}(\mathfrak{d}_{t})\left(\frac{C_{t}^{j}}{C_{t}^{i}}\right)^{-\sigma}Y_{t}^{j},\\ Y_{t}^{i}&=\displaystyle A^{i}_{t}\left[\Psi^{i}(\mathfrak{d}_{t})(C_{t}^{i})^{-\sigma}Y_{t}^{i}\right]^{\frac{\psi^{i}}{1+\varphi}}\prod_{j\in\mathcal{I}}\left[\Lambda^{ji}(\mathfrak{d}_{t})\left(\frac{C_{t}^{i}}{C_{t}^{j}}\right)^{-\sigma}Y_{t}^{i}\right]^{\boldsymbol{\lambda}^{ji}},\end{array}\right.$ (B.12)
where $\Psi$ and $\Lambda$ are defined in (3.11), and $\mathfrak{d}_{t}$ is defined in (3.8).
Proof.
Let $i,j\in\mathcal{I}$ and $t\geq 0$. Combining Proposition B.3 and Proposition B.4, we obtain
$Z_{t}^{ji}=\boldsymbol{\lambda}^{ji}\frac{1-\tau_{t}^{i}\delta_{t}}{1+\zeta^{ji}_{t}\delta_{t}}\frac{1+\kappa_{t}^{j}\delta_{t}}{1+\kappa_{t}^{i}\delta_{t}}\left(\frac{C_{t}^{i}}{C_{t}^{j}}\right)^{-\sigma}Y_{t}^{i}.$ (B.13)
From Propositions B.3 and B.4 again, we also have
$N_{t}^{i}=\psi^{i}\frac{1-\tau_{t}^{i}\delta_{t}}{1+\kappa_{t}^{i}\delta_{t}}(H_{t}^{i})^{-\varphi}(C_{t}^{i})^{-\sigma}Y_{t}^{i}.$
The labor market clearing condition in Definition B.5 yields
$N_{t}^{i}=\left[\psi^{i}\frac{1-\tau_{t}^{i}\delta_{t}}{1+\kappa_{t}^{i}\delta_{t}}(C_{t}^{i})^{-\sigma}Y_{t}^{i}\right]^{\frac{1}{1+\varphi}}.$ (B.14)
Then, by inserting the expression of $N_{t}^{i}$ given in (B.14)and $Z_{t}^{ji}$ given in (B.13) into the production function $F$, we obtain the second equation in (B.12). The first equation in (B.12) is obtained by combining the market clearing condition with (B.13) (at index $(i,j)$ instead of $(j,i)$). ∎
B.4 Output and consumption dynamics and associated growth
For each time $t\geq 0$ and noise realization, the system (B.12) is nonlinear with $2I$ equations and $2I$ variables, and its well-posedness is hence relatively involved. Moreover, it is computationally heavy to solve this system for each price trajectory and productivity scenario. We thus consider a special value for the parameter $\sigma$ which allows to derive a unique solution in closed form. From now on, and following (Golosov et al., , 2014, page 63), we assume that $\sigma=1$, namely $U(x,y):=\log(x)-\frac{y^{1+\varphi}}{1+\varphi}$ on $(0,\infty)^{2}$.
Theorem B.7.
Assume that
  1. 1.
    $\sigma=1$,
  2. 2.
    $\mathbf{I}_{I}-\boldsymbol{\lambda}$ is not singular,
  3. 3.
    $\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{t})^{\top}$ is not singular for all $t\in\mathbb{R}_{+}$.
Then for all $t\geq 0$, there exists a unique $({C}_{t},{Y}_{t})$ satisfying (B.12). Moreover, with $\mathfrak{e}_{t}^{i}:=\frac{Y^{i}_{t}}{C^{i}_{t}}$ for $i\in\mathcal{I}$, we have
$\mathfrak{e}_{t}=\mathfrak{e}(\mathfrak{d}_{t}):=(\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{t})^{\top})^{-1}\mathbf{1},$ (B.15)
and using $\mathcal{B}_{t}=(\mathcal{B}^{i}_{t})_{i\in\mathcal{I}}:=\left[\mathcal{A}_{t}^{i}+\mathfrak{v}^{i}(\mathfrak{d}_{t})\right]_{i\in\mathcal{I}}$ with
$\mathfrak{v}^{i}(\mathfrak{d}_{t}):=\log\left((\mathfrak{e}_{t}^{i})^{-\frac{\varphi\psi^{i}}{1+\varphi}}\left(\Psi^{i}(\mathfrak{d}_{t})\right)^{\frac{\psi^{i}}{1+\varphi}}\prod_{j\in\mathcal{I}}\left(\Lambda^{ji}(\mathfrak{d}_{t})\right)^{\boldsymbol{\lambda}^{ji}}\right),$ (B.16)
we obtain
${C}_{t}=\exp{\left((\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}{\mathcal{B}_{t}}\right)}.$ (B.17)
Proof.
Let $t\geq 0$. When $\sigma=1$, the system (B.12) becomes for all $i\in\mathcal{I}$,
$\left\{\begin{array}[]{rl}Y_{t}^{i}&=\displaystyle C_{t}^{i}+\sum_{j\in\mathcal{I}}\Lambda^{ij}(\mathfrak{d}_{t})\left(\frac{C_{t}^{i}}{C_{t}^{j}}\right)Y_{t}^{j},\\ Y_{t}^{i}&=\displaystyle A_{t}^{i}\left[\Psi^{i}(\mathfrak{d}_{t})\mathfrak{e}^{i}_{t}\right]^{\frac{\psi^{i}}{1+\varphi}}\prod_{j\in\mathcal{I}}\left[\Lambda^{ji}(\mathfrak{d}_{t})C_{t}^{j}\mathfrak{e}_{t}^{i}\right]^{\boldsymbol{\lambda}^{ji}}.\end{array}\right.$ (B.20)
For any $i\in\mathcal{I}$, dividing the first equation in (B.20) by $C_{t}^{i}$, we get
$\mathfrak{e}^{i}_{t}=1+\sum_{j\in\mathcal{I}}\Lambda^{ij}(\mathfrak{d}_{t})\mathfrak{e}^{j}_{t},$
which corresponds to (B.15), thanks to (3.10). Using $\sum_{j\in\mathcal{I}}\boldsymbol{\lambda}^{ji}=1-\psi^{i}$ and $Y_{t}^{i}=\mathfrak{e}_{t}^{i}C_{t}^{i}$ in the second equation in (B.20), we compute
$C_{t}^{i}=A_{t}^{i}(\mathfrak{e}_{t}^{i})^{-\frac{\varphi\psi^{i}}{1+\varphi}}\left[\Psi^{i}(\mathfrak{d}_{t})\right]^{\frac{\psi^{i}}{1+\varphi}}\prod_{j\in\mathcal{I}}\left[\Lambda^{ji}(\mathfrak{d}_{t})\right]^{\boldsymbol{\lambda}^{ji}}\prod_{j\in\mathcal{I}}(C_{t}^{j})^{\boldsymbol{\lambda}^{ji}}.$
Applying log and writing in matrix form, we obtain $(\mathbf{I}_{I}-\boldsymbol{\lambda})\log({C}_{t})={\mathcal{B}_{t}}$, implying (B.17). ∎
Remark B.8.
The matrix $\boldsymbol{\lambda}$ is generally not diagonal, and therefore, from (B.17), the sectors (in output and in consumption) are linked to each other through their respective productivity process. Similarly, an introduction of price in one sector affects the other ones.
Remark B.9.
For any $t\geq 0$, $i\in\mathcal{I}$, we observe that
$\mathcal{B}^{i}_{t}=\mathcal{A}^{i}_{t}+v^{i}(\mathfrak{d}_{t}),$ (B.21)
where $v^{i}(\cdot)$ is defined using (B.16). Namely, $\mathcal{B}_{t}$ is the sum of the (random) productivity term and a term involving the price. The economy is therefore subject to fluctuations of two different natures: the first one comes from the productivity process while the second one comes from the price processes.
We now look at the dynamics of production and consumption growth.
Theorem B.10.
For any $t\geq 0$ and for $\varpi\in\{{Y},{C}\}$. With the same assumptions as in Theorem B.7,
$\mathrm{d}\log{\varpi}_{t}\sim\mathcal{N}\left({m}_{t}^{\varpi},\widehat{\Sigma}_{t}\right),\qquad\text{for }\varpi\in\{{Y},{C}\},$ (B.22)
with
$\widehat{\Sigma}_{t}$ $=\varsigma^{2}(\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}\overline{\Sigma}(\mathbf{I}_{I}-\boldsymbol{\lambda}^{\top})^{-1}(\mathrm{d}t)^{2},$ (B.23)
$m^{C}_{t}$ $=(I-\boldsymbol{\lambda})^{-1}\left[\mu\mathrm{d}t+\mathrm{d}\mathfrak{v}(\mathfrak{d}_{t})\right],$ (B.24)
$m^{Y}_{t}$ $=(I-\boldsymbol{\lambda})^{-1}\left[\mu\mathrm{d}t+\mathrm{d}\mathfrak{v}(\mathfrak{d}_{t})\right],$ (B.25)
and
$v(\mathfrak{d}_{t}):=\mathfrak{v}(\mathfrak{d}_{t})+(\mathbf{I}_{I}-\boldsymbol{\lambda})\log(\mathfrak{e}(\mathfrak{d}_{t})),$ (B.26)
where $\overline{\mu}$ and $\varsigma^{2}\overline{\Sigma}$ are the mean and the variance of the stationary process $\mathcal{Z}$ (Remark 3.2), $v$ is defined in (B.16) and $\mathfrak{e}$ in (B.15).
Proof.
Let $t\geq 0^{*}$, from (B.21), we have, for $i\in\mathcal{I}$,
$\mathrm{d}\mathcal{B}^{i}_{t}=(\mu^{i}+\varsigma\mathcal{Z}_{t}^{i})\mathrm{d}t+\mathrm{d}v^{i}(\mathfrak{d}_{t}).$
Combining the previous equality with (B.17), we get
$\mathrm{d}\log{{C}}_{t}=(\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}\left[(\mu+\varsigma\mathcal{Z}_{t})\mathrm{d}t+\mathrm{d}v(\mathfrak{d}_{t})\right].$ (B.27)
Applying Remark 3.2 leads to $\mathrm{d}\log{{C}}_{t}\sim\mathcal{N}\left({m}_{t}^{{C}},\widehat{\Sigma}_{t}\right)$. Using (B.15), we observe that, for $i\in\mathcal{I}$,
$(\mathrm{d}\log{{Y}_{t}})^{i}=(\mathrm{d}\log{{C}_{t}})^{i}+\mathrm{d}\log(\mathfrak{e}^{i}(\mathfrak{d}_{t})),$ (B.28)
which, using the previous characterization of the law of $\mathrm{d}\log{{C}_{t}}$, allows to conclude. ∎
From the previous result, we observe that output and consumption growth processes have a stationary variance but a time-dependent mean.
Proof.
of Proposition 3.7.
Let $t\geq 0$, $n\in\{1,\ldots,N\}$, and $T>t_{\star}$.
  1. 1.
    we also introduce,
    $\mathcal{V}^{n,K}_{t,\mathfrak{d}}:=F^{n}_{t,\mathfrak{d}}\int_{t}^{+\infty}e^{-r(s-t)}\mathbb{E}_{t}\left[\exp\left((s-t)\mathfrak{a}^{n\cdot}\mu+\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)+\sigma_{n}(\mathcal{W}^{n}_{s}-\mathcal{W}^{n}_{t})\right)\mathrm{d}s\right].$ (B.29)
    Similar computations as (in fact easier than) the ones performed in the proof of Proposition 2.5. in Bouveret et al., (2025) show that $\mathcal{V}^{n}_{t,\mathfrak{d}}=\lim_{K\to+\infty}\mathcal{V}^{n,K}_{t}$ is well defined in $\mathcal{L}^{q}(\mathcal{H},\mathbb{E})$ for any $q\geq 1$. Furthermore,
    $\mathcal{V}^{n,K}_{t,\mathfrak{d}}=F^{n}_{t,\mathfrak{d}}\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{t+s})-v(\mathfrak{d}_{t})\right)\right)}\mathrm{d}s=F^{n}_{t,\mathfrak{d}}e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t})}\int_{s=0}^{K}e^{\varrho_{n}s}\exp{(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s}))}\mathrm{d}s,$
    where $\varrho_{n}$ is defined in the lemma, and from Assumption 3.6 and Corollary 3.5,
    $F^{n}_{t}=F^{n}_{0}\exp\left(\int_{u=0}^{t}\mathfrak{a}^{n\cdot}(\Theta_{u}\mathrm{d}u+\mathrm{d}v(\mathfrak{d}_{u}))+\sigma_{n}\mathrm{d}\mathcal{W}^{n}_{t}\mathrm{d}u\right)=F^{n}_{0}e^{\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{t})-v(\mathfrak{d}_{0})\right)}\exp\left(\mathfrak{a}^{n\cdot}\mathcal{A}^{\circ}_{t}+\sigma_{n}\mathcal{W}^{n}_{t}\right).$
    We then have
    $F^{n}_{t,\mathfrak{d}}e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t})}\int_{s=0}^{K}e^{\varrho_{n}s}\exp{(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s}))}\mathrm{d}s=F^{n}_{0}e^{-\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{0})}\exp\left(\mathfrak{a}^{n\cdot}\mathcal{A}^{\circ}_{t}+\sigma_{n}\mathcal{W}^{n}_{t}\right)\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s.$
  2. 2.
    Moreover,
    • (a)
      If $t<t_{\circ}$, then
      $\begin{split}\mathfrak{R}^{n,K}_{t}(\mathfrak{d})&:=\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s\\ &=\int_{s=0}^{t_{\circ}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s+\int_{s=t_{\circ}-t}^{t_{\star}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s+\int_{s=t_{\star}-t}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s\\ &=e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\circ}})}\frac{1-e^{\varrho_{n}(t_{\circ}-t)}}{-\varrho_{n}}+\int_{s=t_{\circ}-t}^{t_{\star}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s+e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})+\varrho_{n}(t_{\star}-t)}\frac{1-e^{\varrho_{n}(K-t_{\star}+t)}}{-\varrho_{n}}.\end{split}$
    • (b)
      If $t_{\circ}\leq t<t_{\star}$, then
      $\begin{split}\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s&=\int_{s=0}^{t_{\star}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s+\int_{s=t_{\star}-t+1}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s\\ &=\int_{s=0}^{t_{\star}-t}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s+e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})+\varrho_{n}(t_{\star}-t+1)}\frac{1-e^{\varrho_{n}(K-t_{\star}+t)}}{-\varrho_{n}}.\end{split}$
    • (c)
      If $t\geq t_{\star}$, then
      $\begin{split}\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t+s})\right)}\mathrm{d}s&=\int_{s=0}^{K}e^{\varrho_{n}s}\exp{\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})\right)}\mathrm{d}s=e^{\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{t_{\star}})}\frac{1-e^{\varrho_{n}(K+1)}}{-\varrho_{n}}.\end{split}$
    Finally, $e^{\varrho_{n}(K+1)}$ and $e^{\varrho_{n}(K-t_{\star}+t)}$ converge to 00 for $\varrho_{n}<0$ as $K$ tends to infinity, and the result follows.
  3. 3.
    We denote
    $V^{n,T}_{t,\mathfrak{d}}:=\mathbb{E}_{t}\left[\int_{t}^{T}e^{-r(s-t)}F^{n}_{s,\mathfrak{d}}\mathrm{d}s\right].$
    As we have from (3.15), $F^{n}_{s,\mathfrak{d}}=F^{n}_{t,\mathfrak{d}}\exp\left(\mathfrak{a}^{n\cdot}(\mathcal{A}_{s}-\mathcal{A}_{t})+\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)+\sigma_{n}(\mathcal{W}^{n}_{s}-\mathcal{W}^{n}_{t})\right)$, and given that for all $h,t\geq 0$,
    $\mathcal{A}_{t+h}=\mathcal{A}_{t}+\mu h+\varsigma\Upsilon_{h}\mathcal{Z}_{t}-\varsigma\Gamma^{-1}\int_{t}^{t+h}\left(e^{-\Gamma(t+h-s)}-\mathbf{I}_{I}\right)\Sigma\mathrm{d}B_{s}^{\mathcal{Z}}.$
    We obtain
    $V^{n,T}_{t,\mathfrak{d}}$ $=\mathbb{E}_{t}\left[\int_{t}^{T}e^{-r(s-t)}F^{n}_{t}\exp\left(\mathfrak{a}^{n\cdot}(\mathcal{A}_{s}-\mathcal{A}_{t})+\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)+\sigma_{n}(\mathcal{W}^{n}_{s}-\mathcal{W}^{n}_{t})\right)\mathrm{d}s\right]$
    $=F^{n}_{t,\mathfrak{d}}\int_{t}^{T}e^{\left(\frac{1}{2}\sigma_{n}^{2}-r\right)(s-t)}\exp\left(\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)\right)\mathbb{E}_{t}\left[\exp\left(\mathfrak{a}^{n\cdot}(\mathcal{A}_{s}-\mathcal{A}_{t})\right)\right]\mathrm{d}s$
    $=F^{n}_{t,\mathfrak{d}}\int_{t}^{T}e^{\left(\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu-r\right)(s-t)}\exp\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)\exp\left(\varsigma\mathfrak{a}^{n\cdot}\Upsilon_{s-t}\mathcal{Z}_{t}+\frac{1}{2}\mathfrak{a}^{n\cdot}\Sigma^{\mathcal{A},h}_{t}(\mathfrak{a}^{n\cdot})^{\top}\right)\mathrm{d}s.$
    Then using Hölder’s inequality (with $1=\frac{1}{p}+\frac{1}{q}$), we have
    $\lVert V^{n,T}_{t,\mathfrak{d}}\rVert_{1}$ $\leq\lVert F^{n}_{t,\mathfrak{d}}\rVert_{q}\left\lVert\int_{t}^{T}e^{\left(\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu-r\right)(s-t)}\exp\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)\exp\left(\varsigma\mathfrak{a}^{n\cdot}\Upsilon_{s-t}\mathcal{Z}_{t}+\frac{1}{2}\mathfrak{a}^{n\cdot}\Sigma^{\mathcal{A},s-t}_{t}(\mathfrak{a}^{n\cdot})^{\top}\right)\mathrm{d}s\right\rVert_{p}$
    $\leq\lVert F^{n}_{t,\mathfrak{d}}\rVert_{q}\int_{t}^{T}e^{\left(\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu-r\right)(s-t)}\exp\left(\mathfrak{a}^{n\cdot}v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)\exp\left(\frac{1}{2}\mathfrak{a}^{n\cdot}\Sigma^{\mathcal{A},s-t}_{t}(\mathfrak{a}^{n\cdot})^{\top}\right)\left\lVert\exp\left(\varsigma\mathfrak{a}^{n\cdot}\Upsilon_{s-t}\mathcal{Z}_{t}\right)\right\rVert_{p}\mathrm{d}s.$
    Observe that under Assumption 3.3, there exists a constant $\mathfrak{C}_{\mathfrak{d}}>0$ such that
    $\sup_{n,s,t}\exp\left(\mathfrak{a}^{n\cdot}\left(v(\mathfrak{d}_{s})-v(\mathfrak{d}_{t})\right)\right)\leq\mathfrak{C}_{\mathfrak{d}}\,.$
    Given that $\mathcal{Z}$ is stationary and $\Upsilon_{s-t}$ is bounded ((A.1)), there exists $\mathfrak{C}_{n,p}>0$ so that $\leq\mathfrak{C}_{n,p}$
    $\left\lVert\exp\left(\varsigma\mathfrak{a}^{n\cdot}\Upsilon_{s-t}\mathcal{Z}_{t}\right)\right\rVert_{p}=\mathbb{E}\left[\exp\left(\varsigma p\mathfrak{a}^{n\cdot}\Upsilon_{s-t}\mathcal{Z}_{t}\right)\right]^{\frac{1}{p}}\leq\mathfrak{C}_{n,p}.$
    Moreover,
    $\exp\left(\frac{1}{2}\mathfrak{a}^{n\cdot}\Sigma^{\mathcal{A},h}_{t}(\mathfrak{a}^{n\cdot})^{\top}\right)$ $=\exp\left(\frac{1}{2}\varsigma^{2}\int_{0}^{s-t}\mathfrak{a}^{n\cdot}\Upsilon_{u}\Sigma\Sigma^{\top}\Upsilon_{u}^{\top}(\mathfrak{a}^{n\cdot})^{\top}\mathrm{d}u\right)$
    $\leq\exp\left(\frac{1}{2}\varsigma^{2}\int_{0}^{s-t}\lVert\mathfrak{a}^{n\cdot}\rVert^{2}\lVert\Sigma\rVert^{2}\lVert\Upsilon_{u}\rVert^{2}\mathrm{d}u\right)$
    $\leq\exp\left(\frac{1}{2}\varsigma^{2}\frac{c_{\Gamma}^{2}}{\lambda_{\Gamma}^{2}}\lVert\mathfrak{a}^{n\cdot}\rVert^{2}\lVert\Sigma\rVert^{2}(s-t)\right).$
    Next, we can write
    $\lVert V^{n,T}_{t,\mathfrak{d}}\rVert_{1}$ $\leq\mathfrak{C}_{\mathfrak{d}}\mathfrak{C}_{n,p}\lVert F^{n}_{t,\mathfrak{d}}\rVert_{q}\int_{t}^{T}\exp{\left(\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu+\frac{1}{2}\varsigma^{2}\frac{c_{\Gamma}^{2}}{\lambda_{\Gamma}^{2}}\lVert\mathfrak{a}^{n\cdot}\rVert^{2}\lVert\Sigma\rVert^{2}-r\right)(s-t)}\mathrm{d}s,$
    and if (3.21) is satisfied and $T\to+\infty$, then $V^{n,K}_{t,\mathfrak{d}}$ converges to $V^{n}_{t,\mathfrak{d}}$. Finally, similar methods must be used to show $\mathbb{E}\left[\left|\frac{V^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}-\frac{\mathcal{V}^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}\right|\right]\leq C\varsigma$.